A Selected Bibliography of Publications by, and about, Freeman J. Dyson

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

15 December 2023 Version 1.80

Title word cross-reference

(p, k) [AAL22]. 1/N [CDM05, FM21, KPS90]. 1/r^{2} [SLA93]. $11.50 [Dys67b]. $17 [Dys60a]. $18.00 [Dys98b]. $18.95 [Sta03]. $19.95 [Dys02c]. 2 [Meu09]. $20.00 [Dys05f]. $22.00 [Dys98b]. $22.95 [Dys03b]. $23.00 [Dys05e]. $23.95 [Dys03c]. $24.95 [Dys11o]. $25.00 [Dys05e, Dys04c, Dys06e, Dys15a]. $25.95 [Dys06d, Dys07f, Dys10d]. $26.00 [Ber02, Dys02b, Dys05l, Dys14a]. $26.95 [Dys08d, Dys03d, Dys09b]. $27.00 [Dys12f, Dys12d]. $27.50 [Dys05g]. $27.95 [Dys07e, Dys12g, Dys12e, Guh07, Hor06a]. $27.99 [Ben13]. $28.00 [Dys08d]. $28.95 [Dys04d]. $29.00 [Dys13b]. $29.95 [CK12, Dys08c, Dys10c]. $29.99 [Dys11o, Dys15c]. $3.50 [Dys58a]. $30.00 [Dys05f, Sta03, Dys11n, Dys11l]. $32.95 [Dys05h]. $34.95 [Dys96j]. $35.00 [Dys02a, Dys08b]. $37.50 [Dys13e]. $4.95 [Dys64c, Dys62b]. $40.00 [Dys09c]. $6.00 [Dys60c]. $6.80 [Dys60b]. $7.00 [Dys62a]. $8.50 [Dys70c]. $8.80 [Dys70a]. $80.00hb/$30.00pb [Cao06]. $9 [Dys64c]. ^{−} [DZO11]. _{2} [DZO11, MND11, MNLD11, TMD96a, TMD96c, TMD96b]. _{2}^{−} [GGMO07]. _{3} [BFW11]. _{4} [KSW13]. _{5} [GGMO07]. _{m}^{−} [GGMO07]. a [BS89]. β [FM21]. c [BS89]. Δ_{3} [GC81]. η [HKR07, KK02, KK98]. η′ [HKR07, KK02, KK98]. FS × W [GS85]. g − 2 [GFW11, GFW12]. G_{2} [Zei87]. γ − 3π [CM03]. γ^{∗}γ [KK02]. I = 1/2 [ST67]. J = (1/2)^{+} [ST67]. l_{1} [LEdAB03]. l_{2} [LEdAB03]. L_{2}, 3 [BS89]. λ(φ^{⋆}φ)^{2} [CLO10]. λφ^{4} [CMD04]. m = 3, 4, 5 [GGMO07]. N [Bui11, Wad81]. N/D [BW69, Dab75, HJ65a]. n = 4 [Kad85]. O(N) [CDM05, KPS90, Meu09]. p [SZ21]. P_{11} [BW69]. ϕ^{2}σ [CJ80]. πN [AH66, ST67]. πNP [Nak87]. ππ [AH66]. q [BG85, GX06, Kad85]. S [Dys49b, Dys58k, Dys03i]. s = 1/2 [Goc76]. σ [AK00]. SU(2) [sXcHmW87]. SU(3) [DG84]. SU(6) [DhX65a, DX64, DhX65b]. Y = 2 [DX64, DhX65b].

-Dyson [BG85, GX06, Kad85]. -Matrix [Dys49b, Dys58k, Dys03i]. -Model [AK00]. -Si [BS89, BS89]. -spin [SZ21].

. [Dys58b].

0 [Ber02, Dys96j, Dys97a, Dys02a, Dys02d, Pei80]. 0-06-011108-9 [Pei80]. 0-19-856479-1 [Dys02d]. 0-231-10138-4 [Dys96j]. 0-309-06987-4 [Sta03]. 0-387-94774-4 [Dys97a]. 0-465-01859-9 [Sta03]. 0-674-00433-7 [Sta03]. 0-691-03327-7 [Bro96c]. 0-691-03685-3 [Bro96c]. 0-691-07021-0 [Sta03]. 0-7382-0532-X [Dys02a]. 0-8050-5985-7 [Ber02]. 0-8176-4030-4 [Sta03].

1 [Dys02d]. 100th [Ano04]. 11 [Dys89h]. 123rd [Kon90]. 12th [Mon94]. 14.99/$25.00 [Ber04]. 16th [Dys72n]. 17 [Aas86]. 18th [FR94, FR02]. 1943 [Dys05f]. 1945 [Dys05f, She82]. 1950s [Mah11]. 1954 [Mey56]. 1962 [Dys62f, Ros63]. 1963 [Dys63f]. 1966 [MB66]. 1972 [Dys72n, Dys74a]. 1978 [JR80]. 1981 [Ano82]. 1985 [Dys85a, Dys88c, Dys90h, Dys04f]. 1986 [Bye87]. 1987 [AAB^{+}88]. 1989 [CM90]. 1990 [Kon90]. 1991 [Dys91i]. 1994 [Ano94]. 1996 [Ano97]. 1998 [Dys11t]. 1999 [GSS^{+}03]. 19th [FR94, FR02].

2 [UM86, Dys63b]. 2000 [Ano00a]. 2003 [Dys06b, Dys09d]. 2007 [PSP^{+}07]. 2008 [Dys09a, Dys10h]. 2010 [KLR13]. 2012 [Jen14]. 2013 [PKpCH14]. 2014 [Jon16]. 2020 [Dys22, Joh20b]. 20th [Str00, Dys10e, Dys11p]. 20th-Century [Dys10e, Dys11p]. 21st [Dys95c, Dys95k, Kid89]. 21st-Century [Dys95c, Dys95k]. 240pp [Ber04]. 25 [Ano01]. 28 [Joh20b]. 28th [Ano20d]. 2nd [Whi80].

3 [Dys83b]. 3-7643-5312-0 [Sta03]. 3-butadiene [DK06, SWFB05]. 3. [Dys79a]. 305 [Dys96i].

4 [Dys96j, Dys97a]. 4.0 [BHJ^{+}00]. 4.0/Buchtipp [BHJ^{+}00]. 423pp [CK12]. 4385-4389 [UTB^{+}50]. 469pp [Cao06]. 480pp [Ber04]. 4th [Tie09a, Tie09b]. 4th-Grade [Tie09a]. 4th-Grader [Tie09b].

5781 [DLV10]. 58 [And91, Bre91, Dom91, Far91].

60th [MB66]. 61 [HMSW00a]. 63 [KTNG23]. 68 [BD95].

7 [Ber02].

8 [Ben13]. 80th [Ano04]. 83 [GFW12]. 86 [DORQ12b].

9 [Dys10h, Pei80]. 90c [Dys89h]. 90th [Bus15, PKpCH14, Ano04]. 96a [BD95]. 978 [Ben13]. 978-0-312-64235-8 [Ben13]. ’99 [Dys00a].

A-Bombs [Dre00]. Aalborg [Jen14]. AAPT [Ano91]. Abdus [Dys99b]. Aberdeen [Dys88c, Dys90h, Dys04f]. Abert [Dys85a]. Abhandlugen [Dys70d]. ability [Lam06]. Ablowitz [Spo22]. Abolition [Dys86b, Dys84a]. Absorption [Dys55c]. Abstract [Dys62b]. Abstracts [Dys00a]. abundance [BM83]. Abundances [DW48]. Academic [Dys60b, Dys67b]. Académie [Ham91]. Academy [Whi80, Ham91]. Acceleration [Hom96]. Acceptance [Dys77a, Dys85i, Dys91i]. account [Alp73, HAS10]. Accreting [RRD^{+}73]. accuracy [GMO99]. Accurate [GMO98, GMO03]. achievements [Woo80]. Acrobat [BHJ^{+}00]. Acta [BD95]. Acting [Dys75b]. Active [Dys75c]. Ad [Dys95f]. Adams [Dys72l]. adapted [MND11, MNLD11]. adaptive [BM83]. Added [Dys98e]. Addendum [Dys96a]. Addison [Dys98b]. Addison-Wesley [Dys98b]. additivity [RY23]. adenine [SOM10]. adequate [Yeu80]. administrator [Dys06k]. admolecule [Ye92]. Adobe [BHJ^{+}00]. Advanced [Bat50, Bat51, BTD^{+}47, BS51, BG49, DC52, Dys50a, DTS51, Dys51a, Dys52b, DM57b, Dys06c, Dys07a, Dys11k, EW50, ES51, FZ52, Gal50, Gal52, GB48, Mit80, MS50, RRS48, San50, Thé47, TL48, TB49, Thé50, TK51a, TK51b, UTB^{+}50, UL51, Wil48, WG50, Wil50, WL51, Dys54e]. adventures [FL06, Dys84p]. Advisors [Dys76a]. aesthetic [Cur82]. Aether [Dys54d]. Affairs [GR63]. After [Dys81c, Dys95d, Joh20a]. Against [Dys64d, Dys64f, For02, Sta03]. Age [Dys98b, Dys05g, Dys09c, GR63, Bro15, Dys82a, Dys84a]. Agee [Dys05f]. Agenda [Dys85a, Fre84, Kid89, ACN85]. AIP [Swe88]. Air [Dys08b]. Airplane [GD10]. Airy [AvM05]. Al [GGMO07]. Alamos [Dys92c, She82]. Alan [Dys05e]. Albert [BHJ^{+}00, Dys58a, Woo80]. Albuquerque [Dys79l]. Alfred [Has79]. Algebra [Per68]. Algebraic [HGE88, HR22, Par87, Dys47a, Dys48b, Dys62k]. algebras [GZ87, dJ94]. Algorithmic [AHS09, HB12]. Alice [Dys83c, She82, Sta03, Dys96m]. Allison [Dys85a]. Almost [FKS97]. along [GS22]. Alpha [Fin17]. Alternative [Dys12f, Dys12d]. Alternatives [Dys88b, Dys88a, GSE^{+}88, Mar88]. Alto [Dys70c]. Am [And91, Bre91, Dom91, Far91]. Amateurs [Dys02b]. Ambiguity [AW69, McL13]. America [Dys80k, Kev87]. American [Ber04, Don02, Dys63f, Dys82a, Dys91i, Ano99, ALW^{+}88, Bea99, Mcc94, Whe91]. Among [Ano20c]. Amount [Dys77b]. Amplitude [ST67, Nak87, WK74]. Amplitudes [AW69, Man78, MNM94]. analog [Kri08]. Analysis [Dys58c, Ros63, SG86, TMD96a, TMD96c, TMD96b, AI02, Dav92, Dys43c, HGE88, Hof11, HAS10, Par87, Woo92]. Analytic [BMS61]. Analytical [Tak89]. analyticity [Sim22]. analyzed [SG06]. Ancient [Don02]. Ando [CL22]. Andreev [GS04b, GS04a]. Andres [Dys05f]. Andrews [Kad85, Mah05]. angular [OK09]. anharmonic [Yeu80]. anion [CCC^{+}10]. Anisotropy [DÜ55, BFW11, RT82]. Ann [Dys83b, Hor06a, Sta03]. Annihilation [Pra64, DhX65a, Dys62e]. Anniversary [Ano04, ALW^{+}88, Rab78]. Annual [Dys70c]. anomalous [CM03]. Answers [Tie09b]. Anthology [BB63, Sta03]. Anthropic [Spe91]. Anti [Dys69i]. Anti-Missiles [Dys69i]. antiproton [DhX65a]. antiproton-proton [DhX65a]. Anton [Fis10, Fis12, Fis10, Fis12]. any [Dys79g, Dys80i, PD12]. Anything [Dys90l]. Apart [Cao06, Kai05a]. APN [Dys80b, Dys81a]. Apparent [LD66]. Appeals [WDW89]. Appendix [Dys63b]. Application [Mat49a, Mat49b, SS99, SS67, BS89, Wüt13]. Applications [Dys66a]. Applied [Dys60j, NHKK94, NMM^{+}94]. appreciation [Dys70a, Lie04]. Approach [Dys62b, ADM23, AFPT10, BCDM01, CLO10, EN12, EF12, EF13, Fed75, GS04b, GFW11, GFW12, HMSW00a, HMSW00b, HKR07, KK02, KK98, LACTFF22, LNA96, LFJ^{+}13, MEBK11, MNBW07, NAW08, NEA10, SS99, Wad81]. Approaches [Dys76e]. Approximate [Dys64b, tL84, DCCS94]. Approximation [DRS^{+}54, Dys07t, Dys11j, NO99, DvL05, Dys47a, HvSA98b, SRBW05]. approximations [Dys47b, MM90]. April [Dys88c, Dys90h, Dys04f, Dys10h, Kon90]. arbitrary [SZ21]. Archaeology [Don02]. Archibald [Dys04g, Dys10h]. arising [Pas22]. Arith [BD95]. Armageddon [Dys05f]. Armenia [Dys71h]. Arms [Dys71b, Dys73a]. Aronowitz [CQ99]. Arsenal [Dys62e, Dys84a]. Art [Dys95g, Dys13b, Ano95, Dys11o]. arte [Ano95]. Arthur [Sta03]. Article [Fai94, Kor94, Dys60c]. Articles [GR63]. Artificial [Dys60k, MASD60]. Aruri [ABG^{+}89, Dys89a, DW89, LD90, WDW89]. Asia [DGWW67]. Aspects [CLO10, SW72, Poo72]. Assessment [DCCS94]. Assignment [DZO13]. Assistant [Sta03]. Associate [Dys80b, Dys81a, Ham91]. Associates [Swe88]. Association [Dys91i, Whe91]. associés [Ham91]. Astronomical [Kon90]. Astronomy [Dys70c, Dys81m, Dys84c, Dys90j, Dys92i, Dys96l, FF91, Gin75, Mor58]. Astrophysics [Dys70c]. Asymptotic [JK67, JK68, Yeu80]. asymptotics [AAL22, EG11]. Athens [Dys80l]. Atlas [Dys62a]. Atmosphere [Dys77b, Dys90c, HMMD94, Dys60l]. Atmospheric [Dys75c, HMMD94]. Atmospheric-Turbulence [HMMD94]. Atom [Kra01, Opp89, Dys05e, Dys10c]. Atomic [Ber02, Blo04, Dys82a, Dys92c, GR63, Kan02, Dys02h, PVW02, PVW03, VPW01]. atomism [Sta03]. atoms [DvL05, LT91, LT97, LT01, LT05, Thi91, TL97, Thi01, Thi05]. August [Dys63f, Han82, Jen14, PKpCH14]. Australian [Whi80]. Austrianization [Dys85j]. Authorized [Dys60a]. authors [Dys60a]. Autobiography [Dys84p, Dys18]. auxiliary [CMD04]. average [sXcHmW87]. Avoiding [Dys70b, Dys85a, ACN85]. Award [Dys85i, Dys96b, Ano94, San94]. Awarded [Dys65e]. Awardees [Han82]. awards [Ano99]. Azeglio [Ano97].

B [Dys62a, Dys78b, San94]. Bach [Gre81]. Back [Ano17b, Dys07d, Dys11a, PSP^{+}07, GM12]. background [AB12, Can01]. Backscattering [HMMD94]. Backyard [Dys02b]. Bacon [Dys07e, BD07]. Badash [She82]. Badoz [Dys97a]. Baked [BB63]. Balance [Dys74c, Dys08d, Guh07]. Ballistic [Dys64d, Dys64f, Dys69n, McM65, Dys72l]. Ban [Dys59b, Dys63f]. Band [Dys69m, Dys70d, För96]. Bang [Spe91, Kra01, Spe91]. Banquet [Dys90a]. Barabashov [Dys62a]. Barbara [CQ99]. Bargain [LD08]. Bargmann [LSW76]. Barrow [Spe91]. Bart [Dys04c]. Bartusiak [Sta03]. baryon [NHKK94, NMM^{+}94]. baryons [MEBK11]. based [MNK^{+}97]. Basel [Sta03]. Basic [Dys05g, Dys05l, Dys09b, Dys10c, Dys15c, Sta03]. basis [Dir38, She89]. Batalin [dJ94]. Battle [Dys05f]. Battling [Dys13b]. Be [Dys97b, Dys13b, You69, Dys02d, Dys03g, Dys15b]. beans [Smi06]. Beat [SGT^{+}95]. Beaten [Dys05l]. Beaulieu [SGT^{+}95]. Beautiful [Dys80n, Dys83e]. Beauty [Dys09c, Sta03]. Became [Blo04, Joh20a]. Becchi [dJ94]. become [ALW^{+}88]. Been [Dys80c]. before [Dys63f]. began [Dys16]. Beginning [Dys11n]. beginnings [Dys87c]. begyndelser [Dys87c]. behalf [Dys63f, DLV10]. Behavior [Dys56f, HT95, Kon97, LFW10, Man78]. Being [Dys90g, Dys09b]. Belief [Dys98b]. Belknap [Dys10d]. benchmark [DK06]. Benefits [Bas11, Ehr75, Dys76c]. Bengt [Dys67a]. Benjamin [Dys64c]. Benoît [Dys78b]. Benson [Dys72l]. Berg [Dys87a]. Berlin [Dys58b, KLR13, Sta03, Dys03f]. Bermuda [DMZ76]. Bernal [Dys72n]. Berndt [Dys96h, Dys90k]. Bernstein [Gre81]. Best [Dys06e, Dys03g, Fey00, Pit11, Pit12]. Bethe [Dys55a, Dys56b, Dys05c, NEA10, BL06, Dys05d, Dys06j, HG87, KK02, KK98, Lee07, Ler71, MB66, Mun95, NO99, SA03]. between [Dys54c, Dys58d, Dys66b, Dys70e, GBB03, LD66, Neu16]. Beyond [Dys80d, Dys96f, Dys08d, Hof11, Bas22, Kon90, Kra01, Lew99]. bez [Dys83d]. Bi [Han82, CGR06]. bi-fermion [CGR06]. Bi-National [Han82]. bien [Ano95]. Bifurcation [CK94, Rem90]. Big [Spe91, Kra01]. bill [DLV10]. billiards [SRBW05]. Binding [Dys71c, Dys72c, GGMO07, McP74]. Biogenesis [Str00]. Biografie [BHJ^{+}00]. Biografie/Kuhn/Försterling [BHJ^{+}00]. biographical [Dys99b, Dys10h]. Biography [Dys91b, Dys02d]. Biohazards [Dys77e]. biologia [Dys83d]. biological [MMJ91]. Biologie [Dys89m, GBS91]. Biology [Gre81, Dys79m, Dys83d, Dys89m, Gre81, Dys87a]. Biosphere [Dys90c, Dys05h, Gol99, MASD60]. Biotech [BHMD07, Dys07r, FD07]. Bird [Pei85]. Birds [Dai10, Dys09a, Dys10b, Dys11b, Jon16]. Birkbeck [Dys72n]. Birkhaüser [Dys83b, Sta03]. Birmingham [PD13]. Birth [Ano04, Str00, ST11]. Birthday [Ano04, Bus15, DPSY64, MB66, PKpCH14]. bis [Fis10, Fis12]. bites [Dys96i]. Bitter [Dys05f, GD05]. bizarre [Ano20b, Rad20]. Black [Ove08, Spe91, Dys80d, ST11]. Blackett [Dys62e]. Bloch [Pra64]. block [DFPR22]. block-diagonalization [DFPR22]. blooded [Dys97h]. Blunder [TD14]. Blunders [Dys14a, KAD14]. body [HH13, MHFW72, NH91, NRT22, SS99]. Bogoliubov [Dys60a]. Boltzmann [CH99]. Bomb [Dys59a, Dys61b, Dys62j, Dys63g, Dys72f, Dys76a, Dys87d, Dys20, Dys92c]. Bomb-Propelled [Dys72f]. Bombing [Dys90l]. Bombs [Dre00, Dys83a]. Bonding [TMD96a, TMD96b, TMD96c]. Book [Ano80, Ano93, Ano96, Bai17, Bar82, Bar66, Ben97, Ben13, Ber19, Ber02, Bet79, Bro78a, Bro96b, Bro96c, Buc87, Cao06, CQ99, Cre01, Cre13, Don02, Dys55a, Dys55b, Dys56b, Dys58b, Dys58a, Dys58c, Dys60c, Dys60a, Dys60b, Dys62a, Dys62b, Dys62e, Dys64c, Dys67b, Dys70d, Dys70c, Dys70a, Dys74a, Dys79d, Dys80c, Dys80b, Dys81b, Dys81a, Dys82a, Dys83b, Dys83c, Dys84a, Dys84b, Dys85a, Dys85i, Dys87a, Dys89d, Dys90b, Dys91b, Dys91a, Dys92b, Dys96h, Dys96g, Dys96j, Dys96i, Dys96m, Dys97a, Dys98b, Dys98c, Dys02a, Dys02d, Dys02b, Dys02c, Dys03c, Dys03d, Dys03b, Dys03f, Dys04c, Dys04d, Dys05f, DG05, Dys05e, Dys05g, Dys05h, Dys06d, Dys06e, Dys07e, Dys07f, Dys08d, Dys08b, Dys08c, Dys09b, Dys09c, Dys10c, Dys10d, Dys11n, Dys11o, Dys11m, Dys11l, Dys12f, Dys12g, Dys13e, Fie79, Fre84, Gom84, Gre81, Guh07, Hag87]. Book [Has79, Hol81, Hor06a, Joh07a, Joh07b, Jon16, Kan02, Lew99, Nav99, Pie84, Rom89, Sax94, Sha87, Sha01, She82, Spe91, Sta03, Ste92, Str00, Tre94, Vil95, Way95, Wig90, dS99, vH84, Dys98d, Lar89]. Books [Ben13, BB63, CK12, Dys54b, Dys58a, Dys05g, Dys05l, Dys09b, Dys10c, Dys15c, Guh07, Sta03, Str00, Sta98]. Books/St [Ben13]. Born [Dys07t, Dys11j]. Borough [Dys77e]. Borsuk [Yan54, Yan55]. Bose [Bas22, NRT22]. boson [BF76, CGR06, DD96c, DG84, GL82, HGE88, HN08, KV85, tL84, Par87, RS77, SG86, She88, She89, SG93, Tak86, Tak88, Tam83, VHK89]. Boston [Dys83b, Sta03]. Both [WD15]. bound [Bas22, DD96a, DD96b, SA03]. boundary [WYL81]. Bounds [BD03, BF76]. Bouquet [Ber18, Ber19]. box [BFW11]. Brain [Gre81, Dys10a]. brane [AI02]. Braun [Dys08b, LD08]. Breach [Dys62b]. bread [Rog97]. Breakdown [Dys76b]. Breaking [DHD06, AI02, CK97, GS22, JRV22, LNA96, Mun95, RNM04, Rem90, Dys06d]. breakthrough [Joh20a]. Bremsstrahlung [DÜ55, MD57]. Bressoud [GX06]. Brian [Dys04d, DG05, Dys05e]. Bridging [Dys66b]. Brief [PD08, RQ11, Spe91, Dys02d]. Brilliant [Ano20b, Rad20, Dys14a]. Bringing [Ano00b]. Brittle [Dys80b, Dys81a]. Broca [Gre81]. Broch [Dys04c]. Broida [She82]. broken [CDM05, GMO00]. broken-symmetry [CDM05]. Brook [GSS^{+}03]. Brookings [Dys08d]. Brotherhood [Dys07e]. Brown [Dys13b, Ros63]. Brownian [ADvM09, Dys62c, NS93]. Brownian-Motion [Dys62c]. Bruce [Dys90k, Dys96h]. Brueckner [Ort04]. Bruno [Dys15c, Dys15c]. Buchtipp [BHJ^{+}00]. Budget [Bas11]. Budushhee [Dai71]. Build [Dys92c]. Building [KP09]. bulk [AI02]. Bulletin [GR63]. Business [LDB77, Noc08, Tie09c]. Bust [Dys91a]. butadiene [DK06, SWFB05]. Butterflies [Dys84q, Dys91c]. Buying [Ros62].

C [Ber84, Ber04, Bus15, Dys67b, Dys87a, Dys90k, Dys91b, Dys96h, Dys06d, TMD96a, TMD96c, TMD96b]. Calaprice [Sta03, Dys96m]. Calculation [BFM97, För96, LFJ^{+}13, MND11, RM06, CM03, DCCS94, IT01]. Calculations [LD66, Per68, Dab75, GMON98, GMO99, MNLD11, NMM^{+}94, She88]. calculus [GZ02, JL86]. Calentamiento [ZDSN08]. Calif [Dys70c]. California [Dys92c, Dys00a]. Cambrian [Str00]. Cambridge [Dys62a, Dys02a, Dys05e, JR80, Sta03]. Can [Dys77b, Dys97b, Dys12g, Dys12e, Dys00d]. CAN$44.95 [Sta03]. Canberra [Whi80]. canceling [Meu01]. Cancelled [Dys80c]. Candid [HH04a]. Candidate [Dys80b, Dys81a]. canoe [Bro78b, Bro78c, Bro78d, Bro96a, Broxx, Bro78a]. canonical [Dys12c]. Can’t [Noc08]. Can’t-Miss [Noc08]. Canty [BHJ^{+}00]. Carbon [Dys77b, Dys90c]. Care [Spe91]. Career [Ano17b, And13, Lie04]. Carl [Gre81, Sag00]. Carlo [Ano97, Ano97, Mar56, Mey56]. Carlos [CQ99]. Carnesale [Dys85a]. Carroll [Gre81]. Cartan [EJ22]. Cascade [DM57a, Kai05c]. Cascades [Dys57c]. Case [Dys69a, Dys11n, Dys14a, HR22, Pas22, Zei87]. Cassidy [Ber04]. Castillejo [AH66, AW69, BW69, Dab75, HJ65a, HJ65b, JK67, JK68, Kri10, McL13, NN86, Nak87, Roc71, ST67, WK74]. Cat [DF92]. Cathcart [Dys05e]. Cathedral [CK12, Dys05e]. causa [Ano97]. Causal [Dys58g, Dys58h]. Causes [Alb75, Sta03]. Caution [Dys59b]. cavities [TP01]. CB [Ano66]. celebrate [Woo80]. celebrates [ALW^{+}88]. Celebrating [Ano04]. Celebration [Rab78]. Cell [Gre81]. Centauri [Fin17]. centenary [AAB^{+}88, ST11]. centennial [Woo80]. Center [Dys08d, Dys13e]. Centuries [FR94, FR02]. Century [Dys02a, Sta03, Str00, BDF00, Ber04, Dys70i, Dys95c, Dys95k, Dys10e, Dys11p, Kid89, LY95]. cerca [Ano95]. CETI [Sag73]. Chain [Dys53d, Dys53e, Dys66c, FM21, For21]. chains [RNM04]. Challenged [Tay69, Joh20a]. Champaign [AAB^{+}88]. Chance [Dys78b]. Chandra [Dys91b]. Chandrasekhar [ST11, Dys10e, Dys11p, Dys91b]. Chandrasekharan [Dys70d]. Chang [Bus15, Bus15]. Change [Dys71b, Dys73a, Gam67a, GR09, Tel48, Bro10, Wil58, Dys05h]. Changed [Dys90l, Dys08c, Dys14a]. Changing [Whi80]. Channel [AH66]. chaos [DJS95]. Chapter [Dys63b]. character [FL06]. characteristics [RM06]. Characterizing [Dys78b]. Charge [Dys67d, Gam67b]. Charged [CDD56a, CDD56b, Dys67c]. Charles [Dys83c, Dys02d, She82, Dys94a]. Charpak [Dys04c]. Cheaper [Dys95j]. Checks [GMO03, GMO98]. Chemical [Dys71c, Dys72c, TMD96a, TMD96b, Wat48, TMD96c]. Chemistry [BHJ^{+}00]. Chemistry/Wiser [BHJ^{+}00]. Chen [Dys84b, LY95]. Chess [Dys97c]. Chicago [Cao06, Dys05f, Dys06e]. Chicken [BD07].

Chickens [Dre00]. children [Dys81k]. China [PSP^{+}07]. Chiral [FM09, HS98, BK10, GS04b, GS04a, LNA96, LEdAB03, Mun95, NO99, Rem90]. chiral-symmetry-breaking [Rem90]. Choose [Ano69]. chuan [Bro01]. Chueca [CQ99]. ci [Sta03]. Ciampi [Ano97, Ano97]. cibercultura [CQ99]. ciencia [Ano95]. científico [Ano95, Dys10f, Dys08e]. Circle [Dys85i, Dys92f, BCDL93, BD94c]. Circlons [Dys12f, Dys12d]. Citation [Dys86a]. Citizen [Dys98b]. City [Rab78]. civ [Sta03]. Civil [Daw09]. civilizacão [Fre92]. civilization [Fre92, KB89]. Civilizations [SW66]. Cl [DZO11]. Class [Dys71d, Dys72d, Dys90i, BLS03]. Classic [Dys86a, FL06, Dys06i]. Classical [Don02, Dys71c, Dys72c, BCDM01, EJ22, FM21]. classification [GS85]. Claus [CQ99]. Climate [GR09, Joh20a, Sol13]. Clocks [Dys03c]. Clockwork [Dys03c]. Close [Dys15c]. closed [PVW03]. closed- [PVW03]. closely [Ano95]. Closing [Dys66e]. cloth [Dys64c, Sta03]. Cloud [Dys69d]. cluster [MND11, MNLD11, OK07, OK09]. cluster-configuration [MNLD11]. clusters [DZO13]. Co [Dys85a]. coefficients [LEdAB03]. Coexistence [SWFB05]. Coherent [BS90]. Coined [Ove08]. Collapse [Dys69e, Dys76b, Dys65d]. collection [Cam63, Dys67b, Joh07b, SBZ21]. Collective [KD12, tL84]. College [Dys72n]. Collide [Dys88a, GSE^{+}88]. Collider [Dys88b]. collisional [YHKO07]. Colloquium [Han82, Kon90, UM86]. Colonists [Dys78e]. Colonizing [Dys89b]. Color [MNBW07, HMSW00a, HMSW00b, NAW08]. color-flavor-locked [NAW08]. Color-spin [MNBW07]. colored [Dys07q, Dys10i]. coloring [Dys11o]. Colossal [Dys14a]. Columbia [Dys96j, Rab78]. combinatory [Dys43c]. Comedy [Dys80p]. Comets [Dys91g, Dys92d]. Comment [And91, Bre91, Dom91, Dys80d, Far91, Fei03, GF04, HGE88, Pom09]. commentaries [Yan13]. commentary [Dys96e, Dys01d, Dys84b, Dys96h]. Comments [Dys61b, Dys69a, Dys69i]. Committee [Dys63f, Dys77e, Dys77g, Dys77e]. Common [Spe91, LDB77]. communication [Cam63, Dys10a, Sag73]. Communications [CM59]. Community [Dys77e, Kev87, Mit80, Opp89]. Commutativity [Dys58d]. Commutator [Dys58f, Dys60h]. Commutators [Dys58g, Dys58h, SS67]. como [Ano95]. Companions [Roc71]. Comparison [Dys45b, Dys78e, Dys07s, Dys11i, IT01, LD90, MM90, YHKO07]. Comparisons [IO08]. Compensation [HMMD94]. Completion [BMS61, Ano97]. Complex [Dys62g, Dys62h, Dys62i, DM63, MD63, MOR95, BBDY22]. complexity [AAL22, BDH04, LACTFF22, SZ21]. components [GBB03]. comprehension [Jud72]. Compton [EF13]. computer [CK12]. Computers [BHJ^{+}00, URR86]. Computers/Canty [BHJ^{+}00]. Computing [Mah11, MH11]. Concentration [SZ21]. Concept [Dys80m]. Conception [Dys74a]. Concepts [Dys84l, RD99]. Concern [Alb75]. Concerning [LCW^{+}72, Dys99a]. Concluding [Dys91d]. condensate [LFW10, MNK^{+}97]. Condition [CK97, Dys97c]. conditions [WYL81]. conductance [TP01]. cone [KL02]. Conference [BM55, Bus15, Bye87, Cur82, KLR13, AAB^{+}88, PKpCH14, RWP78]. Conferimenti [Ano97]. Conferment [Ano97]. Confidence [Ano54]. Configuration [TC92, MND11, MNLD11]. Configuration-interaction [TC92, MND11]. Confining [ERS07, MM97]. confluent [Dys60l]. conflux [Dys79f]. conformal [AAK^{+}12]. conformers [SWFB05]. Congress [Jen14]. Congruences [AO05, Mah05]. Conjecture [And80, Goo70, Gun62, SZ06, Wil62, Zei87, Kad85]. conjectures [GZ02]. Connection [Dys58d, Ano00b, Sag00]. conocio [Ano95]. conscience [Dys02d]. Conscious [Ben97, GBS91]. Consciousness [CCL^{+}11, JR80, JR80]. Consent [Mav75]. Consequences [Dys69g, Dys69h]. Conservative [Dys03e, Dys99a]. conserving [DvL05]. considerations [Dys68b]. Consilience [Dys98d]. consistency [CLO10, Mun95]. consistent [BEKW94, DvL05, MNBW07, PVW02, PVW03, SS99, VPW01]. Constance [Dys70a]. Constant [Bek86, BG85, AMN02, Alp73, DD96a, DD96b, Kon97, LFW10]. Constants [Gam68, Tel48, Wil58, Alp73, Dir37, Dys72a, Dys78a, Dys78f]. Construction [GS04a]. Containment [Dys59a]. contains [PD12]. Contestation [IT23]. context [BK10, VWS08]. Continued [Dys00f, Dys43b]. Continuous [BD59, Dys51b, GS22, vN98]. continuum [Rem90]. Contrarian [Gol08]. Contrasting [DZO13]. contributions [Cam63, CCC^{+}10]. Control [Dys71b, Dys73a, Dys77b]. Conventional [Dys62e]. Convergence [Hom96, Man77, Man82]. converging [Maa06]. Conversation [Dre00]. conversations [Dre01, HH04a, Tip10]. Conversion [BHJ^{+}00]. Conway [Dys05g]. cool [Dys98a]. Copenhagen [Dys07f]. Cornell [Dys51a, Dys51g, Dys89e, Dys93d]. Corp [Dys62a]. Corporate [Swe88]. Corrected [För96, SS98]. correction [PS04]. Correlation [För96, Guh91, LFJ^{+}13, Ort04, Sim22, SLA93, SS98]. Correlations [DJS95, Dys70e, CH99]. correspondence [Lee07, Neu16]. cosmic [Sag00]. Cosmo [GF04, GFD04]. Cosmological [Spe91, Dir37]. Cosmology [Alp73, CCL^{+}11, Spe91, BDH04, Dir38]. Cosmos [Dys04d, DG05, RD99, Gre81, Sha01]. Cost [Dys75b]. Costs [Dys74b, Dys76c, Dys77c]. could [Dys15b]. Coulomb [Dys63h, Dys71c, Dys72c, Dys95e, ERS07, Lew22, RQ11, WR07, WR10]. Council [Dys69c, Dys77e]. coupled [Det03, HN08, HvSA98a, KK02, KSW13, KK98, OK07, OK09]. coupled-cluster [OK07, OK09]. coupling [AMN02, Kon97, RS77, SG06, Tak89]. course [Dys51a, Dys51g]. covariant [CP93]. Craft [Dys98f]. Crank [AO05, Mah05]. CrasyDSE [HM12]. cream [Dys92e]. Creation [BH34, Gre81]. creators [Ano08]. Crichton [Dys03d, BHJ^{+}00]. Crisis [Olt73b, DARM83]. Criterion [Dys58a]. Crítica [CQ99]. Critical [BM87, HT95, GMO99, Zwa03, Dys98c]. Critics [Dys85i, Dys88a, Dys94a, GSE^{+}88]. critique [Lar89]. Cross [Dys54c, Dys55f, Dys55g, OK09, DCCS94]. Cross-sections [Dys54c].

Crossing [AW69]. Crossing-Symmetric [AW69]. Crossover [AF95, FKS97, RNM04]. crusade [Dys81k]. crystal [Ye92]. Crystals [DÜ55, HG87]. Cubic [Dys04a, BBDY22]. cuerda [DG05]. Culture [Joh07b, Lew99, For02, Sta03]. curious [FL06]. Current [Dys69a, Per68, LW78]. Current-Algebra [Per68]. curve [McP74]. cut [BBDY22]. Cuts [Bas11, HJ65a]. Cybernetics [Dys05g]. Cytosine [SPZ^{+}12]. Czas [Dys83d]. czółno [Bro78b].

D [DORQ12b, Dys60a, Dys72l, Dys72n, Dys08c, GFW12, HMSW00a, Spe91, Meu09]. Daedalus [Dro95, Dys95d]. dai [Bro01]. daily [Ano00b]. Dalitz [AH66, AW69, BW69, Dab75, HJ65a, HJ65b, JK67, JK68, Kri10, NN86, Nak87, Roc71, ST67, WK74, McL13]. Dancing [Gre81, Joh07a]. Dancoff [DD55, Dys53b, Dys53g, DRS^{+}54]. Dangers [Dys58a]. Daniel [Dys67b, Dys98c, Dys06d, Dys11l]. Danish [Dys87c]. dans [Dys09e]. Dark [Dys02b, Dys16, Dys05g]. Darwin [Dys14a]. Darwinian [Dys05i, Dys06f, Dys07b]. David [Ber04, Cao06, Dys70c, Dys11n, Dys13b]. Davies [Dys90b]. Day [Dys81c, Gre81]. DC [Sta03]. Dead [Ove08]. Dear [Neu16, Bai17]. Death [Dys65a, Dys82c, Dys05a, Hei80, Dys81b]. Debate [Str00]. Debunked [Dys04c]. Decade [Coh71]. decay [LFW10]. decays [MEBK11]. December [Aas86, Dys11t, Dys22, RWP78]. Decision [Dys76a]. decomposition [EJ22]. deconfinement [FM09]. decouplings [MNM94]. Découverte [Dys80k]. Dee [Ber04]. Deep [Dys02b, Gol99, Pas22]. defect [IT01]. defects [HG87]. Defense [Dys64d, Dys64f, Dys69a, Dys69n, McM65, DK70, Dys72l]. Defensive [Dys79a]. defined [Dys51b, Dys54a]. Definition [Pra64]. Degree [Dys96b]. Degrees [Ste04]. del [CQ99, Nav99, ZDSN08]. delivered [Dys72n]. della [Cor97, Dys87e, Dys02e]. Delocalization [DZO11, DZO13]. Delocalized [CCC^{+}10]. Delta [NEA10, EN12]. Delusion [Dys82a]. Dementi [Dys98c]. Democracy [PSP^{+}07]. Demystifying [Dys87d]. Denmark [Jen14]. Dennett [Dys06d]. denominators [Meu01]. dense [HMSW00a, HMSW00b]. Densities [BD59, BS89, Dys45a]. density [DCCS94, HS98, NHKK94, NMM^{+}94, NRT22, ST12, VWS08]. density-functional [DCCS94, VWS08]. dependent [VWS08]. Depletion [Olt73b]. dérangeurs [Dys86d, Lar89]. Derivation [Eps55, AHS09, HB12]. derived [SLA93]. Deriving [ABP08]. Dery [CQ99]. DeSantis [Lew99]. description [EF12, tL84, SG86]. Desire [Dys80h, Dys82d]. Destiny [Dys06e]. detachment [CCC^{+}10]. Details [Dys97d]. Detectability [SW66]. Detectable [Dys13c, Dys14b, Dys14c]. Detecting [Dys16]. Detective [Dys12g, Dys12e]. Determinants [Dys63h, Dys72k, Dys73b, Dys76d]. Determination [YHKO07]. Deutsch [Dys11n]. Development [Ano60, Dys60f, Dys61a, Wüt13]. Developments [Dys50d, Dys50e]. Deviations [Dys05l]. Device [Dys99g]. Devil [Dys72n, Dys73d]. diagonal [PS04]. diagonalization [DFPR22]. Diagram [MMJ91]. Diagrams [Wüt11, Cao06, JL86, Kai05a]. Dick [Str00]. died [Ano20d, Joh20b]. Diego [Dys00a]. Dies [Ano20c, Joh20a, Ach20, ST20, Sto20]. different [MNM94, SGT^{+}95]. diffraction [HG87]. diffractive [SRBW05]. Diffusion [Dys55c]. Digital [CK12, CQ99]. dilated [MNM94]. dilema [Fre92]. Dilemma [IT23, Fre92, PD12]. Dimension [Dys67b, LM66, Cur82, Dys78b]. Dimensional [Dys69k, Dys71f, Dys72e, Dys69f, GLS22, HR22, Kon97, Oko91, SLA93]. dimensions [ADM23, LACTFF22, OO23]. d’Inverno [SGT^{+}95]. Diophantine [Dys47b]. Dioxide [Dys77b, Dys90c]. diplomas [Ano97]. diplomi [Ano97]. Dirac [Dys10c, Dys74a, Dys87f, Dys07g, Dys11c, GD10, LW78]. direct [Dys10a, SVT09]. Directions [Dys81f, Dys88c, Dys89f, Dys89g, Dys90h, Dys04f, Spe91, Rom89]. direzione [Dys89g]. Dirichlet [Dys54a]. Disarmament [Dys63c]. Disaster [Dys64g]. discontinuous [Dys71g]. Discovered [Dys09c]. Discovery [CCL^{+}11, Dys80k, Dys16, Dys97a, Gre81]. Discrepancy [LD66]. Discrete [DF92, GLS22]. discrete-time [GLS22]. discretization [Spo22]. Discussion [Dys69b]. Disks [Dys77f]. disorder [SR11]. Disordered [Dys53d, Dys53e, Dys66c, FM21, For21]. Dispel [Dys11l]. Dispersion [Cao06, Nak87, Kai05a, Dys62b]. Dissenting [ABG^{+}89, Dys89a]. Distances [Dys92f]. Distant [Dys79d]. Distribution [BCDL93, Dys71d, Dys53a, Pas22]. distributions [AvM05, OK09, TP01]. disturbers [Lar89]. Disturbing [Dys79e, Dys79i, Dys79j, Dys79k, Dys81d, Dys84d, Dys84e, Dys93c, SZ06, Bet79, Gre81, She82, Ano80, Bar82, Dys79c, Dys79a, Dys79b, Fie79, Has79, Hol81, Joh07a, Pei80]. Divergence [Dys52a]. Diversity [Dys85c]. Divided [Dys15c]. DM [Dys58b, Sta03]. DNA [Dys77g, Dys76c]. Do [Cor10, Spe91, Wil58, Alp73]. Does [Dys12g, Dys12e, Gam67a, Ten69]. domain [RBB95]. dominate [PD12]. Don [Sta03]. Doomsday [Dys80c]. Double [AvM23, Dys58f, Dys60h, IO08]. Doubleday [Dys13e]. Douglas [Gre81]. Doves [Dys85a, ACN85]. Draft [Dys69c]. Dragon [Dys92c]. Dramatic [Dys11o]. Drawing [Kai05a, Cao06]. Dream [Dys07e]. Dreamer [Dys08b]. Dreams [Dys15e, Dys00b, Dys02d]. Dreamt [Ano20c]. Dress [Dys80p]. dried [Dys97h]. Driven [Dys12i, SVT09]. Drum [SGT^{+}95]. Ducher [CQ99]. Duckworth [Ber04]. Duckworth/Ivan [Ber04]. Dukas [Dys82e]. d’un [Dys09e]. d’univers [Lar89]. Dunne [Ben13]. Dwight [Bai17]. Dynamic [Dys05b]. Dynamical [Mun95, AI02, HG87, Yos82]. Dynamics [Dys50e, Dys53d, Dys53e, Dys66c, Dys69d, Dys05h, BF96, CH99, NS93, OO23]. Dynamo [Ler71]. Dyson [Aas86, And91, Ano97, Ano04, Ano13, Bar66, Ber19, Ber02, Bet79, BHJ^{+}00, Bre91, BG85, Bro78a, Bro96b, Bro96c, CK12, CQ99, Dom91, Don02, DORQ12b, Far91, GX06, GFW12, GBS91, Guh07, HGE88, Has79, HMSW00a, Kad85, Lew99, Man76a, Man82, Pei80, PKpCH14, Sax94, Sha01, She82, SGT^{+}95, Spe91, Sta03, Str00, Tre94, Vil95, Way95, Wig90, dJ94, vH84, AI02, Ach20, AvM05, ADvM09, AMN02, ABP08, ABP12, AIP13, AB12, Alb94, AFPT10, AHS09, AF95, AK00, And80, AO05, Ano69, Ano70, Ano80, Ano82, Ano91, Ano93, Ano94, Ano95, Ano97, Ano00a, Ano01, Ano16, Ano17a, Ano17b, Ano20a, Ano20b, Ano20c, Ano20d, AAK^{+}12, AH66, AW69, BMM98, BK73, Bar82, BW69, BS90, BCS89, BMS00, BLS03, Ben97, Ber18, BF95]. Dyson [BF96, BP08, BCDM01, BK10, BM87, BFW11, BB03, BEKW94, BMS61, Bro78a, BS89, Buc87, Bui11, BFM97, BCCM87, CJ80, CH99, CCC^{+}10, CBPG84, CLO10, CBBS11, CBBS12, CK94, CK97, CGR06, CMD04, CDM05, Cor10, CM03, Cre01, CP90, CP93, CK74, Dab75, DvL05, Dav78, Dav92, Daw09, DK06, DD96c, DJS95, Det03, DZO11, DZO13, Dre00, DORQ12a, DCCS94, DG84, Dys51a, Dys65b, Dys88a, Dys91i, Dys96e, EN12, EF12, EF13, ERS07, Eps55, Fai94, Fed00, Fei03, Fie79, FM09, Fol78, Fol80, För96, FM21, For21, FKS97, Gan00, GR09, GS85, GC81, GL82, GZ87, GZ02, GS04b, GS04a, Goc76, GMO98, GMON98, GMO99, GMO00, GMO03, GFW11, GSE^{+}88, Gom84, Goo70, Gre81, GF04, GBB03, GGMO07, Guh91, Gun62, Hag87, HS98, HN08, HH04b]. Dyson [HJ65a, HJ65b, HvSA98a, HvSA98b, HH13, Hof11, HT95, HG87, Hol81, Hom96, HMSW00b, Hor93, HKR07, HBKK11, Hsu03, HAS10, HB12, HM12, Hur52, IO08, IT01, JK67, JK68, JL86, Joh07a, Joh07b, Joh20a, Joh20b, Joh20c, Jon16, Kai05c, Kai05b, Kal64, Kan02, KK02, KV85, KC73, KL02, KP09, KSW13, KK98, Kle89, Kon97, Kor94, KPS90, Kri08, Kri10, Lam06, Lar89, Ler71, LNA96, tL84, LLW^{+}09, LFJ^{+}13, Lie04, LCW^{+}72, Liv54, LEdAB03, LD66, LFW10, Maa06, MEBK11, MMJ91, Mah05, Man76b, Man77, MNBW07, Mar88, Mat49a, Mat49b, MM97, McL13, Mcc94, MNM94, Met73, MOR95, Meu01, Meu09, MND11, MNLD11, MNK^{+}97, Mon94, MM90, Mun95, NO99, NN86, Nak87, NH91, NHKK94, NMM^{+}94, NS93, Nar86, Nav99, Neu16, NAW08, NEA10, OK07, OK09]. Dyson [Oko91, Olt73a, Olt74a, Ort04, OO23, PR09, Par87, PVW02, PVW03, Per68, Pie84, PS04, Pom09, Pra64, RNM04, Rad20, RBB95, RT82, Rem90, RM06, RS77, Roc71, RQ11, Rom89, SW66, SWFB05, San94, Sar05, ST12, SA03, SS99, Sch13, SS67, Sch94b, Sch94a, Sha87, SG86, She88, She89, SG93, ST20, ST67, SZ06, SLA93, SOM10, Smi06, SG06, SBZ21, SPZ^{+}12, Sta98, SRBW05, Ste92, Sto20, SVT09, SS98, Swe88, SR11, TKH62, Tak86, Tak88, TP01, Tak89, Tam83, TC92, Tay69, Tie09a, Tie09b, Tie09c, TMD96a, TMD96c, TMD96b, VPW01, VWS08, VHK89, Wad81, Wal99, WR07, WR10, Whe91, WK74, WKH13, Wil62, WYL81, Woo92, Wüt13, sXcHmW87, Xue88, YHKO07, Yan10, Yan54, Yan55, Ye92, Yeu80, Yos82, You69, YxQCxL10, Zei87, Zho97, Zwa03]. Dyson [dS99, Bai17, Ben13, Bus15, Cre13]. Dyson-corrected [SS98]. Dyson-equation [IT01]. Dyson-like [Nar86].

E-Flat [Dys79h]. Early [SGT^{+}95, Joh20a, Wüt13]. Earth [Dys02b, Dys05h, Kon90, Sha01, Str00, Dys15e, Dys69m, Kra01]. Eating [Rog97]. Economic [Dys78e]. Economy [Dys90i]. Economy-Class [Dys90i]. ed [Dys62b, Sta03]. ed. [Dys70d]. Eddington [Bek86]. Edited [Dys58b, JR80, Bai17, BB63, Bus15, Dys83c, Dys90k, Dys92c, Dys05l, Dys08d].

Edition [Dys06c, Dys60a]. Editor [Dys60i, FD93, Dys89h, ABG^{+}89, BD07, BCD^{+}65, CDB^{+}73, DHD06, Dys59b, DK70, Dys71i, Dys81i, DGB^{+}01, FD07, GD03, GD05, GSE^{+}88, GFD04, LD90, LD08, MD95b, MD08, Mor62, PD08, Ros62, SD06, SD98, SDL^{+}01, SDGK06, Tay69, Ten69, TD05, WDW89, You69]. Editorial [Sta03]. Editors [Dys85a, Dys65d, Dys98c]. Edmund [Dys54d]. eds [Dys62a, Dys67b, Dys70c, Sta03]. Edward [Dys58a, Dys02a, Dys98d, Dys06g, Dys07h, Dys09d]. Edwars [CQ99]. Effect [Dys55c, Dys56g, DTO78, ST12, ST67]. effective [CGR06]. Effects [BFW11, NMM^{+}94]. Efficiency [Dys69e]. Eigenvalue [FKS97, Pas22]. Eigenvalues [Dys62c, Dys71d, Dys70e]. Eighth [Gre81]. einem [Dys89m, GBS91]. Einführung [SGT^{+}95, Dys55b]. Einstein [BHJ^{+}00, Dys09a, Dys14a, Sta03, Ano95, BGA^{+}13, Cal96, Cal00, Cal05, Cal10, Dys80e, Dys96m, Dys03f, Dys15a, Tip10, TD14, Woo80, Dys03c, Dys03f, Dys15a]. Ekeland [Dys06e]. elastic [Kle89]. Electricity [Dys54d, SGT^{+}95]. Electricity/Zell [SGT^{+}95]. Electro [Dys50e]. Electro-Dynamics [Dys50e]. Electrodynamics [Dys49b, Dys50b, Dys52c, Dys58k, Dys03i, Dys05d, Dys06j, Sch58, Sch03, SGT^{+}95, CBPG84, Dys51c, Dys51d, Dys51e, Dys51f, Dys52a, Kon97]. Electromagnetic [Bet47, Dys48c, DM57a, EN12, NEA10]. Electron [Bet33, Dys55c, GGMO07, MMJ91, MNLD11, CCC^{+}10, DK06, DCCS94, IO08, MNM94, MND11, SS98]. electron-momentum-spectroscopy [DCCS94]. Electronic [SGT^{+}95, DK06]. electronically [OK07]. Electrons [BH34, BF32]. Electrostatic [Fol78]. Elektronen [BF32]. Elektronenprobleme [Bet33]. Elementary [Gam67b, EG11]. Elements [BD59, Dys54c, Wat48]. Eleuthera [DMZ76]. elite [Fin06, Hor06a]. Elliot [Dys67b]. Elliott [Dys04a, LT91, LT97, LT01, LT05, Thi91, TL97, Thi01, Thi05]. elliptic [ADM23]. embedding [IT01]. Emergence [Str00]. eminent [Fad90]. emission [BS89]. Emmeche [CQ99]. emocionales [CQ99]. Empires [Dys03c]. Empirical [BW69, BS89]. employing [WYL81]. Empty [Dys95j]. Encouragement [Dys82b]. End [Dys96i, Dys02c, Dys02g, Dys05f, Dys07b, GD05, Sta03, Dys79m, Dys83d, Dys89m, Joh20c]. Ende [Dys89m, GBS91]. Ending [BM99]. Ends [Dys02c, Dys02g]. Enemies [SD06]. Energies [MD57, GBB03, GGMO07, HH13, NMM^{+}94, PS04, SWFB05, SOM10, SZ21, SS98, WKH13]. Energy [Ano71, Bet47, BHJ^{+}00, Dys48c, Dys50d, Dys57a, Dys62g, Dys62h, Dys62i, DM63, Dys67c, Dys69e, Dys71e, Dys71a, Dys75a, Gun62, MD63, Bas22, BM83, BF95, BDL93, BF76, Dys71i, Man78, McP74, NRT22, Zho97]. Engineer [Dys08b]. Engineering [BDF00]. English [Dys60a]. enigma [Ber04, Ber04]. enlarged [Dys60a]. enquiry [Alp73]. Enrico [Dys04b, San94]. ensemble [ADM23, BBDY22, Dys12c, SLA93]. Ensembles [Dys72d, Dys62k, EJ22, FM21, JKM21]. Enterprise [KD12]. entitled [BM83]. Enumeration [Hur52]. environment [LDB77]. envision [BDF00].

Enz [Dys02d, SGT^{+}95]. Eörs [Str00]. epsilon [GBB03]. Equal [SS67]. Equal-Time [SS67]. Equation [CDD56a, CDD56b, Dys44a, Dys57b, För96, Hom96, Kal64, Ler71, AB12, AAK^{+}12, BK10, BS89, CJ80, CH99, CK94, CK97, CP93, DvL05, Dys51f, GFW11, GFW12, HS98, HvSA98b, HT95, HBKK11, IT01, KL02, Kle89, Kon97, LNA96, MM97, MM90, NN86, Nar86, OK07, OK09, PR09, PVW02, PVW03, Rem90, RM06, RQ11, ST12, SS99, Spo22, SRBW05, Tak89, TC92, VPW01, WR10, WYL81, Woo92, sXcHmW87, Xue88, Ye92, Zho97]. equation-of-motion [OK07, OK09]. Equations [BS90, BD59, BFM97, Dys84g, Pom09, AMN02, ABP08, ABP12, AIP13, AHS09, And91, BMM98, BCS89, BMS00, BP08, BFW11, BEKW94, Bre91, Bui11, BCCM87, CBPG84, CMD04, CDM05, CP90, DJS95, Det03, Dom91, DORQ12a, DORQ12b, Dys90e, ERS07, Far91, FM09, FM21, HN08, HvSA98a, HG87, HAS10, HB12, HM12, KP09, KSW13, KPS90, LFW10, Maa06, MMJ91, Mun95, NO99, Oko91, RBB95, Spo22, SR11, VWS08, Wad81, Wal99, WR07, Yos82, YxQCxL10, Zwa03]. equivalence [DD96c, Man76a, Man76b]. Erbium [LCW^{+}72]. Erforschung [Fis10, Fis12]. Ergodicity [OO23]. Erich [Dys05f]. Erinnerungen [Dys81g]. Ernesto [Dys08d]. Eros [Ano93, Ste92, Dys92a, Dys93b, Dys94b, Dys13a, Fre92]. Erratum [BD95, DhX65b, Dys73b, GFW12, HMSW00a, KTNG23, Man76a, Man82, Mat49b]. error [BCDL93, BD94c]. Erwin [Sta03]. escape [CQ99]. Escher [Gre81]. ESO [RWP78]. ESP [Dys04c]. essays [BR01, BM99, For02, Joh07b, LSW76, Man07, Opp89, MB66]. eternities [Dys79f]. Ether [Dys09b]. Ethical [Dys97b, Fre92]. Ethics [Dys96i, Lam06]. ético [Fre92]. étrangers [Ham91]. Euclidean [BFM97]. EUR83.00 [Sta03]. EUR98.00 [Sta03]. Europe [Ell86, IT23]. evaluation [GBB03, LEdAB03, WKH13]. Every [Dys44a]. everything [Bro10, Dys12f, Dys12d]. Evidence [LCW^{+}72, MOR95]. Evolution [BHJ^{+}00, Dys05h, Dys07b]. evolutionary [PD12]. evolutions [BM83]. Exact [SLA93, Zho97]. Exactly [Dys67b, Dav92, EG11, LM66]. Examples [Dys07i, Dys11d, OK07]. Exchange [NZSD08, BHMD07]. excited [OK07]. exclusion [GLS22]. Exist [Dys12g, Dys12e]. Existence [Dys69f, Dys69k, Dys71f, Dys72e]. Existential [Dys12g, Dys12e]. exotic [GLSS21]. Expanded [Sta03, Cal00]. Expanding [Vog99]. Expansion [MOR95, BCCM87, CDM05, KPS90, Nar86, RS77, Tam83, Yeu80]. expansions [HH13]. Experiencing [Gre81]. Experiment [Dys59c]. Experimental [Dys07s, Dys11i]. Experiments [Dys72f, Dys04g, YHKO07]. experts [BDF00]. Explanations [Dys11n]. Exploding [Olt73b]. Exploration [Dys69g, Dys69h, Joh20a]. explore [Dys00d]. explores [Lam06]. Exploring [Sta03]. Explosion [Str00]. Explosions [Dys59a]. exponential [BD94b]. Exponents [BM87, MOR95, GMO99, Zwa03]. Exposition [Pom09]. Express [Ano54]. Extending [BG85]. Extension [VWS08, LS22]. external [LNA96]. Extinction [Dys96i]. Extra [Dys60j]. Extra-Terrestrial [Dys60j]. Extraordinary [May07]. Extraterrestrial [Dys66g, KB89, Sag73, Str00, Cam63, Dys89l, Sag00]. Extraterrestrials [HZ82, ZH95]. Exxon [Noc08]. Eyes [Dys63c].

F [Ben13, Bre91, Dom91, Dys51a, Dys56b, Dys64c, Dys76a, Dys82a, Gre81, Lar89, DZO11]. F. [Ano95, Bet79, Liv54]. Fabián [CQ99]. Fabric [Dys04d, DG05, GF04, GFD04]. Face [Dys90d, Fre92]. facilitated [GLS22]. factor [CM03]. Factors [BFM97, KK02, NEA10]. Facts [Dys58a]. faculty [Mit80]. Failure [Dys06h, Dys06l]. Fairchild [Dys94a]. Faith [Ano96, Dur00, Fra08]. fall [Dys51a, PSP^{+}07]. family [Pei80]. famous [HH04a]. Fano [Dys60b]. Far [Dys62a, Dys11n, Ano20b, Rad20]. far-fetched [Ano20b, Rad20]. Far-Out [Dys11n]. Farmelo [Dys10c]. Farms [BCC^{+}08]. Farrar [Dys05e, Dys11l]. fashions [Dys65c]. Fast [BH34, Dys11l]. Faster [Dys95j]. Fate [Dys80h, Dys82d]. Father [Dys05g, FD93]. Fathers [Dys78e]. Faust [Dys07f]. fazovye [Daj73]. Feathered [Dys85a]. February [Dys06b, Dys22, Joh20b, Ano20d]. Federation [Dys63f]. Fejnman [Dai67]. Fellowship [Dys07e]. Fend [Bas11]. Fermat [Dys00f]. Fermi [Ano94, San94, Dys71j, Dys04b]. Fermion [She88, CGR06, CP93, HN08, KP09, Tak89, VHK89, Zho97]. fermion-boson [HN08]. fermion-photon [KP09]. Ferris [Dys02b, Dys05l]. Ferromagnet [Dys56f, Dys69k, Goc76, LD66, Dys69f, Dys71g, Hof11]. Ferromagnetism [TKH62]. Ferromagnets [Dys69l, Dys71f, Dys71l, Dys72e, RT82]. Ferromagnon [Pra64]. Ferromagnons [Pra64]. fetched [Ano20b, Rad20]. few [ADvM09]. Feynman [And91, Bre91, Bro96b, Bro96c, Cao06, Dai67, Dom91, Dys91a, Dys98b, Dys05l, Dys11o, Far91, KC73, Sax94, SGT^{+}95, Spe91, Tre94, Vil95, Way95, BR93, Cao06, CK74, Dys49a, Dys58j, Dys65e, Dys89e, Dys90e, Dys93d, Dys01d, Dys03h, Dys06i, Dys11o, Dys11u, Eps55, FD93, Fey00, FL06, GZ87, GZ02, Gle92, Hal17, Hur52, JL86, Kai05a, Kai05b, Kra11, Man78, MNM94, Pom09, Sch94b, Sch94a, Sta98, Wüt11, Wüt13, Yeu80, Dys92b, Dys05l]. Feynman-Dyson [KC73]. Feynman/Miller [SGT^{+}95]. fiction [Cas03, Dys85i]. Field [Dys53f, Dys53g, Dys60c, Dys62b, Dys07i, Dys07j, Dys11d, Dys11e, Ler71, BMS00, BCDM01, Bui11, CH99, CMD04, DSW51, Dys65c, Dys96c, GMON98, LNA96, McL13, Oko91, SPZ^{+}12]. Fields [Dys48d, Dys55a, Dys56b, JK67, JK68, Kri08, LW78, SHD91, Dys56b, Dys60a]. fifth [ALW^{+}88, Dys95e]. fiftieth [Rab78]. Fifty [Dys73c, Dys12h]. Fighting [Dys82c, Sol13]. filosofico [Ano95]. Final [Dys91f]. finding [Fey00]. Fine [Bek86, Guh07, DD96a, DD96b]. Fine-Structure [Bek86, DD96a, DD96b]. Finite [CBPG84, Dys67c, Mar88, RBB95, HS98, HN08, HKR07, NH91, NHKK94, NMM^{+}94, RS77, ST12, Ye92]. Finkbeiner [Hor06a]. First [Dys84f, Dys86c, Dys92c, Dys11o, Nee83, BB03, WR07, Dys70i]. Fish [Dys97h]. Fitch [Dys98c]. five [ALW^{+}88, DG84]. Fixed [Dys55e]. fizika [Dai65]. fizike [Dai67, Dai10]. fiziki [Dai71]. fizyka [Dys83d]. FKG [LS22]. Flat [Dys79h]. flavor [NAW08]. flesh [Dys72n, Dys73d]. Flo [Dys05g]. Florida [Mey56]. flow [Wal99]. flows [ST11]. Fluctuation [Ler71]. fluctuations [TP01]. fluff [Sol13]. Flügge [Dys58b]. fluid [Dys95e, ST11]. fluoride [CCC^{+}10]. Fly [Dys05e]. Fock [YHKO07]. force [BM99]. forces [Dys84m, Dys09b]. Foreign [Dys63f, Ham91]. Foreword [Dys60d, Dys60e, Dys62d, Dys80f, Dys80g, Dys87b, Dys96k, Dys96m, Dys99c, Dys99d, Dys99e, Dys00c, Dys04e, Dys06i, Dys07k, Dys07l, Dys10g, Dys11q, Dys12a, Dys12b, Dys05f, Dys05l, Sta03]. Form [BFM97, Dys78b, Dys95g, Ano95, CM03, KK02, NEA10]. forma [Ano95]. formalism [MNK^{+}97, NH91, NHKK94, NMM^{+}94, OK07, SA03, WR07]. Formation [Wat48]. Formative [Sta03]. forms [ADH88, Dys48a, SOM10]. Formula [BMS61]. formulation [CMD04, GLSS21, TC92, NHKK94]. Försterling [BHJ^{+}00]. forties [Dys01a, Dys93e]. Forty [Dys98e]. Forum [Dys63a]. Foundation [Has79, Hor06b]. Foundations [GZ02]. four [Dys48a, HT95, KV85, Tak89]. four-fermion [Tak89]. four-particle [KV85]. four-point [HT95]. Fourier [Dys58c, Dys53a]. Fourth [BDS52]. Fourth-Order [BDS52]. Fractals. [Dys78b]. fraction [Dys43b]. Fractions [Dys00f]. fractons [GLSS21]. Fragile [Dys97a]. framework [HM12, Kon97, LFJ^{+}13]. Francis [BD07]. Frank [Dys09b, Dys15c, Spe91, Dys99a]. Frederic [Dys55a]. Fredholm [Dys76d]. Free [DW89, Dys07m, Dys07o, Dys11f, Dys11g]. Freedom [Dys56e, Dys56a, Dys82c, Dys07c, Ste04]. Freeman [And91, Ano80, Ano93, Ano97, Ano04, Bar82, Bar66, Ben97, Ben13, Bro78a, Buc87, Bus15, CQ99, Cre01, Cre13, Far91, Fie79, Gom84, Gre81, Guh07, Hag87, Has79, Hol81, Joh07a, Jon16, Lew99, Nav99, Pei80, PKpCH14, Pie84, Rom89, Sha87, Sha01, She82, Spe91, Sta03, Ste92, Str00, Whe91, Wig90, dS99, Aas86, Ach20, Alb94, Ano97, Ano00a, Ano01, Ano13, Ano16, Ano17a, Ano17b, Ano20a, Ano20b, Ano20c, Ano20d, Cor10, Dav78, Daw09, Dre00, Dys77g, Dys91i, Dys96e, Dys22, Fai94, Fei03, Gan00, GR09, GF04, HH04b, Hor93, Hsu03, Joh07b, Joh20a, Joh20b, Joh20c, Kai05c, Kor94, Lam06, Lie04, Mcc94, Met73, Mon94, Neu16, Olt73a, Olt74a, Rad20, San94, Sar05, Sch13, ST20, Smi06, SBZ21, Sta98, Sto20, Tie09a, Tie09b, Tie09c]. Freeman [vH84]. Freeze [Dys97h]. Freeze-dried [Dys97h]. Freezing [AMN02]. French [Dys04c, Dys80k, Dys91e, Dys09e, Ham91, Lar89]. Friends [Dys09c]. Fringe [Dys12f, Dys12d]. Frogs [Dys09a, Dai10, Dys10b, Dys11b, Jon16]. Front [Dys07n, SGT^{+}95]. Frontier [Dys91f]. frontiere [Cor97]. Frontiers [Dys80o, Cor95, Cor97, Gin75]. Frozen [BD07]. FRS [BM99]. Fry [Str00]. fu [Bro01]. Fugue [Gre81]. full [KP09]. Function [Dys53c, Dys53h, BD94c, DK06, Dys60l, IT01, JKM21, She89, Yeu80]. functional [DCCS94, HB12, VWS08]. Functions [Dys58d, TKH62, Dys51b, Dys53a, Dys54a, Guh91, LS22, RM06, SS99, Sim22, Dys58c]. fundamental [Alp73, Dys72a]. funding [DLV10]. Further [Mor62, Dys00d]. Fusion [RRD^{+}73]. Future [Ano60, BHMD07, BGA^{+}13, Dys58a, Dys60f, Dys61a, Dys75b, Dys78d, Dys80h, Dys80e, Dys82d, Dys90f, Dys03d, Dys07r, FD07, Ano20b, Bro10, Dai71, Dys70f, Dys72b, Dys72g, Dys72h, Dys81g, DB02, DBB09, Dys11r, Ham91, Rad20, RWP78, Lew99, Lew99]. Futuro [Nav99, CQ99, CQ99, Nav99].

G [Dys60c, Dys60a, Dys60b, Dys70c]. Gaia [Dys90d, Dys92a, Dys93b, Dys94b, Dys13a, Fre92, Ano93, Ste92]. Galápagos [Dys08c]. Galison [Dys03c]. gamma [Mor58]. gamma-ray [Mor58]. Gamow [Alp73]. Gap [Dys66b]. gapped [DFPR22]. garden [Dys88d, Dys01f, Dys13d]. Garvan [Mah05]. Gary [Gre81, Dys89d]. Gas [Dys57a, Dys69d, Bas22, NRT22, SLA93]. Gases [Dys63h, Lew22]. Gauge [BP08, Dys50c, Rem90, ABP08, ERS07, HAS10, Kon97, WR07, WR10, sXcHmW87, dJ94]. Gauge-independent [Rem90]. Gauge-invariant [BP08]. gauges [CP93, RQ11]. Gauss [Sim22]. Gaussian [BF95, BDL93]. General [Dys56c, Dys80b, Dys81a, Dys86a, Dys07o, Dys11g, Dys85i]. Generalised [Dys58c]. Generalization [Man76b, Man76a]. Generalized [GS85, JL86, Man77, Man82, Yan10]. Generation [Dys09c, Yos82]. generators [BMM98]. Genesis [Wüt11]. Genetic [Dys81i]. Geneticist [San94]. Geneva [RWP78, Yor87]. Genius [Ben13, Cre13, Dys05g, Dys10c, Dys13e, Dys15c, Gle92, Hor06a, Joh20a, Sar05, Sch13, Dys92b]. Gennes [Dys97a]. Genom [DB02]. Genome [BHJ^{+}00, DB02, Dys99i, Dys01b]. Gentile [CQ99]. Geoffrey [Sha01]. geometry [vN98]. Georg [Odi96]. George [Ber02, Bro78a, CK12, Dys82a, Dys96c, Kan02, Alp73, Can01, Dys93a, Dys93e, Dys01a]. Georges [Dys04c]. Gerald [Ber04, Sta03]. Gerard [Dys85b, Dys93f]. German [Dys60c, Dys05f, Ano69, BF32, Bet33, Dys70g, Dys81g, Dys89m, DB02, DBB09, Fis10, Fis12]. Germany [Dys58b, Dys05f, Dys03f]. Gesammelte [Dys70d]. Gets [Fed00]. Getting [GR09]. Ghost [AIP13, ABP08, DORQ12a, DORQ12b, RQ11, WR10]. ghost-gluon [AIP13, DORQ12a, DORQ12b]. Giants [Dys13b]. Gibbs [CM90, Dys90a]. Gifford [Dys88c, Dys90h, Dys04f]. Gilbert [Dys07e]. Gilles [Dys97a]. Gimbel [Dys15a]. Ginibre [ADM23, FKS97]. Gino [Dys07f]. Giroux [Dys05e, Dys11l]. giugno [Ano97]. given [Dys51a, Dys51g, Dys88c, Dys90h, Dys04f]. Gladwell [Dys13b]. Glasgow [BM55]. glass [Dys07q, Dys10i]. Gleick [Dys92b, Dys03b, Dys11m]. Global [Dys08d, Gol08, NZSD08, ZDSN08, ADM23]. Globalization [Dys08d]. Globalization/Brookings [Dys08d]. globally [Maa06]. Glory [Dys10d]. Gluon [ABP08, ABP12, AIP13, DORQ12a, DORQ12b, HvSA98b]. Go [SD15]. Goals [Dys95j]. God [Dys98b, Dys02c, Dys02g, Dre00, Dys97d, Dys98b, Tip10]. Gödel [Gre81]. Gofman [DK70]. Goldberg [Dys70c]. Golden [CL22]. Goldstone [DD96c, Mun95]. Goldstone-boson [DD96c]. Goliath [Dys13b]. Good [BR93, SZ06, BB63]. Goodchild [She82]. Gordon [Dys62b]. Got [SD15]. Grade [Tie09a]. graded [Guh91]. Grader [Tie09b]. Graham [Dys85a, Dys10c]. Grand [Dys09b, Dys12c]. Graphs [Hur52, GS04b, GS04a]. grating [RM06]. Gravitational [Dys63d, Dys65d, Dys69e, Dys69m, Dys76b, Dys16]. Graviton [Dys13c, Dys14b, Dys14c]. Gravity [Dys98a, Gam67a]. Great [Dys14a, LY95]. Greatest [Dys06e]. Green [Can01, Dys93a, Dys93e, Dys01a, RM06, SS99, TKH62, Yeu80]. Greenbelt [Kon90]. Greene [Dys04d, DG05]. greening [Dys96c]. Green’s [IT01]. Green’s-function [IT01]. Gribbin [Dys07e]. Gribov [DORQ12b, DORQ12a, HAS10]. Gross [Bas22]. Ground [Dys57a, Dys67c, NRT22, Bas22, LDB77, OK07]. Ground-State [Dys57a, Dys67c]. Group [Bar66, Dys56g, Dys66a, Dys05e, HB12, Kri08]. Groups [Bar66, Dys62k, Dys66i]. Growth [Olt73b, Dys54a, Olt74b]. Guantanamo [SDGK06]. Guardian [Ano20b]. Guarding [Dys02b]. Guesses [Dys44b]. Guide [GMON98].

H [Bus15, DZO11, Dys56b, Dys67b, GGMO07, LT91, LT97, LT01, LT05, Thi91, TL97, Thi01, BS89, MND11, Thi05]. H.R. [DLV10]. Habitable [Vog99]. Hadron [BFM97, EF12]. hadron-meson [EF12]. hadron-photon [EF12]. Hadronic [GFW11, MEBK11, GFW12]. Haifa [Dys77a]. Haldane [Dro95]. Half [Dys15c]. Half-Life [Dys15c]. halide [DZO13]. Hall [Ber04]. Halo [DTO78]. Hamburg [Dys05f]. Hamiltonian [Mar88, SG06, Wal99]. Hamiltonians [CGR06, DFPR22]. Hamish [Lew99]. Hancock [SGT^{+}95]. Hancock-Beaulieu [SGT^{+}95]. handbook [ABD11]. Handbuch [Dys58b]. Hans [Dys05f, Sta03, BL06, Dys55a, Dys05c, Dys05d, Dys06j, MB66]. Hard [Dys57a, Dys97a, Dys99j]. Hard-Sphere [Dys57a]. hardback [Sta03]. hardcover [Ber04, CK12, Dys96j]. harmonic [HR22]. Harper [Pei80]. HarperCollins [Dys03d]. Hartree [YHKO07]. Harvard [Dys10d, Sta03]. Harvey [Dys07e, Dys77a]. Hastings [Dys05f]. having [Ano95]. Havoc [Sta03]. Hawking [Spe91]. Hawks [ACN85, Dys85a]. Hearings [Dys77g, Dys63f]. Heat [Dys54f]. Heated [GR09]. heaven [Cas03]. Heims [Dys81b]. Heisenberg [Dys51c, Dys51d, DLS76, LD66]. Held [JR80, Dys77g, Kon90, Mey56, Ros63]. Helen [Dys82e]. help [Lam06]. Henri [Dys04c]. Henry [Ber02, Sta03]. Hentschel [Sta03]. Heppenheimer [Dys79d]. Herald [Joh07b]. Herbert [Dys76a, She82]. Heretic [Daw09, Dys08f]. Heretical [Dys05j]. Heretics [Dys13f]. Hermann [Dys56d, Dys70a, Bar66, Dys70d]. Hermitian [FKS97, Mar88, Tak86, Tak88]. Hero [Dys05g]. Herrgott [BHJ^{+}00]. Hertz [Dys69m]. Hiai [CL22]. Hidden [Dys74b, Dys75b, Dys77c, Dys10c, Dys89d]. Hiebert [Sta03]. Hierarchical [BM87, GMO03, MOR95, GMO98, GMON98, GMO99, GMO00, Meu01]. High [BCCM87, Dys50d, GMO99, Man78, MOR95, Dys71i, LACTFF22]. High-accuracy [GMO99]. High-energy [Man78, Dys71i]. High-Temperature [MOR95]. higher [ADM23]. Highest [Ano91]. Hilary [Dys11o]. Hilbert [Dys70a]. Hilbert. [Dys70a]. Hindsight [Dys81e]. Hintertreppe [Fis10, Fis12]. hipporoo [Dys06a]. Hiroshima [Yor87]. Hirschfelder [She82]. Historical [Dys11u, Sta03, Str00]. Histories [MH11]. History [Dys54d, Dys81e, Hor06a, Spe91, Sta03, UM86, Aas86, Fin06, KLR13, Kev87]. Hoffmann [Dys55a, Dys56b, Dys96j]. Hofstadter [Gre81]. Hole [Ove08, Dys80d]. holes [ST11, Spe91]. Holland [Dys04c]. Holmes [Dys09c]. Holt [Ber02, Dys12g, Dys12e]. Holton [Sta03]. Homage [Dys93e, Dys01a]. Home [Dys80o]. homogeneous [Dys48a]. honey [Rog97]. Honor [Ano91, yWHH98, Dys04g, LW78, LSW76, MB66, SBZ21]. Honorary [Dys96b, Ano97]. Honored [Swe88]. Honoring [Dys74a]. honoris [Ano97]. Honour [Bus15, BM99, PKpCH14]. Hope [Dys89j, Dys02c, Dys02g, Dys84i, Dys84j, Dys84k, Dys84l, Dys84s, Dys03a, Buc87, Gom84, Pie84, vH84, Fre84]. Hopkins [Dys04c]. Horace [Gre81]. horizon [HAS10]. Hot [Gol99]. House [Dys77g, DLV10]. Howard [Sta03, Dys87a]. HQ [KLR13]. HQ-3 [KLR13]. HST [Dys90m]. Hubbard [SS99]. Hugh [Lew99]. Human [Dys69g, Dys69h, Dys96i, Dys97c, Dys05e, May07, KNGT22, KTNG23, PM^{+}83, Tip10]. Humboldt [Han82]. Hunting [Dys91g, Dys92d]. Hybrid [CBBS11, CBBS12]. hydrated [CCC^{+}10]. Hydrodynamic [Spo22]. Hydrogen [Dys76a]. Hydrogen-Bomb [Dys76a]. Hydrostatic [Dys60g]. hypergeometric [Dys60l]. Hyperscaling [GMO00, AAK^{+}12].

I. [Han82, Has79]. Ian [Dys96g]. IBM [Swe88]. Ice [Dys04a]. iconoclast [And13]. Ideal [Dys56f, Hof11]. idealized [Dys60l]. Ideas [Dys12i, Ano20b, Bro10, Rad20, BB63]. Identities [BG85, DD96c, Dys43c]. identity [CP90]. ideología [Nav99]. If [Dys75b]. Ignorance [Dys63e]. II [Dys55a, Dys70d, Dys51d, Dys55c, Dys60l, Dys62h, Dys79j, Dys84j, Dys90l, Dys06l, Fed75, GS04b, KK02, LD68, NMM^{+}94, TKH62, Yan55, Yan13, Dys54d, Dys90k]. IIASA [BM83]. III [Dys62a, Dys70d, Dys62i, Dys79k, Dys84k]. Illinois [AAB^{+}88]. Illus [Dys60a, Dys64c, Dys70a]. Illusions [Dys11l]. Image [Dys78c]. imaginación [CQ99]. imaginary [ST12]. imagination [Cor95, Cor97]. Imagined [Dys97e, Lew99, Ben97, Cre01, dS99]. immigrant [Pup60]. Implementation [SVT09, OK07]. implications [Dys77g]. Importance [Dys90g, Dys91e]. imposing [KP09]. imprévisible [Dys91e]. improved [Jud72, Man77, Man82]. improvement [Man76a, Man76b]. Inadequacy [KV85]. Inconvenient [Rev09]. Incorporating [Dab75]. indefinite [ADH88]. Independent [Eps55, Rem90, Zwa03]. Industrial [Dys71k, Dys77d]. Industry [Dys98f]. Inelastic [HJ65a, HJ65b]. inequality [CL22, LS22]. inertial [Woo92]. inertial-range [Woo92]. Inexplicable [May07]. Infinite [Ano20d, Dys81f, Dys88c, Dys89f, Dys89g, Dys90h, Dys04f, OO23, Spe91, Rom89]. Infinito [Dys89g]. Infinity [Dys11n]. informática [CQ99]. Information [Dys81i, Dys05g, LLW^{+}09, Yan10, GBS91, Dys11m]. Infrared [Dys60k, HAS10, Kon97, SW66, AB12, RBB95, RQ11, Zwa03]. inhomogeneous [SHD91]. Innenansichten [Dys81g]. Innovation [Dys58e, Dys64a, Dys98e, DB02]. input [Dab75]. Inquiry [Dys07c]. instabilities [Fol80]. Instability [Dys60g, Meu09]. Institute [DLV10, Mit80]. Institution [Dys08d]. instruments [Dys01b]. integers [BD94b, Dys45a]. integrability [RNM04]. integrable [BLS03, BDL93, JRV22]. Integral [Dys58g, Dys58h, Dys58f, Dys60h, JL86, SVT09]. integral-driven [SVT09]. integrals [EG11]. integrated [Jud72]. intellectual [Joh20b]. Intelligence [Dys06h, Dys06l, Sag73, Bro15]. Intelligencer [Dys89h]. Intelligences [Dys60j]. Intelligent [Dys72i]. Intensity [Ler71]. Interacting [Dys67b, GL82, LM66]. interaction [BF32, MND11, MNLD11, TC92]. Interactions [Dys48d, Dys56c, Dys86a, Mat49a, Mat49b, DLS78, NRT22]. Interdisciplinary [JR80, Sta03]. Interesting [DFG10, DFG13]. Interfaces [SGT^{+}95]. Interfaces/Meadows [SGT^{+}95]. Interior [Dys81g]. interlacing [AvM23]. Interlude [Dys05i, Dys06f]. intermediate [Wal99]. Internal [GBS91, DMZ76]. International [Dys62b, Dys05e, Jen14, KLR13, Kon90, Joh07b]. Internet [BHJ^{+}00, DB02, Dys99i, Dys01b, DB02]. Interplanetary [Dys02b]. interplay [CJ80]. Interpretation [DMZ76, GBB03]. Interpretive [Lew99]. Interscience [Dys60a, Dys62a]. Interstellar [Cam63, CM59, Dys68a, Dys69b, Dys82f, Dys85d]. Interview [Dav78, Olt73a, Olt74a, Smi06, Aas86, Daw09, Met73]. interviews [Dre01]. Intrigues [Dys95a]. Intrinsic [She89]. Introduction [Dys58c, Dys81h, Dys07p, Dys11h, Dys05l, Dys60a]. introductory [Mar56]. Invariance [Dys50c, HZ01]. invariant [BP08, JKM21, Kri08]. inventor [Dys99f, Pup60]. Inverse [Dys76e, För96, Hom96, Dys76d]. Investigation [BBDY22]. ion [RT82]. ionization [DK06, GBB03, IO08, OK07, SWFB05, SOM10, SPZ^{+}12, SVT09, YHKO07]. ions [OK09]. Iris [Str00]. irreducible [CDM05, GS85, Dys60b]. Irregularity [Dys78b]. Isaac [Dys03b]. ISBN [Ben13, Ber02, Bro96c, Dys96j, Dys97a, Dys02a, Dys02d, Pei80, Sta03]. ISBN-13 [Ben13]. Ising [BMM98, Dys69f, Dys69k, Dys71f, Dys71g, Dys72e, MOR95]. Islands [Dys08c]. Isotopes [LCW^{+}72]. isotropic [DLS78]. Israel [UM86]. ISSOL [Dys00a]. Issue [Mcc94, Dys77g]. Italian [Ano97, Dys87e, Dys89g, Dys02e]. Italy [Dys74a]. Iterated [PD12]. Iterative [För96, Hom96, DFPR22]. IV [DM63, Dys84l, HH04a, Dys70d]. Ivan [Ber04]. Ivar [Dys06e].

J [Aas86, And91, Ano97, Ano20b, Ano20c, Bar66, Ben97, BB63, Bre91, Dom91, Dre00, Dys51a, Dys58c, Dys77g, Dys81b, Dys91i, Dys08b, Far91, Gan00, HH04b, Hor93, Jon16, KTNG23, Liv54, Nav99, ST20, Spe91, Str00, Whe91, Wig90, Ber04]. J. [Ano97, Ber84, DPSY64, Dys72n, GDH^{+}04, PC06, She82]. Jacques [Dys97a]. Jagdish [Dys74a]. James [Dys03b, Dys11m, Wig90, Dys92b]. January [Dys91i, Dys99b, Dys09d, JR80]. Japanese [Broxx, FR94, FR02]. JASON [IT23, CDB^{+}73]. Jasons [Hor06a, Fin06]. Jellium [Dys04a]. Jeremy [Ber04, Gre81]. Jew [Dys15a]. Jews [BCD^{+}65]. Jim [Dys05g, Dys11o, Dys12g, Dys12e]. JingShin [yWHH98]. Joandomènee [CQ99]. Joel [Dys05f]. John [Dys70c, Dys96i, Dys98b, Dys02c, Dys02g, Dys07e, Dys22, Spe91, Sta03, Str00, Dys80d, Dys04g, Dys10h, Hal17, Hei80, Ove08, Dys81b]. Johnny [Dys13d]. Johns [Dys04c]. joint [AvM05]. Jonathan [Dys84a]. Joseph [Dys85a, She82, Spe91, Sta03, BM99]. Jost [BK73, BMS61, SS67]. Journey [Dys02a, Dys91a]. Jr [Dys85a]. Jr. [Wig90]. ju [Bro96a]. Juan [CQ99]. Jubilee [Whi80]. Judith [Dys02a]. Judson [Gre81]. Julian [Dys50c, Ng96]. July [Dys00a, Dys10h, KLR13, MB66]. June [AAB^{+}88, Ano97, KLR13, Ros63]. just [Guh07]. Justice [Ano01, Dys98g].

ka-nu [Bro96a]. Kafatos [GBS91]. Kahneman [Dys11l]. Kaiser [Cao06]. Kaku [CQ99]. Kakutani [Yan54, Yan55]. Kalnajs [Dys77f]. Kameshwar [Dys91b]. kanū [Broxx]. Keldysh [AK00]. Kemmer [Dys11t]. Kennan [Dys82a]. Kill [Dys62e]. Kimball [She82]. Kinematic [Ler71]. Kinetics [BS90]. King [Wig90]. Kitaev [JRV22]. Kitchen [Dys93f]. Klaus [Sta03]. Klein [Dys62b]. kleinsten [Fis10, Fis12]. Knopf [Dys04d, Dys05f, Dys08b]. Knopf/Smithsonian [Dys08b]. Know [Dys11m, Dys11s, Dys12g, Dys12e, GD11]. Knowing [May07]. knowledge [Bea99, Cas03]. known [Ano95, Lew22]. Koblitz [Dys83b]. Kohn [DCCS94, GBB03, YHKO07]. Kónca [Dys83d]. Konfliktlösungen [BHJ^{+}00]. kong [Bro01]. Korean [Bro96a]. Kornfeld [Dys94a]. Kosmolot [Bro78b]. Krauss [Dys11o]. Kreisel [Odi96]. Kreiseliana [Odi96]. Kuhn [BHJ^{+}00]. kurtosis [Dys43a]. kwa [Bro96a]. Kwek [Bus15]. Kyoto [Dys08d].

L [Bus15, Don02, Dys58a, Dys89h, Dys98c]. Lab [Dys98b]. Laboratories [Rab78]. Labyrinth [Hal17]. Ladder [CM03]. Ladik [Spo22]. Lagrangian [LEdAB03, dJ94]. Lahav [Str00]. Lake [Dys10d]. L’Amérique [Dys80k]. Landau [ABP08, HAS10, RQ11]. landscape [LACTFF22]. Langevin [CH99]. Language [Str00]. Laplace [PS04]. Laplace-transformed [PS04]. Lapp [Dys62e]. Large [Alp73, Dys62e, Ler71, Bui11, Pas22, Wad81]. large- [Bui11, Wad81]. Large-Scale [Ler71]. Lars [Ano04]. Laser [Cor10]. Last [Dys91a, Alp73]. Later [Dys98e]. Latter [Dys58a]. lattice [ABP08, BCDL93, BD94a, BD95, BCCM87, DORQ12a, DORQ12b, sXcHmW87]. Lattices [Dys71c, Dys72c, Fol78, Fol80]. Laura [CQ99]. laureaty [Dai67]. L’avenir [Ham91]. law [BB03]. Lawrence [Dys11o, She82]. Laxenburg [BM83]. Layer [HMMD94]. Layzer [Dys70c]. leading [Bro15]. leap [Fis10, Fis12]. Leaping [Dys09b]. Leaps [Dys16]. Leben [DBB09]. Lecture [Dys51g, Dys52b, Dys66i, Dys72n, Dys09a]. lecture-note [Dys66i]. Lectures [DSW51, Dys92c, Dys06c, PM^{+}83, Dys51a, Dys71j, DARM83, Dys88c, Dys90h, Dys04f, Ano08]. Lee [Sim22]. Legendary [Ano17b, Sto20]. Legislators [Dys75b]. Lehmann [BK73, BMS61, SS67]. Lehmer [DFG10, DFG13]. Leighton [Dys91a]. Leland [Dys11o]. lemon [Dys92e]. Leo [Dys70c]. Leslie [Dys96i]. Letter [BD07, BCD^{+}65, DHD06, Dys59b, Dys71h, Dys71i, Dys81i, Dys89h, Dys11s, FD07, GD03, GD05, GD11, GFD04, LD08, MD95b, MD08, Mor62, PD08, Ros62, SD06, SD98, SDGK06, Tay69, Ten69, TD05, You69]. Letters [Alb75, ABG^{+}89, CDB^{+}73, Dys60i, DK70, Dys80a, Dys83c, Dys96h, DGB^{+}01, Dys05l, FD93, GSE^{+}88, LD90, Mav75, She82, SDL^{+}01, WDW89, Dys18]. leurs [Dys01b]. Level [HR22, LCW^{+}72, BDL93, RNM04, Bus15]. Levels [Bet47, Dys48c, Dys62g, Dys62h, Dys62i, DM63, Gun62, MD63, GC81]. Levenson [Dys03f]. Lewis [Dys62b, Gre81]. Li [Gre81]. Liane [San94]. Lie [Bar66]. Lieb [Dys67b, Thi05, BB03, LT91, LT97, LT01, LT05, Thi91, TL97, Thi01]. Life [Dys59c, Dys70a, Dys72i, Dys80j, Dys81b, Dys81j, Dys82g, Dys82h, Dys84h, Dys85g, Dys85h, Dys85e, Dys87c, Dys87e, Dys92b, Dys99h, Dys02e, Dys08g, DBB09, Dys10c, Dys11o, Dys13e, Dys14a, Dys15c, Gle92, PD08, Sha01, SGT^{+}95, Str00, Ano00b, Cam63, Can01, Dys85f, Dys02d, Dys03g, Dys07q, Dys10i, Hei80, Kra01, Kra11, Neu16, PC06, Dys09e, Str00, Hag87, Sha87, Sha01, Dys15c, Str00]. Life-Time [Dys59c]. Lifshitz [AFPT10]. Lifshitz-type [AFPT10]. Light [BD59, CCL^{+}11, Fin17, KL02, GFW11, GFW12, Joh20a]. light-by-light [GFW11, GFW12]. Light-cone [KL02]. Lighthill [Dys58c]. Lightman [Dys05e]. Lightness [Dys09b]. Like [Ano17c, Nar86]. limit [AB12, BD94a, BD95, SG06, Wad81]. Limitations [Dys92h]. limits [Cas03, SV^{+}08]. L’immaginazione [Cor97]. Linda [CQ99]. line [GS22]. Linear [Dys53d, Dys53e, Dys66c, Dys48a, For21]. linearized [Zho97]. Liouville [Nar86]. Lipsky [Dys62a]. Listening [Sta03]. Little [Dys13b]. Livable [Dys69c]. Live [Dys78c]. Liveright [Dys12g, Dys12e]. Lives [Gre81]. Livets [Dys87c]. Living [Fad90]. Livio [Dys14a]. ljagushki [Dai10]. Local [Dys58d, RY23, ADM23, DFPR22]. locked [NAW08]. locking [MNBW07]. London [Cur82, Dys60c, Dys72n, Sta03]. Long [Cor10, MD08, Dys71g, Lie04]. long-range [Dys71g]. Long-Term [Cor10]. Longitudinal [Dys50b, DM57a, MD57]. look [Dys03g]. Looked [Dys93e, Dys01a]. Looking [Dys03g, Dys08d]. Looks [Ano17b]. loop [DJS95, FM21, HBKK11, PR09, sXcHmW87, Xue88]. loops [Kri08]. Lorentz [Dys66d, HZ01]. Loss [Dys45b]. Low [BD59, NN86, NRT22, CDD56a, CDD56b, Dys57b]. Lucas [Sim22]. Lunik [Dys62a].

M [Dys60a, Dys62e, Dys87f, Dys11o, LW78, Sta03]. M. [Dys58c]. MA [Dys02a, Dys80k]. Macdonald [Zei87]. machine [Bro15]. Machines [Dys63d, Gre81, Bro15]. Made [Bro96b, Bro96c, Sax94, Tre94, Vil95, Way95, Sch94a, SGT^{+}95]. magnetic [LNA96, SHD91]. Magnetization [Dys69k, Hof11]. magnitude [Dys43b]. Magnon [Goc76]. Mainstream [Hor93]. Major [Dys90i]. Make [Dys75b, Dys06a, Lam06]. Maker [Dys18, She82, SM04]. Makers [Gre81]. Making [Yor87, Bro96c]. Malcolm [Dys13b]. Maldonado [CQ99]. Maleev [Goc76, RT82]. Mall [PSP^{+}07]. Malthusian [Dys60j]. Man [Dys05l, Dys08b, Dys10c, Dys11o, RL69, Kra11, Ng96]. Manchester [Dys80l]. Mandelbrot [Dys78b]. Mandelstam [HvSA98b]. Manin [Dys83b, Man07]. Mankind [Dys70g]. manuscript [Dys50c]. Many [Dys10i, Dys07q, HH13, MHFW72, NH91, SS99]. many-body [HH13, NH91, SS99]. Many-colored [Dys10i, Dys07q]. Mapping [DF92, CGR06, GL82, HGE88, KV85, Par87, She88, She89, SG93, SG06]. Mappings [Dys89i, VHK89]. Maps [Dys03c]. Marbles [Don02]. March [Mey56, Whi80]. Marcia [Sta03]. Margaret [Dys12f, Dys12d]. Margit [Dys98c]. Mario [Dys14a]. Mark [CQ99]. Markvart [SGT^{+}95]. Marlow [Dys98c]. Mars [Dys96f]. Marshall [Dys06b]. Martin [Ben13]. Martinsons [CQ99]. Maryland [Kon90]. mas [Ano95]. Mass [Dys53b, Dys62a, Dys09b, LFW10, Yos82]. Mass-Renormalization [Dys53b]. Massacres [WD15]. Masters [Gre81]. match [DORQ12a, DORQ12b]. Matematika [Dai65]. matematike [Dai10]. Materials [SGT^{+}95]. Materials/Mönch [SGT^{+}95]. Math [Dys89h, Joh20a, KTNG23, Tie09a]. Mathematica [BHJ^{+}00]. Mathematical [Jen14, KNGT22, KTNG23, LM66, Dys70a, Joh20b, LSW76, Lie04, ST20, DV63, Poo72, Dys67b]. Mathematician [Alb94, Dys70a, Ano20a, Ano20b, Can01, Rad20]. Mathematicians [Dys89h]. Mathematics [Dys64e, Dys66b, Dys79g, Dys80i, Dys81b, Dys06e, Man07, Ber84, Dai10, Dys10b, Dys12b, FF91, FR94, FR02, Hei80, Mah11, Pit11, Pit12, URR86, Dai65, Dys83b]. Mathematische [SGT^{+}95]. Matrices [Dys71d, FKS97, BBDY22, Pas22]. Matrix [Dys49b, Dys54c, Dys58k, Dys62c, Dys72d, Dys03i, ABD11, DJS95, Dys70e, EJ22, For21, Kri08, LACTFF22]. Matter [Dys66f, Dys66h, DL67, Dys97a, Dys07d, Dys07n, Dys11a, LD68, MD95b, MD95a, BF76, Fed75, Joh20a, LT91, LT97, LT01, LT05, MNK^{+}97, Thi91, TL97, Thi01, Thi05]. Mattis [Dys67b]. Maverick [Sch13, Ben13, Cre13]. Max [Dys70g, Dys05f, Fis10, Fis12, Dys70g, Fis10, Fis12]. Max-Planck-Vortrag [Dys70g]. Maximum [Dys55f, Dys55g, Dys58i]. Maxwell [And91, Ber84, Bre91, Dom91, Far91, Dys84g, Dys90e, Dys99j, Pom09]. May [BM83, CM90, Dys72n, Dys77g, GSS^{+}03, You69, Dys03g]. Maynard [Str00]. Mccutchen [Dys94a]. McRae [Lew99]. me [Dys06a]. Meadows [SGT^{+}95]. Mean [BD94a, BD94b, BD95, Ten69]. Meaning [Dys57b, Dys98b]. means [LLW^{+}09, Wüt13]. Measure [Dys59c]. Measuring [GD03]. Mechanic [Dys51g]. mechanical [Yeu80]. Mechanics [Ano04, Bet33, Coh71, Dys52b, Dys63e, Dys06c, Dys07a, Dys11k, Kal64, Dys51a, Dys51g, DM57b, Dys62k, Poo72, Dys54e]. medal [Dys91i, Ano69, Ano91, Dys66d]. Medalist [Whe91]. medium [NMM^{+}94]. Meet [Swe88, Lam06]. Meeting [Ano04, Noc08, Dys70g, Dys04b]. Mehra [Dys74a, SGT^{+}95]. Mehta [GC81]. Mem [Dys56d]. members [Ham91, Mit80]. membranes [Kle89]. membres [Ham91]. memoir [Dys99b, Dys10h, GM12]. Memoirs [Dys06g, Dys02a]. Memorial [Ano08]. Memoriam [Pat18, Fre08]. Memories [BR93, Dys81g]. memory [KNGT22, KTNG23]. Men [Bro96b, Bro96c, Sax94, SGT^{+}95, Tre94, Vil95, Way95, Fad90, Sch94a]. Menschheit [Dys70g]. Menser [CQ99]. Meson [BM55, Dys48d, DRS^{+}54, Dys55d, Mat49a, Mat49b, BEKW94, DD55, EF12, NMM^{+}94]. Meson-Nucleon [DRS^{+}54, DD55]. Meson-proton [Dys55d]. Mesons [Dys55e, DhX65a, HKR07, MEBK11, Dys55a, Dys56b, Dys55a]. Mesoscopic [AF95]. metabolism [BF95, BF96]. Metals [Dys55c]. metaphor [Man07]. Metaphorical [Gre81]. Method [Dys53b, Dys53g, Ano95, Dys51e, Maa06, Mar56, MND11, SVT09]. Methods [Mat49a, Mat49b, Mey56, IT01]. metodo [Ano95]. metric [YxQCxL10]. Meÿenn [SGT^{+}95]. Michael [CQ99, Dys03d, Dys08b, Wig90]. Michelle [Dys05l]. Michio [CQ99]. Microscopic [GL82]. microverse [Dys00d]. Mikhailov [Dys62a]. Millennium [Bea99]. Miller [SGT^{+}95, Sta03]. Million [Dys04c]. Mills [DORQ12b, AI02, DORQ12a, Wad81, WR07]. Mind [May07]. Minds [Gre81]. Minor [Dys79h]. Miracles [Dys99g]. MIRV [Dys69c]. Misfits [Dys13b]. Miss [Noc08]. Missed [Daĭ80, Dys72j]. Missile [Dys69a, Dys69n, Dys72l, McM65]. Missiles [Dys64d, Dys64f, Dys69i]. missing [GC81]. Mission [Fin17]. Mistake [Dys75b, Dys90m]. Mistakes [Dys14a]. Mitte [BHJ^{+}00]. mixed [LLW^{+}09]. mobile [BB03]. Mode [RS77, Ye92]. Model [AK00, BM87, Dys62c, DLS76, Dys82g, GD10, GMO03, MOR95, Wad81, AFPT10, AAL22, BMM98, BF95, BF96, BCCM87, CBBS11, CBBS12, CDM05, Dav92, DJS95, Det03, GL82, GMO98, GMO99, GMO00, HBKK11, JRV22, KPS90, Kri08, MM97, Meu01, OO23, SS99]. Models [Dys63e, Dys67b, Dys72f, AvM23, BLS03, DV63, KNGT22, KTNG23, LM66, McL13, Meu09, SZ21]. Modern [Dys80p, Ano08, Dys99f, For02, GSS^{+}03, Kev87, MB66]. modified [GLSS21]. molecular [DK06, SVT09, YHKO07]. Molecule [PD08, Dys77g]. molecules [DvL05, MMJ91]. momentum [DK06, DCCS94, MND11, MNLD11]. Monastyrsky [Wig90]. Mönch [SGT^{+}95]. Money [BHJ^{+}00]. Monk [Dys13e]. monotonicity [CL22]. Monte [Mar56, Mey56]. Montoya [CQ99]. Moon [Dys62a, RL69]. Mormon [Dys78e]. Morrison [Fre08]. Morson [Dys89d]. Moss [BHJ^{+}00]. Most [BR93]. Mostly [Dys12i]. Motion [Dys62c, tL84, NS93, OK07, OK09]. motions [ADvM09]. MR0994963 [Dys89h]. MR1307454 [BD95]. mu [Bro01]. multi [HR22]. multi-dimensional [HR22]. multipath [DMZ76]. multiplicative [CP90, KP09]. Munczek [Det03]. mundo [DG05]. Mundos [CQ99, Nav99]. muon [GFW11, GFW12]. Murders [WD15]. Museum [Dys08b]. My [Dys80k]. Myrick [Dys11o]. mysteries [RD99]. Mysterious [Dys05e]. Mystic [Dys10c]. myths [Bro96c]. myths-in-the-making [Bro96c].

N [Bus15, Dys60a, Dys62a, Dys06b, LS22, MND11, MNLD11]. Nadeau [GBS91]. Nadezhda [Dys89j]. Name [Ano70]. Named [San94]. Nanyang [PKpCH14]. NASA [DLV10]. National [Dys08b, GR63, Han82, Sta03, Dys85i]. natura [Cor97]. Natural [Dys06d]. Nature [Cor95, Cor97, Dys71f, Dys72e, Dys79c, Dys06e, Gam68, Wil58, Alp73, SG93, Dys74a, Dys96g]. Neal [Dys83b]. Near [Fin17]. Near-Light-Speed [Fin17]. Nearest [Dys92f]. nearing [SV^{+}08]. Need [Dys13f]. Needs [Dys03d]. Negative [Dys55f, Dys55g, OK09]. Neighbor [Dys92f]. Nejtronnye [DC73]. Nemirovsky [Det03]. Networks [Sta03, Pas22]. Netze [SGT^{+}95]. Netze/Enz [SGT^{+}95]. Neuenschwander [Bai17]. Neufeld [Dys08b]. Neumann [Dys81b, CK12, Dys13d, Hei80]. neural [Pas22]. Neuronaler [SGT^{+}95]. Neutral [CDD56a, CDD56b]. Neutrality [NAW08]. Neutron [Dys56g, Dys59c, Dys61b, Dys69p, Dys71j, LCW^{+}72, RRD^{+}73, BM83, Dys20, DC73]. Neutron-Resonance [LCW^{+}72]. Newton [Dys07e, Dys03b, Fei03, Dys03b]. Next [Cor10, Dys71k, Dys73c, Dys77d, BDF00]. Nicholas [Dys11t]. NII [Dys80b, Dys81a]. nineteen [Dys01a, Dys93e]. nineteen-forties [Dys01a, Dys93e]. Ning [Dys84b, LY95]. Nishimori [GS22]. Nishina [Ano08]. NJ [Han82]. NL [BM99]. No [Dys74b, Dys75b, Dys77c, Dys84n, Dys02c, Dys02g, Ano95, BD95, Broxx, Dys89h, Dys02d]. Noam [Str00]. Nobel [Cur82, Dai67, Dys65e]. nobelevskoj [Dai67]. Noise [Dys75c]. Non [BDL93, Dys69k, Dys95b, Fre84, Dys48a, Dys84m, Dys85i, SVT09]. non-Dyson [SVT09]. Non-Existence [Dys69k]. non-fiction [Dys85i]. non-homogeneous [Dys48a]. Non-Nuclear [Fre84, Dys84m]. Non-Use [Dys95b]. Non-Violence [Dys95b]. Nonequilibrium [Kal64, BCDM01]. nonintersecting [ADvM09]. nonisotropic [DLS78]. nonlinear [Bui11, Meu09, Spo22]. Nonperturbative [CP93, CBPG84, RBB95]. Nontrivial [DORQ12a, DORQ12b]. Nonunique [BCS89]. Nonunitary [SG93]. Norbert [Dys81b, Dys05g, Hei80, TD05]. Nordhaus [Dys08d]. normale [Dys71j]. norms [SPZ^{+}12]. Norton [Dys85a, Dys03c, Dys11o]. Nossack [Dys05f]. Note [DORQ12b, Dys45b, Dys58i, Dys98e, Dys11u, Lar89, Ano93, BF76, Dys43a, Dys66i, Mar56]. Notebook [Dys08a]. Notebooks [Dys90k]. Noted [Gol08]. Notes [And80, Dys50c, Dys52b, Gre81, Dys51a, Dys51g, Dys60c]. Nottingham [Dys96b]. November [Dys88c, Dys90h, Dys99b, Dys04f, Pei80]. nu [Bro96a]. Nuclear [Ano60, Bar66, BM55, Blo04, Dys58a, Dys60f, Dys61a, Dys62e, Dys63f, Dys64b, DGWW67, Dys70b, Dys82a, Dys85a, Dys85j, Dys86b, Dys95a, Fre84, MHFW72, NHKK94, NMM^{+}94, Pei80, RRD^{+}73, ACN85, BM83, Dys66i, DARM83, Dys84m, Ell86, Joh20a, tL84, MNK^{+}97, BM83, Dys64g, Dys84a]. Nuclear-Age [Dys84a]. Nuclei [DW48, NH91, RS77]. Nucleon [BW69, DRS^{+}54, EN12, EF13, BEKW94, DD55, NN86]. Nucleons [Dys48d]. Number [Dys00f, BMM98, BCDL93]. Numbers [Alp73, Dys47a, Dys96g]. numerators [Meu01]. numerical [GS04a, Tak89, NMM^{+}94]. Numerological [Dys68b]. Numerology [Gam68]. Nye [Dys85a].

O [DZO11, Dys98d, GGMO07, She82, Wig90, Fre92, GGMO07, MND11, MNLD11]. Obituary [Ach20, Ano20b, Ano20c, Dys93f, Sto20, Ano20a, Dys56d, Rad20]. Objects [Dys68b, Dys97a, Gam67b]. Observatories [Dys90i, Kon90]. occasion [MB66]. Occultation [Dys90j]. Occurrence [BHJ^{+}00]. Oct [Ano01]. October [Cur82, Dys11t]. Odyssey [Ben13, Cre13, Lew99, Kra01, Sch13, Yor87]. Oersted [Dys91i, Whe91, Ano91]. Off [Bas11, GM12]. offenen [Dys89m, GBS91]. Offensive [Dys79a]. ogni [Dys89g]. Oguchi [LD66]. ohne [Dys89m, GBS91]. Oklo [DD96a, DD96b]. Okubo [CL22]. Old [Dys65c, Dys76e]. Oliveira [BHJ^{+}00]. omega [NEA10]. One [AH66, Bet33, Dys67b, Dys69k, Dys69m, Dys71f, Dys72e, Dys04c, LM66, Cas03, Dys69f, GLS22, SLA93]. One- [Bet33, Dys71f, Dys72e]. One-Channel [AH66]. One-Dimensional [Dys69k, Dys69f, GLS22, SLA93]. One-Hertz [Dys69m]. O’Neill [Dys85b, Dys93f]. Onsager [Ano04]. Opacity [BD59, BD03, Dys58i]. Open [DD04, Dys79m, Dys83d, Dys89m, PVW02, PVW03, SRBW05, dJ94]. open-shell [PVW02, PVW03]. Opening [Dys66e]. operational [JL86]. operator [GZ02]. Operators [Pra64, DG84, Dys51c, Dys51d, GZ87, Kri08]. Opinions [ABG^{+}89, Dys89a]. Oppenheimer [Ber04, Dys76a, Dys83c, Dys13e, She82, Ano54, Ano70, Ano95, Ber04, DPSY64, Dys79c, Dys80a, Dys06k, GDH^{+}04, PC06, Ber04]. Opponent [Blo04, PD12]. opportunities [Daĭ80, Dys72j, Dys58a]. Optical [Dys75c, HMMD94, RM06, RWP78]. optical-waveguide [RM06]. Optics [Dys78c]. optimistic [Dys08f]. Options [Dys08d]. Oral [Aas86, Dys84o]. orbit [Kon90]. orbital [DK06, GBB03, MNM94, SS98]. Orbitals [TMD96a, TMD96b, CCC^{+}10, DZO11, DZO13, DCCS94, GGMO07, MND11, MNLD11, OK07, OK09, Ort04, SWFB05, SOM10, TMD96c, YHKO07]. ordenador [CQ99]. ordenadores [CQ99]. Order [BDS52, Bas22, Dys43b, Dys71g, HH13, IO08, PVW02, PVW03, VPW01, WR07, WKH13]. ordered [GZ87]. Origin [Dys81i, Dys82g, Dys82h, Str00, BF95, BF96, Kri10, Str00]. original [Cam63]. Origini [Dys87e, Dys02e]. Origins [CK12, Dys84h, Dys85g, Dys85h, Dys85e, Dys85f, Dys99h, Dys08g, Str00, Dys87e, Dys02e, DBB09, Sha01, Str00, Hag87, Sha87, Sha01]. Orion [Ber02, Kan02, Dys02h, Pei80]. Orthogonal [Pas22]. Oruzhie [Dys89j]. oscillator [HR22, Yeu80]. Other [Ber18, Dys63c, Dys04c, Spe91, Sta03, Str00, Ber19, Dys84q, GLSS21, Man76a, Man76b]. Ottaviani [Dys11o]. otwartym [Dys83d]. Our [Bek86, Dys62e, Dys74c, Dys07r, Dys14a, FD07, RD99, SD06, DBB09, Fad90, BHMD07, Dys58a]. outer [YHKO07]. outliers [ADvM09]. Overkill [Dys62e]. Overlook [Dys07e]. Overview [Gre81, Str00]. OW [Ano13]. Owls [Dys85a, ACN85]. Own [Ano96, Bek86, Fra08]. Oxford [Dys02d, Sta03, ABD11].

P [Bus15, Dys62a, Dys62e, Dys94a, Dys98b, Dys02d, Dys05l, Fey00, Has79, LW78, She82, Spe91]. Page [SGT^{+}95]. Painlevé [Dys95e]. Pair [DÜ55, MD57, SLA93]. Pais [BHJ^{+}00]. Palindromic [Dys00f]. Palo [Dys70c]. pantalla [CQ99]. Pantheon [CK12, Dys03b, Dys05e, Dys09c]. Paper [Dys64c, Dys69b, Dys08d, Dys08c]. paperback [Bro96c, Sta03].

Papers [Dys84b, Sch58, Sch03, BM83, Dys96e, Jon16, Yan13]. Para [Ano95]. Parametric [KC73, CK74]. parametrized [YxQCxL10]. Part [Dys58b, Dys63b, Dys84d, Dys84e, Dys85g, Dys85h, Gar61, ST12, Dys06l]. partial [Dys43b]. Partially [BS90, SVT09]. Particle [Bar66, Dys64b, Dys66a, Dys07m, Dys07o, Dys11f, Dys11g, BCTG84, CDM05, Dys66i, KV85, MHFW72]. Particles [BH34, Dys67c, Fis10, Fis12, LM66, Dys67b]. Partition [Mah05]. Partitions [ADH88, Dys44b, Dys12c, Dys69j, Dys89i]. Partly [BB63]. Partly-Baked [BB63]. Parts [Dys90k]. Passage [Pei85]. Passions [Sta03]. Past [Dys92h]. pathways [MMJ91]. Patrick [CQ99]. patterns [DZO13, Dys18]. Paul [Dys08c, Dys10c, Dys87f, Dys90b]. Pauli [Dys60c, SGT^{+}95, DGB^{+}01, Dys02d, Dys02d]. PCT [Dys64c]. PDEs [AvM05]. Peace [Dys80c, Yor87]. Pedagogical [Dys80b, Dys81a]. Peierls [Lee07]. Penguin [Dys06d, Gar61]. Pennies [Dys46]. People [Dys84k, Spe91, URR86]. perechody [Daj73]. Perfectly [Dys05l]. perfezionamento [Ano97]. Pergamon [Dys60c]. perhaps [Guh07]. Peril [Dys02b]. Period [DF92]. periodic [PS04]. Perpendicular [Hor93]. perpetuum [BB03]. Perseus [Dys02a]. Persistence [AW69, DM57a]. personal [GM12]. Perspective [Dys81k, Ano00b, BM83, Sag00]. Perspectives [MB66]. perturbation [DFPR22, Dys52a, HH13]. perturbative [SS98]. Peter [Dys80c, Dys03c, She82]. pg [Dys03a]. Phänomene [DBB09]. Phase [Daj73, Dys69l, Dys71f, Dys71l, Dys72e, DLS76, DLS78, BBDY22, CJ80, Dys69f, GMO00, HS98, NAW08]. phase-transition [Dys69f]. Phenomena [BM87, DBB09]. Phenomenon [Dys06d]. Philip [Fre08]. Phillip [Ben13]. Phillips [Dys70c]. Philosopher [Dys15a]. Philosophical [Bea99, Ano95]. philosophies [Fad90]. Philosophy [Rus12, SGT^{+}95, UM86]. Philosophy/Mehra [SGT^{+}95]. photodetachment [OK09]. photoelectron [DZO13, OK09]. photographs [Dys05f]. Photon [Dys75c, EF12, KP09]. Photons [Dys50b]. Phua [Bus15]. Phys [And91, Bre91, Dom91, DORQ12b, Far91, GFW12, HMSW00a, KTNG23]. Physical [Ano99, BHJ^{+}00, Dys64e, GBB03, JR80, Tel48, DV63, For02, HGE88, tL84, Par87, SG86, Whi80]. physicien [Dys09e]. Physicist [Alb94, Dys68d, Dys74a, Dys15c, Gol08, IT23, Lam06, Ove08, Pei85, San94, Ach20, Ano20a, Ano20b, Can01, Dur00, Dys09e, Joh20b, LY95, Ng96, Rad20, ST20, Sto20, Yor87]. Physicists [Ano69, Dys70g, Dys71i, HH04a, Kev87, Bar66]. Physics [Bar66, BM55, BGA^{+}13, BM87, CCL^{+}11, Dai65, Dys50d, Dys58e, Dys64b, Dys64a, Dys66a, Dys66b, Dys67b, Dys76b, Dys78d, Dys79g, Dys79m, Dys80e, Dys80i, Dys83b, Dys89m, Dys91i, Dys93a, Dys93e, Dys98e, Jen14, LM66, Met73, RW64, Sax94, SGT^{+}95, Sta03, Whe91, yWHH98, Ano08, BL06, Dai71, Dai10, Dys66i, Dys70f, Dys72b, Dys72g, Dys72h, Dys83d, Dys01a, Dys02d, Dys10b, FF91, GSS^{+}03, Kai05a, KLR13, LSW76, Lie04, MB66, Mon87, Mon99, MHFW72, Dai67, Dys65e, Dys79c, Gre81, BCTG84, Bar66, Cao06, Dys83b, Dys90b, Dys98c, Dys07f, Wig90, Dys12f, Dys12d]. Physics/Markvart [SGT^{+}95]. Physik [Dys58b, Dys89m, GBS91, SGT^{+}95]. Physik/Sutton [SGT^{+}95]. Physikertagung [Dys70g]. Pi [Ber04]. Picard [CQ99]. Picasso [Sta03]. Picture [Dys11o, MNM94]. Pierre [Dys97a]. Pierre-Gilles [Dys97a]. Pilgrim [Dys78e]. Pion [BW69, Dys55f, Dys55g, LFW10, NN86]. Pion-Nucleon [BW69, NN86]. Pioneering [Ben13, Cre13, Sch13]. Pioneers [Dys78e]. Pisa [Ano97, Ano97, Dys71j]. Pitaevskii [Bas22]. place [Dys07q, Dys10i]. placed [HG87]. places [Dys03g]. Plain [Dys89d]. planar [BCCM87, RT82]. Planck [Dys70g, Fis10, Fis12, Ano69, Dys70g, Fis10, Fis12]. plane [LACTFF22]. Planetary [Vog99]. Planets [Dys91g, Dys92d, Dys03g]. Plants [Dys97h]. Plaquette [sXcHmW87]. platonic [Cas03]. pleasure [Fey00]. poet [Dys06k]. Poetry [Dys83a, Dys09c]. Poincaré [Dys03c, HZ01]. Point [Met73, CP90, HT95]. points [BCDL93, BD94a, BD95]. Poisson [AF95, RNM04, TP01]. Polaritons [BS90]. Polarization [BDS52, Dys57c, DM57a, MD57, Dys50c]. Pole [ST67, NN86, Nak87]. Poles [AH66, BW69, JK67, JK68, Roc71, Dab75, Kri10]. Policies [Dys08d]. Policy [McM65, Dys77g]. Polish [Bro78b, Dys83d]. Politics [Dys02a]. Polkinghorne [Dys98b, Dys02c, Dys02g, Sta98]. Polyakov [HBKK11]. Polyakov-loop [HBKK11]. polyatomics [SPZ^{+}12]. Polymers [För96]. Pomeranchukon [WK74]. pondered [Joh20a]. Pontecorvo [Dys15c, Dys15c]. Population [Olt73b]. por [Ano95]. Portas [Dys95f]. Portrait [Ber04]. Portuguese [Fre92]. Positive [BH34]. Possibilities [Dys11n, Ano20d]. Possible [AH66, Bas11, Dys06e]. Postdoc [Kai05c]. postgraduate [Bus15]. Postscript [Sch94b]. Postwar [Cao06, Hor06a, Fin06, Kai05a]. Potential [BW69, Dys07s, Dys11i]. potentials [GBB03, IO08, Ort04]. Power [Lew99, Ano71]. Powers [May07]. Pp [Dys96i, Sta03, Ben13, Ber02, Bro96c, Dys58b, Dys58a, Dys60c, Dys60a, Dys60b, Dys62a, Dys62b, Dys64c, Dys67b, Dys70c, Dys70a, Dys96j, Dys98b, Dys02a, Dys02d, Dys02b, Dys02c, Dys03c, Dys03d, Dys03b, Dys04c, Dys04d, Dys05f, Dys05e, Dys05g, Dys05h, Dys05l, Dys06d, Dys06e, Dys07e, Dys07f, Dys08d, Dys08b, Dys08c, Dys09b, Dys09c, Dys10c, Dys10d, Dys11n, Dys11o, Dys11l, Dys12f, Dys12g, Dys12d, Dys12e, Dys13e, Dys13b, Dys14a, Dys15a, Dys15c, Hor06a, Pei80, Dys80c]. Practical [Dys95j]. Practicing [MOG^{+}13]. Praise [Dys85c, Dys02b, DD04]. precision [GMON98]. Predictable [Dys75b]. Preface [Dys87g, Dys89k, Dys91h, Dys92g, Dys94c, Dys97f, Dys01c, Dys02f, Dys05k, Dys11v]. Preliminary [Dys96b]. Prelude [Dys79h]. premii [Dai67]. Prentice [Ber04]. presentation [Dys70g]. presented [Dys91i, Dys00a]. Press [Ben13, Ber04, Bro96c, Cao06, Dys60c, Dys60b, Dys67b, Dys80c, Dys98b, Dys02d, Dys02c, Dys04c, Dys05f, Dys05h, Dys06e, Dys08d, Dys08c, Dys10d, Dys15a, Sta03, Dys81b, Dys92c]. Press/Harvard [Dys10d]. Press/Prentice [Ber04]. Prey [Dys03d]. Price [Ber02, Dys02a, Dys10d, Dys58b, Dys02d]. Primary [Sta03]. Primer [Dys92c]. Princeton [Bro96c, Dys77e, Han82, Sta03, Ano20c, Dys77e, Mah11]. Principle [Spe91]. Principles [BHJ^{+}00, Dys60j, Dys58b]. Prism [UM86]. Prisoner [PD12]. Private [Dys84c]. Prize [Dai67, Dys65e, Dys77a, Fed00, Gan00, Ano70, Ano82, Ano00a, Mon94, SDL^{+}01]. Prizes [Dys89h, Ano99]. Probing [Dys02b]. Problem [BMS61, Dys46, Dys76e, Goc76, BD94c, Fed75, RW21, ZDSN08]. Problema [ZDSN08]. Problems [Bat50, Bat51, BTD^{+}47, Bet33, BS51, BG49, DC52, Dys50a, DTS51, Dys64f, Dys98c, Dys07m, Dys07t, Dys11f, Dys11j, EW50, ES51, FZ52, Gal50, Gal52, GB48, HJ65a, MS50, RRS48, San50, Thé47, TL48, TB49, Thé50, TK51a, TK51b, UL51, Wil48, WG50, Wil50, WL51, Dys76d, FR94, FR02, Kap73, UTB^{+}50]. Proceedings [BM55, Dys74a, JR80, AAB^{+}88, Bus15, Bye87, Kon90, Whi80, CM90, PKpCH14, Ros63, DV63, Han82, RWP78, Bus15]. process [GLS22]. Processes [Dys63b, AvM05, Bui11]. Processing [Dys78c]. product [Dys48a]. Production [BHJ^{+}00, DÜ55, MD57]. Prof. [Dys56d]. Professor [Bai17, yWHH98, Dys50c, Dys51a, Neu16]. Profile [Hor93]. Profiles [Gre81]. Program [Ano04, Has79, Jud72, Bea99]. Progress [Dys83b, Dys00e, Gan00, Lew99, RD99]. Project [Dys65a, Dys02h, Ell86, Ber02, Kan02]. projection [tL84]. Proof [Dys44a, GX06, Goo70, Gun62, Kad85, Pom09, Wil62, And91, Bre91, Dom91, Dys90e, Far91, Zei87]. Propagator [WR07, ABP12, AIP13, CP93, Det03, IO08, MM97, MNM94, RQ11, SVT09]. propagators [ABP08, BEKW94, DORQ12a, DORQ12b, KP09, KSW13]. Propelled [Dys72f]. Properties [För96]. property [CL22]. proportion [Woo80]. proportional [FM21]. Proposal [Dys59c]. Propulsion [Dys82f, Dys85d, Cor10]. prospects [Kap73, RD99]. Prosperity [Lew99]. Protestors [CDB^{+}73]. Proton [Dys67d, Dys55d, DhX65a]. protoneutron [CBBS12]. Pseudopath [SRBW05]. Pseudoscience [Dys04c]. Psychology [Dys80b, Dys81a]. PT [BMS00]. PT-symmetric [BMS00]. Pticy [Dai10]. Public [Dys81m, Joh20b]. Publisher [DORQ12b]. Publishers [Dys62a, Dys62b]. Publishing [Dys62a]. Pugwash [Dys62f]. Pulsars [DC73, Dys69o, Dys71j]. pul’sary [DC73]. Pupin [Rab78]. pure [SZ21]. Pursuits [Dys81n, Dys82b, Dys82i, Dys84r, Dys16, Dys83g]. Puzzle [Tie09a, Tie09b].

QCD [HMSW00a, AMN02, BP08, BK10, HS98, HvSA98a, HMSW00b, HBKK11, MM90, NO99, Zwa03]. QED [SGT^{+}95, BFW11, Bro96c, CLO10, CP90, CP93, KP09, KSW13, RBB95, Rem90, Sch94a, Bro96b, Sax94, Tre94, Vil95, Way95]. qi [Bro01]. quadratic [ADH88]. Quantenelektrodynamik [Dys55b]. Quantenmechanik [Bet33]. Quantensprung [Fis10, Fis12]. quantization [Zwa03, dJ94]. Quantized [Dys60a, Dys07i, Dys11d]. Quantum [Dys49b, Dys50b, Dys50e, Dys51g, Dys52b, Dys52c, Dys54e, Dys58b, Dys58k, Dys63e, DLS76, Dys92h, Dys03i, Dys05d, Dys06j, Dys06c, Dys07a, Dys10c, Dys11k, Hal17, Kra11, Sch58, Sch03, SGT^{+}95, BDH04, BMS00, Bui11, CH99, CBPG84, DJS95, Dys51a, Dys51c, Dys51d, Dys51e, Dys51f, Dys52a, DM57b, Dys62k, DLS78, Dys96c, Fis10, Fis12, GZ02, HR22, KLR13, Kon97, NS93, Nar86, Oko91, SW72, SLA93, SRBW05, TP01, Yeu80, Bet33, Dys11o]. Quark [LEdAB03, MNK^{+}97, CBBS11, CBBS12, KL02, LFW10, MM97]. quartic [EG11]. Quasi [Dys68b, EG11, Dys65d]. Quasi-exactly [EG11]. Quasi-Stellar [Dys68b, Dys65d]. Quasifermion [HN08]. quasiparticle [PS04]. Quasistellar [Gam67b]. Quaternion [Dys72k, Dys73b]. quenched [CP93]. Quest [Dys80m]. Question [Dys08d, NZSD08]. Questions [Dys63b, Dys84i]. Quick [Dys80n, Dys83e]. quien [Ano95]. Quotable [Sta03, Cal96, Cal00, Cal05, Cal10]. quotients [Dys43b].

R [Ber04, Dys64c, Dys98c, Wig90]. R. [Dys94a]. R.S. [Dys56d]. Racah [Dys60b]. Race [Dys97g, Dys05e]. Radar [BCC^{+}08]. Radiation [Dys60k, Dys01d, Dys49a, Dys58j, Dys03h]. Radioactivity [PDMS80]. Radiotelepathy [Dys10a]. Raffiniert [BHJ^{+}00]. Ralph [Dys62e, Dys91a]. Ramanujan [AAB^{+}88, AO05, BR01, Dys88d, Dys01f, Dys90k, Dys96h]. Ramón [CQ99]. Rampage [Dys12f, Dys12d]. Random [Dys62c, FKS97, AvM23, ABD11, BMM98, BBDY22, BF95, Bui11, Dys70e, EJ22, For21, LACTFF22, Pas22, Dys87a]. randomly [HG87]. range [Dys71g, Woo92]. Rankin [Dys96h]. Rare [Dys99g]. rate [Dys54a]. Rates [Dys45b]. rather [Ano95]. Rationales [McM65]. rationals [Dys47a]. Ray [Dys13e, BS89, HG87, Mor58]. razón [CQ99]. Reaching [Dys80o]. Readable [Dur60, Dur62]. Reader [Spe91]. reading [Jud72]. Real [Dys71d, Dys79g, Dys80i]. Reality [Dys96i, Hal17, BDH04, DBB09, Dys04d, DG05]. Really [Dys12g, Dys12e]. Reappraisal [Dys76a]. reason [BM99]. Reasonable [Dys05l]. reasons [ALW^{+}88, Dys03g]. Rebel [Dys89c, Dys95h, Dys95i, Dys06m, Dys08h, Ano95, Dys96d, Dys06n, DS07, Guh07, Joh07b]. rebelde [Ano95, Dys08e, Dys10f]. Rebellion [Sta03]. Receive [Gan00]. Receives [Ano91, Ano94, Mon94]. Recollections [Dys80a, Fie79, Pei85, She82, Dys83c]. recombinant [Dys76c, Dys77g]. Reconnaissance [Dys62a]. Recreational [Dys12b]. rectangles [RW21]. reducible [dJ94]. Refined [DORQ12a, DORQ12b]. Refined-Gribov [DORQ12a]. Reflections [Ben97, Ber18, Dys61b, Dys70h, Dys79i, Dys79j, Dys79k, Dys84i, Dys84j, Dys84k, Dys84l, Gre81, Ber19, Dys07q, Dys10i, Fad90]. reflexions [Dys09e, Dys09e]. Reformulation [Tam83]. Refusenik [AF81]. Regime [AF95, Bas22]. region [BBDY22]. Regularity [Dys58d]. Reid [Dys70a]. Reise [BHJ^{+}00]. Related [BMS61, Dys63b, BD94b]. Relation [Dys54c, AAK^{+}12, LLW^{+}09, Yan10]. Relations [Dys62b, Dys63f, Dys82a, GBB03, Nak87]. Relative [DW48]. Relativistic [AW69, Dys51g, Dys53c, Dys53h, MD57, NH91, Dys51a, Dys55d, GZ02, SA03]. Relativitätstheorie [SGT^{+}95]. Relativitätstheorie/Wong [SGT^{+}95]. relativity [Dur60, Dur62, HZ01, Dys60c]. Relaxation [BF96, SHD91].

Release [Dys69e]. Relevant [Dys68b]. Religion [Bye87, Dys86f, Dys87h, Dys00e, Dys02c, Dys02g, Dys06d, Fed00, SDL^{+}01, Bye87, Gan00]. Religious [Dur00]. Remarks [Dys66e, Dys66f, Dys85i, Dys91d, Dys96b, Kri10, SS67, RQ11]. Remembering [GDH^{+}04]. Reminiscences [She82]. Removal [LD66]. renaissance [Ach20]. renormalizability [CP90, KP09]. renormalizable [CP93]. Renormalization [DD55, Dys53b, Dys51e, HB12, Man76a, Man76b, Man77, Man82, Wal99]. Renormalized [CDM05, Man78]. Renormalizing [CMD04]. Rent [Hor06a]. Replica [JRV22]. Replies [Dys65b]. Reply [Dys88a, Dys89a, Dys94a]. Report [Dys77e]. Representation [BK73, Dys58f, Dys60h, Per68, SS67, BD94b, DG84, KL02, Wad81, Wal99, Wüt13]. Representations [Dys58g, Dys58h, GS85]. Representatives [Dys77g]. Reprint [Dys67b, Dys66i]. reprints [Cam63]. repulsion [SLA93]. research [Dys76c, Dys77g]. researcher [Bus15]. Researchers [Bas11]. Reshaping [Sax94]. resolved [VWS08]. Resonance [Dys55c, LCW^{+}72]. Resonances [HJ65b, MNM94]. Resonant [HMMD94]. Resource [Olt73b]. Resources [BHJ^{+}00]. Respect [SD06]. Respectable [KAD14]. Response [Dys69m, Dys96b, Fai94, Kor94, Sta98]. rest [DhX65a]. restrict [CP90]. Results [Dys07s, Dys11i, ABP08, GS04a]. Reta [CQ99]. Retirement [GSS^{+}03]. Return [Dys81l]. Rev [DORQ12b, GFW12, HMSW00a]. Review [Ano80, Ano96, Bai17, Bar82, Bar66, Ben97, Ben13, Ber19, Ber02, Ber04, Bet79, Bro78a, Bro96b, Bro96c, Buc87, CK12, Cao06, CQ99, Cre01, Cre13, Don02, Dys54d, Dys55a, Dys55b, Dys56b, Dys58b, Dys58a, Dys58c, Dys60c, Dys60a, Dys60b, Dys62a, Dys62b, Dys62e, Dys64c, Dys64g, Dys65d, Dys67b, Dys70d, Dys70c, Dys70a, Dys72l, Dys74a, Dys76a, Dys78b, Dys79d, Dys80c, Dys80b, Dys81b, Dys81a, Dys82a, Dys83b, Dys83c, Dys84a, Dys84b, Dys85a, Dys87a, Dys89d, Dys90b, Dys90k, Dys91b, Dys91a, Dys92b, Dys92c, Dys96h, Dys96g, Dys96j, Dys96i, Dys97a, Dys98b, Dys98c, Dys02a, Dys02d, Dys02b, Dys02c, Dys03c, Dys03d, Dys03b, Dys03f, Dys04c, Dys04d, Dys05f, DG05, Dys05e, Dys05g, Dys05h, Dys06d, Dys06e, Dys07e, Dys07f, Dys08d, Dys08b, Dys08c, Dys09b, Dys09c, Dys10c, Dys10d, Dys11n, Dys11m, Dys11l, Dys12f, Dys12g, Dys13e]. Review [Fie79, Fre84, Gom84, Gre81, Guh07, Hag87, Has79, Hol81, Hor06a, Joh07a, Joh07b, Jon16, Kan02, Lew99, Nav99, Pie84, Rom89, Sax94, Sha87, Sha01, Spe91, Ste92, Str00, Tre94, Vil95, Way95, Wig90, dS99, vH84, Dys98d, Fei03, GF04, Lar89, Sta98]. Reviews [Dys11o, She82, Sta03]. revised [Dys60a]. revisited [AAB^{+}88, DD96a, DD96b, Dro95, RY23, SG93]. Revolution [Dys07f, Gre81, Dys71k, Dys77d, Dys07e]. Revolutionary [Dys03e, Dys99a]. Revolutionized [Hal17]. Revolutions [BHJ^{+}00, Dys92i, Dys96l, Dys99i, Dys01b]. Revolutions/Pais [BHJ^{+}00]. Rezensionen [GBS91]. Rhenium [Dys68c]. Rhodes [Dys92c]. Richard [Dys62a, Dys91a, Dys92b, Dys92c, Dys98b, Dys05l, Dys09c, Dys11o, SGT^{+}95, Spe91, BR93, Dys11o, Dys11u, Fey00, Gle92, Hal17, Kra11]. Richardson [SG06]. Ricolfi [GBS91]. ride [Dys79l]. Riemann [Mon87, Mon99, Wig90]. Riesz [Lew22]. Right [You69]. Rights [PD13]. Rings [Dys12f, Dys12d]. Riordan [Swe88]. ripples [Dys16]. Risk [Mav75]. risky [Dys16]. Roald [Dys96j]. Robert [Bar66, Ber04, DPSY64, Dys92c, Dys96h, GDH^{+}04, She82, Dys80a, Dys94a, She82, Dys83c, Dys13e]. Robin [RW21]. Robinson [Dys65d]. Rocket [Dys08b]. Rodman [Dys62a]. Role [Dys84m, Dys10e, Dys11p]. Romance [Gre81]. Romantic [Dys09c]. Root [Dys44a, Zei87]. Roots [Dys86e]. Ros [CQ99]. Rosalind [CQ99]. Rosenbluth [Dys06b]. Rotating [DTO78]. rotator [BCCM87]. Rotblat [BM99]. Rouet [dJ94]. Routledge [Dys96i]. Row [Pei80]. RPA [SS99]. Rules [Eps55, KC73, CK74]. Running [Kon97]. Russell [San94]. Russia [BCD^{+}65]. Russian [Daj73, Dys60a, Dys62a, Dai65, Dai67, Dai71, Dai10, DC73, Dys89j].

S [Bro96b, Bro96c, Dys56b, Dys58b, Dys62e, Dys64c, Dys85a, Dys91b, ST11, Tre94, Vil95, Way95]. S. [Dys56b]. Sachdev [JRV22]. Sagan [Gre81, Sag00]. Sage [Ano16, Guh07]. Sake [Dys81m]. Sakharov [LD90]. Salam [Dys99b, Man76a, Man82, Man76b, Man77]. Salpeter [HG87, KK02, KK98, Ler71, Mun95, NO99, NEA10, SA03]. SALT [Kap73]. Salzburg [Dys70g]. Same [Dys96j, Dys96j]. San [Dys00a]. Saul [Dys89d]. Saying [Dys74b, Dys75b, Dys77c, Dys84n, Ehr75]. Scalar [CDD56a, CDD56b, Roc71, Det03, GMON98, Oko91, SA03]. Scale [Ler71]. Scaling [BK73]. Scatterer [Dys55e]. Scattering [AH66, CDD56a, CDD56b, DRS^{+}54, Dys54c, Dys55e, Dys55f, Dys55g, Dys57b, Dys76e, Dys07m, Dys07o, Dys07t, Dys07s, Dys11f, Dys11g, Dys11j, Dys11i, DD55, DCCS94, Dys55d, Dys76d, EF12, EF13, GFW11, GFW12, McL13, NN86, Nak87, SRBW05, WK74]. Schell [Dys84a]. scheme [BP08, Man76a, Man76b, Man77, Man82]. schemes [Man76b, SG86, Man76a]. Schewe [Ben13]. Schild [Dys65d]. Scholar [Mcc94]. scholars [Mit80]. Scholz [GBS91]. Schrödinger [Dys51f, Spo22]. Schucking [Dys65d]. Schuster [Dys02b, Dys14a]. Schweber [Bro96c, Dys56b, SGT^{+}95, Bro96b, Tre94, Vil95, Way95]. Schwinger [Bro96b, Bro96c, Dai67, DORQ12b, Dys50c, GFW12, NEA10, Sax94, SGT^{+}95, Tre94, Vil95, Way95, AI02, AMN02, ABP08, ABP12, AIP13, AB12, AFPT10, AHS09, AAK^{+}12, BMM98, BS90, BCS89, BMS00, BP08, BCDM01, BK10, BFW11, BEKW94, Bui11, BFM97, BCCM87, CJ80, CH99, CBPG84, CLO10, CBBS11, CBBS12, CK94, CK97, CMD04, CDM05, CM03, CP90, CP93, Dav92, DJS95, Det03, DORQ12a, Dys49a, Dys58j, Dys65e, Dys96b, Dys96c, Dys01d, Dys03h, EN12, EF12, EF13, ERS07, FM09, GFW11, HS98, HN08, HvSA98a, HvSA98b, HT95, HMSW00a, HMSW00b, HKR07, HBKK11, HAS10, HB12, HM12, KK02, KL02, KP09, KSW13, KK98, Kon97, KPS90, Kri08, LNA96, LFJ^{+}13, LEdAB03, LFW10, Maa06, MEBK11, MMJ91, MNBW07, MM97, MNK^{+}97, MM90, Mun95, NO99, NH91, NHKK94, NMM^{+}94]. Schwinger [Ng96, NAW08, Oko91, RBB95, Rem90, RQ11, ST12, SA03, Sch94b, Sch94a, SR11, Tak89, Wad81, WR07, WR10, sXcHmW87, Xue88, Yos82, YxQCxL10, Zho97, Zwa03, dJ94]. Schwinger-equation [BK10]. Schwinger/Bethe [NEA10]. Science [Ano16, Ano17b, BDH04, Ber18, Dys56e, Dys56a, Dys62b, Dys73c, Dys81m, Dys83f, Dys84o, Dys86f, Dys87h, Dys90a, Dys90f, Dys92b, Dys93g, Dys94a, Dys95g, Dys96i, Dys97a, Dys97b, Dys98f, Dys02a, Dys02c, Dys02g, Dys03c, Dys05e, Dys05j, Dys08a, Dys09c, Dys12f, Dys12i, Dys12d, Fai94, Gle92, Gre81, Hor06a, Kor94, May07, Mcc94, SGT^{+}95, Sta98, Sta03, SDL^{+}01, URR86, UM86, Ano95, Ber84, Ber19, Bye87, Cur82, Dre01, Dur00, Dys77g, Dys99f, Dys10e, Dys11p, Dys11r, Fin06, Gar61, Ham91, HH04a, Joh07b, Joh20a, Kra11, Lam06, Neu16, Opp89, SV^{+}08, Tip10, Whi80, Dys98b, Dys09c, Dys11o, Gre81, Dys84p]. Sciences [Dys64e, Dys80b, Dys81a, Ham91, DV63, Ham91]. Scientific [BHJ^{+}00, Dre01, Dys80b, Dys81a, Dys84p, Dys07e, Str00, Alp73, Cas03, Cor95, Cor97, Dys99i, Dys02d, DB02, Kev87]. scientifica [Cor97]. scientifiques [Dys01b]. Scientist [Ano17c, Ano20c, Ben97, Blo04, BB63, Dys79a, Dys89c, Dys95h, Dys95i, Dys06m, Dys08h, Dys15c, Guh07, Joh07b, Ano95, ALW^{+}88, Dys96d, Dys06k, Dys06n, DS07, ALW^{+}88, Dys79c, Dys79b, Dys98b]. Scientists [Ano54, Ano66, Ano96, Ber18, Dys63f, Dys05e, Dys14a, GR63, SDGK06, Fra08, GR63, Ber19]. scintillations [DMZ76]. Scope [Bus15]. Scotland [Dys88c, Dys90h, Dys04f]. Scuola [Dys71j]. Sea [GD03]. Search [Dys60k, Dys66g, Dys05g, Bye87, Cam63, Dys89l]. Searching [CM59]. Second [Dys55f, Dys55g, Dys06c, Dys11o, HH13, IO08, Bas22, PVW02, PVW03, VPW01, WKH13]. Second-order [HH13, IO08, WKH13]. Secret [Dys14d, Hor06a, Fin06]. Section [Dys55f, Dys55g]. sections [DCCS94, Dys54c, OK09]. See [Bas11]. Seeing [Dys05e, Dys02b]. seen [Ham91]. Segrè [Dys07f]. Seismic [Dys69m]. Seismicity [Dys69p]. Selecta [TL97, LT91, LT97, LT01, LT05, Thi91, Thi01, Thi05]. Selected [Dys96e, Sch58, Sch03, Yan13, BM83, DARM83, Jon16, Man07, Dys84b]. Self [BEKW94, DvL05, PVW02, PVW03, VPW01, CLO10, HH13, MNBW07, NMM^{+}94, SS99, WKH13, Zho97]. self-consistency [CLO10]. Self-consistent [BEKW94, DvL05, PVW02, PVW03, VPW01, MNBW07, SS99]. self-energies [HH13, NMM^{+}94, WKH13]. self-energy [Zho97]. Sell [Dys70h, Dys72m]. Sell-Out [Dys72m, Dys70h]. semiclassical [SRBW05, GS04b]. Semiconductor [SGT^{+}95]. Semiconductors [BS90]. sen [Bro01]. Senate [Dys63f]. Senior [Dys80b, Dys81a]. Sense [Dys05e, Lam06]. Sensors [GBS91]. Separated [LCW^{+}72]. Sept [Dys74a, Ham91]. September [Bye87, Dys06b, Dys09d]. Serber [Dys92c]. Series [Dys62b, DFG10, DFG13, Ros63, Dys54a, JL86, Yeu80]. sesquicentennial [Ber84]. set [HvSA98a, Maa06]. sets [Dys45a, Dys60b]. Seventy [ALW^{+}88, Dys95d]. seventy-fifth [ALW^{+}88]. Seventy-five [ALW^{+}88]. SFR [Sta03]. Sham [DCCS94, GBB03, YHKO07]. Shape [Dys13e, Bro10, MNM94]. share [Ano82]. Sharp [AAL22]. Shatterer [She82]. shell [PVW02, PVW03]. Shelters [Dys62j, Dys63g, Mor62]. Sherry [CQ99]. Shift [Bet47, Dys48c, BD94c]. shifted [BCDL93, BD94c]. Shirkov [Dys60a]. Shoolery [Dys02a]. Shopping [PSP^{+}07]. Shores [Don02]. Short [GX06, Goo70, Fey00, Lie04]. Shvinger [Dai67]. Si [BS89, BS89]. Side [Dys62a]. Sides [WD15]. Siegelman [Dys05g]. Siegert [WYL81]. Sight [Dys02c, Dys02g]. sigma [Meu09]. significance [For02]. Signos [CQ99]. Sigrid [Dys67a]. Silence [Dys80b, Dys81a]. Silent [Dys10c]. Silk [Spe91]. Silvan [Bro96b, Tre94, Vil95, Way95, Bro96c]. Silver [Whi80]. similarities [RQ11]. Similarity [Wal99]. Simon [Dys02b, Dys14a]. simulada [CQ99]. Simulation [SGT^{+}95]. simultaneous [Dys47b]. Sine [AvM05]. Singapore [PKpCH14]. Single [DÜ55, RT82]. singular [DFPR22]. Singularities [HJ65b]. sino [Ano95]. Sir [Dys54d]. site [JRV22]. Sitter [AB12]. Six [Dys69i, VHK89]. Six-fermion [VHK89]. Sixth [Dys00f]. Sixtieth [DPSY64]. Skeptic [Hor06b]. Skepticism [May07]. skew [Yan10]. Skin [Dys55c]. Sky [Dys62a, Dys15e]. Skyriders [Dys01e]. Sloan [Has79]. Slow [Dys11l]. Small [Meu01]. Smaller [Dys95j]. smallest [Fis10, Fis12]. Smil [Dys05h]. Smith [Dys83c, She82, Str00]. Smithsonian [Dys08b]. Smoke [Dys12f, Dys12d]. sobre [DG05]. Social [Ano01, Dys98g]. Socialism [Sta03]. Socially [Ben97]. Society [Bea99, Dys05j, Neu16, Ano99]. Sociology [UM86]. Sodium [HMMD94]. Soft [Dys97a, BS89]. soft-X-ray-emission [BS89]. Solar [SGT^{+}95]. soleil [Dys01b]. solid [VWS08]. solids [IT01]. Soluble [Dys67b, LM66]. Solution [AB12, Bat50, BTD^{+}47, BMS00, Det03, Dys50a, För96, Gal50, Hom96, Kle89, Thé47, Thé50, Wil48, Wil50, WYL81, BCS89, BEKW94, CBPG84, CK97, DvL05, ERS07, PVW02, PVW03, Rem90, VPW01, UTB^{+}50]. Solutions [Bat50, Bat51, BTD^{+}47, BS51, BG49, DC52, Dys50a, DTS51, Dys57b, EW50, ES51, FZ52, Gal50, Gal52, GB48, MS50, RRS48, San50, Thé47, TL48, TB49, Thé50, TK51a, TK51b, UL51, Wil48, WG50, Wil50, WL51, AMN02, CJ80, CK94, MM97, Nar86, RBB95, RQ11, Zho97, UTB^{+}50]. solvable [Dav92, EG11, Meu01]. Solving [HvSA98a, HvSA98b, Maa06, Oko91, HM12, MMJ91]. Some [Bas11, Bro96c, Dys44b, Rev09, Woo80, BDL93, Fad90, MM90]. Sonne [DB02]. Soul [Dys07f]. Sounds [Sta03]. Sourcebook [SGT^{+}95]. Sourcebook/Schweber [SGT^{+}95]. Sources [Dys60k, Dys65d, Sta03]. Southeast [DGWW67]. Soviet [Dys82a, AF81]. Space [Cor10, Dre00, Dys68d, Dys69g, Dys69h, Dys78e, Dys81l, Dys83f, Dys84q, Dys84o, Dys86g, Dys90i, Dys02b, Dys04d, DG05, Dys08b, DLV10, Dys15a, KC73, Sta03, CK74, Dys84m, Joh20a, Nee83, She88]. spacecraft [Dys95c, Dys95k]. Spaceflight [Fin17]. Spaceship [Dys72f, Dys02h, Ber02, Kan02]. spacetime [Dys16]. spacing [HR22]. Spacings [LCW^{+}72]. Spanish [Ano95, ZDSN08]. spatial [LACTFF22]. Speak [Ano66, Ano01, SDGK06]. Special [SBZ21]. spectra [DZO13]. Spectral [AF95, GS04b, GS04a]. Spectroscopy [LCW^{+}72, DK06, DCCS94, MND11, MNLD11]. spectrum [DK06, HN08].

Speculation [Dys63a]. Speculations [Dys84q]. Speculator [BB63]. Speech [Dys66d, Dys77a, Dys90a, Dys91i, Dys95f]. Speed [Fin17, Jud72]. Spell [DHD06, Dys06d]. Sphere [Dys57a, Dys84c, BD94a, BD95]. spheres [Dys51b]. spherical [SZ21]. spiked [AAL22]. spills [Smi06]. Spin [Dys55c, Dys56c, Dys64c, Dys86a, TKH62, DLS78, LFJ^{+}13, MNBW07, RNM04, SHD91, SZ21]. Spin-Wave [Dys56c, Dys86a]. Spinning [Dys69d]. Spirit [Gre81, Tip10, Dys05e]. spiritual [Ano00b]. Split [Dys05e, Kai05b]. Spontaneous [Dys69k, Hof11]. Springer [Dys58b, Dys70a]. Springer-Verlag [Dys58b, Dys70a]. spurious [HGE88, Par87]. Spy [Dys15c, Dys15c]. Square [Dys04a, BD94a, BD94b, BD95]. squares [BD94b]. SSC [Dys88a, GSE^{+}88]. SSSR [Dys80b, Dys81a]. St [Ben13]. Stability [Dys60l, Dys66f, Dys66h, DL67, Dys74c, Dys77f, DTO78, LD68, Fed75, Fol78, LT91, LT97, LT01, LT05, Thi91, TL97, Thi01, Thi05, Daj73]. Stachel [Sta03]. staggered [LFJ^{+}13]. staircase [Fis10, Fis12]. Stanley [CQ99]. Star [Dys60g, Dys85j, SM04, Dys05a, GS04b, GS04a]. Stardust [Dys15d]. Stargazers [Dys02b]. Stars [Ano20c, DC73, Dys69p, Dys80o, Joh07a, RRD^{+}73, Wat48, CBBS11, CBBS12, Dys71j, LT91, LT97, LT01, LT05, Thi91, TL97, Thi01, Thi05]. Starship [Bro78b, Bro96a, Broxx, Bro78c, Bro78d, Bro78a]. State [BW69, BD59, Dys57a, Dys67c, GSS^{+}03, Bas22, Joh07b, NRT22]. Statement [Dys69c, Dys69n, DLV10]. States [DX64, DhX65b, BS89, GLS22, HGE88, tL84, LLW^{+}09, OK07, Par87, SA03, SG86, Don02]. Static [Dys07s, Dys11i]. Stationary [GLS22]. statistic [GC81]. Statistical [Ano04, BM87, Coh71, Dys62g, Dys62h, Dys62i, DM63, Gun62, Kal64, MD63, Poo72]. Statistics [AF95, AK00, Dys64c, FKS97, ADM23, BLS03, BDL93, GS04b, GS04a, HR22, RNM04]. Stauffer [BHJ^{+}00]. Stay [MD08]. Stellar [Dys60k, Dys68b, Dys65d]. Stephen [Spe91, Don02]. Sternglass [Dys69i]. Steve [Dys81b]. Steven [Dys10d, Dys15a, Str00]. Stewart [Dys08c, Dys96g]. Stochastic [CH99, WKH13, Zwa03]. Stonier [Dys64g, GBS91]. Stony [GSS^{+}03]. Stopping [BH34]. Stora [dJ94]. Story [Ber02, Dys07e, Kan02, Cao06, Dys02h, Dys12g, Dys12e]. Strain [Rus12]. strangeness [Woo80]. Strangest [Dys10c]. Strategic [Dys90l, Dys84m]. strategies [PD12]. Strategy [Dys62e]. Straus [Dys05e, Dys11l]. Streater [Dys64c]. String [Dys04d, Xue88, Wad81]. strings [Dys91c]. Strömgren [Dys67a]. Strong [SG06, Sim22, SPZ^{+}12]. Strong-coupling [SG06]. strong-field [SPZ^{+}12]. Strongly [KSW13]. Strongly-coupled [KSW13]. Structure [Bek86, GBS91, SGT^{+}95, DD96a, DD96b, DK06, Dys62k, GL82]. Structures [För96]. Struggle [Dys07f, Dys08c, Wüt13]. students [Neu16]. Studies [DLV10, LSW76, Sta03, UM86, Dys62e]. Study [AAK^{+}12, DK06, Dys08d, MM97, SA03, CP93, Fis10, Fis12, Tak89, Mit80]. Stuff [BR93, Dys00b]. Subcommittee [Dys77g]. Subleading [MOR95]. subscription [Dys58b]. subversion [Ano95]. sum [BD94b]. summation [DD96c]. sums [BD94b]. Sun [BHJ^{+}00, Bro96a, Dys99i, DB02]. Suns [Dys79d]. Super [Dys88b, Dys91c]. Superbomb [Dys76a].

Superconducting [Dys88b]. superconductivity [HMSW00a, HMSW00b]. superiore [Dys71j]. Superposition [LACTFF22]. Supplement [Dys54e]. supplementary [Dys60c]. Support [Dys81m]. surface [Ye92]. Surfaces [SGT^{+}95]. surprise [Wüt13]. survey [Gar61]. surveys [BR01]. Sustainability [MOG^{+}13]. Sustainable [Dys13f, BM83]. Sutton [SGT^{+}95]. sweep [Dys89h]. Sycamore [Dys11o]. Symmetric [AW69, Dys71d, JKM21, BMS00, GLS22]. Symmetries [Dys64b, Dys89i, BCTG84]. Symmetry [Dys66i, AI02, CK97, CDM05, Dys62k, Dys69j, GS22, GMO00, Guh91, JRV22, LNA96, MND11, MNLD11, Mun95, NO99, Rem90, dJ94, GSS^{+}03, Bar66]. symmetry-adapted-cluster [MND11]. symmetry-breaking [CK97]. Symphony [Sta03]. sympletic [SLA93]. Symposium [CM90, GSS^{+}03, JR80, Mey56, Ros63, Whi80, yWHH98, Ber84, Dys74a, LW78, Woo80, Nee83]. System [Dys53c, Dys53h, Dys67c, BM83, HN08, Zei87]. system-Dyson [Zei87]. Systems [Dys62g, Dys62h, Dys62i, DM63, Dys75c, DTO78, Dys82f, Dys85d, MD63, Vog99, BDL93, DLS78, NS93, PVW02, PVW03, PS04, VPW01, VWS08, Wad81]. Szathmáry [Str00].

T [Dys79d, Dys85a]. Tactical [DGWW67]. Tai [Bro01]. Take [Hor06b]. taken [Dys50c]. Taking [GM12, HAS10]. Tale [Dys05g, Sar05]. Talking [Noc08, Yor87]. Tallahassee [LW78]. Tamm [DD55, Dys53b, Dys53g, DRS^{+}54]. Tamplin [DK70]. Tanner [DARM83, PM^{+}83]. tanto [Ano95]. Tautomeric [SOM10]. Taylor [Blo04]. Tayseer [ABG^{+}89, Dys89a, WDW89]. Taysir [DW89]. teach [Dys91i]. teacher [Ng96]. Teachers [Dys91i, Whe91]. Teaching [Dys90a]. Technique [Hur52, Kal64]. techniques [Jud72, MMJ91]. Technological [Dys71b, Dys73a, Joh20a, PKpCH14]. Technologien [DB02]. Technologies [Dys81b, Dys82c, BM83, DB02, Hei80]. Technology [Ano01, Dys66g, Dys73c, Dys77g, Dys98g, BDF00, Dys89l, Fre92, Neu16, Ber84]. Tecnociencia [CQ99]. tecnologia [Fre92]. Teeth [Dys92c]. Teilchen [Fis10, Fis12]. Telekinesis [Dys04c]. telescopes [RWP78]. Teller [Dys58a, Dys76a, Dys02a, Dys06g, Dys09d, Dys07h]. Temperament [MD95b, MD95a]. Temperature [MOR95, BK10, BCCM87, HS98, HN08, HKR07, ST12, Ye92]. Templeton [Ano00a, Gan00, Hor06b, SDL^{+}01]. tension [Xue88]. tensor [AAL22]. Tensorial [Dys60b]. Term [BG85, Cor10, Ove08, BCDL93]. terms [DG84, DMZ76, GL82]. Terraforming [Cor10]. Terrestrial [Dys60j]. Terror [Dys09c]. Test [BMM98, Dys63f, GC81]. Testimony [Dys63f, Dys77g]. tests [KV85, VHK89]. Texture [Dys04d, DG05]. Their [Ano96, Dys72a, Fra08]. Theology [Dys82h, Sta98]. Theorem [Liv54, DD96c, Dys45a, Dys48b, Mun95, Sim22, BG85, GX06]. Theorems [Yan54, Yan55]. theoretic [McL13]. Theoretical [yWHH98, Ano20a, Ano20b, Rad20, Sto20]. Theoreticians [Ano70]. Theories [Cao06, CDD56a, CDD56b, Dys54d, Dys01d, Dys07i, Dys11d, Dys12f, Dys12d, Str00, Bui11, CLO10, DK06, Dys49a, Dys58j, Dys03h, GLSS21, Kai05a, SA03, dJ94]. Theory [Bar66, Dys44b, Dys53f, Dys53g, Dys56c, Dys58b, Dys60a, Dys62g, Dys62h, Dys62i, DM63, DX64, DhX65b, Dys66a, Dys69o, Dys86a, Dys92h, Dys07g, Dys07j, Dys07o, Dys11c, Dys11e, Dys11g, Eps55, Gun62, Ler71, LCW^{+}72, MD63, OK07, Pra64, TKH62, AI02, ABD11, BDH04, BMS00, BCDM01, CH99, CMD04, DD55, DFPR22, DSW51, Dys52a, Dys55d, Dys65c, DhX65a, Dys96c, Dys99j, For21, GZ02, GMON98, IO08, JKM21, Kon97, LW78, NH91, Oko91, SW72, Tak86, Tak88, Tam83, VWS08, Wad81, WR07, sXcHmW87, Dys55c, Dys62b, Dys60c]. therefore [Ano95]. Thermal [GBS91]. Thermodynamic [Dys56f]. thermodynamics [BB03]. Thesis [Dys69i]. Things [SD98, Fey00]. Think [Ano17c, Spe91, Bro15]. Thinkers [Rev09, Bro15]. Thinking [DD04, Rus12, Cor10, Dys11l]. Third [KLR13, Dys72n]. Thirring [Dys55b]. Thomas [Ben13, Dys03f, Gre81]. Thompson [CL22]. t’Hooft [Ano82]. Thought [Dys80o, Dys04g]. Thought-experiments [Dys04g]. Thoughts [Dys62j, Dys63g, Dys98b, Dys05j, Mor62]. Three [Dys43c, Goc76, CP90, Kon97, NRT22]. three-body [NRT22]. three-dimensional [Kon97]. Three-Magnon [Goc76]. three-point [CP90]. threefold [Dys62k]. Thrill [Dys97a]. Thriving [KD12]. tides [DMZ76]. tiling [AvM23]. Time [Bek86, Dys59c, Dys67d, Dys79m, Dys83d, Dys04d, DG05, Eps55, Gam67a, GZ87, Hal17, Ros62, Ros63, SS67, Spe91, Sta03, VWS08, Wil58, Zwa03, Alp73, DD96a, DD96b, Dys72a, Dys02d, Fad90, GLS22, Dys89m, Dys03c]. time-dependent [VWS08]. Time-Independent [Eps55, Zwa03]. Time-ordered [GZ87]. Time-resolved [VWS08]. Timeline [BHJ^{+}00]. Times [Dre01, Dys15a]. Timothy [Dys02b, Dys05l]. Tipler [Spe91]. Today [Dys90l, BDF00, Bro15]. Toeplitz [Dys63h]. Tom [Dys64g]. Tomás [CQ99]. Tomonaga [Bro96c, Dai67, Dys49a, Dys58j, Dys65e, Dys01d, Dys03h, Sch94b, Sch94a, SGT^{+}95, Bro96b, Tre94, Vil95, Way95, Sax94]. tomorrow [BDF00]. Too [Dys62e]. tool [DFPR22]. Tools [BHJ^{+}00, Dys84j, Dys12i, Dys99i]. topic [Dys80d]. topological [AAL22]. Topology [Wig90, Dys48b, Mon87, Mon99]. Topping [Dys14d]. Total [Dys63e]. Township [Dys77e]. trace [CL22]. Track [Dys05l]. Traditional [FR94, FR02]. Traditions [KLR13]. Trafí [CQ99]. Tragedy [Dys80p, Tie09c]. Tragic [Dys05g, Sar05]. transcendence [For02]. transcendent [Dys95e]. transfer [MMJ91]. Transform [Dys11n]. Transformation [Goc76, RT82]. transformations [KLR13]. transformed [PS04]. transforms [Dys53a]. transition [BK10, BS89, CJ80, Dys69f, EN12, FM09, HS98, HBKK11, KK02]. Transitions [Daj73, Dys69l, Dys71f, Dys71l, Dys72e, DLS76, DLS78]. Transl [Dys62a]. Translated [Dys60c, Dys60a, Dys83b, Dys04c, Dys05f]. Translator [Sta03]. Transport [Dys69b, Dys68a, SRBW05]. Transportation [Dys95k]. Transverse [SHD91]. trapped [AF81]. treasury [FF91]. Treatment [Dys56g, SS98, Tak86, Tak88]. Treaty [Dys63f]. tree [URR86]. trends [LW78]. triangular [BCCM87]. Tribue [Joh07b]. tribute [ST11]. Trieste [Dys74a]. Trinity [Dys81c]. trivialization [AAL22]. Trouble [Dys93g, Dys94a, Fai94, Kor94, Mcc94, She82]. Trouble-Maker [She82]. True [Ber02, Kan02, Cas03, Dys02h]. truly [Alp73]. truncated [HvSA98a, Maa06]. Truncating [CP90]. Truncation [KPS90, BP08, CP93, Oko91, SG86]. Truth [FD93]. Tsipis [Dys84a]. Tsirul’nikov [Dys80b, Dys81a]. Turbulence [HMMD94, Woo92]. Turing [CK12]. Turkle [CQ99]. Turn [Coh71]. Turned [Joh20a]. Turning [Met73]. Tuva [Dys91a]. Twentieth [Dys02a, LY95, Sta03]. Twentieth-Century [Dys02a, Sta03]. twenty [Neu16, Dys70i]. Twenty-First [Dys70i]. Two [Bet33, Dys56g, Dys96l, Ste04, BBDY22, BF32, BD94b, CDM05, DJS95, DhX65a, Dys87c, IO08, JRV22]. two-cut [BBDY22]. Two-Electron [Bet33, IO08]. Two-Group [Dys56g]. two-matrix [DJS95]. two-particle [CDM05]. two-site [JRV22]. type [AFPT10]. Tyutin [dJ94].

U [Dys60b, DGWW67, GBS91, SGT^{+}95]. U.P [Dys96j]. U.S [Dys77g]. Uchūsen [Broxx]. Ulam [Yan54, Yan55]. ultimate [BDH04, Cal10]. Unapt [LD90]. Uncertainty [LLW^{+}09, Yan10]. Underdog [Dys13b]. Underdogs [Dys13b]. undergraduate [Neu16]. understand [Dys99j]. Understanding [Dys84a, Dys14a]. Unevolved [Rus12]. Unfashionable [Dys81n, Dys82b, Dys82i, Dys83g, Dys84r, Dys16]. Unfinished [Sta03, Dys50c]. Unification [Dys09b]. Unified [EF12]. Uniformly [DTO78]. unifying [ADM23]. Union [Kon90, AF81]. Unitarization [Roc71]. unitary [JKM21]. United [Don02]. univers [Dys86d, Dys09e]. Universal [BM87, GS04b, GS04a, NS93]. Universality [GMO03, GMO98]. Universe [DW48, Dys79m, Dys89b, Dys89m, Dys07q, Dys10i, Dys70g, Dys71e, Dys71a, Dys72i, Dys75a, Dys79e, Dys79i, Dys79j, Dys79k, Dys80j, Dys81d, Dys81j, Dys84d, Dys84e, Dys93c, Dys05b, Dys09e, GBS91, Gre81, She82, Ano80, Bar82, Bet79, CK12, Dys79c, Dys79a, Dys79b, Dys10d, Dys14a, Fie79, Has79, Hol81, Joh07a, Lar89, Pei80]. Universe/Dyson [GBS91]. Universe/Kafatos [GBS91]. Universities [Bas11]. University [AAB^{+}88, Bro96c, CM90, Cao06, Dys92c, Dys98b, Dys02d, Dys02c, Dys04c, Dys05f, Dys06e, Dys08d, Dys08c, Dys10d, Dys15a, GSS^{+}03, JR80, PKpCH14, Rab78, Ros63, Sta03, Dys51a, Dys51g, Mey56]. Universum [Dys89m, GBS91]. Universum/Scholz [GBS91]. Unknown [Dys09b, Dys08i, Lew22]. unlabeled [OO23]. unlikely [Dys03g]. unorthodox [Joh20c]. Unpredictable [Dys90g, Dys91e]. unquenched [KSW13]. Unquenching [ABP12]. Unrenormalized [Yos82, CBPG84]. Unseen [Dys05e]. unserer [DBB09]. Upper [HMMD94, Bas22]. Upper-Atmosphere [HMMD94]. uracil [SPZ^{+}12]. Urals [PDMS80]. Urbana [AAB^{+}88]. Urbana-Champaign [AAB^{+}88]. Urged [Dys59b]. Ursprünge [DBB09]. US$60.00 [Dys02d]. USA [Dys00a]. Use [BHJ^{+}00, Dys53g, Dys79g, Dys80i, Dys95b, GC81, Per68, Pra64, TKH62, SVT09]. Use/Moss [BHJ^{+}00]. Useful [Bea99]. Using [TMD96a, TMD96b, CP93, DK06, KSW13, MNM94, MND11, OK09, RM06, She88, SG06, TMD96c]. Ustojcivost [Daj73].

V [Dys60a, Dai10, MD63]. Vaclav [Dys05h]. Vacuum [BDS52, Dys50c]. Vajk [Dys80c]. Val [Dys98c]. Valentine [LSW76]. value [BD94b]. Values [DARM83, PM^{+}83]. Variability [Gam67b]. variance [BD94c]. Variation [Dys67d, Dys78a, Dys78f, DD96a, DD96b, Dys72a]. vary [Alp73]. vector [Kri08]. Velocidad [CQ99]. Venerated [ST20]. venture [Dys16]. verged [Ano20b, Rad20]. Verlag [Dys58b, Dys70a]. Version [Dys84o]. versus [Dys90i, GS85]. vertex [AIP13, CP90, Det03, DORQ12a, DORQ12b, KP09]. vertical [IO08, SOM10]. via [BCCM87, Sim22]. vibrational [HH13, Ye92]. Victoria [Wig90]. vida [CQ99, CQ99]. vie [Dys09e]. Vietnam [IT23]. View [Gol08, Dys03f, Dys89d]. viewpoint [For21]. Views [Dys69i, Dys10d, Tay69, Dys81g, Whi80]. VIII [LCW^{+}72]. Viking [Dys06d, Dys07f, Dys11n, Hor06a]. Vilkovisky [dJ94]. Villain [GLSS21]. vingt [Ham91]. vingt-sept [Ham91]. violating [DD96c]. Violence [Dys95b]. Vision [Lew99, Cor95, Cor97]. Visionary [Joh20a, Ach20, Ano20a]. visione [Cor97]. Visiones [CQ99]. Visions [CCL^{+}11, Fie79]. Visit [DGB^{+}01]. Vista [Ell86]. vistas [Ber84]. vita [Dys87e, Dys02e]. Vitale [CQ99]. void [Opp89]. Vol [Dys56b, Dys58b, Dys62b, Dys03a]. Volcano [Dys69o]. Volkoff [Dys60a]. Volume [Dys55a, LFW10, Dys66i, UM86]. Vortrag [Dys70g]. Voting [PD13]. vs [CDB^{+}73]. Vu [Dys62a, Ham91]. Vulcanism [Dys69p].

W [CQ99, Dys60c, Dys85a, Dys94a, SGT^{+}95, Spe91, Dys83d]. Wali [Dys91b]. walk [Dys88d, Dys01f, Dys13d]. Walker [Dys12f, Dys12d]. Walks [Dys87a]. Wall [PSP^{+}07]. Walter [Dys55b]. War [BHJ^{+}00, Dys62e, Dys70b, Dys90l, ACN85, BM99, DARM83, IT23, Dys85a, Dys08b]. Ward [CP90, DD96c]. warfare [Joh20a]. Warm [Dys56g, Dys97h]. Warm-blooded [Dys97h]. Warming [Dys08d, Gol08, NZSD08, ZDSN08]. Wars [Dys85j]. Was [Dys90m, Dys15c, DBB09]. Washington [Sta03]. watch [GM12]. Watcher [Gre81]. water [CCC^{+}10, DZO13]. Wave [Dys53c, Dys53h, Dys56c, Dys69m, Dys86a, DK06, Nak87]. waveguide [RM06]. Waves [Dys16, TKH62, DMZ76, LACTFF22]. Way [SZ06, Dys62k]. Wczechswiecie [Dys83d, Dys83d]. Weapons [Ano60, Ano66, Dys60f, Dys61a, DGWW67, Dys84a, Dys84i, Dys84j, Dys84k, Dys84l, Dys84s, Dys86b, Dys89j, Dys03a, Ell86, Yor87, Dys79a, vH84, Buc87, Fre84, Gom84, Pie84]. Wechselwirkung [BF32]. Weighing [Dys08d]. Weinberg [Dys10d]. Weiner [Dys83c, She82]. Weisskopf [Ano82]. Wells [Wig90]. Weltall [Dys70g]. Were [Dys75b, Dys09c]. Wertheim [Dys12f, Dys12d]. Wesley [Dys98b]. Weyl [Dys56d, Dys70d, Dys70a]. Wheeler [Dys80d, Dys04g, Dys10h, Hal17, Ove08]. Where [Lam06, HZ82, ZH95]. Whittaker [Dys54d]. Who [Ano20c, Bro96b, Bro96c, Dys15c, Ove08, Sax94, Tre94, Vil95, Way95, Joh20a, Sch94a, SGT^{+}95]. whose [Ano20b, Rad20]. Width [HBKK11]. Wiener [Dys81b, Dys05g, Hei80, TD05]. Wightman [Dys64c, Dys58d]. Wigner [AF95, AK00, BLS03, Fol78, Fol80, FKS97, GS04b, GS04a, LLW^{+}09, RNM04, TP01, Yan10]. Wilczek [Dys09b]. wild [BD94c]. Will [Ano01, MD08, Bro10]. Willard [Dys90a]. William [Dys08d]. Wilson [Dys98d, PR09, sXcHmW87, Xue88]. Win [Dys13b]. Winchester [Pei80]. Wind [BCC^{+}08]. Winner [Ano70, Dys81o]. Winners [San94]. wins [Ano00a]. Winter [Dys85j, Mcc94]. Wires [AF95]. Wirklichkeit [DBB09]. wisdom [Bye87, Dys85a]. Wise [Dys05l, Guh07]. Wiser [BHJ^{+}00]. wissenschaftliche [BHJ^{+}00, DB02]. within [DvL05, OK07, WR07]. without [DD96c, Dys79m, Dys81e, Dys83d, Dys89m]. Wolf [Ano82, Dys89h]. Wolfgang [Dys02d, DGB^{+}01, Dys02d]. women [Fad90]. Won [Dys05e]. Wonder [Dys09c]. Wong [SGT^{+}95]. Woo [Bro96a]. Woo-ju-sun-kwa [Bro96a]. Word [Dys84f, Dys86c]. Words [Dys95j, SD98, Fra08, Ano96]. work [Can01, Dys70a, Dys02d]. Working [Dys07f]. works [Fey00]. Workshop [BM83]. World [Dys69c, Dys79c, Dys79a, Dys79b, Dys96i, Dys04d, Dys04a, Dys05h, Dys10d, Dys12g, Dys12e, GR63, JR80, Lew99, BM83, Dys72n, Dys73d, Dys84m, FF91, Lam06, Whi80, Dys90l, Dys02c, Dys02g, Dys08c, Dys11n]. Worlds [Dys06e, Str00, Dys97e, Lew99, She82, Ben97, Cre01, dS99]. Wren [Dys07e]. Wright [Mon94]. Writer [Alb94, Ano20c, Joh20a]. Writing [Dys96m, Dys06e, Gre81, Pit11, Pit12]. Writings [SGT^{+}95]. wrong [Dys15b]. Wu [Gre81, yWHH98]. Wu-Li [Gre81]. Wussing [Sta03].

x [Dys64c, Sta03, BS89, Dys02a, HG87]. X-ray [HG87]. xi [Sta03]. xii [Dys70a, Sta03]. xliii [Sta03]. XUV [SPZ^{+}12]. xvi [Dys60a]. XVIIth [Jen14]. xxvii [Bro96c].

Yale [CM90, Dys98b, Dys02c, Dys08d, Dys08c, Dys15a]. Yamabe [Yan54, Yan55]. Yanase [LLW^{+}09, Yan10]. Yang [DORQ12b, Dys84b, GSS^{+}03, LY95, AI02, DORQ12a, Dys99a, Sim22, Wad81, WR07]. Ye [JRV22]. Years [Dys73c, Dys95d, Dys98e, Dys12h, Sta03, HZ01, Nee83, Neu16]. Yes [Dys84n, Ehr75]. York [Ben13, Ber02, Dre01, Dys58a, Dys60c, Dys60a, Dys60b, Dys62a, Dys62b, Dys64c, Dys67b, Dys70a, Dys96j, Dys02d, GSS^{+}03, Guh07, Sta03, Dys76a, Dys85i, Rab78]. Yu [Dys83b, Bro01]. Yujobo [Yan54, Yan55]. Yukawa [AFPT10]. Yuri [Man07].

Zalcman [Dys89h]. Zedillo [Dys08d]. Zeilberger [GX06]. Zeilinger [Fis10, Fis12]. Zeit [BHJ^{+}00, Dys89m, GBS91]. Zeit/Dyson [BHJ^{+}00]. Zell [SGT^{+}95]. zero [Oko91]. zero-dimensional [Oko91]. zeros [WK74, Dys60l, Dys63b]. zhou [Bro01]. zi [Bro01]. Zone [Vog99]. Zubay [Sha01]. Zukav [Gre81]. Zukunft [DBB09, Dys81g, DB02]. zum [Fis10, Fis12]. zvezdy [DC73]. Zwanziger [DORQ12b, DORQ12a]. zwei [BF32, Bet33]. Zwei-Elektronenprobleme [Bet33]. Zweipersonenspiele [BHJ^{+}00]. Zweipersonenspiele/Adobe [BHJ^{+}00].

References

Andrews:1988:RRP

[AAB^{+}88] George E. Andrews, Richard A. Askey, Bruce C. Berndt, K. G. Ramanathan, Robert A. Rankin, et al., editors. Ramanujan revisited: proceedings of the centenary conference, University of Illinois at Urbana-Champaign, June 1–5, 1987. Academic Press, New York, USA, 1988. ISBN 0-12-058560-X. LCCN QA1 .R26 1987.

Aoki:2012:SCH

[AAK^{+}12] Yasumichi Aoki, Tatsumi Aoyama, Masafumi Kurachi, Toshihide Maskawa, Kei ichi Nagai, Hiroshi Ohki, Akihiro Shibata, Koichi Yamawaki, and Takeshi Yamazaki. Study of the conformal hyperscaling relation through the Schwinger–Dyson equation. Physical Review D (Particles and Fields), 85(??):074502, April 2, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.85.074502. The authors are part of the

LatKMI Collaboration.

Auffinger:2022:SCA

[AAL22] Antonio Auffinger, Gerard Ben Arous, and Zhehua Li. Sharp complexity asymptotics and topological trivialization for the (p, k) spiked tensor model. Journal of Mathematical Physics, 63(4): 043303, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL https://pubs.aip.

org/aip/jmp/article/63/4/043303/2843034/Sharp-complexity-
asymptotics-and-topological. Special collection in honor of

Freeman Dyson.

Aaserud:1986:OHI

[Aas86] Finn Aaserud. Oral history interview with Freeman J. Dyson, 1986 December 17. Technical report, American Institute of Physics, Woodbury, NY, USA, 1986. Two hours of interview tapes.

Akhmedov:2012:SDS

[AB12] E. T. Akhmedov and Ph. Burda. Solution of the Dyson–Schwinger equation on a de Sitter background in the infrared limit. Physical Review D (Particles and Fields), 86(??):044031, August 17, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.

org/doi/10.1103/PhysRevD.86.044031.

Akemann:2011:OHR

[ABD11] Gernot Akemann, Jinho Baik, and Philippe Di Francesco, editors. The Oxford handbook of random matrix theory. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2011. ISBN 0-19-957400-6. xxxi + 919 pp. LCCN QA188 .O94 2011. URL http:/

/www.loc.gov/catdir/enhancements/fy1402/2011029624-b.
html;   http://www.loc.gov/catdir/enhancements/fy1402/
2011029624-d.html;   http://www.loc.gov/catdir/enhancements/
fy1402/2011029624-t.html.

Aumann:1989:LED

[ABG^{+}89] Moshe Aumann, Yehuda B. Band, Howard D. Greyber, Lawrence Cranberg, Reuben Rudman, Edward Witten, Freeman Dyson, and Richard Wilson. Letters to the Editor: Dissenting opinions about Tayseer Aruri. Physics Today, 42(8):13–15, 85–88, August 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/42/13/1.

Aguilar:2008:GGP

[ABP08] A. C. Aguilar, D. Binosi, and J. Papavassiliou. Gluon and ghost propagators in the Landau gauge: Deriving lattice results from Schwinger–Dyson equations. Physical Review D (Particles and Fields), 78(??):025010, July 9, 2008. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
78.025010.

Aguilar:2012:UGP

[ABP12] A. C. Aguilar, D. Binosi, and J. Papavassiliou. Unquenching the gluon propagator with Schwinger–Dyson equations. Physical Review D (Particles and Fields), 86(??):014032, July 27, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.86.014032.

Achenbach:2020:OFD

[Ach20] Joel Achenbach. Obituary: Freeman Dyson, a visionary and renaissance physicist, dies at 96. Washington Post, ??(??):??,

February 28, 2020. URL https://www.washingtonpost.com/
local/obituaries/freeman-dyson-a-visionary-and-renaissance-
physicist-dies-at-96/2020/02/28/0ba462e0-5a58-11ea-ab68-
101ecfec2532_story.html.

Allison:1985:HDO

[ACN85] Graham T. Allison, Albert. Carnesale, and Joseph S. Nye, editors. Hawks, doves, and owls: an agenda for avoiding nuclear war. W. W. Norton & Co., New York, NY, USA, 1985. ISBN 0-393-01995-

0. xii + 282 pp. LCCN ???? URL http://catalog.hathitrust.
org/api/volumes/oclc/11621599.html.

Aaserud:1986:FD

[AD86] Finn Aaserud and Freeman Dyson. Freeman dyson. AIP Oral History Interview., December 17, 1986. URL https://

www.aip.org/history-programs/niels-bohr-library/oral-
histories/4585.

Andrews:1988:PIQ

[ADH88] George E. Andrews, Freeman J. Dyson, and Dean Hickerson. Partitions and indefinite quadratic forms. Inventiones Mathematicae, 91(3):391–407, 1988. CODEN INVMBH. ISSN 0020-9910 (print), 1432-1297 (electronic).

Akemann:2023:EGE

[ADM23] G. Akemann, M. Duits, and L. D. Molag. The elliptic Ginibre ensemble: a unifying approach to local and global statistics for higher dimensions. Journal of Mathematical Physics, 64(2):023503, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.aip.

org/aip/jmp/article/64/2/023503/2874065/The-elliptic-
Ginibre-ensemble-A-unifying-approach. Special collection in

honor of Freeman Dyson.

Adler:2009:DNB

[ADvM09] Mark Adler, Jonathan Delépine, and Pierre van Moerbeke. Dyson’s nonintersecting Brownian motions with a few outliers. Communications on Pure and Applied Mathematics (New York), 62(3):334–395, March 2009. CODEN CPAMAT, CPMAMV. ISSN 0010-3640 (print), 1097-0312 (electronic).

Azbel:1981:RTS

[AF81] M. Ya. (Mark Yakovlevich) Azbel and Grace Pierce Forbes, editors. Refusenik, trapped in the Soviet Union. Houghton Mifflin, Boston, MA, USA, 1981. ISBN 0-395-30226-9. xiii + 513 + 8 pp.

LCCN DS135.R95 A937. URL http://catalog.hathitrust.
org/api/volumes/oclc/6735540.html.

Altland:1995:SSM

[AF95] Alexander Altland and Dirk Fuchs. Spectral statistics of mesoscopic wires: Crossover from Wigner–Dyson to Poisson regime. Physical Review Letters, 74(??):4269–??, May 22, 1995. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://link.aps.org/doi/10.1103/PhysRevLett.
74.4269.

Alexandre:2010:SDA

[AFPT10] J. Alexandre, K. Farakos, P. Pasipoularides, and A. Tsapalis. Schwinger–Dyson approach for a Lifshitz-type Yukawa model. Physical Review D (Particles and Fields), 81(??):045002, February 1, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.81.045002.

Atkinson:1966:POC

[AH66] D. Atkinson and M. B. Halpern. Possible one-channel Castillejo– Dalitz–Dyson poles in πN and ππ scattering. Physical Review, 150(4):1377–1386, October 28, 1966. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.150.1377.

Alkofer:2009:ADD

[AHS09] R. Alkofer, M. Q. Huber, and K. Schwenzer. Algorithmic derivation of Dyson–Schwinger equations. Computer Physics Communications, 180(6):965–976, June 2009. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465508004268.

Abe:2002:SDA

[AI02] Hiroyuki Abe and Tomohiro Inagaki. Schwinger–Dyson analysis of dynamical symmetry breaking on a brane with bulk Yang–Mills theory. Physical Review D (Particles and Fields), 66(??):085001, October 1, 2002. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

link.aps.org/doi/10.1103/PhysRevD.66.085001.

Aguilar:2013:GPG

[AIP13] A. C. Aguilar, D. Ibáñez, and J. Papavassiliou. Ghost propagator and ghost-gluon vertex from Schwinger–Dyson equations. Physical Review D (Particles and Fields), 87(??):114020, June 27, 2013. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.87.114020.

Altland:2000:WDS

[AK00] Alexander Altland and Alex Kamenev. Wigner–Dyson statistics from the Keldysh σ-model. Physical Review Letters, 85(??): 5615–??, December 25, 2000. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.85.5615.

Albrecht:1975:LCC

[Alb75] Herbert O. Albrecht. Letters: Causes for concern. Bulletin of the Atomic Scientists, 31(8):3, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys75b].

Albers:1994:FDM

[Alb94] Donald J. Albers. Freeman Dyson: Mathematician, physicist, and writer. College Mathematics Journal, 25(1):2–21, January 1994. CODEN ???? ISSN 0746-8342 (print), 1931-1346

(electronic). URL http://www.jstor.org/stable/2687079;
 http://www.tandfonline.com/doi/abs/10.1080/07468342.
1994.11973574.

Alpher:1973:LNC

[Alp73] Ralph A. Alpher. Large numbers, cosmology and Gamow: Are the fundamental constants of nature truly constant, or do they vary with time? An account of George Gamow’s last scientific enquiry. American Scientist, 61(1):52–58, January/February 1973. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://adsabs.harvard.edu/abs/1973AmSci..
61...52A;   http://www.jstor.org/stable/27843563.

Arnheim:1988:SFR

[ALW^{+}88] Rudolf Arnheim, Thomas E. Lovejoy, David Gordon Wilson, Freeman Dyson, Jane Goodall, Ian Shelton, Kenneth H. Olsen, Irene C. Peden, Richard W. Hamming, Thomas Eisner, Preston Cloud, Matt Cartmill, Samuel C. Florman, Jeremy Bernstein, George A. Miller, Robert M. May, G. Evelyn Hutchinson, Jerome Bruner, Priscilla C. Grew, William Bevan, Elisabeth S. Vrba, Myrdene Anderson, Kevin Padian, Harry Shipman, Victor F. Weisskopf, Walter A. Hill, Patricia D. Moehlman, Melvin Kranzberg, Malak Kotb, Raymond Kurzweil, Marcia McNutt, Masakazu Konishi, Miriam Rothschild, Edward Teller, Alison Jolly, H. Jane Brockmann, Keith Stewart Thomson, Peter J. Denning, Benoît B. Mandelbrot, Abraham Pais, Paul MacCready, Kip S. Thorne, Ruth Sager, Gerald J. Wasserburg, Neal E. Miller, Rita Levi-Montalcini, Stephen Jay Gould, Edwin H. Land, Michel Boudart, Anne Kernan, Douglas R. Hofstadter, Rosalyn S. Yalow, Bruce H. Tiffney, Mimi Koehl, Walter E. Massey, David P. Billington, John A. W. Kirsch, Abner Shimony, J. Donald Fernie, Brian J. Skinner, Lynn Margulis, Sheldon Lee Glashow, Michael LaBarbera, J. Tuzo Wilson, E. R. Ward Neale, Rudolf Peierls, Roald Hoffmann, Mary L. Good, Donald R. Griffin, Vaclav Smil, Michael S. Turner, Sarah Ann Woodin, Luis Alvarez, George A. Bartholomew, and George B. Schaller. Seventy-five reasons to become a scientist: American Scientist celebrates its seventyfifth anniversary. American Scientist, 76(5):450–463, September 1988. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786

(electronic). URL http://www.jstor.org/stable/27855384;
 http://www.jstor.org/stable/pdfplus/27855384.pdf.

Aguilar:2002:FQC

[AMN02] A. C. Aguilar, A. Mihara, and A. A. Natale. Freezing of the QCD coupling constant and solutions of Schwinger–Dyson equations. Physical Review D (Particles and Fields), 65(??):054011, January 30, 2002. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

link.aps.org/doi/10.1103/PhysRevD.65.054011.

Andrews:1980:NDC

[And80] George E. Andrews. Notes on the Dyson conjecture. SIAM Journal on Mathematical Analysis, 11(5):787–792, September 1980. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Anderson:1991:CFP

[And91] James L. Anderson. Comment on “Feynman’s proof of the Maxwell equations,” by Freeman J. Dyson [Am. J. Phys. 58, 209– 211 (1990)]. American Journal of Physics, 59(1):86, 1991. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/59/86/1.

Anderson:2013:IC

[And13] Philip Anderson. An iconoclast’s career. Physics World, 26(3): 62–63, March 2013. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://iopscience.iop.org/pwa/
full/pwa-pdf/26/03/phwv26i03a40.pdf. Review of [Sch13].

Anonymous:1954:SEC

[Ano54] Anonymous. Scientists express confidence in Oppenheimer. Bulletin of the Atomic Scientists, 10(7):283, 286, September 1954. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1960:FDN

[Ano60] Anonymous. The future development of nuclear weapons. Bulletin of the Atomic Scientists, 16(5):167, May 1960. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Response to [Dys60f].

Anonymous:1966:SSC

[Ano66] Anonymous. Scientists speak out on CB weapons. Bulletin of the Atomic Scientists, 22(9):39–40, November 1966. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). The letter is signed by Felix Bloch (Nobel 1952), Konrad E. Bloch (Nobel 1964), James F. Crow, William Doering, Paul Doty, Freeman J. Dyson, John T. Edsall, Bernard Feld, Irwin C. Gunsalus, Robert Hofstadter (Nobel 1961), Arthur Kornberg (Nobel 1959), Fritz Lipmann (Nobel 1953), Robert B. Livingston, Matthew Meselson, Severo Ochoa (Nobel 1959), Ray D. Owen, Keith R. Porter, Charles Price, Eugene Rabinowitch, E. L. Tatum (Nobel 1958), George Wald, and Paul Dudley White.

Anonymous:1969:GPC

[Ano69] Anonymous. German physicists choose Dyson for Planck Medal. Physics Today, 22(7):112, 1969. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/?PTO/22/112/1.

Anonymous:1970:TND

[Ano70] Anonymous. Theoreticians name Dyson as winner of Oppenheimer Prize. Physics Today, 23(3):97, 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/?PTO/23/97/1.

Anonymous:1971:EP

[Ano71] Anonymous. Energy and power. W. H. Freeman, New York, NY, USA, 1971. ISBN 0-7167-0938-4 (paperback), 0-7167-0939-2. viii + 144 pp. LCCN TJ153 .E478. Originally appeared as articles in the September 1971 issue of Scientific American.

Anonymous:1980:BRB

[Ano80] Anonymous. Book review: Disturbing the Universe by Freeman Dyson. The Wilson Quarterly (1976–2012), 4(1):167, Winter

1980. ISSN 0363-3276. URL http://www.jstor.org/stable/
40255778.

Anonymous:1982:DTW

[Ano82] Anonymous. Dyson, t’Hooft and Weisskopf share 1981 Wolf Prize. Physics Today, 35(2):87–88, 1982. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/?PTO/35/87/1.

Anonymous:1991:DRO

[Ano91] Anonymous. Dyson receives Oersted Medal, AAPT’s highest honor. Physics Today, 44(4):125, 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/?PTO/44/125/1.

Anonymous:1993:BNB

[Ano93] Anonymous. Book note: From Eros to Gaia, by Freeman Dyson. Bulletin of the Atomic Scientists, 49(4):54–55, May 1993. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Anonymous:1994:DRF

[Ano94] Anonymous. Dyson receives 1994 Fermi Award. Physics Today, 47(11):104, 1994. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
47/104/1.

Anonymous:1995:CCR

[Ano95] Anonymous. El científico como rebelde — para F. Dyson, quien conocio de cerca a Einstein y a Oppenheimer, la ciencia no es un metodo filosofico sino mas bien una forma de arte y, por tanto, de subversion. (Spanish) [The scientist as rebel — for F. Dyson, having closely known Einstein and Oppenheimer, science is not a philosophical method but rather an art form and therefore subversion]. Quimera, ??(140–141):21–??, 1995. ISSN 0211-3325.

Anonymous:1996:BRB

[Ano96] Anonymous. Book review: The Faith of Scientists: In Their Own Words. BMJ: British Medical Journal, 312(7038):1095, April 27,

1996. ISSN 0959-8138. URL http://www.jstor.org/stable/
29731464.

Anonymous:1997:CDP

[Ano97] Anonymous. Conferimenti dei diplomi di perfezionamento honoris causa a Freeman J. Dyson e Carlo Azeglio Ciampi: Pisa, 4 giugno 1996. (Italian) [Conferment of honorary diplomas of completion on J. Freeman Dyson and Carlo Azeglio Ciampi: Pisa, June 4, 1996]. Report, Scuola Normale Superiore, Pisa, Italy, June 4, 1997. 60 pp.

Anonymous:1999:APS

[Ano99] Anonymous. American Physical Society prizes and awards. The Chronicle of Higher Education, ??(??):??, March 12, 1999. ISSN 0009-5982 (print), 1931-1362 (electronic). URL http://

chronicle.com/article/American-Physical-Society/4938/.

Anonymous:2000:FDW

[Ano00a] Anonymous. Freeman Dyson wins 2000 Templeton Prize.

Web site, 2000. URL http://www.templetonprize.org/
previouswinner.html#dyson.

Anonymous:2000:WCB

[Ano00b] Anonymous. What’s the connection? Bringing a spiritual perspective to daily life. The Christian Science Monitor, ??(??):??, June 12, 2000. ISSN 0882-7729 (print), 1540-4617 (electronic).

URL http://www.csmonitor.com/2000/0612/p23s1.html.

Anonymous:2001:FDW

[Ano01] Anonymous. Freeman Dyson will speak on “Technology and Social Justice” Oct. 25. Web news story., October 10,

2001. URL https://www.amherst.edu/aboutamherst/news/
news_releases/2001/10_2001/node/18906.

Anonymous:2004:PSM

[Ano04] Anonymous. Program of the 90th Statistical Mechanics Meeting Celebrating the 80th Birthday of Freeman Dyson and the 100th Anniversary of the Birth of Lars Onsager. Journal of Statistical Physics, 115(3–4):1139–1145, May 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1023/
B%3AJOSS.0000022576.38044.55.

Anonymous:2008:NML

[Ano08] Anonymous. Nishina Memorial Lectures: creators of modern physics, volume 746 of Lecture notes in physics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. ISBN 4-431-77055-0. xiv + 402 pp. LCCN QC71

.N615 2008. URL http://www.loc.gov/catdir/toc/fy0805/
2007940477.html.

Anonymous:2013:FDO

[Ano13] Anonymous. Freeman Dyson OW. Web document, 2013. URL

http://www.winchestercollege.org/freemason-dyson-ow.

Anonymous:2016:FDS

[Ano16] Anonymous. Freeman Dyson, sage of science. NPR To The Best of Our Knowledge radio broadcast., December 8, 2016. URL https:

//www.ttbook.org/show/freeman-dyson-sage-science.

Anonymous:2017:FD

[Ano17a] Anonymous. Freeman Dyson. NPR To The Best of Our Knowl-

edge Web site., 2017. URL https://www.ttbook.org/people/
freeman-dyson.

Anonymous:2017:FDL

[Ano17b] Anonymous. Freeman Dyson, at 93, looks back on a legendary career in science. NPR To The Best of Our Knowledge radio broadcast., April 16, 2017.

Anonymous:2017:HTL

[Ano17c] Anonymous. How to think like a scientist. NPR To The Best of Our Knowledge radio broadcast., April 16, 2017. URL https:/

/www.ttbook.org/show/how-think-scientist#. Guests: Hope

Jahren; Manu Prakash; Neil deGrasse Tyson; Freeman Dyson; Nick Lantz.

Anonymous:2020:FDV

[Ano20a] Anonymous. Freeman Dyson, visionary theoretical physicist and mathematician — obituary. Telegraph, ??(??):??, March 1, 2020.

URL https://www.telegraph.co.uk/obituaries/2020/03/
01/freeman-dyson-visionary-theoretical-physicist-mathematician/

.

Anonymous:2020:FJDa

[Ano20b] Anonymous. Freeman J. Dyson (1923–2020): Guardian obituary: Brilliant theoretical physicist and mathematician whose farfetched ideas for the future verged on the bizarre. Web site, 2020.

URL https://mathshistory.st-andrews.ac.uk/Obituaries/
Dyson_Guardian/.

Anonymous:2020:FJDb

[Ano20c] Anonymous. Freeman J. Dyson (1923–2020): Princeton obituary: Scientist and writer, who dreamt among the stars, dies at 96.

Web site, 2020. URL https://mathshistory.st-andrews.ac.
uk/Obituaries/Dyson_Princeton/.

Anonymous:2020:IPF

[Ano20d] Anonymous. Infinite possibilities: Freeman Dyson died on February 28th. The Economist, ??(??):??, March 12,

2020. URL https://www.economist.com/obituary/2020/03/
12/freeman-dyson-died-on-february-28th.

Andrews:2005:RCD

[AO05] George E. Andrews and Ken Ono. Ramanujan’s congruences and Dyson’s crank. Proceedings of the National Academy of Sciences of the United States of America, 102(43):15277, October 25, 2005. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/4143422.

Adler:2005:PJD

[AvM05] Mark Adler and Pierre van Moerbeke. PDEs for the joint distributions of the Dyson, Airy and sine processes. Annals of Probability, 33(4):1326–1361, July 2005. CODEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1120224583;
http://www.jstor.org/stable/3481731.

Adler:2023:DIR

[AvM23] Mark Adler and Pierre van Moerbeke. Double interlacing in random tiling models. Journal of Mathematical Physics, 64(3): 033509, March 2023. CODEN JMAPAQ. ISSN 0022-2488 (print),

1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/
article/64/3/033509/2881737/Double-interlacing-in-random-

tiling-models. Special collection in honor of Freeman Dyson.

Atkinson:1969:PCD

[AW69] D. Atkinson and Robert Lee Warnock. Persistence of the Castillejo–Dalitz–Dyson ambiguity in relativistic crossing-symmetric amplitudes. Physical Review, 188(??):2098–??, December 25, 1969. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.188.2098.

Bailin:2017:BRD

[Bai17] David Bailin. Book review: Dear Professor Dyson, edited by Dwight E. Neuenschwander. Contemporary Physics, 58(1):91–92, 2017. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Barut:1966:BRG

[Bar66] A. O. Barut. Book review: Group theory in physics: Symmetry Groups in Nuclear and Particle Physics by Freeman J. Dyson. Lie Groups for Physicists by Robert Hermann. Science, 152(3725): 1048–1049, May 20, 1966. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1718198.

Barnes:1982:BRB

[Bar82] Barry Barnes. Book review: Disturbing the Universe by Freeman Dyson. Leonardo (Oxford, England), 15(3):246, Summer 1982. CODEN LEONDP. ISSN 0024-094X (print), 1530-9282

(electronic). URL http://www.jstor.org/stable/1574709.

Basken:2011:UFB

[Bas11] Paul Basken. As universities fend off budget cuts, some researchers see possible benefits. The Chronicle of Higher Education, ??(??):??, April 3, 2011. ISSN 0009-5982 (print), 1931-

1362 (electronic). URL http://chronicle.com/article/As-
Universities-Fend-Off/127000/.

Basti:2022:SOU

[Bas22] Giulia Basti. A second order upper bound on the ground state energy of a Bose gas beyond the Gross–Pitaevskii regime. Journal of Mathematical Physics, 63(7):071902, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print),

1089-7658 (electronic). URL https://pubs.aip.org/aip/
jmp/article/63/7/071902/2843530/A-second-order-upper-

bound-on-the-ground-state. Special collection in honor of Freeman Dyson.

Bateman:1950:APS

[Bat50] Paul Bateman. Advanced problems and solutions: Problems for solution: 4387. American Mathematical Monthly, 57(3):188–189, March 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Bat51, DTS51, ES51, TK51b, UL51].

Bateman:1951:APSa

[Bat51] Paul Bateman. Advanced problems and solutions: 4387. American Mathematical Monthly, 58(9):638–639, November 1951. CO-DEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UTB^{+}50].

Born:1963:BBS

[BB63] Hedwig Born and Max Born. Books: The Scientist Speculator: An Anthology of Partly-Baked Ideas, edited by I. J. Good. Bulletin of the Atomic Scientists, 19(5):30–32, May 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). This is a strongly critical review of the book. The review begins: “The publication of Books: The Scientist Speculator: An Anthology of Partly-Baked Idea is nothing less than a crime against the ethical code, unwritten but vital, of the community of scientists.” See rebuttal in [Dys63a].

Borchers:2003:FLT

[BB03] H. J. Borchers and Ch. Borchers. The first law of thermodynamics and Dyson–Lieb’s perpetuum mobile. Journal of Mathematical Physics, 44(10):4866–4867, October 2003. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Barhoumi:2022:ITC

[BBDY22] Ahmad Barhoumi, Pavel Bleher, Alfredo Deaño, and Maxim Yattselev. Investigation of the two-cut phase region in the complex cubic ensemble of random matrices. Journal of Mathematical Physics, 63(6):063303, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

URL https://pubs.aip.org/aip/jmp/article/63/6/063303/
2846121/Investigation-of-the-two-cut-phase-region-in-

the. Special collection in honor of Freeman Dyson.

Brenner:2008:WFR

[BCC^{+}08] Michael Brenner, Shelly Cazares, Michael J. Cornwall, Freeman Dyson, Douglas Eardley, Paul Horowitz, Darrell Long, Jeremiah Sullivan, John Vesecky, and Peter J. Weinberger. Wind farms and radar. Report AD-jsr-08-125; AD-a480 068, JASON Program Office, Mitre Corporation, McLean, VA, USA, 2008.

Butera:1987:HTE

[BCCM87] P. Butera, R. Cabassi, M. Comi, and G. Marchesini. High temperature expansion via Schwinger–Dyson equations: The planar rotator model on a triangular lattice. Computer Physics Communications, 44(1–2):143–156, April/May 1987. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
0010465587900245.

Bloch:1965:LEJ

[BCD^{+}65] Felix Bloch, S. Chandrasekhar, Freeman J. Dyson, Yuval Ne’eman, Ivor Robinson, and Edwin E. Salpeter. Letter to the Editor: Jews in Russia. New York Times, ??(??):46, June 9, 1965. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/116741645/.

Bleher:1993:DET

[BCDL93] Pavel M. Bleher, Zheming Cheng, Freeman J. Dyson, and Joel L. Lebowitz. Distribution of the error term for the number of lattice points inside a shifted circle. Communications in Mathematical Physics, 154(3):433–469, 1993. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic). URL http://

projecteuclid.org/getRecord?id=euclid.cmp/1104253074.

Blagoev:2001:SDA

[BCDM01] Krastan B. Blagoev, Fred Cooper, John F. Dawson, and Bogdan Mihaila. Schwinger–Dyson approach to nonequilibrium classical field theory. Physical Review D (Particles and Fields), 64(??): 125003, November 26, 2001. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.64.125003.

Bender:1989:NSS

[BCS89] Carl M. Bender, Fred Cooper, and L. M. Simmons, Jr. Nonunique solution to the Schwinger–Dyson equations. Physical Review D (Particles and Fields), 39(??):2343–??, April 15, 1989. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.39.2343.

Bars:1984:SPP

[BCTG84] Itzhak Bars, Alan Chodos, Chia-Hsiung Tze, and Feza Gürsey, editors. Symmetries in particle physics. Plenum Press, New York, NY, USA; London, UK, 1984. ISBN 0-306-41801-0. LCCN QC793.3.S9 S93 1984.

Bernstein:1959:COE

[BD59] Jeremy Bernstein and Freeman J. Dyson. The continuous opacity and equations of state of light elements at low densities. Report, Project Orion, General Atomic Division of General Dynamics Corporation, ????, July 13, 1959. URL http://www.

daviddarling.info/encyclopedia/O/OrionProj.html.

Bleher:1994:MSL

[BD94a] Pavel M. Bleher and Freeman J. Dyson. Mean square limit for lattice points in a sphere. Acta Arithmetica, 68(4):383–393, 1994. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic). See erratum [BD95].

Bleher:1994:MSV

[BD94b] Pavel M. Bleher and Freeman J. Dyson. Mean square value of exponential sums related to representation of integers as sum of two squares. Acta Arithmetica, 68(1):71–84, 1994. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic).

Bleher:1994:VEF

[BD94c] Pavel M. Bleher and Freeman J. Dyson. The variance of the error function in the shifted circle problem is a wild function of the shift. Communications in Mathematical Physics, 160(3):493–505, 1994. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916

(electronic). URL http://projecteuclid.org/getRecord?id=
euclid.cmp/1104269707.

Bleher:1995:EMS

[BD95] Pavel M. Bleher and Freeman J. Dyson. Erratum: “Mean square limit for lattice points in a sphere” [Acta Arith. 68 (1994), no. 4, 383–393. MR1307454 (96a:11104)]. Acta Arithmetica, 73(2):199, ???? 1995. CODEN AARIA9. ISSN 0065-1036 (print), 1730-6264 (electronic). See [BD94a].

Bernstein:2003:OB

[BD03] Jeremy Bernstein and Freeman Dyson. Opacity bounds. Publications of the Astronomical Society of the Pacific, 115(815): 1383–1387, December 2003. CODEN PASPAU. ISSN 0004-6280 (print), 1538-3873 (electronic). URL http://labs.adsabs.

harvard.edu/ui/abs/2003PASP..115.1383B;   http://www.
jstor.org/stable/10.1086/380420.

Beecroft:2007:LEF

[BD07] Timothy Beecroft and Freeman J. Dyson. Letter to the Editor: Francis Bacon & the frozen chicken. New York Review of Books, 54(9):62, May 31, 2007. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2007/may/31/francis-bacon-the-frozen-chicken/. Com-

ment on [Dys07e].

Bell:2000:ETT

[BDF00] Trudy E. Bell, David Dooling, and Janie McLawhorn Fouke, editors. Engineering tomorrow: today’s technology experts envision the next century. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7803-5360-9 (Platinum ed.), 0-7803-5361-7 (Member hardcover ed.), 0-7803-5362-5 (Trade hardcover ed.). xiv + 308 pp. LCCN T174 .B451 2000.

Barrow:2004:SUR

[BDH04] John D. Barrow, P. C. W. Davies, and Charles L. Harper, editors. Science and ultimate reality: quantum theory, cosmology, and complexity. Cambridge University Press, Cambridge, UK, 2004. ISBN 0-521-83113-X. xx + 721 pp. LCCN QC174.12 .S4 2004.

URL ftp://uiarchive.cso.uiuc.edu/pub/etext/gutenberg/
;   http://www.loc.gov/catdir/description/cam051/2003055903.
html;   http://www.loc.gov/catdir/toc/cam051/2003055903.
html.

Bleher:1993:NGE

[BDL93] Pavel M. Bleher, Freeman J. Dyson, and Joel L. Lebowitz. Non–Gaussian energy level statistics for some integrable systems. Physical Review Letters, 71(19):3047–3050, November 8, 1993.

CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.71.3047.

Baranger:1952:FOV

[BDS52] M. Baranger, F. J. Dyson, and E. E. Salpeter. Fourth-order vacuum polarization. Physical Review, 88(3):680, November 1, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.
88.680.

Bearn:1999:UKA

[Bea99] Alexander G. Bearn, editor. Useful knowledge: the American Philosophical Society Millennium Program, volume 234 of Memoirs of the American Philosophical Society. American Philosophical Society, Philadelphia, PA, USA, 1999. ISBN 0-87169-234-1 (hardcover). ISSN 0065-9738. xxvi + 307 pp. LCCN Q11.P612 vol. 234.

Bekenstein:1986:FSC

[Bek86] Jacob D. Bekenstein. The fine-structure constant: From Eddington’s time to our own. In Ullmann-Margalit [UM86], pages 209–224. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. LCCN Q174 .B67 vol. 95

Q175. URL http://www.springerlink.com/content/978-94-
009-4566-1;   https://link.springer.com/book/10.1007/
978-94-009-4566-1.

Bracco:1994:SCS

[BEKW94] M. E. Bracco, A. Eiras, G. Krein, and L. Wilets. Self-consistent solution of the Schwinger–Dyson equations for the nucleon and meson propagators. Physical Review C (Nuclear Physics), 49 (??):1299–??, March 1, 1994. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.49.1299.

Benford:1997:BRR

[Ben97] Gregory Benford. Book review: The reflections of a socially conscious scientist: Imagined Worlds, by Freeman J. Dyson. Science Fiction Studies, 24(3):506–508, November 1997. ISSN 0091-7729 (print), 2327-6207 (electronic). URL http://www.jstor.org/

stable/4240652.

Benford:2013:BRB

[Ben13] Gregory Benford. Book review: Maverick Genius: The Pioneering Odyssey of Freeman Dyson, Phillip F. Schewe, Thomas Dunne Books/St. Martin’s Press, New York, 2013. $27.99 (339 pp.). ISBN-13 978-0-312-64235-8. Physics Today, 66(6):52, June 2013. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://www.physicstoday.org/resource/
1/phtoad/v66/i6/p52_s1.

Berger:1984:JCM

[Ber84] Melvyn S. (Melvyn Stuart) Berger, editor. J. C. Maxwell, the sesquicentennial symposium: new vistas in mathematics, science, and technology. North-Holland Publishing Co., Amsterdam, The Netherlands, 1984. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Bernstein:2002:BRB

[Ber02] Jeremy Bernstein. Book review: Project Orion: The True Story of the Atomic Spaceship, by George Dyson. 345 pp. Henry Holt, New York, 2002. Price: $26.00 ISBN 0-8050-5985-7. American Journal of Physics, 70(8):876–879, August 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://ajp.aapt.org/resource/1/ajpias/v70/i8/p876_s1;
 http://www.daviddarling.info/encyclopedia/O/OrionProj.
html.

Bernstein:2004:EJR

[Ber04] Barton J. Bernstein. The enigma of J Robert Oppenheimer. [review of Oppenheimer: Portrait of an Enigma, Jeremy Bernstein, 2004 Gerald Duckworth/Ivan R Dee 240pp, £14.99/$25.00 hardcover, J Robert Oppenheimer: The American Century, David C Cassidy 2004 Pi Press/Prentice Hall 480pp]. Physics World, 12 (12):36–37, December 2004. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://iopscience.

iop.org/pwa/full/pwa-pdf/17/12/phwv17i12a32.pdf.

Bernstein:2018:BDO

[Ber18] Jeremy Bernstein. A Bouquet of Dyson and Other Reflections on Science and Scientists. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2018. ISBN 981-323-192-0 (hardcover), 981-323-828-3 (paperback). xi + 175 pp. LCCN QC16.D95 A4 2018.

Berkshire:2019:BRB

[Ber19] Frank Berkshire. Book review: A bouquet of Dyson — and other reflections on science and scientists. Contemporary Physics, 60 (1):73–74, 2019. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bethe:1933:QZE

[Bet33] Hans Bethe. Quantenmechanik der Ein und Zwei-Elektronenprobleme. (German) [Quantum mechanics of one- and two-electron problems]. In Adolf Gustav Smekal, editor, Handbuch der Physik, volume 24-1, pages ix + 853. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1933. LCCN ????

Bethe:1947:ESE

[Bet47] H. A. Bethe. The electromagnetic shift of energy levels. Physical Review, 72(4):339–341, August 15, 1947. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

prola.aps.org/abstract/PR/v72/i4/p339_1.

Bethe:1979:BRF

[Bet79] Hans Bethe. Book review: F. Dyson, Disturbing the Universe. Physics Today, 32(12):51–52, December 1979. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/phtoad/v32/i12/p51/s1.

Bethe:1932:WZE

[BF32] H. Bethe and E. Fermi. Über die Wechselwirkung von zwei Elektronen. (German) [On the interaction of two electrons]. Zeitschrift für Physik, 77(5–6):296–306, May 1932. CODEN ZEPYAA.

ISSN ???? URL http://www.springerlink.com/content/
l561955527vv4513/.

Brydges:1976:NEB

[BF76] David Brydges and Paul Federbush. A note on energy bounds for boson matter. Journal of Mathematical Physics, 17(12):2133– 2134, December 1976. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v17/i12/p2133_s1.

Bicout:1995:GRE

[BF95] Dominique J. Bicout and Martin J. Field. Gaussian random energy model and Dyson’s model for the origin of metabolism. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 52(??):4209–??, October 1, 1995. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRevE.52.4209.

Bicout:1996:RDD

[BF96] D. J. Bicout and M. J. Field. Relaxation dynamics in Dyson’s model for the origin of metabolism. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 54(??):726–??, July 1, 1996. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.54.726.

Burkardt:1997:CHF

[BFM97] M. Burkardt, M. R. Frank, and K. L. Mitchell. Calculation of hadron form factors from Euclidean Dyson–Schwinger equations. Physical Review Letters, 78(??):3059–??, April 21, 1997. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://link.aps.org/doi/10.1103/PhysRevLett.
78.3059.

Bonnet:2011:EAQ

[BFW11] Jacqueline A. Bonnet, Christian S. Fischer, and Richard Williams. Effects of anisotropy in QED_{3} from Dyson–Schwinger equations in a box. Physical Review B: Condensed Matter and Materials Physics, 84(??):024520, July 15, 2011. CODEN PRB-

MDO. ISSN 1098-0121. URL http://link.aps.org/doi/10.
1103/PhysRevB.84.024520.

Butchart:1949:APS

[BG49] J. H. Butchart and R. Goormaghtigh. Advanced problems and solutions: 4245. American Mathematical Monthly, 56(3):189–190, March 1949. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Thé47].

Bressoud:1985:CTI

[BG85] D. M. Bressoud and I. P. Goulden. Constant term identities extending the q-Dyson Theorem. Transactions of the American Mathematical Society, 291(1):203–228, September 1985. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL

http://www.jstor.org/stable/1999904. See [Dys62g].

Bethe:2013:EPF

[BGA^{+}13] Hans Bethe, Marvin L. Goldberger, Stephen L. Adler, Steven Weinberg, Freeman Dyson, and C. N. Yang. Einstein and the physics of the future. In Yang [Yan13], pages 21–34. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4

(e-book). LCCN QC21.3. URL http://www.worldscientific.
com/doi/abs/10.1142/9789814449021_0006.

Bethe:1934:SFP

[BH34] H. A. Bethe and W. Heitler. On the stopping of fast particles and on the creation of positive electrons. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 146(856):83–112, August 1, 1934. CODEN PRLAAZ. ISSN 0080-

4630. URL http://www.jstor.org/stable/2935479.

Birkl:2000:CTR

[BHJ^{+}00] Gerhard Birkl, Stephan Hartmann, Manfred Jacobi, Hans Christoph Wolf, Erhard Keppler, Armin Bunde, Jens Eisert, and Max Rauner. Crichton: Timeline. Eine Reise in die Mitte der Zeit/ Dyson: The Sun, The Genome, and The Internet. Tools of Scientific Revolutions/Pais: Raffiniert ist der Herrgott Albert Einstein. Eine wissenschaftliche Biografie/Kuhn/Försterling: Principles of Physical Chemistry/Wiser: Energy Resources. Occurrence, Production, Conversion, Use/Moss de Oliveira, de Oliveira und Stauffer: Evolution, Money, War, and Computers/Canty: Konfliktlösungen mit Mathematica. Zweipersonenspiele/Adobe Acrobat 4.0/Buchtipp. Physikalische Blätter, 56(10):69–73, October 2000. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.20000561016/abstract.

Berry:2007:OBF

[BHMD07] Wendell Berry, James P. Herman, Christopher B. Michael, and Freeman J. Dyson. ‘Our biotech future’: An exchange. New York Review of Books, 54(14):101–103, September 27, 2007. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/archives/2007/sep/27/our-biotech-future-

an-exchange/. Response to [Dys07r].

Baragar:1973:JLD

[BK73] F. A. Baragar and A. N. Kamal. Jost–Lehmann–Dyson representation and scaling. Physical Review D (Particles and Fields), 7(??):3025–??, May 15, 1973. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.7.3025.

Blank:2010:QCT

[BK10] M. Blank and A. Krassnigg. QCD chiral transition temperature in a Dyson–Schwinger-equation context. Physical Review D (Particles and Fields), 82(??):034006, August 6, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.82.034006.

Brown:2006:HBHa

[BL06] G. E. (Gerald Edward) Brown and Chang-Hwan Lee, editors. Hans Bethe and his physics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2006. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). xiii + 314 pp. LCCN QC16.B46 H36 2006.

URL http://www.worldscibooks.com/physics/5989.html;
http://www.worldscientific.com/worldscibooks/10.1142/
5989.

Block:2004:AST

[Blo04] Melissa Block. Atomic scientist Taylor became nuclear opponent. National Public Radio report, November 2, 2004. URL http://

www.npr.org/templates/story/story.php?storyId=4139566.

Benet:2003:WDS

[BLS03] L. Benet, F. Leyvraz, and T. H. Seligman. Wigner–Dyson statistics for a class of integrable models. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 68(??):045201, October 21, 2003. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRevE.68.045201.

Bellamy:1955:PGC

[BM55] E. H. Bellamy and R. G. Moorehouse, editors. Proceedings of the 1954 Glasgow Conference on Nuclear and Meson Physics. Pergamon Press, Oxford, UK, 1955. LCCN ????

Bauer:1983:NTS

[BM83] G. S. (Günter Siegfried) Bauer and Alan McDonald, editors. Nuclear technologies in a sustainable energy system: selected papers from an IIASA Workshop (entitled “a perspective on adaptive nuclear energy evolutions: towards a world of neutron abundance”, Laxenburg, May 25–27, 1981). Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 3-540-12154-4, 0-387-12154-4. LCCN ????

Bleher:1987:CPU

[BM87] P. M. Bleher and P. Major. Critical phenomena and universal exponents in statistical physics. On Dyson’s hierarchical model. Annals of Probability, 15(2):431–477, April 1987. CO-DEN APBYAE. ISSN 0091-1798 (print), 2168-894X (electronic).

URL http://projecteuclid.org/euclid.aop/1176992155;
http://www.jstor.org/stable/2244058.

Bruce:1999:EWF

[BM99] Maxwell Bruce and Tom Milne, editors. Ending war: the force of reason: essays in honour of Joseph Rotblat, NL, FRS. Macmillan Press, Houndmills, Basingstoke, Hampshire, UK, 1999. ISBN 0-333-76070-0 (Macmillan Press: hardcover), 0-333-77482-5 (Macmillan Press: paperback), 0-312-22570-9 (St. Martin’s Press: hardcover). xxiii + 179 pp. LCCN JZ5665 .E53 1999.

URL http://www.loc.gov/catdir/bios/hol057/99026118.
html;   http://www.loc.gov/catdir/description/hol056/
99026118.html;   http://www.loc.gov/catdir/toc/hol053/
99026118.html.

Ballesteros:1998:TRN

[BMM98] H. G. Ballesteros and V. Martín-Mayor. Test for random number generators: Schwinger–Dyson equations for the Ising model. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 58(5):6787–6791, November 1, 1998. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRevE.
58.6787.

Bros:1961:PAC

[BMS61] J. Bros, A. Messiah, and R. Stora. A problem of analytic completion related to the Jost–Lehmann–Dyson formula. Journal of Mathematical Physics, 2(5):639–651, May 1961. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v2/i5/
p639_s1.

Bender:2000:SSD

[BMS00] Carl M. Bender, Kimball A. Milton, and Van M. Savage. Solution of Schwinger–Dyson equations for PT-symmetric quantum field theory. Physical Review D (Particles and Fields), 62(??): 085001, September 8, 2000. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.62.085001.

Binosi:2008:GIT

[BP08] D. Binosi and J. Papavassiliou. Gauge-invariant truncation scheme for the Schwinger–Dyson equations of QCD. Physical Review D (Particles and Fields), 77(??):061702, March 19, 2008. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.77.061702.

Brown:1993:MGS

[BR93] Laurie M. Brown and John S. Rigden, editors. Most of the Good Stuff: Memories of Richard Feynman. American Institute of Physics, Woodbury, NY, USA, 1993. ISBN 0-88318-870-8. 181 + 16 pp. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Berndt:2001:RES

[BR01] Bruce C. Berndt and Robert A. (Robert Alexander) Rankin. Ramanujan: essays and surveys, volume 22 of History of mathematics. American Mathematical Society, Providence, RI, USA, 2001. ISBN 0-8218-2624-7. xvi + 347 pp. LCCN QA29.R3 B47 2001.

Brehme:1991:CFP

[Bre91] Robert W. Brehme. Comment on “Feynman’s proof of the Maxwell equations,” by F. Dyson [Am. J. Phys. 58, 209–211 (1990)]. American Journal of Physics, 59(1):85–86, 1991. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/59/85/2.

Brower:1978:SCD

[Bro78a] K. Brower. Book review: Starship and Canoe: Freeman Dyson and George Dyson. Omni (New York), 1(3):16, ???? 1978. ISSN 0149-8711.

Brower:1978:KCP

[Bro78b] Kenneth Brower. Kosmolot i czółno. (Polish) [The Starship and the canoe]. Editions Spotkania, Warszawa, Poland, 1978. ISBN 83-85195-70-X. 262 pp. LCCN ????

Brower:1978:SCa

[Bro78c] Kenneth Brower. The starship and the canoe. Howard W. Sams, Indianapolis, IN 46268, USA, 1978. ISBN 0-03-039196-2. 270 pp. LCCN TL789.85.D9 B76 1978.

Brower:1978:SCb

[Bro78d] Kenneth Brower. The starship and the canoe. Whizzard Press, London, UK, 1978. ISBN 0-233-97240-4. 270 pp. LCCN ????

Brower:1996:WJS

[Bro96a] Kenneth Brower. Woo-ju-sun-kwa ka-nu. (Korean) [The Starship and the canoe]. Chang-Jak-Kwa, Bi-Pyung-Sa, Seoul, Korea, 1996. ISBN 89-364-7040-X. 396 pp. LCCN ???? Korean translation by Kyo-Sun Lee of [Bro78c].

Brown:1996:BRB

[Bro96b] Laurie M. Brown. Book review: QED and the Men Who Made It: Dyson, Feynman, Schwinger, and Tomonaga by Silvan S. Schweber. Isis, 87(1):204–205, March 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/235809.

Brown:1996:BRS

[Bro96c] Laurie M. Brown. Book review: Some QED myths-in-the-making?: Silvan S. Schweber, QED and the Men Who Made It: Dyson, Feynman, Schwinger and Tomonaga (Princeton University Press, 1994), xxvii + 732 pp., ISBN 0-691-03685-3, 0-691-03327-7 (paperback). Studies in History and Philosophy of Modern Physics, 27(1):81–90, March 1996. CO-DEN ???? ISSN 1355-2198 (print), 1879-2502 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
1355219895000232. See [Sch94a].

Brower:19xx:UKJ

[Broxx] Kenneth Brower. Uchūsen no to kanū. (Japanese) [The Starship and the canoe]. ????, ????, 19xx. ISBN ???? 349 pp. LCCN ???? Translation to Japanese by Takashi Serizawa of [Bro78c].

Brower:2001:TKC

[Bro01] Kenneth Brower. Tai kong chuan yu du mu zhou: dai sen fu zi chuan qi, volume 16 of Ke xue ren wen. Shi bao wen hua chu ban gong si, Tai bei shi, Republic of China (Taiwan), 2001. ISBN 957-13-3381-6. 300 pp. LCCN ???? Mandarin Chinese translation by Anqi yi Zhuang of [Bro78c].

Brockman:2010:WCE

[Bro10] John Brockman, editor. This will change everything: ideas that will shape the future. Harper Perennial, New York, NY, USA, 2010. ISBN 0-06-189967-4 (paperback). xxiii + 390 pp. LCCN Q175.5 .T48 2010; HM901 .T45 2010.

Brockman:2015:WTA

[Bro15] John Brockman, editor. What to think about machines that think: today’s leading thinkers on the age of machine intelligence. HarperPerennial, New York, NY, USA, 2015. ISBN 0-06-242565-X (paperback). xxvi + 541 pp. LCCN Q335 .W445 2015.

Browne:1951:APS

[BS51] D. H. Browne and N. T. Seely, Jr. Advanced problems and solutions: Solutions: 4396. American Mathematical Monthly, 58(9): 642–644, November 1951. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gal50].

Bruhwiler:1989:EAD

[BS89] P. A. Bruhwiler and S. E. Schnatterly. Empirical application of Dyson’s equation to the L_{2}, 3 soft-X-ray-emission transition densities of states of c-Si and a-Si:H. Physical Review B: Condensed Matter and Materials Physics, 39(17):12649, June 15, 1989. CO-DEN PRBMDO. ISSN 1098-0121. URL http://link.aps.org/

doi/10.1103/PhysRevB.39.12649.

Beloussov:1990:KPC

[BS90] I. V. Beloussov and Yu. M. Shvera. Kinetics of partially coherent polaritons in semiconductors. Schwinger–Dyson equations. Physica Status Solidi. B, Basic Research, 159(1):91–99, May 1, 1990. CODEN PSSBBD. ISSN 0370-1972 (print), 1521-3951 (electronic).

Bellman:1947:APS

[BTD^{+}47] Richard Bellman, Victor Thébault, F. J. Dyson, L. A. Santalo, and Howard Eves. Advanced problems and solutions: Problems for solution: 4259–4263, 4248. American Mathematical Monthly, 54(7 (Part 1)):418–419, August/September 1947. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Thé47, TB49].

Buck:1987:BRB

[Buc87] John Buck. Book review: Weapons and Hope by Freeman Dyson. International Journal on World Peace, 4(4):97–99, October 1987.

URL http://www.jstor.org/stable/20751183.

Buividovich:2011:SDE

[Bui11] P. V. Buividovich. Schwinger–Dyson equations in large-N quantum field theories and nonlinear random processes. Physical Review D (Particles and Fields), 83(??):045021, February 23, 2011. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.83.045021.

Bussey:2015:PCH

[Bus15] Peter J. Bussey. Proceedings of the Conference in Honour of the 90th Birthday of Freeman Dyson, edited by K. K. Phua, L. C. Kwek, N. P. Chang, and A. H. Chang, Scope: conference proceedings. Level: postgraduate, researcher. Contemporary Physics, 56 (4):508–509, 2015. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Bart:1969:EPC

[BW69] George R. Bart and Robert L. Warnock. Empirical N/D potential and Castillejo–Dalitz–Dyson poles in the pion-nucleon P_{11} state. Physical Review Letters, 22(??):1081–??, May 19, 1969. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://link.aps.org/doi/10.1103/PhysRevLett.
22.1081.

Byers:1987:RSS

[Bye87] David M. Byers, editor. Religion, science, and the search for wisdom: proceedings of a Conference on Religion and Science, September 1986, volume 157-1 of Publication / Office of Publishing Services, United States Catholic Conference. Bishops’ Committee on Human Values, National Conference of Catholic Bishops, Washington, DC, USA, 1987. ISBN 1-55586-157-1 (paperback). LCCN BL240.2 .C65 1986.

Calaprice:1996:QE

[Cal96] Alice Calaprice. The quotable Einstein. Princeton University Press, Princeton, NJ, USA, 1996. ISBN 0-691-02696-3 (hardcover). xxxiv + 269 pp. LCCN QC16.E5 A25 1996. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/description/prin031/96003543.html;
 http://www.loc.gov/catdir/toc/prin031/96003543.html.

Calaprice:2000:EQE

[Cal00] Alice Calaprice. The expanded quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2000. ISBN 0-691-07021-0 (hardcover). xliii + 407 pp. LCCN QC16.E5

A25 2000. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/bios/prin051/
00026873.html;   http://www.loc.gov/catdir/description/
prin022/00026873.html;   http://www.loc.gov/catdir/samples/
prin031/00026873.html;   http://www.loc.gov/catdir/toc/
prin031/00026873.html.

Calaprice:2005:NQE

[Cal05] Alice Calaprice, editor. The new quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2005. ISBN 0-691-12074-9 (hardcover), 0-691-12075-7 (paperback). xxxvii + 407 pp. LCCN QC16.E5 A25 2005. URL http://press.princeton. edu/titles/7921.html. Foreword by Freeman Dyson.

Calaprice:2010:UQE

[Cal10] Alice Calaprice, editor. The ultimate quotable Einstein. Princeton University Press, Princeton, NJ, USA, 2010. ISBN 0-691-13817-6. xxviii + 578 pp. LCCN QC16.E5 A25 2010. URL http://press.

princeton.edu/titles/9268.html.

Cameron:1963:ICS

[Cam63] Alastair Graham Walter Cameron. Interstellar communication: the search for extraterrestrial life. A collection of reprints and original contributions. Physical investigations of the universe. W. A. Benjamin, Inc., New York, NY, USA, 1963. xii + 320 pp. LCCN ????

Cannell:2001:GGM

[Can01] D. M. (Doris Mary) Cannell. George Green: mathematician and physicist, 1793–1841: the background to his life and work. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 2001. ISBN 0-89871-463-X. xxxiv + 316 + 10 pp. LCCN QC16.G64

C36 2001. URL http://epubs.siam.org/ebooks/siam/other_
titles_in_applied_mathematics/ot73;   http://www.loc.
gov/catdir/enhancements/fy0726/00041938-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0726/00041938-t.html.

Cao:2006:BRS

[Cao06] Tian Yu Cao. Book review: The story of Feynman diagrams: Drawing Theories Apart: The Dispersion of Feynman Diagrams in Postwar Physics, by David Kaiser, 2005. University of Chicago Press, 469pp $80.00hb/$30.00pb. Physics World, 19(1):36–37, January 2006. CODEN PHWOEW. ISSN 0953-8585 (print),

2058-7058 (electronic). URL http://iopscience.iop.org/pwa/
full/pwa-pdf/19/1/phwv19i1a38.pdf.

Casti:2003:OTP

[Cas03] J. L. Casti. The one true platonic heaven: a scientific fiction on the limits of knowledge. Joseph Henry Press, Washington, DC, USA, 2003. ISBN 0-309-08547-0 (hardcover), 0-309-09510-7 (paperback). xviii + 160 pp. LCCN Q175 .C4339 2003. URL http://

www.loc.gov/catdir/toc/fy037/2003002279.html;   http:
//www.nap.edu/catalog.php?record_id=10533.

Chen:2011:HSD

[CBBS11] H. Chen, M. Baldo, G. F. Burgio, and H.-J. Schulze. Hybrid stars with the Dyson–Schwinger quark model. Physical Review D (Particles and Fields), 84(??):105023, November 16, 2011. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.84.105023.

Chen:2012:HPS

[CBBS12] H. Chen, M. Baldo, G. F. Burgio, and H.-J. Schulze. Hybrid protoneutron stars with the Dyson–Schwinger quark model. Physical Review D (Particles and Fields), 86(??):045006, August 2, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.86.045006.

Cartier:1984:FUN

[CBPG84] J. F. Cartier, A. A. Broyles, R. M. Placido, and H. S. Green. Finite, unrenormalized, nonperturbative solution to the Schwinger– Dyson equations of quantum electrodynamics. Physical Review D (Particles and Fields), 30(??):1742–??, October 1, 1984. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.30.1742.

Canuto:2010:DWF

[CCC^{+}10] Sylvio Canuto, Kaline Coutinho, Benedito J. C. Cabral, V. G. Zakrzewski, and J. V. Ortiz. Delocalized water and fluoride contributions to Dyson orbitals for electron detachment from the hydrated fluoride anion. Journal of Chemical Physics, 132(21): 214507, 2010. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/?JCP/132/
214507/1.

Chiao:2011:VDN

[CCL^{+}11] Raymond Y. Chiao, Marvin L. Cohen, Anthony J. Leggett, William D. Phillips, and Charles L. Harper Jr., editors. Visions of Discovery: New Light on Physics, Cosmology, and Consciousness. Cambridge University Press, Cambridge, UK, 2011. ISBN 0-521-88239-7 (hardback). xxxii + 794 pp. LCCN Q162 .V48

2011. URL http://assets.cambridge.org/97805218/82392/
cover/9780521882392.jpg.

Cini:1973:LEP

[CDB^{+}73] Marcello Cini, Gianfausto Dell’Antonio, Michel Le Bellac, Jean-Marc Levy-Leblond, Daniel Schiff, Jacques Treiner, Freeman J. Dyson, and Sidney D. Drell. Letters to the Editor: Protestors vs Jason. Physics Today, 26(4):11–13, April 1973. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/?PTO/26/11/1;
http://link.aip.org/link/phtoad/v26/i4/p11/s1.

Castillejo:1956:LSEa

[CDD56a] L. Castillejo, R. H. Dalitz, and F. J. Dyson. Low’s scattering equation for the charged and neutral scalar theories. Physical Review, 101(1):453–458, January 1, 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.101.453.

Castillejo:1956:LSEb

[CDD56b] L. Castillejo, R. H. Dalitz, and F. J. Dyson. Low’s scattering equation for the charged and neutral scalar theories. Series of Selected Papers in Physics [Butsurigaku rombun senshu], 77(??): 103–108, ???? 1956.

Cooper:2005:RBS

[CDM05] Fred Cooper, John F. Dawson, and Bogdan Mihaila. Renormalized broken-symmetry Schwinger–Dyson equations and the twoparticle irreducible 1/N expansion for the O(N) model. Physical Review D (Particles and Fields), 71(??):096003, May 9, 2005. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.71.096003.

Civitarese:2006:DBM

[CGR06] O. Civitarese, H. B. Geyer, and M. Reboiro. Dyson boson mapping of effective bi-fermion Hamiltonians. Physical Review C (Nuclear Physics), 73(??):034306, March 13, 2006. CODEN PRV-CAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http:

//link.aps.org/doi/10.1103/PhysRevC.73.034306.

Calzetta:1999:SDC

[CH99] Esteban Calzetta and B. L. Hu. Stochastic dynamics of correlations in quantum field theory: From the Schwinger–Dyson to Boltzmann–Langevin equation. Physical Review D (Particles and Fields), 61(??):025012, December 27, 1999. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
61.025012.

Cahill:1980:PTI

[CJ80] R. T. Cahill and R. T. Janus. ϕ^{2}σ phase transition and interplay of Schwinger–Dyson equation solutions. Physical Review D (Particles and Fields), 22(??):1979–??, October 15, 1980. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.22.1979.

Cvitanovic:1974:FDR

[CK74] Predrag Cvitanovic and T. Kinoshita. Feynman–Dyson rules in parametric space. Physical Review D (Particles and Fields), 10 (??):3978–??, December 15, 1974. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.10.3978.

Cheng:1994:BSS

[CK94] G. Cheng and T. K. Kuo. Bifurcation solutions of the Schwinger– Dyson equation. Journal of Mathematical Physics, 35(12):6270– 6290, December 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v35/i12/p6270_s1.

Cheng:1997:CSB

[CK97] G. Cheng and T. K. Kuo. Condition on the symmetry-breaking solution of the Schwinger–Dyson equation. Journal of Mathematical Physics, 38(12):6119–6125, December 1997. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Campbell-Kelly:2012:NCR

[CK12] Martin Campbell-Kelly. Von Neumann’s computer: [review of Turing’s Cathedral: the Origins of the Digital Universe, George Dyson, 2012 Pantheon Books £25.00 / $29.95 hardcover 423pp]. Physics World, 25(12):44–45, December 2012. CO-DEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic).

URL http://iopscience.iop.org/pwa/full/pwa-pdf/25/12/
phwv25i12a42.pdf.

Carlen:2022:TIA

[CL22] Eric A. Carlen and Elliott H. Lieb. A trace inequality of Ando, Hiai, and Okubo and a monotonicity property of the Golden–Thompson inequality. Journal of Mathematical Physics, 63(6):062203, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:

//pubs.aip.org/aip/jmp/article/63/6/062203/2846030/A-
trace-inequality-of-Ando-Hiai-and-Okubo-and-a. Special

collection in honor of Freeman Dyson.

Casalbuoni:2010:ASC

[CLO10] Roberto Casalbuoni, Massimo Ladisa, and Valerio Olevano. Aspects of self-consistency in the Dyson–Schwinger approach to QED and λ(φ^{⋆}φ)^{2} theories. Physical Review D (Particles and Fields), 82(??):094023, November 17, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
82.094023.

Cocconi:1959:SIC

[CM59] Giuseppe Cocconi and Philip Morrison. Searching for interstellar communications. Nature, 184(4690):844–846, September 19, 1959. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687

(electronic). URL http://www.nature.com/nature/journal/
v184/n4690/pdf/184844a0.pdf.

Caldi:1990:PGS

[CM90] D. G. Caldi and George D. Mostow, editors. Proceedings of the Gibbs Symposium: Yale University, May 15–17, 1989. American Mathematical Society, Providence, RI, USA, 1990. ISBN 0-8218-0157-0. LCCN QC310.15 .G53 1989. URL http://www.gbv.de/

dms/bowker/toc/9780821801574.pdf.

Cotanch:2003:LDS

[CM03] Stephen R. Cotanch and Pieter Maris. Ladder Dyson–Schwinger calculation of the anomalous γ − 3π form factor. Physical Review D (Particles and Fields), 68(??):036006, August 29, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.68.036006.

Cooper:2004:RSD

[CMD04] Fred Cooper, Bogdan Mihaila, and John F. Dawson. Renormalizing the Schwinger–Dyson equations in the auxiliary field formulation of λφ^{4} field theory. Physical Review D (Particles and Fields), 70(??):105008, November 11, 2004. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
70.105008.

Cohen:1971:SMT

[Coh71] E. G. D. Cohen, editor. Statistical Mechanics at the Turn of the Decade. Marcel Dekker, New York, NY, USA, 1971. ISBN 0-8247-1111-4. LCCN QC175 .S7.

Cornwell:1995:NIF

[Cor95] John Cornwell, editor. Nature’s imagination: the frontiers of scientific vision. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-851775-0. xii + 212 pp. LCCN ????

Cornwell:1997:NIF

[Cor97] John Cornwell. Nature’s imagination: the frontiers of scientific vision, L’immaginazione della natura: le frontiere della visione scientifica. Saggi scientifici / [Bollati Boringhieri]. Boringhieri, Torino, Italia, 1997. ISBN 88-339-1039-3. 291 pp. LCCN ????

Cordell:2010:FDW

[Cor10] Bruce Cordell. Freeman Dyson on what to do next in space: Laser propulsion? Terraforming? Thinking longterm? Web review of Dyson talk at the International Space Development Conference (ISDC 2010) in Chicago, IL,

USA., June 1, 2010. URL http://21stcenturywaves.com/
2010/06/01/freeman-dyson-on-what-to-do-next-in-space-
laser-propulsion-terraforming-thinking-long-term/;   http:
//isdc.nss.org/2010/.

Curtis:1990:TSD

[CP90] D. C. Curtis and M. R. Pennington. Truncating the Schwinger– Dyson equations: How multiplicative renormalizability and the Ward identity restrict the three-point vertex in QED. Physical Review D (Particles and Fields), 42(??):4165–??, December 15, 1990. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.42.4165.

Curtis:1993:NSF

[CP93] D. C. Curtis and M. R. Pennington. Nonperturbative study of the fermion propagator in quenched QED in covariant gauges using a renormalizable truncation of the Schwinger–Dyson equation. Physical Review D (Particles and Fields), 48(??):4933–??, November 15, 1993. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

link.aps.org/doi/10.1103/PhysRevD.48.4933.

Castanares:1999:BRI

[CQ99] Wenceslao Castañares and José Luis González Quirós. Book review: La imaginación digital: Signos del futuro: Tecnociencia y cibercultura by Stanley Aronowitz, Barbara Martinsons, Michael Menser, and Patrick Ducher. Los ordenadores emocionales by Rosalind W. Picard and Linda Edwars. Velocidad de escape by Mark Dery and Ramón Montoya. Mundos de futuro by Freeman Dyson and Joandomènee Ros. Vida simulada en el ordenador by Claus Emmeche and Carlos de la Reta. Visiones by Michio Kaku by Fabián Chueca. Crítica de la razón informática by Tomás Maldonado and Juan Carlos Gentile Vitale. La vida en la pantalla by Sherry Turkle and Laura Trafí. Revista de libros, 193(30):33–36,

June 1999. URL http://www.jstor.org/stable/30228987.

Crease:2001:BRBb

[Cre01] Robert P. Crease. Book review: Imagined Worlds by Freeman Dyson. Isis, 92(4):755, December 2001. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.

jstor.org/stable/3080349.

Crease:2013:MGP

[Cre13] Robert P. Crease. Book review: Maverick Genius: The Pioneering Odyssey of Freeman Dyson. Nature, 494(7437):311, February 20, 2013. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic).

Curtin:1982:ADS

[Cur82] Deane W. Curtin, editor. The aesthetic dimension of science: 1980 Nobel Conference [7–8 October, London, 1980]. Philosophical Library, New York, NY, USA, 1982. ISBN 0-8022-2393-1. LCCN Q174 .N6 1980. US$12.50.

Daboul:1975:ICD

[Dab75] Jamil Daboul. Incorporating Castillejo–Dalitz–Dyson poles as input in N/D calculations. Physical Review D (Particles and Fields), 12(??):888–??, August 1, 1975. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

URL http://link.aps.org/doi/10.1103/PhysRevD.12.888.

Daison:1965:MFR

[Dai65] F. Daison. Matematika i fizika. (Russian) [Mathematics and physics]. Uspekhi Fizicheskikh Nauk, 85(2):351–366, February 1965. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652

(electronic). URL http://ufn.ru/ru/articles/1965/2/e/.

Daison:1967:TSF

[Dai67] F. Daison. Tomonaga, Shvinger i Fejnman — laureaty nobelevskoj premii po fizike. (Russian) [Tomonaga, Schwinger and Feynman — Nobel Prize in Physics]. Uspekhi Fizicheskikh Nauk, 91(1):71–73, 1967. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://ufn.ru/ru/

articles/1967/1/f/.

Daison:1971:BFR

[Dai71] F. Daison. Budushhee fiziki. (Russian) [The future of physics]. Uspekhi Fizicheskikh Nauk, 103(3):529–538, 1971. CODEN UF-NAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL

http://ufn.ru/ru/articles/1971/3/e/. Russian translation

of [Dys70f].

Daison:1980:MO

[Daĭ80] F. Dž. Daĭson. Missed opportunities. Russian Mathematical Surveys, 35(1(211)):171–191, 1980. ISSN 0042-1316. Translated from the English by M. I. Monastyrskiĭ.

Daison:2010:PLV

[Dai10] F. Daison. Pticy i ljagushki v matematike i fizike. (Russian) [birds and frogs in mathematics and physics]. Uspekhi Fizicheskikh Nauk, 180(8):859–870, August 2010. CODEN UFNAAG. ISSN 0042-1294 (print), 1996-6652 (electronic). URL http://

ufn.ru/ru/articles/2010/8/f/.

Dajson:1973:UFP

[Daj73] F. Dajson. Ustojcivost’ i fazovye perechody. (Russian) [Stability and Phase Transitions]. Izdatelsvo Mir, Moscow, USSR, 1973. 373 pp. LCCN ???? Translation to Russian by S. P. Malysenko and E. G. Skrockoj.

Dyson:1983:VWS

[DARM83] Freeman J. Dyson, Raymond Aron, Joan Robinson, and Sterling M. McMurrin, editors. Values at war: selected Tanner lectures on the nuclear crisis. University of Utah Press, Salt Lake City, UT, USA, 1983. ISBN 0-87480-226-1. x + 132 pp. LCCN U263 .D95 1983. US$6.00.

Davis:1978:IFD

[Dav78] Monte Davis. Interview: Freeman Dyson. Omni (New York), 1(1):100–107, October 1978. ISSN 0149-8711. URL http:/

/archive.org/details/omni-magazine-1978-10;   http://
omnimagindex.files.wordpress.com/2012/02/78-10-004.jpg;
 http://omnimagindex.wordpress.com/author-indices/authors-

Davis:1992:SDA

[Dav92] John E. Davis. Schwinger–Dyson analysis of an exactly solvable model. Physical Review D (Particles and Fields), 46(??):1750–??, August 15, 1992. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

link.aps.org/doi/10.1103/PhysRevD.46.1750.

Dawidoff:2009:CHI

[Daw09] Nicholas Dawidoff. The civil heretic: interview with Freeman Dyson. The New York Times Magazine, ??(??):32–39, 54–59, March 29, 2009. ISSN 0362-1308. URL http://www.nytimes.

com/2009/03/29/magazine/29Dyson-t.html. Dyson appears on

the magazine cover.

Dyson:2002:SGI

[DB02] Freeman J. Dyson and Michael Bischoff. Die Sonne, das Genom und das Internet: wissenschaftliche Innovation und die Technologien der Zukunft. (German) [The Sun, the Genome and the Internet: scientific innovation and technologies future], volume 15239 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, lizenzausg edition, 2002. ISBN 3-596-15239-9, 3-10-015335-9. 221 pp. LCCN ???? EUR 8.90. URL http:/

/www.gbv.de/dms/bsz/toc/bsz087284472inh.pdf;   http:/
/www.gbv.de/dms/faz-rez/FR220001028640893.pdf.

Dyson:2009:LWU

[DBB09] Freeman Dyson, John Brockman, and Kurt Beginnen. Leben, was ist das?: Ursprünge, Phänomene und die Zukunft unserer Wirklichkeit. (German) [What is Life?: Origins, phenomena, and the future of our reality]. Fischer, Frankfurt am Main, 2009. ISBN 3-596-18240-9 (paperback). 164 pp. LCCN ???? sfr 18.60 (freier Pr.), EUR 9.95. URL http:/

/deposit.d-nb.de/cgi-bin/dokserv?id=3107511%26prov=M%
26dok_var=1%26dok_ext=htm;   http://www.gbv.de/dms/faz-
rez/FD1N200812312101998_1.pdf.

Droussent:1952:APSa

[DC52] Lucien Droussent and E. D. Camier. Advanced problems and solutions: 4398. American Mathematical Monthly, 59(1):47–48, January 1952. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gal50].

Dajson:1973:NZP

[DC73] F. Dajson and D. Ter Chaar. Nejtronnye zvezdy i pul’sary. (Russian) [Neutron Stars and Pulsars]. Izdatelsvo Mir, Moscow, USSR, 1973. 191 pp. LCCN ????

Duffy:1994:AKS

[DCCS94] Patrick Duffy, Delano P. Chong, Mark E. Casida, and Dennis R. Salahub. Assessment of Kohn–Sham density-functional orbitals as approximate Dyson orbitals for the calculation of electronmomentum-spectroscopy scattering cross sections. Physical Review A (Atomic, Molecular, and Optical Physics), 50(??):4707–??, December 1, 1994. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.aps.org/

doi/10.1103/PhysRevA.50.4707.

Dalitz:1955:RNT

[DD55] R. H. Dalitz and F. J. Dyson. Renormalization in the new Tamm– Dancoff theory of meson-nucleon scattering. Physical Review (2), 99(1):301–314, July 1, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.99.301.

Damour:1996:OBTa

[DD96a] Thibault Damour and Freeman Dyson. The Oklo bound on the time variation of the fine-structure constant revisited. arxiv.org, ??(??):1–23, June 27, 1996. URL http://arxiv.org/pdf/hepph/9606486. IHES/P/96/40. IASSNS-HEP-96/62.

Damour:1996:OBTb

[DD96b] Thibault Damour and Freeman Dyson. The Oklo bound on the time variation of the fine-structure constant revisited. Nuclear Physics B, 480(1–2):37–54, November 25, 1996. CODEN NUPBBO. ISSN 0550-3213 (print), 1873-1562 (electronic).

Denner:1996:DSV

[DD96c] Ansgar Denner and Stefan Dittmaier. Dyson summation without violating Ward identities and the Goldstone-boson equivalence theorem. Physical Review D (Particles and Fields), 54 (??):4499–??, October 1, 1996. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.54.4499.

Dyson:2004:POT

[DD04] Freeman J. Dyson and George Dyson. In praise of open thinking. 42 minute audio recording., July 29, 2004. URL http://itc.

conversationsnetwork.org/shows/detail170.html.

Detmold:2003:SCV

[Det03] W. Detmold. Solution of coupled vertex and propagator Dyson– Schwinger equations in the scalar Munczek–Nemirovsky model. Physical Review D (Particles and Fields), 67(??):085011, April 21, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.67.085011.

Dyson:1992:PDC

[DF92] Freeman J. Dyson and Harold Falk. Period of a discrete cat mapping. American Mathematical Monthly, 99(7):603–614, August/ September 1992. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
2324989.

Dyson:2010:LIS

[DFG10] F. J. Dyson, N. E. Frankel, and M. L. Glasser. Lehmer’s interesting series. arxiv.org, ??(??):1–14, September 22, 2010. URL

http://arxiv.org/abs/1009.4274.

Dyson:2013:LIS

[DFG13] Freeman J. Dyson, Norman E. Frankel, and M. Lawrence Glasser. Lehmer’s interesting series. American Mathematical Monthly, 120 (2):116–130, February 2013. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://www.jstor.org/

stable/10.4169/amer.math.monthly.120.02.116.

DelVecchio:2022:LIB

[DFPR22] Simone Del Vecchio, Jürg Fröhlich, Alessandro Pizzo, and Stefano Rossi. Local iterative block-diagonalization of gapped Hamiltonians: a new tool in singular perturbation theory. Journal of Mathematical Physics, 63(7):073503, July 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
073503/2843679/Local-iterative-block-diagonalization-

of-gapped. Special collection in honor of Freeman Dyson.

Dundarer:1984:DRT

[DG84] Resit Dündarer and Feza Gürsey. Dyson representation of SU(3) in terms of five boson operators. Journal of Mathematical Physics, 25(3):431–432, March 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v25/i3/p431_s1.

Dyson:2005:BRM

[DG05] Freeman Dyson and Luis Gago. Book review: El mundo sobre una cuerda. The Fabric of the Cosmos: Space, Time, and the Texture of Reality by Brian Greene. Revista de libros de la Fundación Caja Madrid, 67(99):24–28, March 2005. URL http://www.jstor.

org/stable/30230488.

Dyson:2001:LEA

[DGB^{+}01] Freeman Dyson, Kurt Gottfried, Luis J. Boya, Karl von Meyenn, and Engelbert L. Schucking. Letters to the Editor: Another visit with Wolfgang Pauli. Physics Today, 54(8):11–13, August 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

labs.adsabs.harvard.edu/ui/abs/2001PhT....54h..11D;
http://link.aip.org/link/phtoad/v54/i8/p11/s1.

Dyson:1967:TNW

[DGWW67] F. J. Dyson, R. Gomer, S. Weinberg, and S. C. Wright. Tactical nuclear weapons in Southeast Asia (U). Study S-266, Institute for Defense Analyses, JASON Division, Washington, DC, USA, March 1967. iv + 55 + 1 pp. URL https://irp.fas.org/

agency/dod/jason/tactical.pdf.

Dennett:2006:LEB

[DHD06] Daniel C. Dennett, Nicholas Humphrey, and Freeman J. Dyson. Letter to the Editor: ‘breaking the spell’. New York Review of Books, 53(13):65, August 10, 2006. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2006/aug/10/breaking-the-spell/. In response to

[Dys06d].

Dyson:1965:TAP

[DhX65a] Freeman J. Dyson and Nguyen huu Xuong. SU(6) theory and the antiproton-proton annihilation at rest into two mesons. Physical Review Letters, 14(16):654–657, April 19, 1965. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-

0145. URL http://link.aps.org/doi/10.1103/PhysRevLett.
14.654.

Dyson:1965:EST

[DhX65b] Freeman J. Dyson and Nguyen huu Xuong. Erratum: Y = 2 states in Su(6) theory. Physical Review Letters, 14(9):339, March 1, 1965. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://link.aps.org/doi/10.
1103/PhysRevLett.14.339.

Dirac:1937:CC

[Dir37] P. A. M. Dirac. The cosmological constants. Nature, 139(3512): 323, February 20, 1937. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://adsabs.harvard.

edu/abs/1937Natur.139..323D;   http://www.nature.com/
nature/journal/v139/n3512/pdf/139323a0.pdf. See related

work in [Dir38, Tel48, Wil58, Dys67d, Gam67a, Gam67b, Gam68, Alp73].

Dirac:1938:NBC

[Dir38] P. A. M. Dirac. A new basis for cosmology. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 165(921):199–208, April 5, 1938. CODEN PRLAAZ. ISSN

0080-4630. URL http://adsabs.harvard.edu/abs/1938RSPSA.
165..199D;   http://www.jstor.org/stable/97105. See re-

lated work in [Dir37, Tel48, Wil58, Dys67d, Gam67a, Gam67b, Gam68, Alp73].

deJonghe:1994:SDB

[dJ94] Frank de Jonghe. Schwinger–Dyson Becchi–Rouet–Stora–Tyutin symmetry and the Batalin–Vilkovisky Lagrangian quantization of gauge theories with open or reducible gauge algebras. Journal of Mathematical Physics, 35(6):2734–2742, June 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v35/i6/
p2734_s1;   http://link.aip.org/link/?JMP/35/2734/1.

Deo:1995:DSL

[DJS95] Nivedita Deo, Sanjay Jain, and B. Sriram Shastry. Dyson– Schwinger loop equations of the two-matrix model: Correlations in quantum chaos. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 52(??):4836–??, November 1, 1995. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.52.4836.

Dyson:1970:LED

[DK70] Freeman J. Dyson and Henry A. Knoll. Letters to the Editor: In defense of Gofman and Tamplin. Physics Today, 23(7):9–11, July 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/23/9/1.

Deleuze:2006:SMS

[DK06] M. S. Deleuze and S. Knippenberg. Study of the molecular structure, ionization spectrum, and electronic wave function of 1,3butadiene using electron momentum spectroscopy and benchmark Dyson orbital theories. Journal of Chemical Physics, 125(10): 104309, 2006. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/?JCP/125/
104309/1.

Dyson:1967:SM

[DL67] Freeman J. Dyson and A. Lenard. Stability of matter. I. Journal of Mathematical Physics, 8(3):423–434, March 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v8/i3/
p423_s1;   http://link.aip.org/link/?JMP/8/423/1.

Dyson:1976:PTQ

[DLS76] Freeman J. Dyson, Elliott H. Lieb, and Barry Simon. Phase transitions in the quantum Heisenberg model. Physical Review Letters, 37(3):120–123, July 19, 1976. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:

//link.aps.org/doi/10.1103/PhysRevLett.37.120.

Dyson:1978:PTQ

[DLS78] Freeman J. Dyson, Elliott H. Lieb, and Barry Simon. Phase transitions in quantum spin systems with isotropic and nonisotropic interactions. Journal of Statistical Physics, 18(4):335– 383, 1978. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dyson:2010:SAH

[DLV10] Freeman Dyson, John S. Lewis, and Lee Valentine. Statement about the house’s funding bill, H.R. 5781 for NASA on behalf of The Space Studies Institute. Letter to US Congress., Septem-

ber 10, 2010. URL http://www.parabolicarc.com/2010/09/
10/freemon-dyson-congress-reject-hr-5781-choose-side-
history/.

Dyson:1957:PLP

[DM57a] Freeman J. Dyson and Kirk McVoy. Persistence of longitudinal polarization in an electromagnetic cascade. Unpublished preprint., 1957.

Dyson:1957:AQM

[DM57b] Freeman J. Dyson and Michael J. Moravcsik. Advanced quantum mechanics. Laboratory of Nuclear Studies, Cornell University, Ithaca, NY, USA, 1957. 213 pp.

Dyson:1963:STE

[DM63] Freeman J. Dyson and Madan Lal Mehta. Statistical theory of the energy levels of complex systems. IV. Journal of Mathematical Physics, 4(5):701–712, May 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i5/
p701_s1;   http://link.aip.org/link/?JMP/4/701/1.

Dyson:1976:IMS

[DMZ76] Freeman Dyson, Walter Munk, and Bernard Zetler. Interpretation of multipath scintillations Eleuthera to Bermuda in terms of internal waves and tides. Journal of the Acoustical Society of America, 59(5):1121–1133, 1976. CODEN JASMAN. ISSN

0001-4966. URL http://link.aip.org/link/?JAS/59/1121/1.

Dombey:1991:CFP

[Dom91] Norman Dombey. Comment on “Feynman’s proof of the Maxwell equations,” by F. J. Dyson [Am. J. Phys. 58, 209–211 (1990)]. American Journal of Physics, 59(1):85, 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://

link.aip.org/link/?AJP/59/85/1.

Donohue:2002:BRS

[Don02] A. A. Donohue. Book review: Stephen L. Dyson: Ancient Marbles to American Shores: Classical Archaeology in the United States. Isis, 93(2):294–295, June 2002. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://www.jstor.org/

stable/10.1086/344983.

Dudal:2012:NGG

[DORQ12a] D. Dudal, O. Oliveira, and J. Rodríguez-Quintero. Nontrivial ghost-gluon vertex and the match of the refined-Gribov– Zwanziger, Dyson–Schwinger equations, and lattice Yang–Mills propagators. Physical Review D (Particles and Fields), 86(??): 105005, November 2, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.86.105005.

Dudal:2012:PNN

[DORQ12b] D. Dudal, O. Oliveira, and J. Rodriguez-Quintero. Publisher’s note: Nontrivial ghost-gluon vertex and the match of the refined Gribov–Zwanziger, Dyson–Schwinger equations, and lattice Yang–Mills propagators [Phys. Rev. D 86, 105005 (2012)]. Physical Review D (Particles and Fields), 86(??):109902, November 13, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.86.109902.

Dyson:1964:JRO

[DPSY64] Freeman Dyson, Abraham Pais, Bengt Stromgren, and Chen Ning Yang. To J. Robert Oppenheimer on his sixtieth birthday. Reviews of Modern Physics, 36(2):507–508, April 1964. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.36.507;   http://rmp.aps.org/abstract/RMP/
v36/i2/p507_1.

Dreifus:2000:CFJ

[Dre00] Claudia Dreifus. A conversation with Freeman J. Dyson: Abombs, space chickens and God. The New York Times Magazine, ??(??):??, August 1, 2000. ISSN 0362-1308. URL http://www.

nytimes.com/2000/08/01/science/a-conversation-with-freeman-
j-dyson-a-bombs-space-chickens-and-god.html.

Dreifus:2001:SCI

[Dre01] Claudia Dreifus. Scientific conversations: interviews on science from the New York Times. Times Books/Henry Holt, New York, NY, USA, 2001. ISBN 0-7167-4661-1. xxii + 250 pp. LCCN Q141 .D74 2001. URL http://www.loc.gov/

catdir/enhancements/fy0667/2001043630-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0667/2001043630-d.

html;   http://www.loc.gov/catdir/enhancements/fy0667/

2001043630-t.html. With a foreword by Natalie Angier.

Dronamraju:1995:HBR

[Dro95] Krishna R. Dronamraju, editor. Haldane’s Daedalus revisited. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1995. ISBN 0-19-854846-X. xiv + 147 pp. LCCN Q171 .H1566 1995. With an introduction by Krishna R. Dronamraju. Foreword by Joshua Lederberg.

Dyson:1954:MNS

[DRS^{+}54] F. J. Dyson, M. Ross, E. E. Salpeter, S. S. Schweber, M. K. Sundaresan, W. M. Visscher, and H. A. Bethe. Meson-nucleon scattering in the Tamm–Dancoff approximation. Physical Review, 95(6):1644–1658, September 15, 1954. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http:

//link.aps.org/doi/10.1103/PhysRev.95.1644;   http://
prola.aps.org/abstract/PR/v95/i6/p1644_1.

deSyon:1999:BRB

[dS99] Guillaume de Syon. Book review: Imagined Worlds by Freeman Dyson. Utopian Studies, 10(2):234–236, ???? 1999. ISSN 1045-991X (print), 2154-9648 (electronic). URL http://www.jstor.

org/stable/20718114.

Dyson:2007:SR

[DS07] Freeman Dyson and Glenn Suter. The scientist as rebel. Integrated Environmental Assessment and Management, 3(2):305– 306, April 2007. ISSN 1551-3777 (print), 1551-3793 (electronic).

Dyson:1947:LFT

[DSW51] Freeman J. Dyson, Robert Serber, and Gregor Wentzel. Lectures on field theory. Report, Cornell University, Ithaca, NY, USA, 1947–1951.

Dyson:1978:EHS

[DTO78] Freeman J. Dyson, S. Tremaine, and J. P. Ostriker. The effect of a halo on the stability of uniformly rotating systems. Submitted to the Astrophysical Journal., 1978.

Dyson:1951:APS

[DTS51] F. J. Dyson, Ernest Trost, and G. Szegő. Advanced problems and solutions: Solutions: 4389. American Mathematical Monthly, 58 (9):640–641, November 1951. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UTB^{+}50].

Dyson:1955:ABP

[DÜ55] F. J. Dyson and H. Überall. Anisotropy of Bremsstrahlung and pair production in single crystals. Physical Review, 99(2):604–605, July 15, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.99.604.

Durell:1960:RR

[Dur60] Clement V. (Clement Vavasor) Durell. Readable relativity. Harper, New York, NY, USA, 1960. xi + 146 pp. LCCN ???? Foreword by Freeman J. Dyson. Reprint of 1926 original edition: London: G. Bell.

Durell:1962:RR

[Dur62] Clement V. (Clement Vavasor) Durell. Readable relativity. Bell, London, UK, 1962. xi + 146 pp. LCCN ???? Foreword by Freeman J. Dyson. Reprint of 1926 original edition.

Durrani:2000:RPF

[Dur00] Matin Durrani. Religious physicist with faith in science. Physics World, 13(6):10, June 2000. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://stacks.iop.

org/2058-7058/13/i=6/a=11.

Drobot:1963:MMP

[DV63] Stefan Drobot and Paul A. Viebrock, editors. Mathematical models in physical sciences: proceedings. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1963. LCCN QA401 .C65 1962.

Dahlen:2005:SCS

[DvL05] Nils Erik Dahlen and Robert van Leeuwen. Self-consistent solution of the Dyson equation for atoms and molecules within a conserving approximation. Journal of Chemical Physics, 122(16): 164102, 2005. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/?JCP/122/
164102/1.

DeToledo:1948:RAN

[DW48] Paulo Saraiva De Toledo and Gleb Wataghin. On the relative abundances of nuclei in the universe. Physical Review, 73 (1):79–80, January 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.73.79.2. See remarks in [Dys93a] about

the relation of this work to [?], and the subsequent incorrect neglect of Wataghin’s work. See also related papers [?, ?, Wat48].

Dyson:1989:FTA

[DW89] Freeman J. Dyson and Edward Witten. Free Taysir Aruri! New York Review of Books, 36(13):60, August 17, 1989. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/archives/1989/aug/17/free-taysir-aruri/.

Dyson:1964:ST

[DX64] Freeman J. Dyson and Nguyen-Huu Xuong. Y = 2 states in Su(6) theory. Physical Review Letters, 13(26):815–817, December 28, 1964. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.aps.org/doi/

10.1103/PhysRevLett.13.815. See erratum [DhX65b].

Dyson:1943:NK

[Dys43a] F. J. Dyson. A note on kurtosis. Journal of the Royal Statistical Society, 106(4):360–361, 1943. ISSN 0952-8385. URL http://

www.jstor.org/stable/2980484.

Dyson:1943:OMP

[Dys43b] F. J. Dyson. On the order of magnitude of the partial quotients of a continued fraction. Journal of the London Mathematical Society, 18:40–43, 1943. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Dyson:1943:TIC

[Dys43c] F. J. Dyson. Three identities in combinatory analysis. Journal of the London Mathematical Society, 18:35–39, 1943. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Dyson:1944:PEE

[Dys44a] Freeman J. Dyson. A proof that every equation has a root. Eureka (Cambridge), 8(??):3–4, ???? 1944.

Dyson:1944:SGT

[Dys44b] Freeman J. Dyson. Some guesses in the theory of partitions. Eureka (Cambridge), 8(??):10–15, ???? 1944. See later work [AO05, Mah05, EG11].

Dyson:1945:TDS

[Dys45a] F. J. Dyson. A theorem on the densities of sets of integers. Journal of the London Mathematical Society, 20:8–14, 1945. CODEN JLMSAK. ISSN 0024-6107 (print), 1469-7750 (electronic).

Dyson:1943:NCL

[Dys45b] Freeman J. Dyson. Note on the comparison of loss rates. Report, Operational Research Section Headquarters, R.A.F. Bomber Command, London, UK, 1943–1945.

Dyson:1946:PP

[Dys46] Freeman J. Dyson. The problem of the pennies. Mathematical Gazette (London), 30(29):231–234, October 1946.

Dyson:1947:AAN

[Dys47a] F. J. Dyson. The approximation to algebraic numbers by rationals. Acta Mathematica, 79:225–240, 1947. CODEN ACMAA8. ISSN 0001-5962 (print), 1871-2509 (electronic).

Dyson:1947:SDA

[Dys47b] F. J. Dyson. On simultaneous Diophantine approximations. Proceedings of the London Mathematical Society, Series 2, 49:409– 420, 1947. ISSN 0024-6115 (print), 1460-244x (electronic).

Dyson:1948:PFN

[Dys48a] F. J. Dyson. On the product of four non-homogeneous linear forms. Annals of Mathematics (2), 49(1):82–109, 1948. CODEN ANMAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Dyson:1948:TAT

[Dys48b] F. J. Dyson. A theorem in algebraic topology. Annals of Mathematics (2), 49(1):75–81, 1948. CODEN ANMAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Dyson:1948:ESE

[Dys48c] Freeman J. Dyson. The electromagnetic shift of energy levels. Physical Review, 73(6):617–626, March 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.73.617.

Dyson:1948:INM

[Dys48d] Freeman J. Dyson. The interactions of nucleons with meson fields. Physical Review, 73(8):929–930, April 15, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.73.929.2.

Dyson:1949:RTT

[Dys49a] F. J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman. Physical Review (2), 75(3):486–502, February 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/

PhysRev.75.486. Reprinted in [Dys58j].

Dyson:1949:MQE

[Dys49b] Freeman J. Dyson. The S-matrix in quantum electrodynamics. Physical Review (2), 75(11):1736–1755, June 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.

75.1736. Reprinted in [Dys58k].

Dyson:1950:APS

[Dys50a] F. J. Dyson. Advanced problems and solutions: Problems for solution: 4389. American Mathematical Monthly, 57(3):188–189, March 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Bat51, DTS51, ES51, TK51b, UL51].

Dyson:1950:LPQ

[Dys50b] F. J. Dyson. Longitudinal photons in quantum electrodynamics. Physical Review, 77(3):420, February 1, 1950. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://

link.aps.org/doi/10.1103/PhysRev.77.420.

Dyson:1950:NTU

[Dys50c] Freeman J. Dyson. Notes taken from an unfinished manuscript of Professor Julian Schwinger, On Gauge Invariance and Vacuum Polarization. Unpublished., 1950.

Dyson:1950:RDH

[Dys50d] Freeman J. Dyson. Recent developments in high energy physics. Report, University of Michigan, Ann Arbor, MI, USA, 1950. Notes of a course of lectures given by Professor Robert Serber at the Summer Physics Symposium of the University of Michigan, June 28–July 17, 1950.

Dyson:1950:RDQ

[Dys50e] Freeman J. Dyson. Recent developments in quantum electrodynamics. Report, University of Michigan, Ann Arbor, MI, USA, 1950. 41 pp. Notes prepared by David Park, Summer Symposium, University of Michigan, 1950.

Dyson:1951:AQM

[Dys51a] F. J. Dyson. Advanced quantum mechanics: lectures notes by Professor F. J. Dyson for a course in relativistic quantum mechanics given at Cornell University in the fall of 1951. Report, Laboratory of Nuclear Studies, Cornell University, Ithaca, NY, USA, 1951. 148 pp.

Dyson:1951:CFD

[Dys51b] F. J. Dyson. Continuous functions defined on spheres. Annals of Mathematics (2), 54(3):534–536, 1951. CODEN ANMAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Dyson:1951:HOQa

[Dys51c] F. J. Dyson. Heisenberg operators in quantum electrodynamics. I. Physical Review (2), 82(3):428–439, May 1, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.82.428.

Dyson:1951:HOQb

[Dys51d] F. J. Dyson. Heisenberg operators in quantum electrodynamics. II. Physical Review (2), 83(3):608–627, August 1, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.83.608.

Dyson:1951:RMQ

[Dys51e] F. J. Dyson. The renormalization method in quantum electrodynamics. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 207(??):395–401, 1951. CO-DEN PRLAAZ. ISSN 0962-8444.

Dyson:1951:SEQ

[Dys51f] F. J. Dyson. The Schrödinger equation in quantum electrodynamics. Physical Review (2), 83(6):1207–1216, September 15, 1951. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.83.1207.

Dyson:1951:QML

[Dys51g] Freeman J. Dyson. Quantum mechanics: Lecture notes for a course in ‘relativistic quantum mechanics’ given at Cornell University, 1951. Technical report, Princeton University, Princeton, NJ, USA, 1951.

Dyson:1952:DPT

[Dys52a] F. J. Dyson. Divergence of perturbation theory in quantum electrodynamics. Physical Review (2), 85(4):631–632, February 15, 1952. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.85.631.

Dyson:1952:LNA

[Dys52b] Freeman J. Dyson. Lecture notes on advanced quantum mechanics. Report, Cornell Laboratory of Nuclear Studies, Cornell University, Ithaca, NY, USA, 1952. Reprinted in [Dys07a, Dys11k].

Dyson:1952:QE

[Dys52c] Freeman J. Dyson. Quantum electrodynamics. Physics Today, 5 (9):6–9, September 1952. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/phtoad/v5/i9/p6/s1.

Dyson:1953:FTD

[Dys53a] F. J. Dyson. Fourier transforms of distribution functions. Canadian Journal of Mathematics = Journal canadien de mathématiques, 5(??):554–558, ???? 1953. CODEN CJMAAB. ISSN 0008-414X (print), 1496-4279 (electronic).

Dyson:1953:MRT

[Dys53b] F. J. Dyson. Mass-renormalization with the Tamm–Dancoff method. Physical Review, 91(2):421–422, July 15, 1953. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.91.421.2.

Dyson:1953:WFRa

[Dys53c] F. J. Dyson. The wave function of a relativistic system. Physical Review (2), 91(6):1543–1550, September 15, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.91.1543.

Dyson:1953:DDLa

[Dys53d] Freeman J. Dyson. The dynamics of a disordered linear chain. Physical Review (2), 92(6):1331–1338, December 15, 1953. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.

92.1331. Reprinted in [Dys66c].

Dyson:1953:DDLb

[Dys53e] Freeman J. Dyson. The dynamics of a disordered linear chain. Series of Selected Papers in Physics [Butsurigaku rombun senshu], 119(19):61–68, ???? 1953. Check year: it is not given in Dyson’s bibliography.

Dyson:1953:FT

[Dys53f] Freeman J. Dyson. Field theory. Scientific American, 188 (4):57–64, April 1953. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v188/n4/pdf/scientificamerican0453-
57.pdf.

Dyson:1953:UTD

[Dys53g] Freeman J. Dyson. The use of the Tamm–Dancoff method in field theory. Physical Review, 90(5):994, June 1, 1953. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.90.994.

Dyson:1953:WFRb

[Dys53h] Freeman J. Dyson. The wave function of a relativistic system. Series of Selected Papers in Physics [Butsurigaku rombun senshu], 67(19):95–102, ???? 1953. Check year: it is not given in Dyson’s bibliography.

Dyson:1954:RGF

[Dys54a] F. J. Dyson. The rate of growth of functions defined by Dirichlet series. Annals of Mathematics (2), 60(3):437–446, 1954. CODEN ANMAAH. ISSN 0003-486x (print), 1939-8980 (electronic).

Dyson:1954:B

[Dys54b] Freeman J. Dyson. Books. Scientific American, 190(3): 92–99, March 1954. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v190/n3/pdf/scientificamerican0354-
92.pdf.

Dyson:1954:RBS

[Dys54c] Freeman J. Dyson. On the relation between scattering matrix elements and cross-sections. Report, L’Université de Grenoble, Grenoble, France, August 1954. Lecture notes mimeographed at L’École d’Été de Physique Théorique, Les Houches, Haute-Savoie, France. Included in [Dys11k].

Dyson:1954:RSE

[Dys54d] Freeman J. Dyson. Review of Sir Edmund Whittaker’s A History of the Theories of Aether and Electricity, II. Scientific American, 190(3):92–99, March 1954. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v190/n3/pdf/scientificamerican0354-92.pdf.

Dyson:1954:SBQ

[Dys54e] Freeman J. Dyson. Supplement to Advanced Quantum Mechanics. Report, L’Université de Grenoble, Grenoble, France, August 1954. Lecture notes mimeographed at L’École d’Été de Physique Théorique, Les Houches, Haute-Savoie, France. Included in [Dys11k].

Dyson:1954:WH

[Dys54f] Freeman J. Dyson. What is heat? Scientific American, 191(3): 58–63, September 1954. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.com/

scientificamerican/journal/v191/n3/pdf/scientificamerican0954-
58.pdf.

Dyson:1955:BRH

[Dys55a] Freeman J. Dyson. Book review: Hans A. Bethe and Frederic de Hoffmann, Mesons and Fields. Volume II: Mesons. Physics Today, 8(6):27, June 1955. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/?PTO/8/27/1;   http://link.aip.org/link/
phtoad/v8/i6/p27/s1.

Dyson:1955:BRW

[Dys55b] Freeman J. Dyson. Book review: Walter Thirring, Einführung in die Quantenelektrodynamik. Physics Today, 8(12):22, December 1955. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/8/22/2.

Dyson:1955:ESR

[Dys55c] Freeman J. Dyson. Electron spin resonance absorption in metals. II. Theory of electron diffusion and the skin effect. Physical Review, 98(2):349–359, April 15, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.98.349.

Dyson:1955:MPS

[Dys55d] Freeman J. Dyson. Meson-proton scattering in relativistic meson theory. In Bellamy and Moorehouse [BM55], pages 263–264. LCCN ????

Dyson:1955:SMF

[Dys55e] Freeman J. Dyson. Scattering of mesons by a fixed scatterer. Physical Review (2), 100(1):344–348, October 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.100.344.

Dyson:1955:SMNa

[Dys55f] Freeman J. Dyson. Second maximum in the negative pion scattering cross section. Physical Review, 99(3):1037, August 1, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.
99.1037.

Dyson:1955:SMNb

[Dys55g] Freeman J. Dyson. Second maximum in the negative pion scattering cross section. Series of Selected Papers in Physics [Butsurigaku rombun senshu], 84(??):154–155, ???? 1955.

Dyson:1956:SFb

[Dys56a] Freeman Dyson. Science and freedom. Science, 124(3219):432– 433, September 7, 1956. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.org/

stable/1750282;   http://www.sciencemag.org/content/124/
3219/432.full.pdf.

Dyson:1956:BRS

[Dys56b] Freeman J. Dyson. Book review: S. S. Schweber, H. A. Bethe, and F. De Hoffmann, Mesons and Fields. Vol. 1: Fields. Physics Today, 9(5):32–34, May 1956. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/?PTO/9/32/1;   http://link.aip.org/link/
phtoad/v9/i5/p32/s1.

Dyson:1956:GTS

[Dys56c] Freeman J. Dyson. General theory of spin-wave interactions. Physical Review (2), 102(5):1217–1230, June 1, 1956. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.102.1217. See

comments in [Dys86a].

Dyson:1956:PHW

[Dys56d] Freeman J. Dyson. Prof. Hermann Weyl, For. Mem. R.S.: [obituary]. Nature, 177(4506):457–458, March 10, 1956. CO-DEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (elec-

tronic). URL http://www.nature.com/nature/journal/v177/
n4506/pdf/177457a0.pdf.

Dyson:1956:SFa

[Dys56e] Freeman J. Dyson. Science and freedom. Baltimore Morning Sun, ??(??):14, June 26, 1956.

Dyson:1956:TBI

[Dys56f] Freeman J. Dyson. Thermodynamic behavior of an ideal ferromagnet. Physical Review (2), 102(5):1230–1244, June 1956. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.102.1230.

Dyson:1956:TGT

[Dys56g] Freeman J. Dyson. Two-group treatment of the warm neutron effect. Unpublished paper written at General Atomic, San Diego, in connection with the TRIGA reactor, 1956.

Dyson:1957:GSE

[Dys57a] F. J. Dyson. Ground-state energy of a hard-sphere gas. Physical Review, 106(1):20–26, April 1, 1957. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.106.20.

Dyson:1957:MSL

[Dys57b] F. J. Dyson. Meaning of the solutions of Low’s scattering equation. Physical Review, 106(1):157–159, April 1, 1957. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.106.157.

Dyson:1957:PC

[Dys57c] Freeman J. Dyson. Polarization in cascades. Unpublished preprint., 1957.

Dyson:1958:BRBb

[Dys58a] Freeman J. Dyson. Book review: Our Nuclear Future: Facts, Dangers and Opportunities. By Edward Teller and Albert L. Latter. 184 pp. Criterion Books, New York, 1958. $3.50. Physics Today, 11(8):38–40, August 1958. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/?PTO/11/38/3;   http://physicstoday.org/
resource/1/phtoad/v11/i8/p38_s3.

Dyson:1958:BRBa

[Dys58b] Freeman J. Dyson. Book review: Principles of Quantum Theory I. Vol. 5, Part 1 of Handbuch der Physik. Edited by S. Flügge. 376 pp. Springer-Verlag, Berlin, Germany, 1958. DM 90.00 (subscription price DM 72.00). Physics Today, 11(7):26, July 1958. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

URL http://link.aip.org/link/?PTO/11/26/1.

Dyson:1958:BRM

[Dys58c] Freeman J. Dyson. Book review: M. J. Lighthill, An Introduction to Fourier Analysis and Generalised Functions. Physics Today, 11(6):28, June 1958. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
11/28/1.

Dyson:1958:CBL

[Dys58d] Freeman J. Dyson. Connection between local commutativity and regularity of Wightman functions. Physical Review (2), 110(2): 579–581, April 15, 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.org/

doi/10.1103/PhysRev.110.579.

Dyson:1958:IP

[Dys58e] Freeman J. Dyson. Innovation in physics. Scientific American, 199(3):74–82, September 1958. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v199/n3/pdf/scientificamerican0958-
74.pdf.

Dyson:1958:IRD

[Dys58f] Freeman J. Dyson. Integral representation of a double commutator. Physical Review (2), 111(6):1717–1718, September 15, 1958. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRev.
111.1717.

Dyson:1958:IRCa

[Dys58g] Freeman J. Dyson. Integral representations of causal commutators. Physical Review (2), 110(6):1460–1464, June 15, 1958. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.110.1460.

Dyson:1958:IRCb

[Dys58h] Freeman J. Dyson. Integral representations of causal commutators. Series of Selected Papers in Physics [Butsurigaku rombun senshu], 104(??):27–31, ???? 1958.

Dyson:1958:NMO

[Dys58i] Freeman J. Dyson. Note on maximum opacity. Report GAMD-469, Project Orion, General Atomic Division of General Dynamics Corporation, ????, July 8, 1958. URL http://www.

daviddarling.info/encyclopedia/O/OrionProj.html.

Dyson:1958:RTT

[Dys58j] Freeman J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman. In Schwinger [Sch58], page ?? ISBN 0-486-60444-6. LCCN QC680 .S35. Reprint of [Dys49a].

Dyson:1958:MQE

[Dys58k] Freeman J. Dyson. The S-matrix in quantum electrodynamics. In Schwinger [Sch58], page ?? ISBN 0-486-60444-6. LCCN QC680 .S35. Reprint of [Dys49b].

Dyson:1959:CBE

[Dys59a] Freeman J. Dyson. Containment of bomb explosions. In Proceedings of the Conference on Scientific Applications of Nuclear Explosions, held July 6–8, 1959, Los Alamos, New Mexico, pages 55–60. Los Alamos National Laboratory, Los Alamos, NM, USA, 1959. LCCN ???? LAMS-2443.

Dyson:1959:LEC

[Dys59b] Freeman J. Dyson. Letter to the Editor: Caution on ban urged. Bulletin of the Atomic Scientists, 15(10):401, December 1959. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1959:PEM

[Dys59c] Freeman J. Dyson. Proposal for an experiment to measure the life-time of the neutron. In Proceedings of the Conference on Scientific Applications of Nuclear Explosions, held July 6–8, 1959, Los Alamos, New Mexico, pages 55–60. Los Alamos National Laboratory, Los Alamos, NM, USA, 1959. LCCN ???? LAMS-2443.

Dyson:1960:BRBb

[Dys60a] Freeman J. Dyson. Book review: Introduction to the Theory of Quantized Fields. N. N. Bogoliubov and D. V. Shirkov. Authorized English edition, revised and enlarged by the authors. Translated from the Russian by G. M. Volkoff. Interscience, New York, 1959. xvi + 720 pp. Illus. $17. Science, 132(3419):84, July 8, 1960. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1706047;
http://www.sciencemag.org/content/132/3419/84.1.full.
pdf.

Dyson:1960:BRBc

[Dys60b] Freeman J. Dyson. Book review: Irreducible Tensorial Sets. By U. Fano and G. Racah. 171 pp. Academic Press Inc., New York, 1959. $6.80. Physics Today, 13(8):38, August 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/?PTO/13/38/1.

Dyson:1960:BRBa

[Dys60c] Freeman J. Dyson. Book review: Theory of Relativity. By W. Pauli. Translated by G. Field from 1921 German article, with 25 pp. of supplementary notes by Pauli. 241 pp. Pergamon Press, London & New York, 1958. $6.00. Physics Today, 13(6):46, June 1960. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/phtoad/
v13/i6/p46/s1.

Dyson:1960:Fa

[Dys60d] Freeman J. Dyson. Foreword. In Readable relativity [Dur60], page ?? LCCN ???? Foreword by Freeman J. Dyson. Reprint of 1926 original edition: London: G. Bell.

Dyson:1960:Fb

[Dys60e] Freeman J. Dyson. Foreword. In From immigrant to inventor [Pup60], page ?? LCCN T40.P8 A3 1960. With a foreword by Freeman J. Dyson.

Dyson:1960:FDN

[Dys60f] Freeman J. Dyson. The future development of nuclear weapons. Foreign Affairs, 38(3):457–464, April 1960. CODEN FR-

NAA3. ISSN 0015-7120. URL http://www.jstor.org/stable/

20029432. Reprinted in [Dys60f].

Dyson:1960:HIS

[Dys60g] Freeman J. Dyson. Hydrostatic instability of a star. Unpublished preprint., 1960.

Dyson:1960:IRD

[Dys60h] Freeman J. Dyson. Integral representation of a double commutator. Physical Review, 117(6):1616, March 15, 1960. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.117.1616.

Dyson:1960:LE

[Dys60i] Freeman J. Dyson. Letters to the editor. Science, 132(3421): 252–253, July 22, 1960. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.sciencemag.

org/content/132/3421/252.2.full.pdf.

Dyson:1960:MPA

[Dys60j] Freeman J. Dyson. Malthusian principles applied to extraterrestrial intelligences. Unpublished preprint., 1960.

Dyson:1960:SAS

[Dys60k] Freeman J. Dyson. Search for artificial stellar sources of infrared radiation. Science, 131(3414):1667–1668, June 3, 1960. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1705101;
http://www.sciencemag.org/content/131/3414/1667.full.

pdf. See comments [SW66].

Dyson:1960:SIA

[Dys60l] Freeman J. Dyson. Stability of an idealized atmosphere. II. Zeros of the confluent hypergeometric function. Physics of Fluids, 3 (2):155–157, 1960. CODEN PHFLE6. ISSN 1070-6631. URL

http://link.aip.org/link/?PFL/3/155/1.

Dyson:1961:FDN

[Dys61a] Freeman J. Dyson. The future development of nuclear weapons. In Garratt [Gar61], page ?? LCCN Q171 .P36. Reprint of [Dys60f].

Dyson:1961:RCN

[Dys61b] Freeman J. Dyson. Reflections and comments: The neutron bomb. Bulletin of the Atomic Scientists, 17(7):271–272, September 1961. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1962:BRBa

[Dys62a] Freeman J. Dyson. Book review: An Atlas of the Moon’s Far Side. The Lunik III Reconnaissance. N. P. Barabashov, A. A. Mikhailov, Vu. N. Lipsky, eds. Transl. from Russian by Richard B. Rodman. 147 pp, Sky Publishing Corp., Cambridge. Mass., and Interscience Publishers, Inc., New York, 1961. $7.00. Physics Today, 15(4):64, April 1962. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/?PTO/15/64/1.

Dyson:1962:BRBb

[Dys62b] Freeman J. Dyson. Book review: Dispersion Relations and the Abstract Approach to Field Theory. Lewis Klein, ed. Vol. 1 of International Science Review Series. 273 pp. Gordon & Breach, Publishers, Inc., New York, 1961. $4.95. Physics Today, 15(7): 52–54, July 1962. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
15/52/2.

Dyson:1962:BMM

[Dys62c] Freeman J. Dyson. A Brownian-motion model for the eigenvalues of a random matrix. Journal of Mathematical Physics, 3(6):1191–1198, November/December 1962. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v3/i6/p1191_
s1;   http://link.aip.org/link/?JMP/3/1191/1.

Dyson:1962:F

[Dys62d] Freeman J. Dyson. Foreword. In Readable relativity [Dur62], page ?? LCCN ???? Foreword by Freeman J. Dyson. Reprint of 1926 original edition.

Dyson:1962:OAT

[Dys62e] Freeman J. Dyson. Is our arsenal too large?: Book review: Studies of War: Nuclear and Conventional, by P. M. S. Blackett. Kill and Overkill: The Strategy of Annihilation, by Ralph E. Lapp. New York Times, ??(??):326, November 11, 1962. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/115600467/.

Dyson:1962:P

[Dys62f] Freeman J. Dyson. Pugwash 1962. Physics Today, 15(11):24– 26, November 1962. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
15/24/1.

Dyson:1962:STEa

[Dys62g] Freeman J. Dyson. Statistical theory of the energy levels of complex systems. I. Journal of Mathematical Physics, 3(1):140– 156, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v3/i1/p140_s1;   http:/
/link.aip.org/link/?JMP/3/140/1.

Dyson:1962:STEb

[Dys62h] Freeman J. Dyson. Statistical theory of the energy levels of complex systems. II. Journal of Mathematical Physics, 3(1):157– 165, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v3/i1/p157_s1;   http:/
/link.aip.org/link/?JMP/3/157/1.

Dyson:1962:STEc

[Dys62i] Freeman J. Dyson. Statistical theory of the energy levels of complex systems. III. Journal of Mathematical Physics, 3(1):166– 175, January/February 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v3/i1/p166_s1;   http:/
/link.aip.org/link/?JMP/3/166/1.

Dyson:1962:TBS

[Dys62j] Freeman J. Dyson. Thoughts on bomb shelters. Bulletin of the Atomic Scientists, 18(3):14–15, March 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Reprinted in [Dys63g]. See responses [Mor62, Ros62].

Dyson:1962:TWA

[Dys62k] Freeman J. Dyson. The threefold way. Algebraic structure of symmetry groups and ensembles in quantum mechanics. Journal of Mathematical Physics, 3(6):1199–1215, November/December 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v3/i6/p1199_s1;   http://link.aip.org/link/?
JMP/3/1199/1.

Dyson:1963:FS

[Dys63a] Freeman Dyson. Forum: On speculation. Bulletin of the Atomic Scientists, 19(7):32, September 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [BB63].

Dyson:1963:ACB

[Dys63b] Freeman J. Dyson. Appendix to chapter 4 of Zeros of Processes and Related Questions, Part 2. In Rosenblatt [Ros63], pages 82– 88. LCCN QA280.

Dyson:1963:DTO

[Dys63c] Freeman J. Dyson. Disarmament through other eyes. Bulletin of the Atomic Scientists, 19(10):37–38, December 1963. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1963:GM

[Dys63d] Freeman J. Dyson. Gravitational machines. In Interstellar communication: the search for extraterrestrial life. A collection of reprints and original contributions [Cam63], page ?? LCCN ???? Submitted as prize essay to Gravity Foundation, April 1962.

Dyson:1963:MTI

[Dys63e] Freeman J. Dyson. Models of total ignorance in quantum mechanics. In Drobot and Viebrock [DV63], pages 49–60. LCCN QA401 .C65 1962.

Dyson:1963:TBF

[Dys63f] Freeman J. Dyson. Testimony on behalf of the Federation of American Scientists before the Senate Committee on Foreign Relations, hearings on the Nuclear Test Ban Treaty, August 12–27, 1963. U.S. Government Printing Office, 1963.

Dyson:1963:TBS

[Dys63g] Freeman J. Dyson. Thoughts on bomb shelters. In Grodzins and Rabinowitch [GR63], pages 313–316. LCCN D842 .B78. Reprint of [Dys62j].

Dyson:1963:TDC

[Dys63h] Freeman J. Dyson. Toeplitz determinants and Coulomb gases. Talk to Eastern Theoretical Physics Conference, Chapel Hill, North Carolina., October 25, 1963.

Dyson:1964:IP

[Dys64a] Freeman Dyson. Innovation in physics. In Rapport and Wright [RW64], pages 256–271.

Dyson:1964:ASN

[Dys64b] Freeman J. Dyson. Approximate symmetries in nuclear and particle physics. Technical report, New York University, New York, NY, USA, 1964. Three lectures given under the Tenth Annual Lectureship in Physics at New York University, December 10–15, 1964. Notes prepared by M. E. Arons. Published as the introduction of [Dys66i].

Dyson:1964:BRB

[Dys64c] Freeman J. Dyson. Book review: PCT, Spin and Statistics, and All That. R. F. Streater and A. S. Wightman. Benjamin, New York, 1964. x + 181 pp. Illus. Paper, $4.95, cloth, $9. Science, 145(3631):475, July 31, 1964. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic).  URL http://www.jstor.org/stable/1714570;
http://www.sciencemag.org/content/145/3631/475.1.full.
pdf.

Dyson:1964:DAB

[Dys64d] Freeman J. Dyson. Defense against ballistic missiles. Bulletin of the Atomic Scientists, 20(6):12–18, June 1964. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1964:MPS

[Dys64e] Freeman J. Dyson. Mathematics in the physical sciences. Scientific American, 211(3):128–146, September 1964. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v211/n3/pdf/scientificamerican0964-128.pdf.

Dyson:1964:PDA

[Dys64f] Freeman J. Dyson. Problems of defense against ballistic missiles. F.A.S. Newsletter, ??(??):??, January 1964.

Dyson:1964:RTS

[Dys64g] Freeman J. Dyson. Review of Tom Stonier’s Nuclear Disaster. Disarmament and Arms Control, 2(4):459–461, ???? 1964.

Dyson:1965:DP

[Dys65a] Freeman J. Dyson. Death of a project. Science, 149(3680):141– 144, July 9, 1965. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.daviddarling.info/
encyclopedia/O/OrionProj.html;   http://www.jstor.org/
stable/1716283;   http://www.sciencemag.org/content/149/
3680/141.full.pdf.

Dyson:1965:DR

[Dys65b] Freeman J. Dyson. Dyson replies. Bulletin of the Atomic Scientists, 21(3):40, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys64d, McM65].

Dyson:1965:ONF

[Dys65c] Freeman J. Dyson. Old and new fashions in field theory. Physics Today, 18(6):21–24, June 1965. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/?PTO/18/21/1.

Dyson:1965:RBS

[Dys65d] Freeman J. Dyson. Review of Quasi-Stellar Sources and Gravitational Collapse [Robinson, Schild and Schucking, editors]. Publications of the Astronomical Society of the Pacific, 77(458):398, November 1965. CODEN PASPAU. ISSN 0004-6280 (print),

1538-3873 (electronic). URL http://www.jstor.org/stable/
40674240.

Dyson:1965:TSF

[Dys65e] Freeman J. Dyson. Tomonaga, Schwinger, and Feynman awarded Nobel Prize for Physics. Science, 150(3696):588–589, October 29, 1965. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1717147;
 http://www.sciencemag.org/content/150/3696/588.full.
pdf.

Dyson:1966:AGT

[Dys66a] Freeman J. Dyson. Applications of group theory in particle physics. SIAM Review, 8(1):1–10, ???? 1966. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic). URL http://

www.jstor.org/stable/2028169. Von Neumann Lecture given

at Cornell University, Ithaca, on September 2, 1965, for the Society of Industrial and Applied Mathematics.

Dyson:1966:BGB

[Dys66b] Freeman J. Dyson. Bridging the gap between physics and mathematics. Essay for the Committee on Support of Research in the Mathematical Sciences, 1966., 1966.

Dyson:1966:DDL

[Dys66c] Freeman J. Dyson. The dynamics of a disordered linear chain. In Lieb and Mattis [LM66], pages 132–139. LCCN QC174.5 .L47. Reprint of [Dys53d].

Dyson:1966:LMS

[Dys66d] Freeman J. Dyson. Lorentz Medal speech. Given to the Netherlands Academy., December 17, 1966.

Dyson:1966:OCR

[Dys66e] Freeman J. Dyson. Opening and closing remarks. CNRS Conference, Gif-sur-Yvette, April 1–5, 1966, 1966.

Dyson:1966:RSM

[Dys66f] Freeman J. Dyson. Remarks on the stability of matter. Talk delivered at the sixtieth birthday celebration of Hans Bethe at Cornell University, Ithaca, October 1966., October 1966.

Dyson:1966:SET

[Dys66g] Freeman J. Dyson. The search for extraterrestrial technology. In Marshak and Blaker [MB66], pages 641–655. LCCN QC774.B4 M3.

Dyson:1966:SM

[Dys66h] Freeman J. Dyson. Stability of matter. Talk given to Eastern Theoretical Physics Conference at Brown University., November 1966.

Dyson:1966:SGN

[Dys66i] Freeman J. Dyson, editor. Symmetry groups in nuclear and particle physics: a lecture-note and reprint volume. The Mathematical physics monograph series. W. A. Benjamin, Inc., New York, NY, USA, 1966. xii + 320 pp. LCCN QC721 .D96.

Dyson:1967:BSS

[Dys67a] Freeman J. Dyson. Bengt and Sigrid Strömgren. Remarks for farewell lunch, April 11, 1967.

Dyson:1967:BRB

[Dys67b] Freeman J. Dyson. Book review: Mathematical Physics in One Dimension: Exactly Soluble Models of Interacting Particles. (Reprint collection) Elliot H. Lieb, Daniel C. Mattis, eds. 565 pp. Academic Press, New York 1966. $11.50. Physics Today, 20(9):81–82, September 1967. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.

aip.org/link/?PTO/20/81/1.

Dyson:1967:GSE

[Dys67c] Freeman J. Dyson. Ground-state energy of a finite system of charged particles. Journal of Mathematical Physics, 8(8):

1538–1545, August 1967. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v8/i8/p1538_s1;   http:
//link.aip.org/link/?JMP/8/1538/1.

Dyson:1967:TVC

[Dys67d] Freeman J. Dyson. Time variation of the charge of the proton. Physical Review Letters, 19(22):1291–1293, November 27, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (elec-

tronic), 1092-0145. URL http://link.aps.org/doi/10.1103/
PhysRevLett.19.1291.

Dyson:1968:IT

[Dys68a] Freeman J. Dyson. Interstellar transport. Physics Today, 21 (10):41–45, October 1968. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.org/

link/?PTO/21/41/1.

Dyson:1968:NCR

[Dys68b] Freeman J. Dyson. Numerological considerations relevant to quasi-stellar objects. Astrophysical Journal, 154(??):37–39, ???? 1968. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Dyson:1968:RA

[Dys68c] Freeman J. Dyson. Rhenium 187 and all that. Unpublished., 1968.

Dyson:1968:SP

[Dys68d] Freeman J. Dyson. Space and the physicist. New Physics (Korean Physical Society), 7(2):5–8, June 1968.

Dyson:1969:CCC

[Dys69a] Freeman J. Dyson. Current comments: A case for missile defense. Bulletin of the Atomic Scientists, 25(4):31–33, April 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ten69, Tay69, You69].

Dyson:1969:DPI

[Dys69b] Freeman J. Dyson. Discussion paper: Interstellar transport. Annals of the New York Academy of Sciences, 163(1):347–357, September 1969. CODEN ANYAA9. ISSN 0077-8923 (print), 1749-6632 (electronic).

Dyson:1969:DSM

[Dys69c] Freeman J. Dyson. Draft statement on MIRV for the council for a livable world. Unpublished., 1969.

Dyson:1969:DSG

[Dys69d] Freeman J. Dyson. Dynamics of a spinning gas cloud. Journal of Mathematics and Mechanics, 18(??):91–102, ???? 1969. CODEN JOMMAN. ISSN 0095-9057 (print), 1943-5274 (electronic).

Dyson:1969:EER

[Dys69e] Freeman J. Dyson. The efficiency of energy release in gravitational collapse. Comments on Astrophysics and Space Physics, 1(??): 75–80, ???? 1969.

Dyson:1969:EPT

[Dys69f] Freeman J. Dyson. Existence of a phase-transition in a onedimensional Ising ferromagnet. Communications in Mathematical Physics, 12(2):91–107, 1969. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic). URL http://projecteuclid.

org/euclid.cmp/1103841344.

Dyson:1969:HCEa

[Dys69g] Freeman J. Dyson. Human consequences of the exploration of space. Bulletin of the Atomic Scientists, 25(7):8–10, 12–13, September 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Carlson Memorial Lecture, April 1968. Reprinted in [Dys69h].

Dyson:1969:HCEb

[Dys69h] Freeman J. Dyson. Human consequences of the exploration of space. In Rabinowitch and Lewis [RL69], page ?? Reprint of [Dys69g].

Dyson:1969:MAM

[Dys69i] Freeman J. Dyson. Missiles and anti-missiles … six views: Comments on Sternglass thesis. Bulletin of the Atomic Scientists, 25 (6):27, June 1969. CODEN BASIAP. ISSN 0096-3402 (print),

1938-3282 (electronic). URL http://en.wikipedia.org/wiki/
Ernest_J._Sternglass.

Dyson:1969:NSP

[Dys69j] Freeman J. Dyson. A new symmetry of partitions. Journal of Combinatorial Theory, 7:56–61, 1969. CODEN JCTHAR. ISSN 0021-9800 (print), 1878-1756 (electronic).

Dyson:1969:NES

[Dys69k] Freeman J. Dyson. Non-existence of spontaneous magnetization in a one-dimensional Ising ferromagnet. Communications in Mathematical Physics, 12(3):212–215, ???? 1969. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic). URL

http://projecteuclid.org/euclid.cmp/1103841388.

Dyson:1969:PTF

[Dys69l] Freeman J. Dyson. Phase transitions in ferromagnets. Talk given at Evanston Statistical Mechanics Conference in honor of G. Uhlenbeck. Published in [Dys71l]., October 31, 1969.

Dyson:1969:SRE

[Dys69m] Freeman J. Dyson. Seismic response of the Earth to a gravitational wave in the one-Hertz band. Astrophysical Journal, 156 (??):529–540, ???? 1969. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic).

Dyson:1969:SBM

[Dys69n] Freeman J. Dyson. Statement on ballistic missile defense. Submitted to the U.S. Senate Committee on Armed Services., May 1969.

Dyson:1969:VTP

[Dys69o] Freeman J. Dyson. Volcano theory of pulsars. Nature, 223(5205): 486–487, August 2, 1969. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v223/n5205/pdf/223486a0.pdf.

Dyson:1969:VSN

[Dys69p] Freeman J. Dyson. Vulcanism and seismicity in neutron stars. Comments on Astrophysics and Space Physics, 1(??):198–206, ???? 1969.

Dyson:1970:BRL

[Dys70a] Freeman Dyson. Book review: Life of a mathematician: Hilbert. Constance Reid. With an appreciation of Hilbert’s mathematical work by Hermann Weyl. Springer-Verlag, New York, 1970. xii, 292 pp., illus. $8.80. Science, 170(3961):965–966, November 27, 1970. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1731559;
http://www.sciencemag.org/content/170/3961/965.2.full.
pdf.

Dyson:1970:ANW

[Dys70b] Freeman J. Dyson. Avoiding nuclear war. Talk given at Winds of Change Seminar, Michigan State University., April 16, 1970.

Dyson:1970:BRBc

[Dys70c] Freeman J. Dyson. Book review: Annual Review of Astronomy And Astrophysics, Leo Goldberg, David Layzer, John G. Phillips, eds. 528 pp. Palo Alto, Calif., 1968. $8.50. Physics Today, 23(4): 68–69, April 1970. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
23/68/1.

Dyson:1970:BRBa

[Dys70d] Freeman J. Dyson. Book review: Hermann Weyl Gesammelte Abhandlugen, Band I, II, III, IV, K. Chandrasekharan, ed. Physics Today, 23(5):73, May 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/?PTO/23/73/1.

Dyson:1970:CBE

[Dys70e] Freeman J. Dyson. Correlations between eigenvalues of a random matrix. Communications in Mathematical Physics, 19(3):235– 250, 1970. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-

0916 (electronic). URL http://projecteuclid.org/euclid.
cmp/1103842703.

Dyson:1970:FP

[Dys70f] Freeman J. Dyson. The future of physics. Physics Today, 23 (9):23–28, September 1970. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip. org/link/?PTO/23/23/1. Lecture given at the dedication of Jadwin and Fine Halls, Princeton University, March 17, 1970. Translations in Czech [Dys72g], Japanese [Dys72b], Russian [Dai71], and Slovene [Dys72h].

Dyson:1970:MWM

[Dys70g] Freeman J. Dyson. Menschheit und Weltall: Max-Planck-Vortrag von der Physikertagung in Salzburg. (German) [Mankind and the Universe: Max Planck presentation of the physicists meeting in Salzburg]. Physikalische Blätter, 26(1):7–14, January 1970. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19700260104/abstract.

Dyson:1970:RS

[Dys70h] Freeman J. Dyson. Reflections: The sell-out. The New Yorker, 46(??):44–??, February 21, 1970. ISSN 0028-792X.

URL http://www.newyorker.com/archive/1970/02/21/1970_
02_21_044_TNY_CARDS_000298991.

Dyson:1970:TFC

[Dys70i] Freeman J. Dyson. The Twenty-First Century. Vanuxem Lecture, Princeton University., February 26, 1970.

Dyson:1971:EUb

[Dys71a] F. J. Dyson. Energy in the Universe. In Energy and power [Ano71], page ?? ISBN 0-7167-0938-4 (paperback), 0-7167-0939-2. LCCN TJ153 .E478. Reprinted in [Dys75a].

Dyson:1971:ACT

[Dys71b] Freeman J. Dyson. Arms control and technological change. Talk prepared for the Committee on International Relations, University of Chicago. Printed in [Dys73a]., April 1971.

Dyson:1971:CBC

[Dys71c] Freeman J. Dyson. Chemical binding in classical Coulomb lattices. Annals of Physics, 63(1):1–11, March 1971. CO-DEN APNYA6. ISSN 0003-4916 (print), 1096-035X (electronic).

URL http://www.sciencedirect.com/science/article/pii/

0003491671902946. Reprinted in [Dys72c].

Dyson:1971:DEC

[Dys71d] Freeman J. Dyson. Distribution of eigenvalues for a class of real symmetric matrices. Revista Mexicana de Fisica, 20(??):231–237, ???? 1971.

Dyson:1971:EUa

[Dys71e] Freeman J. Dyson. Energy in the Universe. Scientific American, 225(3):50–59, September 1971. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic). URL http://www.nature.

com/scientificamerican/journal/v225/n3/pdf/scientificamerican0971-
50.pdf.

Dyson:1971:ENP

[Dys71f] Freeman J. Dyson. Existence and nature of phase transitions in one- dimensional Ising ferromagnets. Talk to American Mathematical Society., April 7, 1971.

Dyson:1971:IFD

[Dys71g] Freeman J. Dyson. An Ising ferromagnet with discontinuous long-range order. Communications in Mathematical Physics, 21(4):269–283, 1971. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic). URL http://projecteuclid.

org/euclid.cmp/1103857374.

Dyson:1971:LA

[Dys71h] Freeman J. Dyson. Letter from Armenia. The New Yorker, 47 (??):126–??, November 6, 1971. ISSN 0028-792X. URL http:/

/www.newyorker.com/archive/1971/11/06/1971_11_06_126_
TNY_CARDS_000302733.

Dyson:1971:LEH

[Dys71i] Freeman J. Dyson. Letter to the Editor: For high-energy physicists. Physics Today, 24(4):11, April 1971. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/?PTO/24/11/1.

Dyson:1971:NSP

[Dys71j] Freeman J. Dyson. Neutron stars and pulsars: Fermi lectures 1970 at the Scuola normale superiore di Pisa. Accademia nazionale dei Lincei. Problemi attuali di scienza e di cultura, Quaderno 152 (??):89–90, 1971.

Dyson:1971:NIR

[Dys71k] Freeman J. Dyson. The next Industrial Revolution. Tinlot Memorial Lecture, Rochester University. Printed in [Dys77d]., February 8, 1971.

Dyson:1971:PTF

[Dys71l] Freeman J. Dyson. Phase transitions in ferromagnets. In Cohen [Coh71], page ?? ISBN 0-8247-1111-4. LCCN QC175 .S7.

Dyson:1972:FCT

[Dys72a] F. J. Dyson. The fundamental constants and their time variation. In Salam and Wigner [SW72], pages 213–236. ISBN 0-521-08600-0. LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.

uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http://
www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Dyson:1971:FP

[Dys72b] Freeman J. Dyson. The future of physics. Shizen, ??(??):35–??, ???? 1971–1972. Japanese translation of [Dys70f].

Dyson:1972:CBC

[Dys72c] Freeman J. Dyson. Chemical binding in classical Coulomb lattices. In Morse et al. [MHFW72], pages 1–11. ISBN 0-12-508201-0 (vol. 1), 0-323-14303-2 (e-book). LCCN QC173 .N8837 1972.

URL http://www.sciencedirect.com/science/article/pii/
B9780125082013500089. Reprint of [Dys71c].

Dyson:1972:CME

[Dys72d] Freeman J. Dyson. A class of matrix ensembles. Journal of Mathematical Physics, 13(1):90–97, January 1972. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v13/i1/
p90_s1;   http://link.aip.org/link/?JMP/13/90/1.

Dyson:1972:ENP

[Dys72e] Freeman J. Dyson. Existence and nature of phase transitions in one- dimensional Ising ferromagnets. In Pool [Poo72], pages 1–12. ISBN 0-8218-1324-2. LCCN QC174.7 .S92 1971.

Dyson:1972:EBP

[Dys72f] Freeman J. Dyson. Experiments with bomb-propelled spaceship models. Adventures in Experimental Physics, B(??):321–??, ????

Dyson:1972:FPa

[Dys72g] Freeman J. Dyson. The future of physics. Casopis pro fysiku, A22(??):74–??, ???? 1972. Czech translation of [Dys70f].

Dyson:1972:FPb

[Dys72h] Freeman J. Dyson. The future of physics. [unknown], ??(??):??, ???? 1972. Slovene translation of [Dys70f].

Dyson:1972:ILU

[Dys72i] Freeman J. Dyson. Intelligent life in the Universe. Lecture given at San Francisco, September 18, 1972, under sponsorship of the Astronomical Society of the Pacific, NASA, and the City College of San Francisco., 1972.

Dyson:1972:MO

[Dys72j] Freeman J. Dyson. Missed opportunities. Bulletin of the American Mathematical Society, 78(5):635–652, 1972. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic). URL

http://projecteuclid.org/euclid.bams/1183533964. Josiah

Willard Gibbs Lecture, given under the auspices of the American Mathematical Society, January 17, 1972; received by the editors January 17, 1972.

Dyson:1972:QD

[Dys72k] Freeman J. Dyson. Quaternion determinants. Helvetica Physica Acta, 45(??):289–302, ???? 1972. CODEN HPACAK. ISSN 0018-0238. See erratum [Dys73b].

Dyson:1972:RBM

[Dys72l] Freeman J. Dyson. Review of Ballistic Missile Defense, by Benson D. Adams. Bulletin of the Atomic Scientists, 28(5):41, May 1972. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Dyson:1972:S

[Dys72m] Freeman J. Dyson. The sell-out. In The techniques of reading: an integrated program for improved comprehension and speed [Jud72], pages 447–456. ISBN 0-15-589690-3. LCCN LB1050.5 .J8 1972. in consultation with William S. Schaill.

Dyson:1972:WFD

[Dys72n] Freeman J. Dyson. The world, the flesh and the Devil: the third J. D. Bernal lecture delivered at Birkbeck College, London, 16th May 1972, volume 3 of The J. D. Bernal lecture. Birkbeck College, London, UK, 1972. ISBN 0-900975-55-5. 15 pp. LCCN ???? Reprinted in [Dys73d].

Dyson:1973:ACT

[Dys73a] Freeman J. Dyson. Arms control and technological change. In Kaplan, comp [Kap73], page ?? LCCN JX1974.7 .K24; JZ5665 .K37

1973x. URL http://catalog.hathitrust.org/api/volumes/
oclc/623340.html.

Dyson:1973:EQD

[Dys73b] Freeman J. Dyson. Erratum: Quaternion determinants. Helvetica Physica Acta, 46(??):274, ???? 1973. CODEN HPACAK. ISSN 0018-0238. See [Dys72k].

Dyson:1973:STN

[Dys73c] Freeman J. Dyson. Science and technology in the next fifty years. Talk at Naval Research Laboratory 50th Anniversary Symposium, Washington, DC., October 3, 1973.

Dyson:1973:WFD

[Dys73d] Freeman J. Dyson. The world, the flesh and the Devil. In Sagan [Sag73], pages 371–389. ISBN 0-262-19106-7. LCCN QB54 .S68 1971a.

Dyson:1974:BRH

[Dys74a] Freeman Dyson. Book review: Honoring Dirac: The Physicist’s Conception of Nature, by Jagdish Mehra. Proceedings of a symposium, Trieste, Italy, Sept. 1972. Science, 185(4157):1160–1161, September 27, 1974. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1738373.

Dyson:1974:HCS

[Dys74b] Freeman J. Dyson. The hidden costs of saying no. Talk to the International Meeting on Scientific Research and Energy Problems, Madrid. Reprinted in [Dys75b, Dys77c], October 16, 1974.

Dyson:1974:OSB

[Dys74c] Freeman J. Dyson. Our stability is but balance. Pauli Memorial Lecture, Zurich, February 18, 1974.

Dyson:1975:EU

[Dys75a] Freeman J. Dyson. Energy in the Universe. In New frontiers in astronomy [Gin75], page ?? ISBN 0-7167-0519-2, 0-7167-0520-6 (paperback). LCCN QB51 .F74 1975.

Dyson:1975:HCS

[Dys75b] Freeman J. Dyson. The hidden cost of saying no!: Legislators make a mistake in acting as if the future were predictable. Bulletin of the Atomic Scientists, 31(6):23–27, June 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See comments [Ehr75, Alb75, Mav75].

Dyson:1975:PNA

[Dys75c] Freeman J. Dyson. Photon noise and atmospheric noise in active optical systems. Journal of the Optical Society of America, 65(5): 551–558, May 1975. CODEN JOSAAH. ISSN 0030-3941.

Dyson:1976:HBD

[Dys76a] F. J. Dyson. The hydrogen-bomb decision: a reappraisal: Review of The Advisors: Oppenheimer, Teller and The Superbomb, by Herbert F. York. Science, 193(4254):668–669, August 20, 1976. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/193/
4254/668.full.pdf.

Dyson:1976:BPG

[Dys76b] Freeman J. Dyson. Breakdown of physics in gravitational collapse. Unpublished., 1976.

Dyson:1976:CBR

[Dys76c] Freeman J. Dyson. Costs and benefits of recombinant DNA research. Science, 193(4247):6, July 2, 1976. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic).  URL http://www.jstor.org/stable/1742296;
http://www.sciencemag.org/content/193/4247/6.1.full.pdf.

Dyson:1976:FDI

[Dys76d] Freeman J. Dyson. Fredholm determinants and inverse scattering problems. Communications in Mathematical Physics, 47(2):171– 183, 1976. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-

0916 (electronic). URL http://projecteuclid.org/euclid.
cmp/1103899727.

Dyson:1976:ONA

[Dys76e] Freeman J. Dyson. Old and new approaches to the inverse scattering problem. In Lieb et al. [LSW76], pages 151–167. LCCN ????

Dyson:1977:ASH

[Dys77a] Freeman J. Dyson. Acceptance speech for Harvey Prize, Haifa. Unpublished., June 1977.

Dyson:1977:CWC

[Dys77b] Freeman J. Dyson. Can we control the amount of carbon dioxide in the atmosphere? Energy, 2(??):287–291, ???? 1977.

Dyson:1977:HCS

[Dys77c] Freeman J. Dyson. The hidden costs of saying no. In Leonard et al. [LDB77], pages 256–264. ISBN 0-89164-046-0. LCCN HD60 B979e.

Dyson:1977:NIR

[Dys77d] Freeman J. Dyson. The next Industrial Revolution. The Key Reporter, 42(3):??, Spring 1977.

Dyson:1977:RPC

[Dys77e] Freeman J. Dyson. Report of the Princeton Community Biohazards Committee — to Princeton Borough Council and Princeton Township Committee, 1977.

Dyson:1977:SKD

[Dys77f] Freeman J. Dyson. Stability of Kalnajs disks. Unpublished., 1977.

Dyson:1977:TFJ

[Dys77g] Freeman J. Dyson. Testimony by Freeman J. Dyson [at] Hearings on the science policy implications of the DNA recombinant molecule research issue, May 5, 1977 [held by] Subcommittee on Science, research and Technology of the Committee on Science and Technology of the U.S. House of Representatives. United States. Congress. House. Committee on Science and Technology. Subcommittee on Science, Research and Technology, 1977.

Dyson:1978:VCa

[Dys78a] F. J. Dyson. Variation of constants. In Lannutti and Williams [LW78], pages 163–167. ISBN 0-88318-147-9. LCCN QC793.3.F5 C87.

Dyson:1978:RCI

[Dys78b] Freeman Dyson. Review: Characterizing irregularity: Fractals. Form, Chance, and Dimension, by Benoît B. Mandelbrot. Science, 200(4342):677–678, May 12, 1978. CO-DEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic).  URL http://www.jstor.org/stable/1746976;
http://www.sciencemag.org/content/200/4342/677.full.pdf.

Dyson:1978:IPL

[Dys78c] Freeman J. Dyson. Image processing and live optics. In Richter et al. [RWP78], pages 439–444. LCCN QB88 .O685

1977. URL http://catalog.hathitrust.org/api/volumes/
oclc/4374966.html.

Dyson:1978:PF

[Dys78d] Freeman J. Dyson. Physics in the future. In Rabi [Rab78], pages 106–112. LCCN ????

Dyson:1978:PFM

[Dys78e] Freeman J. Dyson. Pilgrim fathers, Mormon pioneers, and space colonists: An economic comparison. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 122(2):63–68, ???? 1978. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www. jstor.org/stable/986622. Lecture to American Philosophical Society, November 10, 1977.

Dyson:1978:VCb

[Dys78f] Freeman J. Dyson. Variation of constants. AIP Conference Proceedings, 48(1):163–168, 1978. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. URL http://

link.aip.org/link/?APC/48/163/1.

Dyson:1979:BUBb

[Dys79a] Freeman Dyson. Disturbing the Universe + Defensive and Offensive Weapons — World of the Scientist. 3. The New Yorker, 55 (27):36, ???? 1979. ISSN 0028-792X.

Dyson:1979:BUW

[Dys79b] Freeman Dyson. Disturbing the Universe. 1. World of the Scientist. The New Yorker, 55(25):37, ???? 1979. ISSN 0028-792X.

Dyson:1979:BUBa

[Dys79c] Freeman Dyson. Disturbing the Universe. Oppenheimer, His Nature and His Physics. 2. World of the Scientist. The New Yorker, 55(26):64, ???? 1979. ISSN 0028-792X.

Dyson:1979:BRB

[Dys79d] Freeman J. Dyson. Book review: Toward Distant Suns by T. A. Heppenheimer. American Scientist, 67(6):710, November 1979. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (elec-

tronic). URL http://www.jstor.org/stable/27849542.

Dyson:1979:DU

[Dys79e] Freeman J. Dyson. Disturbing the Universe. Harper & Row, New York, NY, USA, 1979. ISBN 0-06-011108-9. x + 283 pp. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1979:CE

[Dys79f] Freeman J. Dyson. In the conflux of eternities. Nature, 282(5735): 177, November 8, 1979. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://www.nature.com/

nature/journal/v282/n5735/pdf/282177a0.pdf.

Dyson:1979:RMU

[Dys79g] Freeman J. Dyson. Is real mathematics of any use to physics? Talk to Australian Academy of Science Silver Jubilee Symposium, Canberra., March 1979.

Dyson:1979:PFM

[Dys79h] Freeman J. Dyson. Prelude in E-flat minor. In Disturbing the Universe [Dys79e], pages 84–93. ISBN 0-06-011108-9. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1979:RDUa

[Dys79i] Freeman J. Dyson. Reflections: Disturbing the Universe: I. The New Yorker, 55(??):37–??, August 6, 1979. ISSN 0028-792X.

URL http://www.newyorker.com/archive/1979/08/06/1979_
08_06_037_TNY_CARDS_000326372.

Dyson:1979:RDUb

[Dys79j] Freeman J. Dyson. Reflections: Disturbing the Universe: II. The New Yorker, 55(??):64–??, August 6, 1979. ISSN 0028-792X.

URL http://www.newyorker.com/archive/1979/08/13/1979_
08_13_064_TNY_CARDS_000330871.

Dyson:1979:RDUc

[Dys79k] Freeman J. Dyson. Reflections: Disturbing the Universe: III. The New Yorker, 55(??):36–??, August 20, 1979. ISSN 0028-792X.

URL http://www.newyorker.com/archive/1979/08/20/1979_
08_20_036_TNY_CARDS_000332666.

Dyson:1979:RA

[Dys79l] Freeman J. Dyson. A ride to Albuquerque. In Disturbing the Universe [Dys79e], pages 58–69. ISBN 0-06-011108-9. LCCN QC16.D95 A33 1979. US$12.95. Commissioned by the Science Book Program of the Alfred P. Sloan Foundation. Parts of the book were published in The New Yorker, August 6, 13, and 20, 1979, and in The Observer, October 28, 1979.

Dyson:1979:TEP

[Dys79m] Freeman J. Dyson. Time without end: Physics and biology in an open universe. Reviews of Modern Physics, 51 (3):447–460, July 1, 1979. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.51.447;   http:
//rmp.aps.org/abstract/RMP/v51/i3/p447_1. James Arthur

Lectures on Time and its Mysteries, New York University, 1978 Series.

Dyson:1980:ROL

[Dys80a] F. J. Dyson. Robert Oppenheimer — letters and recollections. New Republic, 182(21):37–39, ???? 1980. ISSN 0028-6583.

Dyson:1980:BRBc

[Dys80b] Freeman J. Dyson. Book review: Brittle Silence, By A. Tsirul’nikov, Candidate in Pedagogical Sciences Senior Scientific Associate, NII of General and Pedagogical Psychology, APN SSSR. Literaturanaya Gazeta, ??(??):??, December 3, 1980.

Dyson:1980:BRBa

[Dys80c] Freeman J. Dyson. Book review: Doomsday has Been Cancelled, by Peter Vajk, Peace Press, 1978. 238 pp. Journal of Social and Biological Structures, 3(1):83, January 1980. ISSN 0140-1750 (print), 1878-2787 (electronic). URL http://www.

sciencedirect.com/science/article/pii/014017508090024X.

Dyson:1980:CTB

[Dys80d] Freeman J. Dyson. Comment on the topic “beyond the black hole” [by John Wheeler]. In Woolf [Woo80], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Dyson:1980:PF

[Dys80e] Freeman J. Dyson. Einstein and the physics of the future. In Woolf [Woo80], page ?? ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Dyson:1980:Fa

[Dys80f] Freeman J. Dyson. Foreword. In Josephson and Ramachandran [JR80], page ?? ISBN 0-08-024695-8, 0-08-024694-X (paperback).

LCCN BF311 .C645 1980. URL http://catalog.hathitrust.
org/api/volumes/oclc/5029030.html. With a foreword by F.

J. Dyson.

Dyson:1980:Fb

[Dys80g] Freeman J. Dyson. Foreword. In Josephson and Ramachandran [JR80], page ?? ISBN 0-08-024695-8, 0-08-024694-X (paperback).

LCCN BF311 .C645 1980. URL http://catalog.hathitrust.
org/api/volumes/oclc/5029030.html. With a foreword by F.

J. Dyson.

Dyson:1980:FDF

[Dys80h] Freeman J. Dyson. The future of desire and the future of fate. Hilldale Lecture, given at the University of Wisconsin, Madison., April 15, 1980.

Dyson:1980:RMU

[Dys80i] Freeman J. Dyson. Is real mathematics of any use to physics? In White [Whi80], pages 1–8. ISBN 0-85847-058-6 (v. 1), 0-85847-060-8 (v. 2), 0-85847-064-0 (v. 3). LCCN QC6.9 .C48 1980.

Dyson:1980:LU

[Dys80j] Freeman J. Dyson. Life in the Universe. Third Albert Einstein Lecture at Rockefeller University., March 5, 1980.

Dyson:1980:MDA

[Dys80k] Freeman J. Dyson. Ma découverte de l’Amérique. (French) [My discovery of America]. La Recherche, 11(107):91–98, January 1980. CODEN RCCHBV. ISSN 0029-5671 (print), 1625-9955 (electronic). Translation into French of extracts from [Dys79e].

Dyson:1980:MA

[Dys80l] Freeman J. Dyson. Manchester and Athens. Talk given to Nobel Conference at Gustavus Adolphus College, St. Peter, Minnesota., October 8, 1980.

Dyson:1980:QC

[Dys80m] Freeman J. Dyson. The quest for concept. Talk to the conference on the Arms Race under the auspices of the Princeton Episcopalian and Presbyterian Churches, September 27–28, 1980.

Dyson:1980:QB

[Dys80n] Freeman J. Dyson. Quick is beautiful. Talk to Monsanto Symposium, given at St. Louis, MO., October 23, 1980.

Dyson:1980:RHS

[Dys80o] Freeman J. Dyson. Reaching for a home in the stars: Towards the frontiers of thought. The Christian Science Monitor, ??(??): ??, May 27, 1980. ISSN 0882-7729 (print), 1540-4617 (electronic).

URL http://www.csmonitor.com/1980/0527/052705.html.

Dyson:1980:TCM

[Dys80p] Freeman J. Dyson. Tragedy and comedy in modern dress. Keynote talk for the 29th Annual Conference of the Mental Health Association in New Jersey given at Morristown, NJ., June 3, 1980.

Dyson:1981:BRBb

[Dys81a] Freeman J. Dyson. Book review: Brittle Silence, By A. Tsirul’nikov, Candidate in Pedagogical Sciences Senior Scientific Associate, NII of General and Pedagogical Psychology, APN SSSR. Outlook, ??(??):??, Summer 1981.

Dyson:1981:BRBa

[Dys81b] Freeman J. Dyson. Book review: John von Neumann and Norbert Wiener: From Mathematics to the Technologies of Life and Death, by Steve J. Heims, MIT Press. Technology Review (M.I.T.), 83(4):16–??, February/March 1981. CODEN TEREAU. ISSN 0040-1692.

Dyson:1981:DAT

[Dys81c] Freeman J. Dyson. The day after Trinity. J. Robert Oppenheimer and the Atomic Bomb. PBS Film., April 29, 1981.

Dyson:1981:DU

[Dys81d] Freeman J. Dyson. Disturbing the Universe. Pan Books, London, UK, 1981. ISBN 0-330-26324-2 (paperback). vii + 280 pp. LCCN QC16.D95.

Dyson:1981:HH

[Dys81e] Freeman J. Dyson. History without hindsight. Technology Review (M.I.T.), 83(2):16–??, February/March 1981. CODEN TEREAU. ISSN 0040-1692. URL http://www.

technologyreview.com/magazine/1981/02/pdf/.

Dyson:1981:IAD

[Dys81f] Freeman J. Dyson. Infinite in all directions. Phi Beta Kappa Public Lecture, given at the American Association for the Advancement of Science Annual Meeting in Toronto., January 6, 1981.

Dyson:1981:IEZ

[Dys81g] Freeman J. Dyson. Innenansichten: Erinnerungen in die Zukunft. (German) [Interior views: Memories in the future]. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1981. ISBN 3-7643-1200-9. 286 pp. LCCN ???? German translation by Jeannette Zehnder.

Dyson:1981:I

[Dys81h] Freeman J. Dyson. Introduction. In Azbel and Forbes [AF81], page ?? ISBN 0-395-30226-9. LCCN DS135.R95

A937. URL http://catalog.hathitrust.org/api/volumes/
oclc/6735540.html.

Dyson:1981:LEO

[Dys81i] Freeman J. Dyson. Letter to the Editor: The origin of genetic information. Scientific American, 244(6):8, June 1981. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dyson:1981:LU

[Dys81j] Freeman J. Dyson. Life in the Universe. Darwin Lecture, given at Darwin College, Cambridge, England., November 10, 1981.

Dyson:1981:PCC

[Dys81k] Freeman J. Dyson. Perspective: The children’s crusade. Bulletin of the Atomic Scientists, 37(10):1, 4, December 1981. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). Excerpt from the author’s book, Disturbing the Universe.

Dyson:1981:RS

[Dys81l] Freeman J. Dyson. Return to space. PBS Film, PTV Publications, Kent, Ohio., October 5, 1981.

Dyson:1981:SSS

[Dys81m] Freeman J. Dyson. Science for science’s sake: Public support of astronomy. Lecture to ‘Saturday at the University’, given at the University of Pennsylvania, Philadelphia., February 7, 1981.

Dyson:1981:UP

[Dys81n] Freeman J. Dyson. Unfashionable pursuits. Lecture given on at Yale University in honor of Feza Gursey’s sixtieth birthday, April 11, 1981.

Dyson:1981:W

[Dys81o] Freeman J. Dyson. Winner. New York Review of Books, 28(7): ??, April 30, 1981. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
1981/apr/30/winner/.

Dyson:1982:BRB

[Dys82a] Freeman J. Dyson. Book review: The Nuclear Delusion: Soviet– American Relations in the Atomic Age, by George F. Kennan. The Christian Science Monitor, ??(??):??, October 5, 1982. ISSN 0882-7729 (print), 1540-4617 (electronic).

Dyson:1982:EUP

[Dys82b] Freeman J. Dyson. The encouragement of unfashionable pursuits. Lecture sponsored by the Program in the History of Science and Technology, University of Minnesota, given at Minneapolis, April 14, 1982.

Dyson:1982:FFT

[Dys82c] Freeman J. Dyson. Fighting for freedom with the technologies of death. Compton Memorial Lecture, given at the Massachusetts Institute of Technology., February 22, 1982.

Dyson:1982:FDF

[Dys82d] Freeman J. Dyson. The future of desire and the future of fate. Priroda, 8(??):60–70, ???? 1982. Russian translation by N. V. Gorskoi.

Dyson:1982:HD

[Dys82e] Freeman J. Dyson. Helen Dukas. Remarks at the Dukas Memorial Ceremony, Institute for Advanced Study, Princeton, NJ., March 15, 1982.

Dyson:1982:IPS

[Dys82f] Freeman J. Dyson. Interstellar propulsion systems. In Hart and Zuckerman [HZ82], pages 41–45. ISBN 0-08-026342-9, 0-08-026341-0 (paperback). LCCN QB54 .E95 1982.

Dyson:1982:MOL

[Dys82g] Freeman J. Dyson. A model for the origin of life. Journal of Molecular Evolution, 18(5):344–350, ???? 1982. CODEN JMEVAU. ISSN 0022-2844 (print), 1432-1432 (electronic).

Dyson:1982:TOL

[Dys82h] Freeman J. Dyson. Theology and the origin of life. Notes for a talk to Graduate Theological Union, Berkeley, CA., November 11, 1982.

Dyson:1982:UP

[Dys82i] Freeman J. Dyson. Unfashionable pursuits. In Hanle [Han82], pages 29–40. LCCN ???? Published in cooperation with the Institute for Advanced Study, Princeton, NJ.

Dyson:1983:BP

[Dys83a] Freeman J. Dyson. Bombs and poetry. In Peterson et al. [PM^{+}83], pages 81–145. ISBN 0-585-19772-5 (e-book), 0-87480-216-4. LCCN BD232 .T24 1983eb. Lectures given at Brasenose College, Oxford, May 1982.

Dyson:1983:BRBa

[Dys83b] Freeman J. Dyson. Book review: Mathematics and Physics, by Yu. I. Manin, translated by Ann and Neal Koblitz (Boston, Birkhaüser, 1981), Progress in Physics, 3. The Mathematical Intelligencer, 5(2):54–57, ???? 1983. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Dyson:1983:BRBb

[Dys83c] Freeman J. Dyson. Book review: Robert Oppenheimer, Letters and Recollections, edited by Alice Smith and Charles Weiner. New Republic, ??(??):??, ???? 1983. ISSN 0028-6583.

Dyson:1983:CBK

[Dys83d] Freeman J. Dyson. Czas bez kónca: fizyka: biologia w otwartym wczechświecie. (Polish) [time without end: physics biology in an open wczechswiecie]. Postepy Fizyki [Progress of Physics], 34(??): 263–291, ???? 1983. Polish translation of [Dys79m].

Dyson:1983:QB

[Dys83e] Freeman J. Dyson. Quick is beautiful. In Bauer and McDonald [BM83], page ?? ISBN 3-540-12154-4, 0-387-12154-4. LCCN ???? After-dinner remarks.

Dyson:1983:SS

[Dys83f] Freeman J. Dyson. Science and space. In Needell [Nee83], pages 90–106. ISBN 0-87474-668-X. LCCN TL787 .F5 1983.

Dyson:1983:UP

[Dys83g] Freeman J. Dyson. Unfashionable pursuits. The Mathematical Intelligencer, 5(3):47–54, 1983. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Dyson:1984:BRB

[Dys84a] F. Dyson. Book review: Arsenal — Understanding Weapons in a Nuclear-Age, by K. Tsipis and The Abolition by Jonathan Schell. Science 84, 5(5):88, ???? 1984. ISSN 0193-4511.

Dyson:1984:BRC

[Dys84b] Freeman Dyson. Book review: Chen Ning Yang, Selected Papers, 1945–1980, with Commentary. American Journal of Physics, 52 (3):286–287, 1984. CODEN AJPIAS. ISSN 0002-9505 (print),

1943-2909 (electronic). URL http://link.aip.org/link/?AJP/
52/286/1.

Dyson:1984:APS

[Dys84c] Freeman J. Dyson. Astronomy in a private sphere. American Scholar, 53(2):169–182, Spring 1984. ISSN 0003-0937 (print), 2162-2892 (electronic).

Dyson:1984:DUPa

[Dys84d] Freeman J. Dyson. Disturbing the Universe, part 1. Priroda, 3 (??):123–128, ???? 1984. Russian translation of chapter 18 of [Dys79e].

Dyson:1984:DUPb

[Dys84e] Freeman J. Dyson. Disturbing the Universe, part 2. Priroda, 4 (??):123–128, ???? 1984. Russian translation of chapter 21 of [Dys79e].

Dyson:1984:FW

[Dys84f] Freeman J. Dyson. First word. Omni (New York), 6(10): 6, July 1984. ISSN 0149-8711. URL http://archive.org/

details/omni-magazine-1984-07;   http://omnimagindex.
files.wordpress.com/2012/02/84-07-004.jpg;   http://omnimagindex.
wordpress.com/author-indices/authors-d/.

Dyson:1984:ME

[Dys84g] Freeman J. Dyson. The Maxwell equations. In Berger [Ber84], pages 17–22. ISBN 0-444-86707-4 (Elsevier). LCCN QC669 .J17 1984.

Dyson:1984:OL

[Dys84h] Freeman J. Dyson. Origins of life. Nishina Memorial Lecture, given at Tokyo and Kyoto, Japan., October 1984.

Dyson:1984:RWHa

[Dys84i] Freeman J. Dyson. Reflections: Weapons and hope I — questions. The New Yorker, 60(??):52–??, February 6, 1984. ISSN 0028-

792X. URL http://www.newyorker.com/archive/1984/02/06/
1984_02_06_052_TNY_CARDS_000337681.

Dyson:1984:RWHb

[Dys84j] Freeman J. Dyson. Reflections: Weapons and hope II — tools. The New Yorker, 60(??):67–??, February 13, 1984. ISSN 0028-

792X. URL http://www.newyorker.com/archive/1984/02/13/
1984_02_13_067_TNY_CARDS_000339938.

Dyson:1984:RWHc

[Dys84k] Freeman J. Dyson. Reflections: Weapons and hope III — people. The New Yorker, 60(??):52–??, February 20, 1984. ISSN 0028-

792X. URL http://www.newyorker.com/archive/1984/02/20/
1984_02_20_052_TNY_CARDS_000340268.

Dyson:1984:RWHd

[Dys84l] Freeman J. Dyson. Reflections: Weapons and hope IV — concepts. The New Yorker, 60(??):54–??, February 27, 1984. ISSN

0028-792X. URL http://www.newyorker.com/archive/1984/
02/27/1984_02_27_054_TNY_CARDS_000338385.

Dyson:1984:RSS

[Dys84m] Freeman J. Dyson. Role of strategic space forces in a non-nuclear world. Physics Today, 37(6):9, 89, June 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/?PTO/37/9/1.

Dyson:1984:SYS

[Dys84n] Freeman J. Dyson. Saying yes and saying no. Convocation speech to Susquehanna University., September 2, 1984.

Dyson:1984:SSO

[Dys84o] Freeman J. Dyson. Science and space, oral version. Evening talk at the Cornell University Symposium in honor of Ed Salpeter’s sixtieth birthday., October 1, 1984.

Dyson:1984:SAB

[Dys84p] Freeman J. Dyson. Scientific autobiography for Adventures in Science. Princeton, NJ., February 1984.

Dyson:1984:SBO

[Dys84q] Freeman J. Dyson. Space butterflies: And other speculations. Science 85: sixth anniversary issue, 85(??):127–130, November

Dyson:1984:UP

[Dys84r] Freeman J. Dyson. Unfashionable pursuits. In Bars et al. [BCTG84], pages 265–285. ISBN 0-306-41801-0. LCCN QC793.3.S9 S93 1984.

Dyson:1984:WH

[Dys84s] Freeman J. Dyson. Weapons and hope. Harper & Row, New York, NY, USA, 1984. ISBN 0-06-039031-X. viii + 340 pp. LCCN U264 .D97 1984. US$17.50. Parts of the book were published in The New Yorker, February, 1984.

Dyson:1985:FWB

[Dys85a] Freeman J. Dyson. Feathered wisdom: Book review: Hawks, Doves and Owls: An Agenda for Avoiding Nuclear War, by Graham T. Allison, Abert Carnesale, Joseph S. Nye, Jr. Editors. W. W. Norton and Co. 1985. Nature, 318(6045):415–416, December 5, 1985. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v318/n6045/pdf/318415a0.pdf.

Dyson:1985:GO

[Dys85b] Freeman J. Dyson. Gerard O’Neill. Remarks at retirement dinner, Princeton University, Princeton, NJ., May 30, 1985.

Dyson:1985:PD

[Dys85c] Freeman J. Dyson. In praise of diversity. Syllabus of Gifford Lectures, University of Aberdeen, Scotland., May 1985.

Dyson:1985:IPS

[Dys85d] Freeman J. Dyson. Interstellar propulsion systems. In Zuckerman and Hart [ZH95], page ?? ISBN 0-521-44335-0 (hardback), 0-521-44803-4 (paperback). LCCN QB54 .E95 1995. URL http:/

/www.loc.gov/catdir/description/cam026/94043739.html;
 http://www.loc.gov/catdir/toc/cam021/94043739.html.

Dyson:1985:OLa

[Dys85e] Freeman J. Dyson. Origins of life. Tanner Lectures, given at Trinity College, Cambridge. Lecture I: Illustrious Predecessors; Lecture II: Experiments and Theories; Lecture III: A Toy Model; Lecture IV: Open Questions., January/February 1985.

Dyson:1985:OLb

[Dys85f] Freeman J. Dyson. Origins of life. Cambridge University Press, Cambridge, UK, 1985. ISBN 0-521-30949-2. ix + 81 pp.

LCCN QH325 .D88 1985. URL http://www.loc.gov/catdir/
description/cam032/85019015.html.

Dyson:1985:OLPa

[Dys85g] Freeman J. Dyson. Origins of life, part 1. Kagaku, 55(??):268–276, ???? 1985. Japanese translation of Nishina Memorial lecture.

Dyson:1985:OLPb

[Dys85h] Freeman J. Dyson. Origins of life, part 2. Kagaku, 55(??):369–377, ???? 1985. Japanese translation of Nishina Memorial lecture.

Dyson:1985:RAN

[Dys85i] Freeman J. Dyson. Remarks at the acceptance of the National Book Critics Circle Award for general non-fiction, New York. Unpublished., January 31, 1985.

Dyson:1985:SWA

[Dys85j] Freeman J. Dyson. Star wars and Austrianization and nuclear winter. Albert Pick Lecture, given at the University of Chicago., April 4, 1985.

Dyson:1986:CCG

[Dys86a] F. J. Dyson. Citation classic: General theory of spin-wave interactions. Current Contents/Engineering Technology & Applied Sciences, ??(36):16, September 8, 1986. ISSN 0011-

3395. URL www.garfield.library.upenn.edu/classics1986/

A1986D731200001.pdf. Letter describing background of [Dys56c], and work that has since simplified, and superseded, that paper.

Dyson:1986:ANW

[Dys86b] Freeman J. Dyson. The abolition of nuclear weapons. New York Times, ??(??):??, November 1986. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095.

Dyson:1986:FW

[Dys86c] Freeman J. Dyson. First word. Omni (New York), 8(5):10, February 1986. ISSN 0149-8711. URL http://archive.org/

details/omni-magazine-1986-05;   http://omnimagindex.
wordpress.com/author-indices/authors-d/.

Dyson:1986:DU

[Dys86d] Freeman J. Dyson. Les dérangeurs de l’univers. Payot, Paris, France, 1986. ISBN 2-228-65010-2 (paperback). ii + 318 pp. LCCN ???? Translation to French by Odile Laversanne of [Dys79e]. Preface by Hubert Reeves.

Dyson:1986:R

[Dys86e] Freeman J. Dyson. Roots. Harvard Magazine, 88(??):17–20, July/ August 1986. Oration for the Phi Beta Kappa Literary Exercises at Harvard University, June 3, 1986.

Dyson:1986:SR

[Dys86f] Freeman J. Dyson. Science and religion. Statement to the Committee on Human Values, National Conference of Catholic Bishops, Detroit, Michigan. Published in [Dys87h]., September 16, 1986.

Dyson:1986:S

[Dys86g] Freeman J. Dyson. Space. Omni (New York), 8(8):22, May 1986. ISSN 0149-8711. URL http://archive.org/

details/omni-magazine-1986-05;   http://omnimagindex.
wordpress.com/author-indices/authors-d/.

Dyson:1987:BRH

[Dys87a] Freeman Dyson. Book review: Howard C. Berg, Random Walks in Biology. Physics Today, 40(3):73–74, March 1987. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/?PTO/40/73/1.

Dyson:1987:F

[Dys87b] Freeman Dyson. Foreword. In Monastyrskiĭ [Mon87], page ?? ISBN 0-8176-3262-X. LCCN QA29.R425 M6613 1987. US$34.95. With a foreword by Freeman J. Dyson. Translated from Russian by James King and Victoria King.

Dyson:1987:LBD

[Dys87c] Freeman Dyson. Livets to begyndelser. (Danish) [Life’s two beginnings]. Nysyn. Munksgaard, København, Danmark, 1987. ISBN 87-16-06950-1. 106 pp. LCCN ????

Dyson:1987:DB

[Dys87d] Freeman J. Dyson. Demystifying the bomb. New York Times, ?? (??):SM52–??, April 5, 1987. CODEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-8095. URL http://search.proquest.

com/hnpnewyorktimes/docview/110801349/.

Dyson:1987:ODV

[Dys87e] Freeman J. Dyson. Origini della vita. (Italian) [Origins of Life]. Saggi. Scienze. Bollati Boringhieri, Torino, Italia, 1987. 97 pp. LCCN ????

Dyson:1987:PMD

[Dys87f] Freeman J. Dyson. Paul A. M. Dirac. American Philosophical Society Year Book for 1986, pages 100–104, 1987. CODEN YAP-SAL. ISSN 0065-9762. Obituary notice.

Dyson:1987:P

[Dys87g] Freeman J. Dyson. Preface. In The physicists: the history of a scientific community in modern America [Kev87], page ?? ISBN 0-674-66655-0 (paperback). LCCN QC9.U5 K48

1987. URL http://catalog.hathitrust.org/api/volumes/
oclc/15860449.html.

Dyson:1987:SR

[Dys87h] Freeman J. Dyson. Science and religion. In Byers [Bye87], pages 47–62. ISBN 1-55586-157-1 (paperback). LCCN BL240.2 .C65 1986.

Dyson:1988:SAC

[Dys88a] F. Dyson. SSC alternatives — critics collide with Dyson — reply. Physics Today, 41(5):132, ???? 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Dyson:1988:ASS

[Dys88b] Freeman Dyson. Alternatives to the Superconducting Super Collider. Physics Today, 41(2):77, February 1988. CODEN PH-TOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/?PTO/41/77/1.

Dyson:1988:IAD

[Dys88c] Freeman J. Dyson, editor. Infinite in all directions: Gifford lectures given at Aberdeen, Scotland, April–November 1985. Harper and Row, New York, NY, USA, 1988. ISBN 0-06-039081-6. viii + 321 pp. LCCN Q175.3 .D97 1988. US$19.95.

Dyson:1988:WTR

[Dys88d] Freeman J. Dyson. A walk through Ramanujan’s garden. In Andrews et al. [AAB^{+}88], pages 7–28. ISBN 0-12-058560-X. LCCN QA1 .R26 1987.

Dyson:1989:DOA

[Dys89a] F. Dyson. Dissenting opinions about Tayseer Aruri — reply. Physics Today, 42(8):87, ???? 1989. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic).

Dyson:1989:CU

[Dys89b] Freeman Dyson. Colonizing the universe. In An Agenda for the 21st Century [Kid89], pages 115–122. ISBN 0-262-11128-4. LCCN CB161 .A35 1987.

Dyson:1989:SR

[Dys89c] Freeman Dyson. The scientist as rebel. New York Review of Books, 42(9):??, May 25, 1989. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1995/may/25/the-scientist-as-rebel/.

Dyson:1989:BRG

[Dys89d] Freeman J. Dyson. Book review: Gary Saul Morson, Hidden in Plain View. Tolstoy Studies Journal, 2(1–3):??, ???? 1989.

Dyson:1989:FC

[Dys89e] Freeman J. Dyson. Feynman at Cornell. Physics Today, 42(2):32– 38, February 1989. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/

42/32/1. Special issue on Richard Feynman.

Dyson:1989:IAD

[Dys89f] Freeman J. Dyson, editor. Infinite in all directions. Pelican books. Penguin, 1989. ISBN 0-14-022873-X (paperback). viii + 32 pp. LCCN Q158.5.

Dyson:1989:IOD

[Dys89g] Freeman J. Dyson. Infinito in ogni direzione. (Italian) [Infinite in all directions]. Rizzoli, Milano, Italia, 1989. ISBN 88-17-85479-4. 290 pp. LCCN ????

Dyson:1989:LEM

[Dys89h] Freeman J. Dyson. Letter to the editor: “Mathematicians sweep 1988 Wolf Prizes” [Math. Intelligencer 11 (1989), no. 2, 39–48, MR0994963 (90c:01052)] by L. Zalcman. The Mathematical Intelligencer, 11(4):3, 1989. CODEN MAINDC. ISSN 0343-6993 (print), 1866-7414 (electronic).

Dyson:1989:MSP

[Dys89i] Freeman J. Dyson. Mappings and symmetries of partitions. Journal of Combinatorial Theory (Series A), 51(2):169–180, 1989. CODEN JCBTA7. ISSN 0097-3165 (print), 1096-0899 (electronic).

Dyson:1989:ONR

[Dys89j] Freeman J. Dyson. Oruzhie i Nadezhda. (Russian) [Weapons and Hope]. Progress Publishers, Moscow, Russia, 1989. ISBN ???? ???? pp. LCCN ????

Dyson:1989:P

[Dys89k] Freeman J. Dyson. Preface. In Atom and void: essays on science and community [Opp89], page ?? ISBN 0-691-08547-1, 0-691-02434-0 (paperback). LCCN Q175 .O65 1989. URL http://www.

loc.gov/catdir/description/prin031/89010413.html. Pref-

ace by Freeman J. Dyson.

Dyson:1989:SET

[Dys89l] Freeman J. Dyson. The search for extraterrestrial technology. In Kuiper and Brin [KB89], page ?? ISBN 0-917853-38-5. LCCN QB54 .E945; QB54 .E97 1989.

Dyson:1989:ZEP

[Dys89m] Freeman J. Dyson. Zeit ohne Ende: Physik und Biologie in einem offenen Universum. (German) [Time without end: Physics and biology in an open universe]. Brinkmann and Bose, Berlin, Germany, 1989. ISBN 3-922660-39-8. 87 pp. LCCN ????

Dyson:1990:BSW

[Dys90a] Freeman J. Dyson. Banquet speech: Willard Gibbs and the teaching of science. In Caldi and Mostow [CM90], pages 269–276. ISBN 0-8218-0157-0. LCCN QC310.15 .G53 1989. URL http://www.

gbv.de/dms/bowker/toc/9780821801574.pdf.

Dyson:1990:BRP

[Dys90b] Freeman J. Dyson. Book review: Paul Davies, The New Physics. American Journal of Physics, 58(3):286–287, 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/58/286/1.

Dyson:1990:CDA

[Dys90c] Freeman J. Dyson. Carbon dioxide in the atmosphere and the biosphere. Radcliffe Lecture given at Green College, Oxford., October 11, 1990.

Dyson:1990:FG

[Dys90d] Freeman J. Dyson. The face of Gaia. In Fadiman [Fad90], chapter 2, pages 7–15. ISBN 0-385-24880-6. LCCN P85.S74 Z68 1990; BJ1571 .L635 1990.

Dyson:1990:FPM

[Dys90e] Freeman J. Dyson. Feynman’s proof of the Maxwell equations. American Journal of Physics, 58(3):209–211, 1990. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/58/209/1.

Dyson:1990:FS

[Dys90f] Freeman J. Dyson. The future of science. Remarks for Symposium at the Académie des Sciences, Paris, France. French translation in [Dys91e]., May 15, 1990.

Dyson:1990:IBU

[Dys90g] Freeman J. Dyson. The importance of being unpredictable. Statement at the Encyclopaedia Britannica award ceremony, New York, NY., February 28, 1990.

Dyson:1990:IAD

[Dys90h] Freeman J. Dyson. Infinite in all directions: Gifford lectures given at Aberdeen, Scotland, April–November 1985. Penguin Science. Penguin Books, Harmondsworth, UK, repr. edition, 1990. ISBN 0-14-014482-X. viii + 321 pp. LCCN ????

Dyson:1990:MOV

[Dys90i] Freeman J. Dyson. Major observatories versus economy-class observatories in space. In Kondo [Kon90], pages 399–405. ISBN 0-7923-1133-7 (hardcover). LCCN QB500.267 .I56 1990.

Dyson:1990:OA

[Dys90j] Freeman J. Dyson. Occultation astronomy. In Kondo [Kon90], pages 413–415. ISBN 0-7923-1133-7 (hardcover). LCCN QB500.267 .I56 1990.

Dyson:1990:RBN

[Dys90k] Freeman J. Dyson. Review of Ramanujan’s Notebooks, Parts I & II, edited by Bruce C. Berndt. Bulletin of the London Mathematical Society, 22(??):607–609, ???? 1990. CODEN LMSBBT. ISSN 0024-6093 (print), 1469-2120 (electronic).

Dyson:1990:SBW

[Dys90l] Freeman J. Dyson. Strategic bombing in World War II and today. Has anything changed? Talk given in the series The Legacy of Strategic Bombing at the National Air and Space Museum, Washington, DC., March 1, 1990.

Dyson:1990:WHM

[Dys90m] Freeman J. Dyson. Was HST a mistake? Sky & Telescope, ?? (??):357, April 1990.

Dyson:1991:BRR

[Dys91a] Freeman Dyson. Book review: Ralph Leighton, Tuva or Bust! Richard Feynman’s Last Journey. Physics Today, 44(10):114– 115, October 1991. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
44/114/1.

Dyson:1991:BRK

[Dys91b] Freeman J. Dyson. Book review: Kameshwar C. Wali, Chandra: A Biography of S. Chandrasekhar. Physics Today, 44(3):65–66, March 1991. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/?PTO/44/
65/1.

Dyson:1991:BSS

[Dys91c] Freeman J. Dyson. Butterflies and super strings. In Ferris and Fadiman [FF91], pages 128–?? ISBN 0-316-28129-8. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Dyson:1991:CR

[Dys91d] Freeman J. Dyson. Concluding remarks. In USA–USSR Joint Conference on The Search for Extraterrestrial Intelligent Life:

University of California, Santa Cruz, August 5–9, 1991, page ?? ????, ????, 1991.

Dyson:1991:IIF

[Dys91e] Freeman J. Dyson. De l’importance de l’imprévisible. (French) [on the importance of the unpredictable]. In L’avenir de la science: vu par vingt-sept membres associés étrangers de l’Académie des Sciences. (French) [The future of science: seen by 27 foreign associate members of the Academy of Sciences] [Ham91], pages 27–32. ISBN 2-10-000115-9 (paperback). LCCN ???? Debates directed by Alain Connes, Jacques-Louis Lions, François Jacob and others.

Dyson:1991:FF

[Dys91f] Freeman J. Dyson. The final frontier. Issues in Science and Technology, 8(1):6, Fall 1991. CODEN ???? ISSN 0748-5492 (print), 1938-1557 (electronic). Missions to Mars.

Dyson:1991:HCP

[Dys91g] Freeman J. Dyson. Hunting for comets and planets. Milne Lecture given at Oxford. Published in [Dys92d]., October 24, 1991.

Dyson:1991:P

[Dys91h] Freeman J. Dyson. Preface. In Thirring [Thi91], page ?? ISBN 3-540-53039-8 (Berlin), 0-387-53039-8 (New York). LCCN QC173.4.T48 L54 1991. With a preface by F. Dyson.

Dyson:1991:TTF

[Dys91i] Freeman J. Dyson. “To teach or not to teach”, Freeman J. Dyson’s acceptance speech for the 1991 Oersted medal presented by the American Association of Physics Teachers, 22 January 1991. American Journal of Physics, 59(6):491–495, 1991. CO-DEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic).

URL http://link.aip.org/link/?AJP/59/491/1.

Dyson:1992:EG

[Dys92a] Freeman Dyson. From Eros to Gaia. A Cornelia and Michael Bessie book. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-41307-3. xi + 371 pp. LCCN Q158.5.

Dyson:1992:BRJ

[Dys92b] Freeman J. Dyson. Book review: James Gleick, Genius: The Life and Science of Richard Feynman. Physics Today, 45(11):

87, November 1992. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/

45/87/1. See retraction of one remark in [FD93].

Dyson:1992:DTR

[Dys92c] Freeman J. Dyson. Dragon’s teeth: Review of The Los Alamos Primer: The First Lectures on How To Build an Atomic Bomb, edited by Robert Serber and Richard Rhodes, University of California Press. Science, 256(5055):388–389, April 17, 1992. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/2877089;
http://www.sciencemag.org/content/256/5055/388.extract.

Dyson:1992:HCP

[Dys92d] Freeman J. Dyson. Hunting for comets and planets. Quarterly Journal of the Royal Astronomical Society, 33(2):45–57, June 1992. CODEN QJRAAK. ISSN 0035-8738. URL http:/

/adsabs.harvard.edu/abs/1992QJRAS..33...45D. Milne Lec-

ture given at Oxford, 24 October 1991.

Dyson:1992:LC

[Dys92e] Freeman J. Dyson. The lemon and the cream. Technology Review (M.I.T.), 95(5):30–36, July 1992. CODEN TEREAU. ISSN 0040-1692.

Dyson:1992:NND

[Dys92f] Freeman J. Dyson. Nearest neighbor distances on a circle. IASSNS-HEP-92/27., 1992.

Dyson:1992:P

[Dys92g] Freeman J. Dyson. Preface. In Against transcendence: essays on the physical significance of modern culture [For02], page ?? ISBN 0-521-43020-8, 0-521-43615-X. LCCN ????

Dyson:1992:QPL

[Dys92h] Freeman J. Dyson. Quantum past: The limitations of quantum theory. Schrödinger Lecture, given at Imperial College. IASSNS-HEP-92/87., May 7, 1992.

Dyson:1992:RA

[Dys92i] Freeman J. Dyson. Revolutions in astronomy. Compton Memorial Lecture, given at Washington University, St. Louis, MO., October 14, 1992.

Dyson:1993:GGP

[Dys93a] Freeman Dyson. George Green and physics. Physics World, 6 (8):33–38, August 1993. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL iopscience.iop.org/pwa/

full/pwa-pdf/6/8/phwv6i8a28.pdf.

Dyson:1993:EG

[Dys93b] Freeman J. Dyson. Da Eros a Gaia. Rizzoli, Milano, Italia, 1993. ISBN 88-17-84257-5. 373 + 1 pp. LCCN ???? Translation to Italian by Libero Sosio of [Dys92a].

Dyson:1993:DU

[Dys93c] Freeman J. Dyson. Disturbing the Universe, page ???? Commonwealth Publishing Co., Taipei, Republic of China, 1993. ISBN ???? New preface.

Dyson:1993:FC

[Dys93d] Freeman J. Dyson. Feynman at Cornell. In Brown and Rigden [BR93], pages 39–52. ISBN 0-88318-870-8. LCCN QC16.F49 A3 1993. Commentary by Joan Feynman, John Wheeler, Hans Bethe, Julian Schwinger, Murray Gell-Mann, Daniel Hillis, David Goodstein, Freeman Dyson, and Laurie Brown.

Dyson:1993:HGG

[Dys93e] Freeman J. Dyson. Homage to George Green: How physics looked in the Nineteen-Forties. Talk given at the George Green Bicentenary Celebrations Nottingham University, England. Mentioned in University of Nottingham Gazette, Autumn 1993, pp. 2327., July 14, 1993.

Dyson:1993:OGK

[Dys93f] Freeman J. Dyson. Obituary: Gerard Kitchen O’Neill. Physics Today, 46(2):97–98, February 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://www.

physicstoday.org/resource/1/phtoad/v46/i2/p97_s2.

Dyson:1993:ST

[Dys93g] Freeman J. Dyson. Science in trouble. American Scholar, 62 (4):513–525, Autumn 1993. ISSN 0003-0937 (print), 2162-2892 (electronic).

Dyson:1994:STR

[Dys94a] F. Dyson. Science in trouble: Reply to critics R. P. Fairchild, Robert Kornfeld and Charles W. Mccutchen. American Scholar, 63(1):158, ???? 1994. ISSN 0003-0937 (print), 2162-2892 (electronic).

Dyson:1994:EG

[Dys94b] Freeman J. Dyson. De Eros a Gaia, volume 35 of Metatemas: libros para pensar la ciencia. Tusquets Editores, Barcelona, España, 1994. ISBN 84-7223-775-3. 387 pp. LCCN Q158.5 .D9718 1994. Translation to Spanish by José Luis Fernández-Villanueva of [Dys92a].

Dyson:1994:P

[Dys94c] Freeman J. Dyson. Preface. In Traditional Japanese mathematics problems of the 18th and 19th Centuries [FR94], page ?? ISBN ???? LCCN ????

Dyson:1995:NI

[Dys95a] F. J. Dyson. Nuclear intrigues. Scientific American, 273(3):10, ???? 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dyson:1995:NUN

[Dys95b] Freeman Dyson. Non-use and non-violence. In Oak Ridge Symposium on Non-Use of Nuclear Weapons, Oak Ridge, Tennessee, May 1995, pages 7–?? ????, ????, 1995.

Dyson:1995:CS

[Dys95c] Freeman J. Dyson. 21st-Century spacecraft. Scientific American, 273(3):114–??, September 1995. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

Dyson:1995:BAS

[Dys95d] Freeman J. Dyson. Daedalus after seventy years. In Dronamraju [Dro95], pages 55–63. ISBN 0-19-854846-X. LCCN Q171 .H1566 1995. With an introduction by Krishna R. Dronamraju. Foreword by Joshua Lederberg.

Dyson:1995:CFF

[Dys95e] Freeman J. Dyson. The Coulomb fluid and the fifth Painlevé transcendent. In Liu and Yau [LY95], pages 131–146. ISBN 1-57146-001-2. LCCN QC16.Y364 A3 1995.

Dyson:1995:APS

[Dys95f] Freeman J. Dyson. Ad Portas speech. The Wykehamist, ??(??): ??, June 1995.

Dyson:1995:SAF

[Dys95g] Freeman J. Dyson. Science as an art form. The Chronicle of Higher Education, ??(??):??, June 16, 1995. ISSN 0009-5982 (print), 1931-1362 (electronic). URL http://chronicle.com/

article/Science-as-an-Art-Form/83630/.

Dyson:1995:SRa

[Dys95h] Freeman J. Dyson. The scientist as rebel. In Cornwell [Cor95], pages 1–11. ISBN 0-19-851775-0. LCCN ????

Dyson:1995:SRb

[Dys95i] Freeman J. Dyson. The scientist as rebel. New York Review of Books, 42(9):31–33, May 25, 1995. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1995/may/25/the-scientist-as-rebel/. Reprint of

[Dys95h].

Dyson:1995:SFC

[Dys95j] Freeman J. Dyson. Smaller, faster, cheaper: Empty words or practical goals. Popular Science, ??(??):??, Fall 1995.

Dyson:1995:TCS

[Dys95k] Freeman J. Dyson. Transportation: 21st-Century spacecraft. Scientific American, 273(3):114–116, 116A, September 1995. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v273/n3/pdf/scientificamerican0995-144.pdf.

Dyson:1996:A

[Dys96a] Freeman Dyson. Addendum. In Odifreddi [Odi96], page 75. ISBN 1-56881-061-X. LCCN QA29.K73 K74 1996.

Dyson:1996:PRS

[Dys96b] Freeman Dyson. Preliminary remarks — Schwinger’s response to the award of an honorary degree at Nottingham. In Ng [Ng96], pages 9–12. ISBN 981-02-2531-8 (hardcover), 981-02-2532-6 (paperback), 981-283-044-8 (e-book). LCCN QC16.S29

J85 1996. URL http://www.worldscientific.com/doi/abs/
10.1142/9789812830449_0002.

Dyson:1996:SGQ

[Dys96c] Freeman Dyson. Schwinger’s “The greening of quantum field theory: George and I”. In Ng [Ng96], page ?? ISBN 981-02-2531-8 (hardcover), 981-02-2532-6 (paperback), 981-283-044-8 (e-book).

LCCN QC16.S29 J85 1996. URL http://site.ebrary.com/id/
10691945;   http://www.myilibrary.com?id=494808.

Dyson:1996:SR

[Dys96d] Freeman Dyson. The scientist as rebel. American Mathematical Monthly, 103(9):800–805, November 1996. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http://

www.jstor.org/stable/2974452.

Dyson:1996:SPF

[Dys96e] Freeman Dyson. Selected papers of Freeman Dyson with commentary, volume 5 of Collected Works. American Mathematical Society, Providence, RI, USA, 1996. ISBN 0-8218-0561-4. xii + 601 pp. With a foreword by Elliott H. Lieb, and commentaries by the author about how the papers were written.

Dyson:1996:BM

[Dys96f] Freeman J. Dyson. Beyond Mars. Ad Astra, the magazine of the National Space Society, 8(??):40–43, May/June 1996.

Dyson:1996:BRBb

[Dys96g] Freeman J. Dyson. Book review: Nature’s Numbers by Ian Stewart. American Mathematical Monthly, 103(7):610–612, August/ September 1996. CODEN AMMYAE. ISSN 0002-9890 (print),

1930-0972 (electronic). URL http://www.jstor.org/stable/
2974684.

Dyson:1996:BRBa

[Dys96h] Freeman J. Dyson. Book review: Ramanujan: Letters and Commentary by Bruce C. Berndt and Robert A. Rankin. Isis, 87(2): 387, June 1996. CODEN ISISA4. ISSN 0021-1753 (print), 1545-

6994 (electronic). URL http://www.jstor.org/stable/236134.

Dyson:1996:BRRb

[Dys96i] Freeman J. Dyson. Book review: Reality bites: The End of the World: The Science and Ethics of Human Extinction, By John Leslie. Routledge: 1996. Pp. 305. Nature, 380(6572):296, March 28, 1996. CODEN NATUAS. ISSN 0028-0836 (print),

1476-4687 (electronic). URL http://www.nature.com/nature/
journal/v380/n6572/pdf/380296a0.pdf.

Dyson:1996:BRRa

[Dys96j] Freeman J. Dyson. Book review: Roald Hoffmann, The Same and Not the Same, Columbia U.P. New York, 1995. 294 pp. $34.95 hardcover, ISBN 0-231-10138-4. Physics Today, 49(1):64–65, January 1996. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-

0699 (electronic). URL http://link.aip.org/link/?PTO/49/
64/2.

Dyson:1996:F

[Dys96k] Freeman J. Dyson. Foreword. In The quotable Einstein [Cal96], page ?? ISBN 0-691-02696-3 (hardcover). LCCN QC16.E5

A25 1996. URL ftp://uiarchive.cso.uiuc.edu/pub/etext/
gutenberg/;   http://www.loc.gov/catdir/description/prin031/
96003543.html;   http://www.loc.gov/catdir/toc/prin031/
96003543.html.

Dyson:1996:TRA

[Dys96l] Freeman J. Dyson. Two revolutions in astronomy. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 140(1):1–9, March 1996. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL

http://www.jstor.org/stable/987272.

Dyson:1996:WFA

[Dys96m] Freeman J. Dyson. Writing a foreword for Alice Calaprice’s new Einstein book. Princeton University Library Chronicle, 57(3): 491–502, ???? 1996. See [Cal96].

Dyson:1997:BRP

[Dys97a] Freeman J. Dyson. Book review: Pierre-Gilles de Gennes and Jacques Badoz, Fragile Objects: Soft Matter, Hard Science, and the Thrill of Discovery, ISBN 0-387-94774-4. American Journal of Physics, 65(7):675, 1997. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.

org/link/?AJP/65/675/1.

Dyson:1997:CSE

[Dys97b] Freeman J. Dyson. Can science be ethical? New York Review of Books, 44(6):46, April 10, 1997. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1997/apr/10/can-science-be-ethical/.

Dyson:1997:CHC

[Dys97c] Freeman J. Dyson. Chess and the human condition. Think [House magazine published by IBM], ??(??):??, Fall 1997.

Dyson:1997:GD

[Dys97d] Freeman J. Dyson. God is in the details. In Eating bread and honey [Rog97], page ?? ISBN 1-57131-406-7. LCCN PS3568.O454 E38 1997.

Dyson:1997:IW

[Dys97e] Freeman J. Dyson. Imagined worlds. Jerusalem–Harvard lectures. Harvard University Press, Cambridge, MA, USA, 1997. ISBN 0-674-53908-7. 216 pp. LCCN AC8 .D97 1997. Expanded version of the 1985 Harvard–Jerusalem lectures.

Dyson:1997:P

[Dys97f] Freeman J. Dyson. Preface. In Thirring and Lieb [TL97], page ?? ISBN 3-540-61565-2. LCCN ???? URL http://www.gbv.de/

dms/goettingen/215927141.pdf.

Dyson:1997:R

[Dys97g] Freeman J. Dyson. The race is over. New York Review of Books, 44(4):4, March 6, 1997. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
1997/mar/06/the-race-is-over/.

Dyson:1997:WBP

[Dys97h] Freeman J. Dyson. Warm-blooded plants and freeze-dried fish. Atlantic Monthly, 280(??):71–80, November 1997. ISSN 1072-7825 (print), 2151-9463 (electronic).

Dyson:1998:GC

[Dys98a] F. J. Dyson. Gravity is cool. Forbes, ??(??):160–16+, November 1998. CODEN FORBA5. ISSN 0015-6914.

Dyson:1998:BRG

[Dys98b] Freeman J. Dyson. Book review: Is God in the lab? The Meaning of It All: Thoughts of a Citizen Scientist, by Richard P. Feynman. Addison-Wesley, 133 pp., $22.00. Belief in God in an Age of Science, by John Polkinghorne. Yale University Press, 138 pp., $18.00. New York Review of Books, 45(9):8–??, May 28, 1998. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1998/may/28/is-god-in-the-lab/. See responses

[SD98, Sta98].

Dyson:1998:BRV

[Dys98c] Freeman J. Dyson. Book review: Val L. Fitch, Daniel R. Marlow, Margit A. E. Dementi, editors, Critical Problems in Physics. American Journal of Physics, 66(9):837–838, September 1998. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909

(electronic). URL http://labs.adsabs.harvard.edu/ui/abs/
1998AmJPh..66..837F;   http://link.aip.org/link/?AJP/
66/837/1.

Dyson:1998:CRB

[Dys98d] Freeman J. Dyson. Consilience: review of book by Edward O. Wilson. The New England Journal of Medicine, ??(??):??, July 16, 1998. CODEN NEJMAG. ISSN 0028-4793 (print), 1533-4406 (electronic).

Dyson:1998:IPN

[Dys98e] Freeman J. Dyson. Innovation in physics (with note added forty years later). In you Wu et al. [yWHH98], pages 73–93. ISBN 981-02-3372-8. LCCN QC20 .J554 1997.

URL http://www.worldscientific.com/doi/abs/10.1142/
9789812816566_0008.

Dyson:1998:SCI

[Dys98f] Freeman J. Dyson. Science as a craft industry. Science, 280 (5366):1014–1015, May 15, 1998. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://labs.adsabs.

harvard.edu/ui/abs/1998Sci...280R1014D;   http://www.
jstor.org/stable/2895427;   http://www.sciencemag.org/
content/280/5366/1014.full.

Dyson:1998:TSJ

[Dys98g] Freeman J. Dyson. Technology and social justice. Technical report, Carnegie Council on Ethics and International Affairs, New York, NY, USA, March 1998.

Dyson:1999:CRC

[Dys99a] Freeman Dyson. A conservative revolutionary [concerning Frank Yang]. Modern Physics Letters A (MPLA), 14(22):1455–1459, 1999. CODEN MPLAEQ. ISSN 0217-7323 (print), 1793-6632

(electronic). URL http://labs.adsabs.harvard.edu/ui/abs/
1999MPLA...14.1455D.

Dyson:1999:ASJ

[Dys99b] Freeman J. Dyson. Abdus Salam (29 January 1926–21 November 1996): biographical memoir. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 143(2):347–350, June 1999. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.

org/stable/3181945.

Dyson:1999:Fa

[Dys99c] Freeman J. Dyson. Foreword. In The Deep Hot Biosphere [Gol99], page ?? ISBN 0-387-98546-8 (hardcover). LCCN TN870.5 .G66 1999. Foreword by Freeman Dyson.

Dyson:1999:Fb

[Dys99d] Freeman J. Dyson. Foreword. In Riemann, topology, and physics [Mon99], pages ix–x. ISBN 0-8176-3789-3, 3-7643-3789-3. LCCN

QA29.R425 M6613 1999. URL http://link.springer.com/
book/10.1007/978-0-8176-4779-7/page/1. Foreword by Free-

man J. Dyson. Translated by Roger Cooke, James King, Victoria King.

Dyson:1999:Fc

[Dys99e] Freeman J. Dyson. Foreword. In Bruce and Milne [BM99], pages ix–xii. ISBN 0-333-76070-0 (Macmillan Press: hardcover), 0-333-77482-5 (Macmillan Press: paperback), 0-312-22570-9 (St. Martin’s Press: hardcover). LCCN JZ5665 .E53 1999.

URL http://www.loc.gov/catdir/bios/hol057/99026118.
html;   http://www.loc.gov/catdir/description/hol056/
99026118.html;   http://www.loc.gov/catdir/toc/hol053/
99026118.html.

Dyson:1999:IMS

[Dys99f] Freeman J. Dyson. The inventor of modern science. Nature, 400(6739):27, July 1, 1999. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

labs.adsabs.harvard.edu/ui/abs/1999Natur.400...27D;

http://www.nature.com/nature/journal/v400/n6739/full/
400027a0.html.

Dyson:1999:MRD

[Dys99g] Freeman J. Dyson. Miracles of rare device. The Sciences (New York), 39(2):32–37, March/April 1999. CODEN SCNCAD. ISSN 0036-861X.

Dyson:1999:OL

[Dys99h] Freeman J. Dyson. Origins of Life. Cambridge University Press, Cambridge, UK, revised edition, 1999. ISBN 0-521-62668-4 (paperback). ix + 100 pp. LCCN QH325 .D88 1999. URL http:/

/www.loc.gov/catdir/description/cam029/99021079.html;
 http://www.loc.gov/catdir/samples/cam032/99021079.html;
 http://www.loc.gov/catdir/toc/cam025/99021079.html.

Dyson:1999:SGI

[Dys99i] Freeman J. Dyson. The Sun, the genome and the Internet: tools of scientific revolutions. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1999. ISBN 0-19-512942-3. xvi + 124 pp. LCCN QC20 .D98 1999. URL http://www.loc.gov/

catdir/enhancements/fy0637/98053830-d.html;   http://
www.loc.gov/catdir/enhancements/fy0724/98053830-b.html.

Dyson:1999:WMT

[Dys99j] Freeman J. Dyson. Why is Maxwell’s theory so hard to understand? In James Clerk Maxwell Commemorative Booket for The Fourth International Congress on Industrial and Applied Mathematics, page ?? James Clerk Maxwell Foundation, Edinburgh, UK, 1999. URL http://

ieeexplore.ieee.org/stamp/stamp.jsp?arnumber=4458240;
 http://www.clerkmaxwellfoundation.org/DysonFreemanArticle.
pdf;   http://www.damtp.cam.ac.uk/user/tong/em/dyson.
pdf.

Dyson:2000:API

[Dys00a] F. J. Dyson. Abstracts presented at ISSOL ’99, July 1999, San Diego, California, USA. Origins of Life and Evolution of the Biosphere, 30(2–4):115–397, August 2000. CODEN OGLFAU. ISSN 0169-6149 (print), 1573-0875 (electronic). URL http://

link.springer.com/article/10.1023/A%3A1017300204711.

Dyson:2000:SD

[Dys00b] F. J. Dyson. The stuff of dreams. Natural History, 109(1):27, ???? 2000. ISSN 0028-0712.

Dyson:2000:F

[Dys00c] Freeman J. Dyson. Foreword. In The pleasure of finding things out: the best short works of Richard P. Feynman [Fey00], pages ix–xiv. ISBN 0-7382-0349-1 (paperback). LCCN Q171 .F385 1999. Edited by Jeffrey Robbins, and foreword by Freeman Dyson.

Dyson:2000:HCW

[Dys00d] Freeman J. Dyson. How can we further explore the ‘microverse’? In Bell et al. [BDF00], pages 196–199. ISBN 0-7803-5360-9 (Platinum ed.), 0-7803-5361-7 (Member hardcover ed.), 0-7803-5362-5 (Trade hardcover ed.). LCCN T174 .B451 2000.

Dyson:2000:PR

[Dys00e] Freeman J. Dyson. Progress in religion. Document on Edge Web

site., May 16, 2000. URL http://www.edge.org/documents/
archive/edge68.html.

Dyson:2000:SFN

[Dys00f] Freeman J. Dyson. The sixth Fermat number and palindromic continued fractions. L’Enseignement mathématique, 46(3–4):385– 389, ???? 2000. CODEN ENMAAR. ISSN 0013-8584.

Dyson:2001:HGG

[Dys01a] Freeman J. Dyson. Homage to George Green: how physics looked in the nineteen-forties. In George Green: mathematician and physicist, 1793–1841: the background to his life and work [Can01], page ?? ISBN 0-89871-463-X. LCCN QC16.G64

C36 2001. URL http://epubs.siam.org/ebooks/siam/other_
titles_in_applied_mathematics/ot73;   http://www.loc.
gov/catdir/enhancements/fy0726/00041938-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0726/00041938-t.html.

Dyson:2001:SGI

[Dys01b] Freeman J. Dyson. Le soleil, le génome et Internet: les révolutions scientifiques et leurs instruments. Flammarion, Paris, France, 2001. ISBN 2-08-211248-9 (paperback). 137 pp. LCCN ???? French translation by Jacqueline Carnaud of [Dys99i].

Dyson:2001:P

[Dys01c] Freeman J. Dyson. Preface. In Thirring [Thi01], page ?? ISBN 3-540-42083-5. LCCN QC173.4.T48 L54 2001. With a preface by F. Dyson.

Dyson:2001:RTT

[Dys01d] Freeman J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman (with commentary). In Hsu and Zhang [HZ01], pages 316–337. ISBN 981-02-4721-4. LCCN

QC174.17.S9 H78 2001. URL http://www.loc.gov/catdir/
toc/fy031/2002280428.html.

Dyson:2001:S

[Dys01e] Freeman J. Dyson. Skyriders. In Scientific conversations: interviews on science from the New York Times [Dre01], pages 21–26. ISBN 0-7167-4661-1. LCCN Q141 .D74 2001. URL http://www.

loc.gov/catdir/enhancements/fy0667/2001043630-b.html;
 http://www.loc.gov/catdir/enhancements/fy0667/2001043630-
d.html;   http://www.loc.gov/catdir/enhancements/fy0667/

2001043630-t.html. With a foreword by Natalie Angier.

Dyson:2001:WTR

[Dys01f] Freeman J. Dyson. A walk through Ramanujan’s garden. In Ramanujan: essays and surveys [BR01], pages 261–275. ISBN 0-8218-2624-7. LCCN QA29.R3 B47 2001. MR0938957 (89f:01065).

Dyson:2002:BRB

[Dys02a] Freeman J. Dyson. Book review: Memoirs: A Twentieth-Century Journey in Science and Politics. Edward Teller with Judith Shoolery. 602 pp. Perseus, Cambridge, MA, 2001. Price: $35.00 ISBN 0-7382-0532-X. American Journal of Physics, 70(4):462–463, April 2002. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://ajp.aapt.org/resource/1/ajpias/v70/
i4/p462_s1.

Dyson:2002:BRP

[Dys02b] Freeman J. Dyson. Book review: In praise of amateurs: Seeing in the Dark: How Backyard Stargazers Are Probing Deep Space and Guarding Earth from Interplanetary Peril, by Timothy Ferris. Simon and Schuster, 379 pp., $26.00. New York Review of Books, 59(19):??, December 5, 2002. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2002/dec/05/in-praise-of-amateurs/http://www.nybooks.
com/articles/archives/2002/dec/05/in-praise-of-amateurs/

.

Dyson:2002:BRS

[Dys02c] Freeman J. Dyson. Book review: Science & religion: No ends in sight: The God of Hope and the End of the World, by John Polkinghorne. Yale University Press, 154 pp., $19.95. New York Review of Books, 49(5):4–??, March 28, 2002. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/2002/mar/28/
science-religion-no-ends-in-sight/.

Dyson:2002:BRC

[Dys02d] Freeman J. Dyson. Book review: The conscience of physics: The life, work and dreams of Wolfgang Pauli [No time to be brief: a scientific biography of Wolfgang Pauli: Charles P. Enz; Oxford University Press, New York, 2002, pp. 581, price US$60.00, £35.00, ISBN 0-19-856479-1]. Nature, 420 (6916):607–608, December 12, 2002. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

labs.adsabs.harvard.edu/ui/abs/2002Natur.420..607D;
http://www.nature.com/nature/journal/v420/n6916/full/
420607a.html.

Dyson:2002:ODV

[Dys02e] Freeman J. Dyson. Origini della vita. (Italian) [Origins of Life]. Saggi. Scienze. Bollati Boringhieri, Torino, Italia, second edition, 2002. ISBN 88-339-1380-5. 141 pp. LCCN ????

Dyson:2002:P

[Dys02f] Freeman J. Dyson. Preface. In Traditional Japanese mathematics problems of the 18th and 19th Centuries [FR02], page ?? ISBN 981-04-2759-X. LCCN ????

Dyson:2002:SRN

[Dys02g] Freeman J. Dyson. Science & religion: No ends in sight: The God of Hope and the End of the World by John Polkinghorne. New York Review of Books, 49(5):??, March 28, 2002. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/2002/mar/28/
science-religion-no-ends-in-sight/.

Dyson:2002:POT

[Dys02h] George Dyson. Project Orion: the true story of the atomic spaceship. Henry Holt and Company, New York, NY, USA, 2002. ISBN 0-8050-5985-7. xv + 345 pp. LCCN TL783.5 .D95

2002. URL http://www.daviddarling.info/encyclopedia/O/
OrionProj.html;   http://www.loc.gov/catdir/bios/hol051/
2001046500.html;   http://www.loc.gov/catdir/description/
hol021/2001046500.html.

Dyson:2003:WHV

[Dys03a] F. Dyson. Weapons and hope (vol 31, pg 257, 2002). Contemporary Sociology — a Journal of Reviews, 32(1):vii, January 2003. ISSN 0094-3061 (print), 1939-8638 (electronic).

Dyson:2003:BRN

[Dys03b] Freeman J. Dyson. Book review: A new Newton: Isaac Newton, by James Gleick. Pantheon, 272 pp., $22.95. New York Review of Books, 50(11):??, July 3, 2003. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2003/jul/03/a-new-newton/. See comment [Fei03].

Dyson:2003:BRC

[Dys03c] Freeman J. Dyson. Book review: Clockwork science: Einstein’s Clocks, Poincaré’s Maps: Empires of Time, by Peter Galison. Norton, 389 pp., $23.95. New York Review of Books, 50(17):42– 44, November 6, 2003. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2003/nov/06/clockwork-science/.

Dyson:2003:BRF

[Dys03d] Freeman J. Dyson. Book review: The future needs us! Prey, by Michael Crichton. HarperCollins, 367 pp., $26.95. New York Review of Books, 50(2):11–13, February 13, 2003. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2003/feb/13/the-future-needs-us/.

Dyson:2003:CR

[Dys03e] Freeman J. Dyson. A conservative revolutionary. In Goldhaber et al. [GSS^{+}03], pages 1–5. ISBN 981-238-503-7 (hardcover), 981-238-563-0 (paperback), 981-279-508-1 (e-book). LCCN

QC174.17.S9 S994 2003. URL http://www.worldscientific.
com/doi/abs/10.1142/9789812795083_0001.

Dyson:2003:EVG

[Dys03f] Freeman J. Dyson. Einstein’s view of Germany: [book review: Einstein in Berlin, by Thomas Levenson]. Nature, 422 (6934):811–812, April 24, 2003. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

labs.adsabs.harvard.edu/ui/abs/2003Natur.422..811D;
http://www.nature.com/nature/journal/v422/n6934/full/
422811a.html.

Dyson:2003:LLU

[Dys03g] Freeman J. Dyson. Looking for life in unlikely places: reasons why planets may not be the best places to look for life. International Journal of Astrobiology, 2(2):103–110, April 2003. CO-DEN IJANFR. ISSN 1473-5504 (print), 1475-3006 (electronic).

URL http://labs.adsabs.harvard.edu/ui/abs/2003IJAsB..
.2..103D.

Dyson:2003:RTT

[Dys03h] Freeman J. Dyson. The radiation theories of Tomonaga, Schwinger, and Feynman. In Selected Papers on Quantum Electrodynamics [Sch03], page ?? ISBN 0-486-60444-6. LCCN QC680 .S35. Reprint of [Dys49a].

Dyson:2003:MQE

[Dys03i] Freeman J. Dyson. The S-matrix in quantum electrodynamics. In Selected Papers on Quantum Electrodynamics [Sch03], page ?? ISBN 0-486-60444-6. LCCN QC680 .S35. Reprint of [Dys49b].

Dyson:2004:EWS

[Dys04a] Freeman Dyson. Elliott’s world: From square ice to cubic jellium. Journal of Statistical Physics, 116(1–4):3–8, August 2004. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://labs.adsabs.harvard.edu/ui/abs/2004JSP...
116....3D;   http://link.springer.com/article/10.1023/
B%3AJOSS.0000037240.68210.60.

Dyson:2004:MEF

[Dys04b] Freeman Dyson. A meeting with Enrico Fermi. Nature, 427 (6972):297, January 22, 2004. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

labs.adsabs.harvard.edu/ui/abs/2004Natur.427..297D;
http://www.nature.com/nature/journal/v427/n6972/full/
427297a.html.

Dyson:2004:BRO

[Dys04c] Freeman J. Dyson. Book review: One in a million: Debunked! ESP, Telekinesis, Other Pseudoscience by Georges Charpak and Henri Broch, translated from the French by Bart K. Holland Johns Hopkins University Press, 176 pp., $25.00. New York Review of Books, 51(5):??, March 25, 2004. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2004/mar/25/one-in-a-million/.

Dyson:2004:BRW

[Dys04d] Freeman J. Dyson. Book review: The world on a string: The Fabric of the Cosmos: Space, Time, and the Texture of Reality, by Brian Greene, Knopf, 569 pp., $28.95. New York Review of Books, 51(8):16–19, May 13, 2004. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2004/may/13/the-world-on-a-string/. See re-

sponses [?, GFD04].

Dyson:2004:F

[Dys04e] Freeman J. Dyson. Foreword. In Stapledon and McCarthy [SM04], page ?? ISBN 0-8195-6692-6 (hardcover), 0-8195-6693-4 (paperback). LCCN PR6037.T18 S7 2004. URL http:/

/www.loc.gov/catdir/toc/fy045/2004299753.html. Original

edition 1937. New edition introduced and Edited by Patrick A. McCarthy. Foreword by Freeman Dyson.

Dyson:2004:IAD

[Dys04f] Freeman J. Dyson, editor. Infinite in all directions: Gifford lectures given at Aberdeen, Scotland, April–November 1985. Perennial, New York, NY, USA, 2004. ISBN 0-06-072889-2. xx + 321 pp. LCCN Q175.3 .D97 2004. URL http://www.loc.gov/

catdir/enhancements/fy0910/2004047706-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0910/2004047706-d.
html.

Dyson:2004:TEH

[Dys04g] Freeman J. Dyson. Thought-experiments in honor of John Archibald Wheeler. In Barrow et al. [BDH04], pages 72–89. ISBN 0-521-83113-X. LCCN QC174.12 .S4 2004. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/description/cam051/2003055903.html;
 http://www.loc.gov/catdir/toc/cam051/2003055903.html.

Dyson:2005:DS

[Dys05a] Freeman Dyson. The death of a star. Nature, 438(7071): 1086, December 21, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://

labs.adsabs.harvard.edu/ui/abs/2005Natur.438.1086D;
http://www.nature.com/nature/journal/v438/n7071/full/
4381086a.html.

Dyson:2005:DU

[Dys05b] Freeman Dyson. Dynamic Universe. Nature, 435(7045): 1033, June 22, 2005. CODEN NATUAS. ISSN 0028-0836 (print), 1476-4687 (electronic). URL http://labs.adsabs.

harvard.edu/ui/abs/2005Natur.435.1033D;   http://www.
nature.com/nature/journal/v435/n7045/full/4351033a.html.

Dyson:2005:HB

[Dys05c] Freeman Dyson. Hans A. Bethe (1906–2005). Science, 308(5719): 219, April 8, 2005. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
3841346;   http://www.sciencemag.org/content/308/5719/
219.full.pdf.

Dyson:2005:HBQ

[Dys05d] Freeman Dyson. Hans Bethe and quantum electrodynamics. Physics Today, 58(10):48–50, October 2005. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/phtoad/v58/i10/p48/s1.

Dyson:2005:BRS

[Dys05e] Freeman J. Dyson. Book review: Seeing the unseen: The Fly in the Cathedral: How a Group of Cambridge Scientists Won the International Race to Split the Atom, by Brian Cathcart. Farrar, Straus and Giroux, 308 pp., $25.00. A Sense of the Mysterious: Science and the Human Spirit, by Alan Lightman. Pantheon, 211 pp., $23.00. New York Review of Books, 52(3):11– 13, February 24, 2005. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2005/feb/24/seeing-the-unseen/.

Dyson:2005:BRB

[Dys05f] Freeman J. Dyson. Book review: The bitter end: Armageddon: The Battle for Germany, 1944–1945, by Max Hastings.

Knopf, 584 pp., $30.00. The End: Hamburg 1943, by Hans Erich Nossack, translated from the German and with a foreword by Joel Agee, and with photographs by Erich Andres. University of Chicago Press, 85 pp., $20.00. New York Review of Books, 52 (7):4–6, April 28, 2005. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2005/apr/28/the-bitter-end/.

Dyson:2005:BRT

[Dys05g] Freeman J. Dyson. Book review: The tragic tale of a genius: Dark Hero of the Information Age: In Search of Norbert Wiener, the Father of Cybernetics, by Flo Conway and Jim Siegelman. Basic Books, 423 pp., $27.50. New York Review of Books, 52(12): 10–??, July 14, 2005. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2005/jul/14/the-tragic-tale-of-a-genius/.

Dyson:2005:BRWa

[Dys05h] Freeman J. Dyson. Book review: What a world! The Earth’s Biosphere: Evolution, Dynamics, and Change by Vaclav Smil. MIT Press, 346 pp., $32.95. New York Review of Books, 50(8): 4–6, May 15, 2005. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2003/may/15/what-a-world/. See response [GD03].

Dyson:2005:DI

[Dys05i] Freeman J. Dyson. The Darwinian interlude. Technology Review (M.I.T.), 108(3):27, March 1, 2005. CODEN TEREAU. ISSN

0040-1692. URL http://www.technologyreview.com/article/
403777/the-darwinian-interlude/.

Dyson:2005:HTA

[Dys05j] Freeman J. Dyson. Heretical thoughts about science and society. Frederick S. Pardee Distinguished Lecture Series, Boston University, 2005.

Dyson:2005:P

[Dys05k] Freeman J. Dyson. Preface. In Thirring [Thi05], page ?? ISBN 3-540-22212-X. LCCN QC173.4.T48 L54 2005. URL http://www.

loc.gov/catdir/enhancements/fy0663/2004108033-d.html.

Dyson:2005:WMB

[Dys05l] Freeman J. Dyson. Wise man: Perfectly Reasonable Deviations from the Beaten Track: The Letters of Richard P. Feynman, edited and with an introduction by Michelle Feynman, with a foreword by Timothy Ferris. Basic Books, 486 pp., $26.00. New York Review of Books, 52(16):??, October 20, 2005. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.

com/articles/archives/2005/oct/20/wise-man/.

Dyson:2006:MMH

[Dys06a] F. Dyson. Make me a hipporoo. New Scientist, 189(2538):36–39, February 11, 2006. CODEN NWSCAL. ISSN 0262-4079 (print), 1364-8500 (electronic).

Dyson:2006:MRF

[Dys06b] Freeman Dyson. Marshall N. Rosenbluth, 5 February 1927–28 September 2003. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 150 (2):403–407, June 2006. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic). URL http://www.jstor.org/

stable/4599001.

Dyson:2006:LAQ

[Dys06c] Freeman J. Dyson. 1951 lectures on advanced quantum mechanics second edition. arxiv.org, ??(??):1–??, August 2006. URL http:/

/arxiv.org/abs/arXiv:quant-ph/0608140;   http://hrst.
mit.edu/hrs/renormalization/dyson51-intro/;   http://
www.sns.ias.edu/~dyson/.

Dyson:2006:BRR

[Dys06d] Freeman J. Dyson. Book review: Religion from the outside: Breaking the Spell: Religion as a Natural Phenomenon by Daniel C. Dennett. Viking Penguin, 448 pp., $25.95. New York Review of Books, 53(11):4–??, June 22, 2006. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2006/jun/22/religion-from-the-outside/. See re-

sponses [DHD06, SD06].

Dyson:2006:BRW

[Dys06e] Freeman J. Dyson. Book review: Writing nature’s greatest book: The Best of All Possible Worlds: Mathematics and Destiny by Ivar Ekeland. University of Chicago Press, 207 pp., $25.00. New York Review of Books, 53(16):53–55, October 19, 2006. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2006/oct/19/writing-natures-greatest-book/.

Dyson:2006:DI

[Dys06f] Freeman J. Dyson. The Darwinian interlude. Technology Review (M.I.T.), 109(??):??, February 16, 2006. CODEN TEREAU.

ISSN 0040-1692. URL http://www.technologyreview.com/
news/405328/the-darwinian-interlude/.

Dyson:2006:ETB

[Dys06g] Freeman J. Dyson. Edward Teller’s Memoirs. In The scientist as rebel [Dys06n], pages 167–172. ISBN 1-59017-216-

7. LCCN Q158.5 .D977 2006. URL http://www.loc.gov/

catdir/enhancements/fy0665/2006022081-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0665/2006022081-d.
html;   http://www.loc.gov/catdir/toc/ecip0616/2006022081.

html. Collection of book reviews.

Dyson:2006:FI

[Dys06h] Freeman J. Dyson. A failure of intelligence. Technology Review (M.I.T.), 109(5):62–71, November 2006. CODEN TEREAU.

ISSN 0040-1692. URL http://www.technologyreview.com/
article/406789/a-failure-of-intelligence/.

Dyson:2006:FCF

[Dys06i] Freeman J. Dyson. Foreword: [Classic Feynman]. In Classic Feynman: all the adventures of a curious character [FL06], pages 5–?? ISBN 0-393-06132-9. LCCN QC16.F49 A3 2006. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/toc/ecip0515/2005018928.html.

Dyson:2006:HBQ

[Dys06j] Freeman J. Dyson. Hans Bethe and quantum electrodynamics. In Brown and Lee [BL06], pages 157–163. ISBN 981-256-609-0, 981-256-610-4 (paperback), 981-277-450-5 (e-book). LCCN QC16.B46

H36 2006. URL http://www.worldscientific.com/doi/abs/
10.1142/9789812774507_0009.

Dyson:2006:OSA

[Dys06k] Freeman J. Dyson. Oppenheimer as scientist, administrator, and poet. In The scientist as rebel [Dys06n], pages 229–242. ISBN 1-59017-216-7. LCCN Q158.5 .D977 2006. URL http://www.

loc.gov/catdir/enhancements/fy0665/2006022081-b.html;
 http://www.loc.gov/catdir/enhancements/fy0665/2006022081-
d.html;   http://www.loc.gov/catdir/toc/ecip0616/2006022081.

html. Collection of book reviews.

Dyson:2006:PIF

[Dys06l] Freeman J. Dyson. Part II: A failure of intelligence. Technology Review (M.I.T.), 109(6):??, December 2006. CODEN TEREAU.

ISSN 0040-1692. URL http://www.technologyreview.com/
news/406948/part-ii-a-failure-of-intelligence/.

Dyson:2006:SRb

[Dys06m] Freeman J. Dyson. The scientist as rebel. New York Review of Books, ??(??):??, November 14, 2006. ISSN 0028-7504 (print),

1944-7744 (electronic). URL http://www.nybooks.com/books/
imprints/collections/the-scientist-as-rebel/.

Dyson:2006:SRa

[Dys06n] Freeman J. Dyson. The scientist as rebel. New York Review Books, New York, NY, USA, 2006. ISBN 1-59017-216-7. xvi + 360 pp. LCCN Q158.5 .D977 2006. URL http://www.loc.gov/

catdir/enhancements/fy0665/2006022081-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0665/2006022081-d.
html;   http://www.loc.gov/catdir/toc/ecip0616/2006022081.

html. Collection of book reviews.

Dyson:2007:AQM

[Dys07a] Freeman Dyson. Advanced Quantum Mechanics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-270-661-5 (paperback), 981-270-622-4. xiv + 220 pp. URL http:/

/www.gbv.de/dms/bowker/toc/9789812706614.pdf;   http:
//www.worldscientific.com/worldscibooks/10.1142/6427.

Transcribed and with a preface by David Derbes.

Dyson:2007:EDE

[Dys07b] Freeman Dyson. The end of Darwinian Evolution. New Perspectives Quarterly: NPQ, 24(4):58–59, Fall 2007. ISSN 0893-7850 (print), 1540-5842 (electronic).

Dyson:2007:FI

[Dys07c] Freeman Dyson. Freedom of inquiry. Bulletin of the Atomic Scientists, 63(1):62–64, January/February 2007. CODEN BASIAP.

ISSN 0096-3402 (print), 1938-3282 (electronic). URL http://

bos.sagepub.com/content/63/1/62.full.pdf+html.

Dyson:2007:BM

[Dys07d] Freeman J. Dyson. Back matter. In Advanced Quantum Mechanics [Dys07a], pages 205–220. ISBN 981-270-661-5 (paperback),

981-270-622-4. URL http://www.worldscientific.com/doi/
abs/10.1142/9789812708212_bmatter. Transcribed and with

a preface by David Derbes.

Dyson:2007:BRD

[Dys07e] Freeman J. Dyson. Book review: The dream of scientific brotherhood: The Fellowship: Gilbert, Bacon, Harvey, Wren, Newton, and the Story of a Scientific Revolution, by John Gribbin. Overlook, 336 pp., $27.95. New York Review of Books, 54(8): 47–??, May 10, 2007. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2007/may/10/the-dream-of-scientific-brotherhood/.

Dyson:2007:BRW

[Dys07f] Freeman J. Dyson. Book review: Working for the revolution: Faust in Copenhagen: A Struggle for the Soul of Physics, by Gino Segrè. Viking, 310 pp., $25.95. New York Review of Books, 54(16): 45–47, October 25, 2007. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2007/oct/25/working-for-the-revolution/.

Dyson:2007:DT

[Dys07g] Freeman J. Dyson. The Dirac theory. In Advanced Quantum Mechanics [Dys07a], pages 5–30. ISBN 981-270-661-5 (paper-

back), 981-270-622-4. URL http://www.worldscientific.com/
doi/abs/10.1142/9789812708212_0002. Transcribed and with

a preface by David Derbes.

Dyson:2007:ET

[Dys07h] Freeman J. Dyson. Edward Teller (1908–2003). National Academy of Sciences, ??(??):1–21, ???? 2007. URL http://www.

nap.edu/html/biomems/eteller.pdf.

Dyson:2007:EQF

[Dys07i] Freeman J. Dyson. Examples of quantized field theories. In Advanced Quantum Mechanics [Dys07a], pages 61–123. ISBN 981-270-661-5 (paperback), 981-270-622-4. URL http://www.

worldscientific.com/doi/abs/10.1142/9789812708212_0005.

Transcribed and with a preface by David Derbes.

Dyson:2007:FT

[Dys07j] Freeman J. Dyson. Field theory. In Advanced Quantum Mechanics [Dys07a], pages 47–60. ISBN 981-270-661-5 (paperback), 981-

270-622-4. URL http://www.worldscientific.com/doi/abs/

10.1142/9789812708212_0004. Transcribed and with a preface by David Derbes.

Dyson:2007:Fa

[Dys07k] Freeman J. Dyson. Foreword. In Mathematics as metaphor: selected essays of Yuri I. Manin [Man07], page ?? ISBN 0-8218-4331-1. LCCN QA7 .M279 2007. With foreword by Freeman J. Dyson.

Dyson:2007:Fb

[Dys07l] Freeman J. Dyson. Foreword. In Extraordinary Knowing: Science, Skepticism, and the Inexplicable Powers of the Human Mind [May07], page ?? ISBN 0-553-80335-2 (hardcover). LCCN BF315 .M36 2007. URL http://www.loc.gov/

catdir/enhancements/fy0738/2006025661-b.html;   http:
//www.loc.gov/catdir/enhancements/fy0738/2006025661-d.
html;   http://www.loc.gov/catdir/enhancements/fy0738/
2006025661-s.html;   http://www.loc.gov/catdir/toc/ecip0619/

2006025661.html. Forewords by Freeman Dyson and Carol Gilligan.

Dyson:2007:FPS

[Dys07m] Freeman J. Dyson. Free particle scattering problems. In Advanced Quantum Mechanics [Dys07a], pages 125–143. ISBN 981-270-661-5 (paperback), 981-270-622-4. URL http://www.

worldscientific.com/doi/abs/10.1142/9789812708212_0006.

Transcribed and with a preface by David Derbes.

Dyson:2007:FM

[Dys07n] Freeman J. Dyson. Front matter. In Advanced Quantum Mechanics [Dys07a], pages i–xiii. ISBN 981-270-661-5 (paperback),

981-270-622-4. URL http://www.worldscientific.com/doi/
abs/10.1142/9789812708212_fmatter. Transcribed and with

a preface by David Derbes.

Dyson:2007:GTF

[Dys07o] Freeman J. Dyson. General theory of free particle scattering. In Advanced Quantum Mechanics [Dys07a], pages 145–182. ISBN 981-270-661-5 (paperback), 981-270-622-4.

URL http://www.worldscientific.com/doi/abs/10.1142/

9789812708212_0007. Transcribed and with a preface by David Derbes.

Dyson:2007:I

[Dys07p] Freeman J. Dyson. Introduction. In Advanced Quantum Mechanics [Dys07a], pages 1–4. ISBN 981-270-661-5 (paperback), 981-

270-622-4. URL http://www.worldscientific.com/doi/abs/

10.1142/9789812708212_0001. Transcribed and with a preface by David Derbes.

Dyson:2007:MCG

[Dys07q] Freeman J. Dyson. A many-colored glass: reflections on the place of life in the universe. University of Virginia Press, Charlottesville, VA, USA and London, UK, July 2007. ISBN 0-8139-2663-7 (hardcover). xi + 162 pp. LCCN QH501 .D97 2007. URL

http://www.loc.gov/catdir/toc/ecip077/2006103104.html.

Dyson:2007:OBF

[Dys07r] Freeman J. Dyson. Our biotech future. New York Review of Books, 54(12):4–??, July 19, 2007. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2007/jul/19/our-biotech-future/.

Dyson:2007:SSP

[Dys07s] Freeman J. Dyson. Scattering by a static potential: Comparison with experimental results. In Advanced Quantum Mechanics [Dys07a], pages 183–203. ISBN 981-270-661-5 (paperback), 981-

270-622-4. URL http://www.worldscientific.com/doi/abs/

10.1142/9789812708212_0008. Transcribed and with a preface by David Derbes.

Dyson:2007:SPB

[Dys07t]  Freeman J. Dyson.  Scattering problems and Born ap-

proximation. In Advanced Quantum Mechanics [Dys07a], pages 31–46. ISBN 981-270-661-5 (paperback), 981-270-622-

4. URL http://www.worldscientific.com/doi/abs/10.1142/

9789812708212_0003. Transcribed and with a preface by David Derbes.

Dyson:2008:SN

[Dys08a] Freeman Dyson. Science notebook. Science News (Washington, DC), 174(13):4, December 20, 2008. CODEN SCNEBK. ISSN 0036-8423 (print), 1943-0930 (electronic). URL http://

www.jstor.org/stable/20465769.

Dyson:2008:BRR

[Dys08b] Freeman J. Dyson. Book review: Rocket man: Von Braun: Dreamer of Space, Engineer of War, by Michael J. Neufeld. Knopf/Smithsonian National Air and Space Museum, 587 pp., $35.00. New York Review of Books, 55(1):8–??, January 17, 2008. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://

www.nybooks.com/articles/archives/2008/jan/17/rocket-
man/.

Dyson:2008:BRS

[Dys08c] Freeman J. Dyson. Book review: Struggle for the islands: Galápagos: The Islands That Changed the World, by Paul D. Stewart and others. Yale University Press, 240 pp., $29.95 (paper). New York Review of Books, 55(16):34–36, October 23, 2008. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/2008/oct/23/
struggle-for-the-islands/.

Dyson:2008:BRQ

[Dys08d] Freeman J. Dyson. Book review: The question of global warming: A Question of Balance: Weighing the Options on Global Warming Policies, by William Nordhaus. Yale University Press, 234 pp., $28.00. Book Review: Global Warming: Looking Beyond Kyoto, edited by Ernesto Zedillo. Yale Center for the Study of Globalization/Brookings Institution Press, 237 pp., $26.95 (paper). New York Review of Books, 55(10): 43–45, June 12, 2008. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2008/jun/12/the-question-of-global-warming/. See re-

sponses [MD08, NZSD08, PD08].

Dyson:2008:CR

[Dys08e] Freeman J. Dyson. El científico rebelde. Debate Editorial, Barcelona, España, 2008. ISBN 84-8306-767-6. 375 + 1 pp. LCCN Q 158.5 D9718.2008. Translation to Spanish by Mercedes García Garmilla of [Dys06n].

Dyson:2008:OH

[Dys08f] Freeman J. Dyson. The optimistic heretic. Physics World, 21 (1):12–13, January 2008. CODEN PHWOEW. ISSN 0953-8585 (print), 2058-7058 (electronic). URL http://iopscience.iop.

org/pwa/full/pwa-pdf/21/01/phwv21i01a21.pdf.

Dyson:2008:OL

[Dys08g] Freeman J. Dyson. Origins of life. In Nishina Memorial Lectures: creators of modern physics [Ano08], pages 71–?? ISBN 4-431-77055-0. LCCN QC71 .N615 2008. URL http://labs.adsabs.

harvard.edu/ui/abs/2008LNP...746...71D.

Dyson:2008:SR

[Dys08h] Freeman J. Dyson. The scientist as rebel. New York Review of Books, ??(??):??, September 9, 2008. ISSN 0028-7504 (print),

1944-7744 (electronic). URL http://www.nybooks.com/books/
imprints/collections/the-scientist-as-rebel-2/.

Dyson:2008:U

[Dys08i] Freeman J. Dyson. [unknown]. In Steiner et al. [SV^{+}08], pages 155–?? ISBN 1-84777-007-X. LCCN ???? URL http://www.

carcanet.co.uk/cgi-bin/indexer?product=9781847770073.

Dyson:2009:BFE

[Dys09a] Freeman J. Dyson. Birds and frogs: 2008 Einstein Lecture. Notices of the American Mathematical Society, 56(2):212–223, February 2009. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/notices/
200902/rtx090200212p.pdf.

Dyson:2009:BRL

[Dys09b] Freeman J. Dyson. Book review: Leaping into the grand unknown: The Lightness of Being: Mass, Ether, and the Unification of Forces by Frank Wilczek. Basic Books, 270 pp., $26.95. New York Review of Books, 56(6):59–61, April 9, 2009. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2009/apr/09/leaping-into-the-grand-
unknown/.

Dyson:2009:BRW

[Dys09c] Freeman J. Dyson. Book review: When science & poetry were friends: The Age of Wonder: How the Romantic Generation Discovered the Beauty and Terror of Science, by Richard Holmes.

Pantheon, 552 pp., $40.00. New York Review of Books, 56(13): 15–18, August 13, 2009. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2009/aug/13/when-science-poetry-were-friends/.

Dyson:2009:ETJ

[Dys09d] Freeman J. Dyson. Edward Teller: January 15, 1908–September 9, 2003. Biographical memoirs — National Academy of Sciences of the United States of America, pages 413–430, 2009. CODEN BMNSAC. ISBN 0-309-12148-5. ISSN 0077-2933.

Dyson:2009:VUR

[Dys09e] Freeman J. Dyson. La vie dans l’univers: réflexions d’un physicien. (French) [Life in the Universe: reflexions of a physicist]. Gallimard, Paris, France, 2009. 256 pp. LCCN ???? French translation of [Dys07q] by Stéphane Schmitt.

Dyson:2010:RDC

[Dys10a] Freeman Dyson. Radiotelepathy: direct communication from brain to brain. In Brockman [Bro10], page ?? ISBN 0-06-189967-4 (paperback). LCCN Q175.5 .T48 2010; HM901 .T45 2010.

Dyson:2010:BFM

[Dys10b] Freeman J. Dyson. Birds and frogs in mathematics and physics. Physics-Uspekhi, 53(8):825–834, ???? 2010. CODEN PHUSEY. ISSN 1063-7869 (print), 1468-4780 (electronic). English version of [Dai10].

Dyson:2010:BRS

[Dys10c] Freeman J. Dyson. Book review: Silent quantum genius: The Strangest Man: The Hidden Life of Paul Dirac, Mystic of the Atom, by Graham Farmelo. Basic Books, 539 pp., $29.95. New York Review of Books, 57(3):20–??, February 25, 2010. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://

www.nybooks.com/articles/archives/2010/feb/25/silent-

quantum-genius/. See response [GD10].

Dyson:2010:BRW

[Dys10d] Freeman J. Dyson. Book review: ‘what price glory?’: Lake Views: This World and the Universe, by Steven Weinberg. Belknap Press/Harvard University Press, 259 pp., $25.95. New York Review of Books, 57(10):8–??, June 10, 2010. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2010/jun/10/what-price-glory/.

Dyson:2010:CRC

[Dys10e] Freeman J. Dyson. Chandrasekhar’s role in 20th-Century science. Physics Today, 63(12):44–48, December 2010. CO-DEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://www.physicstoday.org/resource/1/
phtoad/v63/i12/p44_s1.

Dyson:2010:CRH

[Dys10f] Freeman J. Dyson. El científico rebelde. Ciencia. Debolsillo, Barcelona, Spain, 2010. ISBN 84-9908-146-0. 376 pp. LCCN ???? Translation to Spanish by Mercedes García Garmilla of [Dys06n].

Dyson:2010:F

[Dys10g] Freeman J. Dyson. Foreword. In Calaprice [Cal10], pages xxviii + 578. ISBN 0-691-13817-6. LCCN QC16.E5 A25 2010. URL

http://press.princeton.edu/titles/9268.html.

Dyson:2010:JAW

[Dys10h] Freeman J. Dyson. John Archibald Wheeler (9 July 1911–13 April 2008): biographical memoir. Proceedings of the American Philosophical Society held at Philadelphia for promoting useful knowledge, 154(??):126–129, March 2010. CODEN PAPCAA. ISSN 0003-049X (print), 2326-9243 (electronic).

Dyson:2010:MCG

[Dys10i] Freeman J. Dyson. Many-colored glass: reflections on the place of life in the universe. University of Virginia Press, Charlottesville, VA, USA and London, UK, 2010. ISBN 0-8139-2973-3. xi + 162 pp. LCCN ???? Page–Barbour Lectures for 2004.

Dyson:2011:BM

[Dys11a] Freeman Dyson. Back matter. In Advanced Quantum Mechanics [Dys11k], pages 205–289. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN QC174.12 .D97

2011. URL http://www.worldscientific.com/doi/abs/10.

1142/9789814383424_bmatter. Translated and transcribed by David Derbes.

Dyson:2011:BF

[Dys11b] Freeman Dyson. Birds and frogs. In Pitici [Pit11], pages 57– 78. ISBN 0-691-14841-4. LCCN ???? Foreword by William P. Thurston.

Dyson:2011:DT

[Dys11c] Freeman Dyson. The Dirac theory. In Advanced Quantum Mechanics [Dys11k], pages 5–30. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN QC174.12

.D97 2011. URL http://www.worldscientific.com/doi/abs/

10.1142/9789814383424_0002. Translated and transcribed by David Derbes.

Dyson:2011:EQF

[Dys11d] Freeman Dyson. Examples of quantized field theories. In Advanced Quantum Mechanics [Dys11k], pages 61–123. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (ebook). LCCN QC174.12 .D97 2011. URL http://www.

worldscientific.com/doi/abs/10.1142/9789814383424_0005.

Translated and transcribed by David Derbes.

Dyson:2011:FT

[Dys11e] Freeman Dyson. Field theory. In Advanced Quantum Mechanics [Dys11k], pages 47–60. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN QC174.12

.D97 2011. URL http://www.worldscientific.com/doi/abs/

10.1142/9789814383424_0004. Translated and transcribed by David Derbes.

Dyson:2011:FPS

[Dys11f] Freeman Dyson. Free particle scattering problems. In Advanced Quantum Mechanics [Dys11k], pages 125–143. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN

QC174.12 .D97 2011. URL http://www.worldscientific.com/
doi/abs/10.1142/9789814383424_0006. Translated and tran-

scribed by David Derbes.

Dyson:2011:GTF

[Dys11g] Freeman Dyson. General theory of free particle scattering. In Advanced Quantum Mechanics [Dys11k], pages 145–182. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN QC174.12 .D97 2011. URL http://www.

worldscientific.com/doi/abs/10.1142/9789814383424_0007.

Translated and transcribed by David Derbes.

Dyson:2011:I

[Dys11h] Freeman Dyson. Introduction. In Advanced Quantum Mechanics [Dys11k], pages 1–4. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN QC174.12

.D97 2011. URL http://www.worldscientific.com/doi/abs/

10.1142/9789814383424_0001. Translated and transcribed by David Derbes.

Dyson:2011:SSP

[Dys11i] Freeman Dyson. Scattering by a static potential: Comparison with experimental results. In Advanced Quantum Mechanics [Dys11k], pages 183–203. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN QC174.12

.D97 2011. URL http://www.worldscientific.com/doi/abs/

10.1142/9789814383424_0008. Translated and transcribed by David Derbes.

Dyson:2011:SPB

[Dys11j] Freeman Dyson. Scattering problems and Born approximation. In Advanced Quantum Mechanics [Dys11k], pages 31– 46. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). LCCN QC174.12 .D97 2011. URL http://www.

worldscientific.com/doi/abs/10.1142/9789814383424_0003.

Translated and transcribed by David Derbes.

Dyson:2011:AQM

[Dys11k] Freeman J. Dyson. Advanced Quantum Mechanics. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, second edition, 2011. ISBN 981-4383-40-6, 981-4383-41-4 (paperback), 981-4383-42-2 (e-book). xxv + 289 pp. LCCN QC174.12 .D97 2011. Translated and transcribed by David Derbes.

Dyson:2011:BRHb

[Dys11l] Freeman J. Dyson. Book review: How to dispel your illusions: Thinking, Fast and Slow , by Daniel Kahneman. Farrar, Straus and Giroux, 499 pp., $30.00. New York Review of Books, 58 (20):40–??, December 22, 2011. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2011/dec/22/how-dispel-your-illusions/.

Dyson:2011:BRHa

[Dys11m] Freeman J. Dyson. Book review: How we know: The Information, by James Gleick. New York Review of Books, 58(4): 8–??, March 10, 2011. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2011/mar/10/how-we-know/. See corrections [Dys11s].

Dyson:2011:BRC

[Dys11n] Freeman J. Dyson. Book review: The case for far-out possibilities: The Beginning of Infinity: Explanations that Transform the World, by David Deutsch. Viking, 487 pp., $30.00. New York Review of Books, 58(17):26–27, November 10, 2011. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http:

//www.nybooks.com/articles/archives/2011/nov/10/case-
far-out-possibilities/.

Dyson:2011:BRD

[Dys11o] Freeman J. Dyson. Book reviews: The ‘dramatic picture’ of Richard Feynman: Quantum Man: Richard Feynman’s Life in Science, by Lawrence M. Krauss. Norton, 350 pp., $24.95. Feynman, by Jim Ottaviani, with art by Leland Myrick and coloring by Hilary Sycamore. First Second, 266 pp., $29.99. New York Review of Books, 58(12):39–40, July 14, 2011. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/2011/jul/14/
dramatic-picture-richard-feynman/.

Dyson:2011:CRC

[Dys11p] Freeman J. Dyson. Chandrasekhar’s role in 20th-Century science. In Saikia and Trimble [ST11], pages x + 299. ISBN 981-4374-76-8 (hardcover), 981-4374-77-6 (e-book). LCCN QB461

.F58 2011. URL http://www.worldscientific.com/doi/abs/
10.1142/9789814374774_0002.

Dyson:2011:F

[Dys11q] Freeman J. Dyson. Foreword. In Akemann et al. [ABD11], pages vii–ix. ISBN 0-19-957400-6. LCCN QA188 .O94

2011.  URL http://www.loc.gov/catdir/enhancements/
fy1402/2011029624-b.html;   http://www.loc.gov/catdir/
enhancements/fy1402/2011029624-d.html;   http://www.loc.
gov/catdir/enhancements/fy1402/2011029624-t.html.

Dyson:2011:FS

[Dys11r] Freeman J. Dyson. The future of science. In Chiao et al. [CCL^{+}11], pages 39–54. ISBN 0-521-88239-7 (hardback).

LCCN Q162 .V48 2011. URL http://assets.cambridge.org/
97805218/82392/cover/9780521882392.jpg.

Dyson:2011:LHW

[Dys11s] Freeman J. Dyson. Letter: How we know. New York Review of Books, 58(7):74, April 28, 2011. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2011/apr/28/how-we-know/. Corrections

to [Dys11m] in response to [GD11].

Dyson:2011:NKD

[Dys11t] Freeman J. Dyson. Nicholas Kemmer: 7 December 1911–21 October 1998. Biographical Memoirs of Fellows of the Royal Society, 57(1):189–204, ???? 2011. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic). URL http://www.jstor.org/

stable/41412878.

Dyson:2011:HNR

[Dys11u] Freeman J. Dyson. Of historical note: Richard Feynman. The Institute Letter, Spring 2011. URL https://www.ias.edu/ideas/

2011/dyson-of-historical-note.

Dyson:2011:P

[Dys11v] Freeman J. Dyson. Preface. In Chiao et al. [CCL^{+}11], pages xxi– xxii. ISBN 0-521-88239-7 (hardback). LCCN Q162 .V48 2011.

URL http://assets.cambridge.org/97805218/82392/cover/
9780521882392.jpg.

Dyson:2012:F

[Dys12a] Freeman Dyson. Foreword. In Gold and Mitton [GM12], page ?? ISBN 3-642-27588-5 (e-book). ISSN 0067-0057. LCCN QB36.G653 G65 2012.

Dyson:2012:FRM

[Dys12b] Freeman Dyson. Foreword: Recreational mathematics. In Pitici [Pit12], pages xi–xvi. ISBN 0-691-15315-9 (paperback). LCCN QA8.6.

Dyson:2012:PGC

[Dys12c] Freeman Dyson. Partitions and the grand canonical ensemble. The Ramanujan Journal, 29(1–3):423–429, 2012. CODEN RA-JOF9. ISSN 1382-4090 (print), 1572-9303 (electronic).

Dyson:2012:SRB

[Dys12d] Freeman Dyson. Science on the rampage: Physics on the Fringe: Smoke Rings, Circlons, and Alternative Theories of Everything by Margaret Wertheim, Walker, 323 pp., $27.00. New York Review of Books, 59(6):??, April 5, 2012. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/2012/apr/05/
science-rampage-natural-philosophy/.

Dyson:2012:WCY

[Dys12e] Freeman Dyson. What can you really know? Why Does the World Exist?: An Existential Detective Story by Jim Holt, Liveright, 307 pp., $27.95. New York Review of Books, 59(17):??, November 8, 2012. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2012/nov/08/what-can-you-really-know/.

Dyson:2012:BRS

[Dys12f] Freeman J. Dyson. Book review: Science on the rampage: Physics on the Fringe: Smoke Rings, Circlons, and Alternative Theories of Everything, by Margaret Wertheim, Walker, 323 pp., $27.00. New York Review of Books, 59(6):38–39, April 5, 2012. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2012/apr/05/science-rampage-natural-philosophy/.

Dyson:2012:BRW

[Dys12g] Freeman J. Dyson. Book review: What can you really know? Why Does the World Exist?: An Existential Detective Story, by Jim Holt, Liveright, 307 pp., $27.95. New York Review of Books, 59(17):18–??, November 8, 2012. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2012/nov/08/what-can-you-really-know/. See crit-

ical commentary [Rus12].

Dyson:2012:FY

[Dys12h] Freeman J. Dyson. Fifty years. Essay from the fiftieth anniversary of the Association of Amateur Astronomers of Princeton, Sidereal Times., November 2012.

Dyson:2012:SMD

[Dys12i] Freeman J. Dyson. Is science mostly driven by ideas or by tools? Science, 338(6113):1426–1427, December 14, 2012. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (elec-

tronic). URL http://www.jstor.org/stable/23332361;
http://www.sciencemag.org/content/338/6113/1426.full.
pdf.

Dyson:2013:EG

[Dys13a] Freeman Dyson. From Eros to Gaia. Random House, New York, NY, USA, 2013. ISBN 1-299-68439-4 (e-book). ???? pp. LCCN

???? URL http://lib.myilibrary.com?id=499689.

Dyson:2013:HUW

[Dys13b] Freeman Dyson. How to be an underdog, and win: David and Goliath: Underdogs, Misfits, and the Art of Battling Giants by Malcolm Gladwell Little, Brown, 305 pp., $29.00. New York Review of Books, 60(18):??, November 21, 2013. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2013/nov/21/how-to-be-underdog-and-

win/. See comments [PD13, Dys14d].

Dyson:2013:GD

[Dys13c] Freeman Dyson. Is a graviton detectable? International Journal of Modern Physics. A, Particles and Fields, Gravitation, Cosmology, 28(??):1330041:1–1330041:14, 2013. CODEN IMPAEF. ISSN

0217-751X. URL http://www.worldscientific.com/doi/abs/
10.1142/S0217751X1330041X.

Dyson:2013:WTJ

[Dys13d] Freeman Dyson. A walk through Johnny von Neumann’s garden. Notices of the American Mathematical Society, 60(2):154– 161, February 2013. CODEN AMNOAN. ISSN 0002-9920 (print),

1088-9477 (electronic). URL http://www.ams.org/notices/

201302/rnoti-p154.pdf. Talk given at Brown University, Providence, Rhode Island, May 4, 2010.

Dyson:2013:BRO

[Dys13e] Freeman J. Dyson. Book review: Oppenheimer: The shape of genius: Robert Oppenheimer: A Life Inside the Center, by Ray Monk, Doubleday, 825 pp., $37.50. New York Review of Books, 60(13):18–19, August 15, 2013. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2013/aug/15/oppenheimer-shape-genius/.

Dyson:2013:NSH

[Dys13f] Freeman J. Dyson. The need for sustainable heretics. In Madhavan et al. [MOG^{+}13], pages 71–76. ISBN 1-4614-4348-2, 1-4614-4349-0 (e-book). LCCN HC79.E5 P6765 2013. URL http://www.

loc.gov/catdir/enhancements/fy1316/2012947954-b.html;
 http://www.loc.gov/catdir/enhancements/fy1316/2012947954-
d.html;   http://www.loc.gov/catdir/enhancements/fy1316/

2012947954-t.html. Foreword by Michael Spence. Editorials by Klaus Schwab, Robert Rubin, and George Whitesides. Afterword by M. S. Swaminathan.

Dyson:2014:CBB

[Dys14a] Freeman Dyson. The case for blunders: Brilliant Blunders: From Darwin to Einstein — Colossal Mistakes by Great Scientists That Changed Our Understanding of Life and the Universe by Mario Livio, Simon and Schuster, 341 pp., $26.00. New York Review of Books, 61(4):??, March 6, 2014. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://

www.nybooks.com/articles/archives/2014/mar/06/darwin-
einstein-case-for-blunders/. See comments [KAD14, TD14].

Dyson:2014:GDa

[Dys14b] Freeman Dyson. Is a graviton detectable? In Phua et al. [PKpCH14], pages 1–14. ISBN 981-4590-10-X (hardcover), 981-4590-70-3 (paperback), 981-4590-12-6 (e-book). LCCN

???? URL http://www.worldscientific.com/doi/abs/10.
1142/9789814590112_0001.

Dyson:2014:GDb

[Dys14c] Freeman Dyson. Is a graviton detectable? In Jensen [Jen14], pages 670–682. ISBN 981-4449-23-7 (hardcover), 981-4449-24-5 (e-book). LCCN QC19.2 .I58 2012; HC445.8.C47 Y66

2014. URL http://www.worldscientific.com/doi/abs/10.
1142/9789814449243_0071.

Dyson:2014:TS

[Dys14d] Freeman Dyson. Topping the secret. New York Review of Books, 61(1):??, January 9, 2014. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2014/jan/09/topping-secret/. See [Dys13b].

Dyson:2015:EJP

[Dys15a] Freeman Dyson. Einstein as a Jew and a philosopher: Einstein: His Space and Times by Steven Gimbel, Yale University Press, 191 pp., $25.00. New York Review of Books, 62 (8):??, May 7, 2015. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2015/may/07/albert-einstein-jew-and-philosopher/. See

comments [SD15, WD15].

Dyson:2015:CW

[Dys15b] Freeman Dyson. I could be wrong. In Brockman [Bro15], page 79. ISBN 0-06-242565-X (paperback). LCCN Q335 .W445 2015.

Dyson:2015:SSG

[Dys15c] Freeman Dyson. Scientist, spy, genius: Who was Bruno Pontecorvo? Half-Life: The Divided Life of Bruno Pontecorvo, Physicist or Spy by Frank Close, Basic Books, 378 pp., $29.99. New York Review of Books, 62(4):??, March 5, 2015. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/2015/mar/05/
scientist-spy-genius-bruno-pontecorvo/.

Dyson:2015:S

[Dys15d] Freeman Dyson. Stardust. New York Review of Books, 62(??): ??, May 27, 2015. ISSN 0028-7504 (print), 1944-7744 (electronic).

URL http://www.nybooks.com/blogs/gallery/2015/may/27/
star-dust/.

Dyson:2015:DES

[Dys15e] Freeman J. Dyson. Dreams of earth and sky. New York Review Books, New York, NY, USA, 2015. ISBN 1-59017-854-8, 1-59017-855-6. xiii + 298 pp. LCCN Q172.5.S47

D97 2015. URL http://www.nybooks.com/books/imprints/
collections/dreams-of-earth-and-sky/.

Dyson:2016:UPL

[Dys16] Freeman Dyson. Unfashionable pursuits, leaps in the dark, and detecting gravitational waves: How the recent discovery of spacetime ripples began as a risky venture. The Institute Letter, ??(??):

1, 3, Spring 2016. URL https://www.ias.edu/ideas/2016/
dyson-gravitational-waves.

Dyson:2018:MPA

[Dys18] Freeman J. Dyson. Maker of patterns: an autobiography through letters. Liveright Publishing Corporation, New York, NY, USA, 2018. ISBN 0-87140-386-2 (hardcover). xvi + 400 pp. LCCN QC16.D95 A3 2018.

Dyson:2020:NB

[Dys20] Freeman Dyson. 1961: The neutron bomb. Bulletin of the Atomic Scientists, 76(6):371–373, 2020. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). URL http:

//www.tandfonline.com/doi/full/10.1080/00963402.2020.
1847491.

Dyson:2022:FJD

[Dys22] George Dyson. Freeman John Dyson. 15 December 1923–28 February 2020. Biographical Memoirs of Fellows of the Royal Society, 73:197–226, June 2022. CODEN BMFRA3. ISSN 0080-4606 (print), 1748-8494 (electronic).

Dolgounitcheva:2011:DDO

[DZO11] O. Dolgounitcheva, V. G. Zakrzewski, and J. V. Ortiz. Delocalization of Dyson orbitals in F^{−} (H_{2}O) and Cl^{−} (H_{2}O). International Journal of Quantum Chemistry, 111(7–8):1701–1708, June/July 2011. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Dolgounitcheva:2013:APS

[DZO13] O. Dolgounitcheva, V. G. Zakrzewski, and J. V. Ortiz. Assignment of photoelectron spectra of halide–water clusters: Contrasting patterns of delocalization in Dyson orbitals. Journal of Chemical Physics, 138(16):164317, 2013. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.

aip.org/link/?JCP/138/164317/1.

Eichmann:2012:UDH

[EF12] Gernot Eichmann and Christian S. Fischer. Unified description of hadron-photon and hadron-meson scattering in the Dyson– Schwinger approach. Physical Review D (Particles and Fields), 85(??):034015, February 9, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.85.034015.

Eichmann:2013:NCS

[EF13] Gernot Eichmann and Christian S. Fischer. Nucleon Compton scattering in the Dyson–Schwinger approach. Physical Review D (Particles and Fields), 87(??):036006, February 8, 2013. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.87.036006.

Eremenko:2011:QES

[EG11] Alexandre Eremenko and Andrei Gabrielov. Quasi-exactly solvable quartic: elementary integrals and asymptotics. Journal of Physics A: Mathematical and Theoretical, 44(31):312001, 2011. CODEN JPAMB5. ISSN 1751-8113 (print), 1751-8121 (elec-

tronic). URL http://stacks.iop.org/1751-8121/44/i=31/a=

312001. See [Dys44b].

Ehrlich:1975:BSY

[Ehr75] Paul R. Ehrlich. The benefits of saying YES! Bulletin of the Atomic Scientists, 31(7):49–51, September 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys75b].

Edelman:2022:CDC

[EJ22] Alan Edelman and Sungwoo Jeong. On the Cartan decomposition for classical random matrix ensembles. Journal of Mathematical Physics, 63(6):061705, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https:/

/pubs.aip.org/aip/jmp/article/63/6/061705/2845972/On-
the-Cartan-decomposition-for-classical-random. Special

collection in honor of Freeman Dyson.

Elliot:1986:PVN

[Ell86] David C. Elliot. Project Vista and nuclear weapons in Europe. International Security, 11(1):163–183, Summer 1986. CODEN ????

ISSN 0162-2889 (print), 1531-4804 (electronic). URL http://

muse.jhu.edu/journals/international_security/summary/
v011/11.1.elliot.html;   http://www.jstor.org/stable/
2538879.

Eichmann:2012:NDE

[EN12] G. Eichmann and D. Nicmorus. Nucleon to delta electromagnetic transition in the Dyson–Schwinger approach. Physical Review D (Particles and Fields), 85(??):093004, May 9, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.85.093004.

Epstein:1955:DFD

[Eps55] Saul T. Epstein. Derivation of the Feynman–Dyson rules from time-independent theory. Physical Review, 98(??):196–??, April 1, 1955. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.98.196.

Epple:2007:CSD

[ERS07] D. Epple, H. Reinhardt, and W. Schleifenbaum. Confining solution of the Dyson–Schwinger equations in Coulomb gauge. Physical Review D (Particles and Fields), 75(??):045011, February 20, 2007. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.75.045011.

Erdos:1951:APSe

[ES51] Paul Erdős and G. Szegő. Advanced problems and solutions: 4388. American Mathematical Monthly, 58(9):639–640, November 1951. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UTB^{+}50].

Erdos:1950:APSc

[EW50] Paul Erdős and J. E. Wilkins, Jr. Advanced problems and solutions: 4319. American Mathematical Monthly, 57(5):346, May 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wil48].

Fadiman:1990:LPR

[Fad90] Clifton Fadiman, editor. Living philosophies: the reflections of some eminent men and women of our time. Doubleday, New York, NY, USA, 1990. ISBN 0-385-24880-6. xii + 290 pp. LCCN P85.S74 Z68 1990; BJ1571 .L635 1990.

Fairchild:1994:STR

[Fai94] R. P. Fairchild. Science in trouble + response to Freeman Dyson article. American Scholar, 63(1):156–157, ???? 1994. ISSN 0003-0937 (print), 2162-2892 (electronic).

Farquhar:1991:CFP

[Far91] I. E. Farquhar. Comment on “Feynman’s proof of the Maxwell equations,” by Freeman J. Dyson [Am. J. Phys. 58, 209–211 (1990)]. American Journal of Physics, 59(1):87, 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL

http://link.aip.org/link/?AJP/59/87/1.

Feynman:1993:LET

[FD93] Joan Feynman and Freeman Dyson. Letters to the editor: The truth about Feynman’s father. Physics Today, 46(5):91, May 1993. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699

(electronic). URL http://link.aip.org/link/?PTO/46/91/1.

Dyson retracts a mistaken in a remark about the father of Joan and Richard Feynman that he made in his review [Dys92b].

Firestone:2007:LEO

[FD07] Raymond A. Firestone and Freeman J. Dyson. Letter to the Editor: ‘our biotech future’. New York Review of Books, 54(15): 49, October 11, 2007. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2007/oct/11/our-biotech-future/. Response to [Dys07r].

Federbush:1975:NASb

[Fed75] Paul Federbush. A new approach to the stability of matter problem. II. Journal of Mathematical Physics, 16(3):706–709, March 1975. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v16/i3/p706_s1.

Feder:2000:DGR

[Fed00] Toni Feder. Dyson gets religion prize. Physics Today, 53(5): 53, May 2000. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://www.physicstoday.org/
resource/1/phtoad/v53/i5/p53_s1.

Feingold:2003:NNC

[Fei03] Mordechai Feingold. ‘A new Newton’? (comment on a review by Freeman Dyson). New York Review of Books, 50(15): 50, October 9, 2003. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2003/oct/09/a-new-newton/. See [Dys03b].

Feynman:2000:PFT

[Fey00] Richard P. Feynman. The pleasure of finding things out: the best short works of Richard P. Feynman. Perseus Publishers, Cambridge, MA, USA, 2000. ISBN 0-7382-0349-1 (paperback). xvi + 270 pp. LCCN Q171 .F385 1999. Edited by Jeffrey Robbins, and foreword by Freeman Dyson.

Ferris:1991:WTP

[FF91] Timothy Ferris and Clifton Fadiman, editors. The world treasury of physics, astronomy, and mathematics. Little, Brown and Co., Boston, MA, USA, 1991. ISBN 0-316-28129-8. xv + 859 pp. LCCN QC71 .W67 1991. Foreword by Clifton Fadiman.

Field:1979:RVB

[Fie79] G. B. Field. Recollections and visions: Book review: Disturbing the Universe, by Freeman Dyson. Science, 206(4419):675–676, November 9, 1979. CODEN SCIEAS. ISSN 0036-8075 (print),

1095-9203 (electronic). URL http://www.jstor.org/stable/
1749044.

Finkbeiner:2006:JSH

[Fin06] Ann K. Finkbeiner. The Jasons: the secret history of science’s postwar elite. Viking, New York, NY, USA, 2006. ISBN 0-670-03489-4. xxx + 304 pp. LCCN Q141 .F536 2006. URL http:

//en.wikipedia.org/wiki/JASON_%28advisory_group%29;
http://www.aip.org/aip/writing/books08.html;   http:/
/www.loc.gov/catdir/enhancements/fy0719/2005043471-b.
html;   http://www.loc.gov/catdir/enhancements/fy0719/
2005043471-d.html.

Finkbeiner:2017:SNL

[Fin17] Ann Finkbeiner. Spaceflight: Near-light-speed mission to Alpha Centauri. Scientific American, 316(3):30–37, March 2017. CO-DEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (electronic).

URL http://www.nature.com/scientificamerican/journal/
v316/n3/full/scientificamerican0317-30.html;   http:/
/www.nature.com/scientificamerican/journal/v316/n3/pdf/
scientificamerican0317-30.pdf.

Fischer:2010:HQE

[Fis10] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger]. Herbig, München, Germany, 2010. ISBN 3-7766-2643-7. 350 pp. LCCN ????

Fischer:2012:HQE

[Fis12] Ernst Peter Fischer. Die Hintertreppe zum Quantensprung: die Erforschung der kleinsten Teilchen; von Max Planck bis Anton Zeilinger. (German) [The staircase to the quantum leap: the study of the smallest particles from Max Planck to Anton Zeilinger], volume 19406 of Fischer. Fischer-Taschenbuch-Verlag, Frankfurt am Main, Germany, 2012. ISBN 3-596-19406-7. 350 pp. LCCN ????

Fyodorov:1997:AHR

[FKS97] Yan V. Fyodorov, Boris A. Khoruzhenko, and Hans-Jürgen Sommers. Almost Hermitian random matrices: Crossover from Wigner–Dyson to Ginibre eigenvalue statistics. Physical Review Letters, 79(??):557–??, July 28, 1997. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http:

//link.aps.org/doi/10.1103/PhysRevLett.79.557.

Feynman:2006:CFA

[FL06] Richard Phillips Feynman and Ralph Leighton. Classic Feynman: all the adventures of a curious character. W. W. Norton & Co., New York, NY, USA, 2006. ISBN 0-393-06132-9. x + 511 pp. LCCN QC16.F49 A3 2006. URL ftp://

uiarchive.cso.uiuc.edu/pub/etext/gutenberg/;   http:/
/www.loc.gov/catdir/toc/ecip0515/2005018928.html.

Fischer:2009:CDT

[FM09] Christian S. Fischer and Jens A. Mueller. Chiral and deconfinement transition from Dyson–Schwinger equations. Physical Review D (Particles and Fields), 80(??):074029, October 23, 2009.

CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.80.074029.

Forrester:2021:CEP

[FM21] P. J. Forrester and G. Mazzuca. The classical β ensembles with β proportional to 1/N: From loop equations to Dyson’s disordered chain. Journal of Mathematical Physics, 62(7):073505, July 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Foldy:1978:ESW

[Fol78] Leslie L. Foldy. Electrostatic stability of Wigner and Wigner– Dyson lattices. Physical Review B (Solid State), 17(??):4889–??, June 15, 1978. CODEN PLRBAQ. ISSN 0556-2805. URL http:

//link.aps.org/doi/10.1103/PhysRevB.17.4889.

Foldy:1980:NIW

[Fol80] L. L. Foldy. New instabilities in Wigner–Dyson lattices. Physical Review B: Condensed Matter and Materials Physics, 22(??):4992– ??, November 15, 1980. CODEN PRBMDO. ISSN 1098-0121.

URL http://link.aps.org/doi/10.1103/PhysRevB.22.4992.

Forner:1996:PIS

[För96] Wolfgang Förner. The properties of the iterative solution of the inverse Dyson equation for the calculation of correlation corrected band structures of polymers. Journal of Computational Physics, 125(2):477–487, May 1996. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0021999196901096.

Forman:2002:ATE

[For02] Paul Forman. Against transcendence: essays on the physical significance of modern culture. Cambridge University Press, Cambridge, UK, 2002. ISBN 0-521-43020-8, 0-521-43615-X. 400 pp. LCCN ????

Forrester:2021:DDL

[For21] Peter J. Forrester. Dyson’s disordered linear chain from a random matrix theory viewpoint. Journal of Mathematical Physics, 62 (10):103302, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Fukagawa:1994:TJM

[FR94] Hidetoshi Fukagawa and John F. Rigby. Traditional Japanese mathematics problems of the 18th and 19th Centuries. Morikita Shuppan Co., Tokyo, Japan, 1994. ISBN ???? ???? pp. LCCN ????

Fukagawa:2002:TJM

[FR02] Hidetoshi Fukagawa and John F. Rigby. Traditional Japanese mathematics problems of the 18th and 19th Centuries. SCT Publ., Singapore, 2002. ISBN 981-04-2759-X. xvi + 191 pp. LCCN ????

Frankenberry:2008:FST

[Fra08] Nancy Frankenberry, editor. The faith of scientists in their own words. Princeton University Press, Princeton, NJ, USA, 2008. ISBN 0-691-13487-1 (hardcover), 1-4008-2980-1 (e-book). xviii + 523 pp. LCCN BL241 .F358 2008.

Frei:1984:BRN

[Fre84] Daniel Frei. Book review: A non-nuclear agenda: Weapons and Hope. Science, 225(4665):918, August 31, 1984. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.jstor.org/stable/1693945.

Freeman:1992:EGD

[Fre92] Dyson Freeman. De Eros a Gaia o dilema ético da civilização em face da tecnologia. (Portuguese) [From Eros to Gaia: the ethical dilemma of civilization in face of technology]. Best Seller, São Paulo, SP, Brazil, 1992. ISBN 85-7123-466-3. 413 pp. LCCN ???? Translation to Portuguese by Luiz Fernando Martins Estevez of [Dys92a].

French:2008:MPM

[Fre08] A. P. French. In memoriam: Philip Morrison. Physics in Perspective (PIP), 10(1):110–122, March 2008. CODEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic). URL http://

link.springer.com/article/10.1007/s00016-007-0343-5.

Fan:1952:APSa

[FZ52] Ky Fan and H. S. Zuckerman. Advanced problems and solutions: Solutions: 4399. American Mathematical Monthly, 59(1):48–49, January 1952. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gal50].

Gale:1950:APS

[Gal50] David Gale. Advanced problems and solutions: Problems for solution: 4395. American Mathematical Monthly, 57(5):342–343, May 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BS51, DC52, FZ52, Gal52, WL51].

Gale:1952:APS

[Gal52] David Gale. Advanced problems and solutions: 4395. American Mathematical Monthly, 59(1):46–47, January 1952. CODEN AM-MYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gal50].

Gamow:1967:DGC

[Gam67a] George Gamow. Does gravity change with time? Proceedings of the National Academy of Sciences of the United States of America, 57(2):187–193, February 15, 1967. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

adsabs.harvard.edu/abs/1967PNAS...57..187G;   http://
www.jstor.org/stable/pdfplus/57929.pdf.

Gamow:1967:VEC

[Gam67b] George Gamow. Variability of elementary charge and quasistellar objects. Physical Review Letters, 19(16):913–914, October 16, 1967. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114

(electronic), 1092-0145. URL http://prl.aps.org/abstract/
PRL/v19/i16/p913_1.

Gamow:1968:NCN

[Gam68] George Gamow. Numerology of the constants of nature. Proceedings of the National Academy of Sciences of the United States of America, 59(2):313–318, February 15, 1968. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic). URL http://

www.jstor.org/stable/pdfplus/58638.pdf.

Ganley:2000:FJD

[Gan00] Susan Ganley. Freeman J. Dyson to receive Templeton Prize for Progress in Religion. The Chronicle of Higher Education, ??(??):??, March 22, 2000. ISSN 0009-5982 (print), 1931-1362

(electronic). URL http://chronicle.com/article/Freeman-
J-Dyson-to-Receive/105102/.

Garratt:1961:PSS

[Gar61] Arthur Garratt, editor. Penguin science survey, Part 1. Pelican books. Penguin Books, Harmondsworth, UK, 1961. ???? pp. LCCN Q171 .P36.

Greenwood:1948:APS

[GB48] J. A. Greenwood and Robert Breusch. Advanced problems and solutions: 4246. American Mathematical Monthly, 55(8):512–513, October 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Thé47].

Gritsenko:2003:PIE

[GBB03] O. V. Gritsenko, B. Braida, and E. J. Baerends. Physical interpretation and evaluation of the Kohn–Sham and Dyson components of the epsilon — I relations between the Kohn–Sham orbital energies and the ionization potentials. Journal of Chemical Physics, 119(4):1937–1950, 2003. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.

aip.org/link/?JCP/119/1937/1.

Gornitz:1991:RSB

[GBS91] Th. Görnitz, H. J. Blome, and E. Schrüfer. Rezensionen: Stonier: Information and the Internal Structure of the Universe/Kafatos u. Nadeau: The Conscious Universe/Dyson: Zeit ohne Ende: Physik und Biologie in einem offenen Universum/Scholz u. Ricolfi: Thermal Sensors. Physikalische Blätter, 47(8):775–776, August 1991. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19910470811/abstract.

Georgopulos:1981:UDM

[GC81] Peter D. Georgopulos and Harry S. Camarda. Use of the Dyson– Mehta Δ_{3} statistic as a test of missing levels. Physical Review C (Nuclear Physics), 24(??):420–??, August 1, 1981. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL

http://link.aps.org/doi/10.1103/PhysRevC.24.420.

Galvin:2003:LEM

[GD03] Cyril Galvin and Freeman J. Dyson. Letter to the Editor: Measuring the sea. New York Review of Books, 50(13):57, August 14, 2003. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2003/aug/14/measuring-the-sea/. In response to [Dys05h].

Gaynes:2005:LEB

[GD05] Martin Gaynes and Freeman J. Dyson. Letter to the Editor: ‘the bitter end’. New York Review of Books, 52(16):55, October 20, 2005. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2005/oct/20/the-bitter-end/. In response to [Dys05f].

Gluck:2010:DMA

[GD10] Malcolm Gluck and Freeman J. Dyson. Dirac’s model airplane. New York Review of Books, 57(5):46, March 25, 2010. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://

www.nybooks.com/articles/archives/2010/mar/25/diracs-

model-airplane/. In response to [Dys10c].

Gezerlis:2011:LHW

[GD11] Alex Gezerlis and Freeman J. Dyson. Letter: How we know. New York Review of Books, 58(6):85, April 7, 2011. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2011/apr/07/how-we-know/. Comments

on [Dys11m].

Gromyko:2004:RJR

[GDH^{+}04] A. Gromyko, F. Dyson, K. C. Hart, P. H. Nitze, A. M. Smith, S. L. Rearden, M. Moore, J. Reston, R. Aron, and A. Mill. Remembering J. Robert Oppenheimer. Bulletin of the Atomic Scientists, 60(2):37–??, March/April 2004. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Green:2004:FCC

[GF04] Brian Green and Solomon Feferman. The ‘fabric of the cosmos’ (comment on a review by Freeman Dyson). New York Review of Books, 51(12):61, July 15, 2004. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2004/jul/15/the-fabric-of-the-cosmos/.

Greene:2004:LEF

[GFD04] Brian Greene, Solomon Feferman, and Freeman J. Dyson. Letter to the Editor: ‘the fabric of the cosmos’. New York Review of Books, 51(12):60, July 15, 2004. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2004/jul/15/the-fabric-of-the-cosmos/. In re-

sponse to [Dys04d].

Goecke:2011:HLL

[GFW11] Tobias Goecke, Christian S. Fischer, and Richard Williams. Hadronic light-by-light scattering in the muon g − 2: a Dyson– Schwinger equation approach. Physical Review D (Particles and Fields), 83(??):094006, May 4, 2011. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.

83.094006. See erratum [GFW12].

Goecke:2012:EHL

[GFW12] Tobias Goecke, Christian S. Fischer, and Richard Williams. Erratum: Hadronic light-by-light scattering in the muon g − 2: a Dyson–Schwinger equation approach [Phys. Rev. D 83, 094006 (2011)]. Physical Review D (Particles and Fields), 86(??): 099901, November 7, 2012. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.86.099901. See

[GFW11].

Guevara-Garcia:2007:EBE

[GGMO07] Alfredo Guevara-Garcia, Ana Martinez, and J. V. Ortiz. Electron binding energies and Dyson orbitals of Al_{5}O_{m}^{−} (m = 3, 4, 5 ) and Al_{5}O_{5}H_{2}^{−}. Journal of Chemical Physics, 127(23):234302, 2007. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (elec-

tronic). URL http://link.aip.org/link/?JCP/127/234302/1.

Gingerich:1975:NFA

[Gin75] Owen Gingerich. New frontiers in astronomy. Readings from Scientic American. W. H. Freeman, New York, NY, USA, 1975. ISBN 0-7167-0519-2, 0-7167-0520-6 (paperback). 369 pp. LCCN QB51 .F74 1975.

Geyer:1982:MSI

[GL82] H. B. Geyer and S. Y. Lee. Microscopic structure of an interacting boson model in terms of the Dyson boson mapping. Physical Review C (Nuclear Physics), 26(??):642–??, August 1, 1982. CO-DEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

URL http://link.aps.org/doi/10.1103/PhysRevC.26.642.

Gleick:1992:GLS

[Gle92] James Gleick. Genius: The Life and Science of Richard Feynman. Pantheon Books, New York, NY, USA, 1992. ISBN 0-679-40836-3. x + 531 pp. LCCN QC16.F49G54 1992. US$27.50.

Goldstein:2022:SSO

[GLS22] S. Goldstein, J. L. Lebowitz, and E. R. Speer. Stationary states of the one-dimensional discrete-time facilitated symmetric exclusion process. Journal of Mathematical Physics, 63(8):083301, August 2022. CODEN JMAPAQ. ISSN 0022-2488 (print),

1089-7658 (electronic). URL https://pubs.aip.org/aip/jmp/
article/63/8/083301/2846043/Stationary-states-of-the-

one-dimensional-discrete. Special collection in honor of Freeman Dyson.

Gorantla:2021:MVF

[GLSS21] Pranay Gorantla, Ho Tat Lam, Nathan Seiberg, and Shu-Heng Shao. A modified Villain formulation of fractons and other exotic theories. Journal of Mathematical Physics, 62(10):102301, October 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Gold:2012:TBW

[GM12] Thomas Gold and Simon Mitton, editors. Taking the back off the watch: a personal memoir, volume 381 of Astrophysics and space science library. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 3-642-27588-5 (ebook). ISSN 0067-0057. xvii + 232 pp. LCCN QB36.G653 G65 2012.

Godina:1998:ACU

[GMO98] J. J. Godina, Y. Meurice, and M. B. Oktay. Accurate checks of universality for Dyson’s hierarchical model. Physical Review D (Particles and Fields), 57(??):R6581–??, June 1, 1998. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.57.R6581.

Godina:1999:HAC

[GMO99] J. J. Godina, Y. Meurice, and M. B. Oktay. High-accuracy calculations of the critical exponents of Dyson’s hierarchical model. Physical Review D (Particles and Fields), 59(??):096002, April 2, 1999. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.59.096002.

Godina:2000:HBS

[GMO00] J. J. Godina, Y. Meurice, and M. B. Oktay. Hyperscaling in the broken symmetry phase of Dyson’s hierarchical model. Physical Review D (Particles and Fields), 61(??):114509, May 9, 2000. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.61.114509.

Godina:2003:UAC

[GMO03] J. J. Godina, Y. Meurice, and M. B. Oktay. Universality: Accurate checks in Dyson’s hierarchical model. AIP Conference Proceedings, 670(1):176–183, 2003. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. URL http://

link.aip.org/link/?APC/670/176/1.

Godina:1998:GPC

[GMON98] J. J. Godina, Y. Meurice, M. B. Oktay, and S. Niermann. Guide to precision calculations in Dyson’s hierarchical scalar field theory. Physical Review D (Particles and Fields), 57(??):6326–??, May 15, 1998. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.57.6326.

Gochev:1976:DMT

[Goc76] I. G. Gochev. Dyson–Maleev transformation and three-magnon problem for an s = 1/2 ferromagnet. Physica Status Solidi. B, Basic Research, 74(1):311–316, March 1, 1976. CODEN PSSBBD. ISSN 0370-1972 (print), 1521-3951 (electronic).

Gold:1999:DHB

[Gol99] Thomas Gold. The Deep Hot Biosphere. Copernicus, New York, NY, USA, 1999. ISBN 0-387-98546-8 (hardcover). xiv + 235 pp. LCCN TN870.5 .G66 1999. Foreword by Freeman Dyson.

Goldstein:2008:NPC

[Gol08] Evan R. Goldstein. A noted physicist’s contrarian view of global warming. The Chronicle of Higher Education, ?? (??):??, June 20, 2008. ISSN 0009-5982 (print), 1931-1362

(electronic). URL http://chronicle.com/article/A-Noted-
Physicists-Contrarian/16289/.

Gomer:1984:BRB

[Gom84] Robert Gomer. Book review: Weapons and Hope, by Freeman Dyson. Bulletin of the Atomic Scientists, 40(8):36–40, October 1984. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Good:1970:SPC

[Goo70] I. J. Good. Short proof of a conjecture by Dyson. Journal of Mathematical Physics, 11(6):1884, June 1970. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v11/i6/p1884_s1.

Grodzins:1963:AAS

[GR63] Morton Grodzins and Eugene I. Rabinowitch, editors. The Atomic Age: Scientists in National and World Affairs. Articles from the Bulletin of the Atomic Scientists 1945–1962. Basic Books, New York, NY, USA, 1963. xviii + 616 pp. LCCN D842 .B78. With the assistance of Harvey Flaumenhaft and Lois Gradner.

Garfield:2009:GH

[GR09] Bob Garfield and Joe Romm. Getting heated: Freeman Dyson on climate change. Radio broadcast, April 10, 2009. URL

http://www.onthemedia.org/2009/apr/10/getting-heated/
transcript/.

Green:1981:BRW

[Gre81] Martin Green. Book review: Writing about science: Broca’s Brain: Reflections on the Romance of Science, by Carl Sagan. Cosmos by Carl Sagan. The Dancing Wu-Li Masters: An Overview of the New Physics by Gary Zukav. Experiencing Science: Profiles in Discovery by Jeremy Bernstein. The Eighth Day of Creation: The Makers of the Revolution in Biology by Horace F. Judson. Gödel, Escher, Bach: A Metaphorical Fugue on Minds and Machines in the Spirit of Lewis Carroll by Douglas Hofstadter. The Lives of a Cell: Notes of a Biology Watcher by Lewis Thomas. Disturbing the Universe by Freeman Dyson. College English, 43(6):569–577, October 1981. CODEN ???? ISSN 0010-0994 (print), 2161-8178 (electronic). URL http://www.jstor.

org/stable/376994.

Garola:1985:GVD

[GS85] Claudio Garola and Luigi Solombrino. Generalized FS×W versus Dyson’s classification of irreducible representations. Journal of Mathematical Physics, 26(8):1889–1901, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/
p1889_s1.

Gnutzmann:2004:USSb

[GS04a] Sven Gnutzmann and Burkhard Seif. Universal spectral statistics in Wigner–Dyson, chiral, and Andreev star graphs. I. Construction and numerical results. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 69(??):056219, May 25, 2004. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.69.056219.

Gnutzmann:2004:USSa

[GS04b] Sven Gnutzmann and Burkhard Seif. Universal spectral statistics in Wigner–Dyson, chiral, and Andreev star graphs. II. Semiclassical approach. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 69(??):056220, May 25, 2004. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (elec-

tronic). URL http://link.aps.org/doi/10.1103/PhysRevE.
69.056220.

Garban:2022:CSB

[GS22] Christophe Garban and Thomas Spencer. Continuous symmetry breaking along the Nishimori line. Journal of Mathematical Physics, 63(9):093302, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

URL https://pubs.aip.org/aip/jmp/article/63/9/093302/
2844382/Continuous-symmetry-breaking-along-the-Nishimori.

Special collection in honor of Freeman Dyson.

Goldwasser:1988:LES

[GSE^{+}88] Edwin L. Goldwasser, Robert Siemann, Martin Einhorn, Gordon Kane, A. Abashian, and Freeman Dyson. Letters to the Editor: SSC alternatives: Critics collide with Dyson. Physics Today, 41 (5):9–15, 132–134, May 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://link.aip.

org/link/?PTO/41/9/1.

Goldhaber:2003:SMP

[GSS^{+}03] Alfred S. Goldhaber, R. Shrock, J. Smith, G. Sterman, P. van Nieuwenhuizen, and W. Weisberger, editors. Symmetry & modern physics: Yang Retirement Symposium: State University of New York, Stony Brook, 21–22 May 1999. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2003. ISBN 981-238-503-7 (hardcover), 981-238-563-0 (paperback), 981-279-508-1 (e-book). LCCN

QC174.17.S9 S994 2003. URL http://www.worldscientific.
com/doi/abs/10.1142/9789812795083_0001;   http://www.
worldscientific.com/worldscibooks/10.1142/5349.

Guhr:1991:DCF

[Guh91] Thomas Guhr. Dyson’s correlation functions and graded symmetry. Journal of Mathematical Physics, 32(2):336–347, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v32/i2/p336_s1.

Guha:2007:BRF

[Guh07] Pathik Guha. Book review: A fine balance — wise and just, but perhaps not a sage: The Scientist as Rebel, by Freeman Dyson, New York Review Books, $27.95. The Telegraph (Calcutta, India), ??(??):??, January 5, 2007. URL http://www.

telegraphindia.com/1070105/asp/opinion/story_7219599.
asp.

Gunson:1962:PCD

[Gun62] J. Gunson. Proof of a conjecture by Dyson in the statistical theory of energy levels. Journal of Mathematical Physics, 3(4): 752–753, April 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v3/i4/p752_s1.

Gessel:2006:SPZ

[GX06] Ira M. Gessel and Guoce Xin. A short proof of the Zeilberger– Bressoud q-Dyson Theorem. Proceedings of the American Mathematical Society, 134(8):2179–2187, August 2006. CODEN PAM-YAR. ISSN 0002-9939 (print), 1088-6826 (electronic). URL http:

//www.jstor.org/stable/4098254. See [Dys62g].

Gill:1987:TOO

[GZ87] Tepper L. Gill and W. W. Zachary. Time-ordered operators and Feynman–Dyson algebras. Journal of Mathematical Physics, 28 (7):1459–1470, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i7/p1459_s1.

Gill:2002:FRQ

[GZ02] Tepper L. Gill and W. W. Zachary. Foundations for relativistic quantum theory. I. Feynman’s operator calculus and the Dyson conjectures. Journal of Mathematical Physics, 43(1):69–93, January 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Hagan:1987:BRBb

[Hag87] William J. Hagan, Jr. Book review: Origins of Life by Freeman Dyson. Isis, 78(2):309–310, June 1987. CODEN ISISA4. ISSN 0021-1753 (print), 1545-6994 (electronic). URL http://

www.jstor.org/stable/231589.

Halpern:2017:QLH

[Hal17] Paul Halpern. The Quantum Labyrinth: How Richard Feynman and John Wheeler Revolutionized Time and Reality. Basic Books, New York, NY, USA, 2017. ISBN 0-465-09758-8 (hardcover), 0-465-09759-6 (e-book). ix + 311 pp. LCCN QC174.12 .H347 2017.

Hamburger:1991:LSV

[Ham91] Jean Hamburger. L’avenir de la science: vu par vingt-sept membres associés étrangers de l’Académie des Sciences. (French) [The future of science: seen by 27 foreign associate members of the Academy of Sciences]. Dunod, Paris, France, 1991. ISBN 2-10-000115-9 (paperback). 110 pp. LCCN ???? Debates directed by Alain Connes, Jacques-Louis Lions, François Jacob and others.

Hanle:1982:BNC

[Han82] Helmut Hanle, editor. I. Bi-National Colloquium for Humboldt Awardees: Princeton, NJ, August 23–26, 1981: proceedings. Alexander von Humboldt Foundation, Bad Godesberg, Bonn, West Germany, 1982. LCCN ???? Published in cooperation with the Institute for Advanced Study, Princeton, NJ.

Haskins:1979:BRB

[Has79] Caryl P. Haskins. Book review: Disturbing the Universe, by Freeman Dyson, Alfred P Sloan Foundation Program. I. American Scientist, 67(6):708–709, November 1979. CODEN AMSCAC. ISSN 0003-0996 (print), 1545-2786 (electronic). URL http://

www.jstor.org/stable/27849538.

Huber:2010:IAD

[HAS10] M. Q. Huber, R. Alkofer, and S. P. Sorella. Infrared analysis of Dyson–Schwinger equations taking into account the Gribov horizon in Landau gauge. Physical Review D (Particles and Fields), 81(??):065003, March 1, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.81.065003.

Huber:2012:ADF

[HB12] Markus Q. Huber and Jens Braun. Algorithmic derivation of functional renormalization group equations and Dyson–Schwinger equations. Computer Physics Communications, 183(6):1290– 1320, June 2012. CODEN CPHCBZ. ISSN 0010-4655 (print),

1879-2944 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S001046551200029X.

Horvatic:2011:WQT

[HBKK11] D. Horvatić, D. Blaschke, D. Klabucar, and O. Kaczmarek. Width of the QCD transition in a Polyakov-loop Dyson–Schwinger equation model. Physical Review D (Particles and Fields), 84(??): 016005, July 21, 2011. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.84.016005.

Heims:1980:JNN

[Hei80] Steve J. Heims. John von Neumann and Norbert Wiener: from mathematics to the technologies of life and death. MIT Press, Cambridge, MA, USA, 1980. ISBN 0-262-08105-9. xviii + 547 pp. LCCN QA28 .H44. From mathematics to the technologies of life and death.

Holy:1987:DBS

[HG87] V. Holý and K. T. Gabrielyan. Dyson and Bethe–Salpeter equations for dynamical X-ray diffraction in crystals with randomly placed defects. Physica Status Solidi. B, Basic Research, 140 (1):39–50, March 1, 1987. CODEN PSSBBD. ISSN 0370-1972 (print), 1521-3951 (electronic).

Hahne:1988:CAA

[HGE88] F. J. W. Hahne, H. B. Geyer, and C. A. Engelbrecht. Comment on “Algebraic analysis of physical and spurious states in Dyson boson mapping”. Physical Review C (Nuclear Physics), 37(??):885–??, February 1, 1988. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.37.885. See [Par87].

Hargittai:2004:CSI

[HH04a] Magdolna Hargittai and István Hargittai, editors. Candid science IV: conversations with famous physicists. Imperial College Press, London, UK, 2004. ISBN 1-86094-414-0, 1-86094-416-7 (paperback). xvi + 711 pp. LCCN QC15 .H295 2004. URL

http://www.worldscibooks.com/physics/p304.html.

Hargittai:2004:FJD

[HH04b] Magdolna Hargittai and István Hargittai. Freeman J. Dyson. In Candid science IV: conversations with famous physicists [HH04a], pages 440–477. ISBN 1-86094-414-0, 1-86094-416-7 (paperback).

LCCN QC15 .H295 2004. URL http://www.worldscibooks.
com/physics/p304.html.

Hermes:2013:SOM

[HH13] Matthew R. Hermes and So Hirata. Second-order many-body perturbation expansions of vibrational Dyson self-energies. Journal of Chemical Physics, 139(3):034111, 2013. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://

link.aip.org/link/?JCP/139/034111/1.

Hartle:1965:IPC

[HJ65a] James B. Hartle and C. Edward Jones. Inelastic N/D problems and Castillejo–Dalitz–Dyson cuts. Physical Review Letters, 14 (??):801–??, May 10, 1965. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://link.

aps.org/doi/10.1103/PhysRevLett.14.801.

Hartle:1965:IRC

[HJ65b] James B. Hartle and C. Edward Jones. Inelastic resonances and Castillejo–Dalitz–Dyson singularities. Physical Review, 140(??):

B90–??, October 11, 1965. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRev.140.B90.

Horvatic:2007:MDS

[HKR07] D. Horvatić, D. Klabucar, and A. E. Radzhabov. η and η′ mesons in the Dyson–Schwinger approach at finite temperature. Physical Review D (Particles and Fields), 76(??):096009, November 29, 2007. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.76.096009.

Huber:2012:CFS

[HM12] Markus Q. Huber and Mario Mitter. CrasyDSE: a framework for solving Dyson–Schwinger equations. Computer Physics Communications, 183(11):2441–2457, November 2012. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465512001889.

Happer:1994:ATC

[HMMD94] W. Happer, G. J. Macdonald, C. E. Max, and F. J. Dyson. Atmospheric-turbulence compensation by resonant optical backscattering from the sodium layer in the upperatmosphere. Journal of the Optical Society of America. A, Optics, image science, and vision, 11(1):263–276, ???? 1994. CODEN JOAOD6. ISSN 0740-3232.

Hong:2000:ESD

[HMSW00a] Deog Ki Hong, V. A. Miransky, I. A. Shovkovy, and L. C. R. Wijewardhana. Erratum: Schwinger–Dyson approach to color superconductivity in dense QCD [Phys. Rev. D 61, 056001 (2000)]. Physical Review D (Particles and Fields), 62(??):059903, August 7, 2000. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.62.059903. See [HMSW00b].

Hong:2000:SDA

[HMSW00b] Deog Ki Hong, V. A. Miransky, I. A. Shovkovy, and L. C. R. Wijewardhana. Schwinger–Dyson approach to color superconductivity in dense QCD. Physical Review D (Particles and Fields), 61 (??):056001, January 25, 2000. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.61.056001. See

erratum [HMSW00a].

Harada:2008:QSF

[HN08] Masayasu Harada and Yukio Nemoto. Quasifermion spectrum at finite temperature from coupled Schwinger–Dyson equations for a fermion-boson system. Physical Review D (Particles and Fields), 78(??):014004, July 3, 2008. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.78.014004.

Hofmann:2011:SMI

[Hof11] Christoph P. Hofmann. Spontaneous magnetization of an ideal ferromagnet: Beyond Dyson’s analysis. Physical Review B: Condensed Matter and Materials Physics, 84(??):064414, August 19, 2011. CODEN PRBMDO. ISSN 1098-0121. URL http://link.

aps.org/doi/10.1103/PhysRevB.84.064414.

Holyoke:1981:BRB

[Hol81] T. C. Holyoke. Book review: Disturbing the Universe by Freeman Dyson. The Antioch Review, 39(1):125, Winter 1981. ISSN 0003-

5769. URL http://www.jstor.org/stable/4638408.

Homeier:1996:CAI

[Hom96] H. H. H. Homeier. On convergence acceleration for the iterative solution of the inverse Dyson equation. Journal of Molecular Structure. Theochem, 368:81–91, 1996. CODEN THEODJ. ISSN 0166-1280 (print), 1872-7999 (electronic).

URL http://www.chemie.uni-regensburg.de/pub/preprint/

preprint.html#TCQM954. Proceedings of the 2nd Electronic Computational Chemistry Conference.

Horgan:1993:PFJ

[Hor93] John Horgan. Profile: Freeman J. Dyson: Perpendicular to the mainstream. Scientific American, 269(2):27–28, August 1993. CODEN SCAMAC. ISSN 0036-8733 (print), 1946-7087 (elec-

tronic). URL http://www.nature.com/scientificamerican/
journal/v269/n2/pdf/scientificamerican0893-27.pdf.

Horgan:2006:BRR

[Hor06a] John Horgan. Book review: Rent a genius: The Jasons: The Secret History of Science’s Postwar Elite, by Ann Finkbeiner. 304 pp. Viking. $27.95. The New York Times Magazine, ??(??):

??, April 16, 2006. ISSN 0362-1308. URL http://www.nytimes.

com/2006/04/16/books/review/16horgan.html.

Horgan:2006:TFS

[Hor06b] John Horgan. The Templeton Foundation: a skeptic’s take. The Chronicle of Higher Education, ??(??):??, April 7, 2006. ISSN 0009-5982 (print), 1931-1362 (electronic). URL http://

chronicle.com/article/The-Templeton-Foundation-a/20080/

.

Haynes:2022:LSS

[HR22] Alan Haynes and Roland Roeder. Level spacing statistics for the multi-dimensional quantum harmonic oscillator: Algebraic case. Journal of Mathematical Physics, 63(1):012102, January 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Harada:1998:CPT

[HS98] Masayasu Harada and Akihiro Shibata. Chiral phase transition of QCD at finite temperature and density from the Schwinger– Dyson equation. Physical Review D (Particles and Fields), 59 (??):014010, November 23, 1998. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.59.014010.

Hsu:2003:FD

[Hsu03] Stephen Hsu. Freeman Dyson. Web blog., March 28, 2003. URL

http://infoproc.blogspot.com/2009/03/freeman-dyson.html.

Comment on [Daw09].

Holdom:1995:CBF

[HT95] B. Holdom and G. Triantaphyllou. Critical behavior of a fourpoint Schwinger–Dyson equation. Physical Review D (Particles and Fields), 51(??):7124–??, June 15, 1995. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
51.7124.

Hurst:1952:EGF

[Hur52] C. A. Hurst. The enumeration of graphs in the Feynman–Dyson technique. Proceedings of the Royal Society of London. Series A, Mathematical and physical sciences, 214(1116):44–61, August 7, 1952. CODEN PRLAAZ. ISSN 0080-4630. URL http://www.

jstor.org/stable/99158.

Hauck:1998:SCS

[HvSA98a] A. Hauck, L. von Smekal, and R. Alkofer. Solving a coupled set of truncated QCD Dyson–Schwinger equations. Computer Physics Communications, 112(2–3):166–182, August 1998. CO-DEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465598000459.

Hauck:1998:SGD

[HvSA98b] A. Hauck, L. von Smekal, and R. Alkofer. Solving the gluon Dyson–Schwinger equation in the Mandelstam approximation. Computer Physics Communications, 112(2–3):149–165, August 1998. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944

(electronic). URL http://www.sciencedirect.com/science/
article/pii/S0010465598000460.

Hart:1982:EWT

[HZ82] Michael H. Hart and Ben Zuckerman, editors. Extraterrestrials — where are they? Pergamon Press, New York, NY, USA, 1982. ISBN 0-08-026342-9, 0-08-026341-0 (paperback). LCCN QB54 .E95 1982.

Hsu:2001:LPI

[HZ01] J. P. (Jong-Ping) Hsu and Yuanzhong Zhang, editors. Lorentz and Poincaré invariance: 100 years of Relativity, volume 8 of Advanced series on theoretical physical science. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2001. ISBN 981-02-4721-4. xxxi + 583 pp. LCCN

QC174.17.S9 H78 2001. URL http://www.loc.gov/catdir/
toc/fy031/2002280428.html.

Ida:2008:SOT

[IO08] T. Ida and J. V. Ortiz. Second-order, two-electron Dyson propagator theory: Comparisons for vertical double ionization potentials. Journal of Chemical Physics, 129(8):084105, 2008. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL

http://link.aip.org/link/?JCP/129/084105/1.

Ishida:2001:CED

[IT01] H. Ishida and M. I. Trioni. Comparison of the embedding and Dyson-equation methods in the Green’s-function calculation of a defect in solids. Physical Review B: Condensed Matter and Materials Physics, 63(??):155108, March 28, 2001. CODEN PRB-

MDO. ISSN 1098-0121. URL http://link.aps.org/doi/10.
1103/PhysRevB.63.155108.

Ienna:2023:JEC

[IT23] Gerardo Ienna and Simone Turchetti. JASON in Europe: Contestation and the physicists’ dilemma about the Vietnam War. Physics in Perspective (PIP), 25(3):??, November 2023. CO-DEN PHPEF2. ISSN 1422-6944 (print), 1422-6960 (electronic).

URL https://link.springer.com/article/10.1007/s00016-
023-00302-5.

Jensen:2014:XIC

[Jen14] A. (Arne) Jensen, editor. XVIIth International Congress on Mathematical Physics: Aalborg, Denmark, 6–11 August 2012. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4449-23-7 (hardcover), 981-4449-24-5 (e-book). LCCN QC19.2 .I58 2012;

HC445.8.C47 Y66 2014. URL http://www.worldscientific.
com/worldscibooks/10.1142/8700.

Jernow:1967:CDD

[JK67] Stanley Jernow and Emil Kazes. Castillejo–Dalitz–Dyson poles and asymptotic fields. Physical Review, 160(??):1428–??, August 25, 1967. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.160.1428.

Jernow:1968:CDD

[JK68] Stanley Jernow and Emil Kazes. Castillejo–Dalitz–Dyson poles and asymptotic fields. Physical Review, 166(??):1862–??, February 25, 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.166.1862.3.

Jonnadula:2021:SFT

[JKM21] Bhargavi Jonnadula, Jonathan P. Keating, and Francesco Mezzadri. Symmetric function theory and unitary invariant ensembles. Journal of Mathematical Physics, 62(9):093512, September 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Johnson:1986:GDS

[JL86] Gerald W. Johnson and Michel L. (Michel Laurent) Lapidus. Generalized Dyson series, generalized Feynman diagrams, the Feynman integral, and Feynman’s operational calculus, volume 351 of Memoirs of the American Mathematical Society. American Mathematical Society, Providence, RI, USA, 1986. ISBN 0-8218-2413-9 (paperback). ISSN 0065-9266. vi + 78 pp. LCCN QA3 .A57 no. 351 QC174.12.

Johnson:2007:BRD

[Joh07a] George Johnson. Book review: Dancing with the stars: Disturbing the Universe by Freeman Dyson. The New York Times Magazine, ??(??):??, January 7, 2007. ISSN 0362-

1308.  URL http://www.nytimes.com/2007/01/07/books/
review/Johnson.t.html.

Johnson:2007:FDC

[Joh07b] George Johnson. From Freeman Dyson, a collection of essays on the state of science — culture – International Herald Tribue: Book review: The Scientist as Rebel. The New York Times Magazine, ??(??):??, January 5, 2007. ISSN

0362-1308. URL http://www.nytimes.com/2007/01/05/arts/
05iht-idside6.4110240.html.

Johnson:2020:FDM

[Joh20a] George Johnson. Freeman Dyson, math genius turned technological visionary, dies at 96. After an early breakthrough on light and matter, he became a writer who challenged climate science and pondered space exploration and nuclear warfare. New York Times, ??(??):??, February 28, 2020. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL https://www.nytimes.com/2020/02/28/science/
freeman-dyson-dead.html.

Johnston:2020:FDM

[Joh20b] Hamish Johnston. Freeman Dyson (1923–2020). The mathematical physicist and public intellectual Freeman Dyson died on 28 February 2020. IOP Web site, 2020. URL http://www.iop.org/

about/obituaries/page_73973.html.

Johnston:2020:FDU

[Joh20c] Hamish Johnston. Freeman Dyson: unorthodox to the end. Physics World, 33(4):??, April 2020. CODEN PHWOEW.

ISSN 0953-8585 (print), 2058-7058 (electronic). URL https:

//physicsworld.com/a/freeman-dyson-unorthodox-to-the-
end/.

Jones:2016:BRB

[Jon16] Derry W. Jones. Book review: Birds and frogs: selected papers, 1990–2014, by Freeman J. Dyson. Contemporary Physics, 57(4): 596–598, 2016. CODEN CTPHAF. ISSN 0010-7514 (print), 1366-5812 (electronic).

Josephson:1980:CPW

[JR80] B. D. (Brian David) Josephson and V. S. Ramachandran, editors. Consciousness and the Physical World: Edited Proceedings of an Interdisciplinary Symposium on Consciousness Held at the University of Cambridge in January 1978. Pergamon Press, Oxford, UK, 1980. ISBN 0-08-024695-8, 0-08-024694-X (paperback).

LCCN BF311 .C645 1980. URL http://catalog.hathitrust.
org/api/volumes/oclc/5029030.html. With a foreword by F.

J. Dyson.

Jia:2022:RSB

[JRV22] Yiyang Jia, Dario Rosa, and Jacobus J. M. Verbaarschot. Replica symmetry breaking for the integrable two-site Sachdev– Ye–Kitaev model. Journal of Mathematical Physics, 63(10): 103302, October 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic). URL https://pubs.

aip.org/aip/jmp/article/63/10/103302/2843256/Replica-
symmetry-breaking-for-the-integrable-two. Special collec-

tion in honor of Freeman Dyson.

Judson:1972:TRI

[Jud72] Horace Judson. The techniques of reading: an integrated program for improved comprehension and speed. Harcourt Brace Jovanovich, New York, third edition, 1972. ISBN 0-15-589690-3. xviii + 487 pp. LCCN LB1050.5 .J8 1972. in consultation with William S. Schaill.

Kadell:1985:PAD

[Kad85] Kevin W. J. Kadell. A proof of Andrews’ q-Dyson Conjecture for n = 4. Transactions of the American Mathematical Society, 290(1):127–144, July 1985. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic). URL http://www.jstor.org/

stable/1999787. See [Dys62g].

Konrad:2014:RB

[KAD14] Michael Konrad, Tonio Andrade, and Freeman Dyson. Respectable blunders. New York Review of Books, 61(6):??, April 3, 2014. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2014/apr/03/respectable-blunders/. See [Dys14a].

Kaiser:2005:DTA

[Kai05a] David Kaiser. Drawing theories apart: the dispersion of Feynman diagrams in postwar physics. University of Chicago Press, Chicago, IL, USA and London, UK, 2005. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). xix + 469 pp. LCCN QC794.6.F4

K35 2005. URL http://www.loc.gov/catdir/bios/uchi051/
2004023335.html;   http://www.loc.gov/catdir/enhancements/
fy0617/2004023335-d.html;   http://www.loc.gov/catdir/
toc/ecip051/2004023335.html.

Kaiser:2005:FDS

[Kai05b] David Kaiser. The Feynman–Dyson split. In DTA [Kai05a], pages 175–194. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). LCCN QC794.6.F4 K35 2005. URL http://www.

loc.gov/catdir/bios/uchi051/2004023335.html;   http:/
/www.loc.gov/catdir/enhancements/fy0617/2004023335-d.
html;   http://www.loc.gov/catdir/toc/ecip051/2004023335.
html.

Kaiser:2005:FDP

[Kai05c] David Kaiser. Freeman Dyson and the postdoc cascade. In DTA [Kai05a], chapter 3, pages 60–111. ISBN 0-226-42266-6, 0-226-42267-4 (paperback). LCCN QC794.6.F4 K35

2005.  URL http://www.loc.gov/catdir/bios/uchi051/
2004023335.html;   http://www.loc.gov/catdir/enhancements/
fy0617/2004023335-d.html;   http://www.loc.gov/catdir/
toc/ecip051/2004023335.html.

Kalman:1964:DET

[Kal64] G. Kalman. Dyson equation technique in nonequilibrium statistical mechanics. Physics of Fluids, 7(6):910–911, 1964. CODEN

PHFLE6. ISSN 1070-6631. URL http://link.aip.org/link/?
PFL/7/910/1.

Kantrowitz:2002:BRB

[Kan02] Arthur Kantrowitz. Book review: Project Orion: The True Story of the Atomic Spaceship, by George Dyson. Physics Today, 55(11):69–70, November 2002. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http:/

/www.daviddarling.info/encyclopedia/O/OrionProj.html;
 http://www.physicstoday.org/resource/1/phtoad/v55/i11/
p69_s2.

Kaplan:1973:SPP

[Kap73] Morton A. Kaplan, comp, editor. SALT: problems and prospects. General Learning Press, Morristown, NJ, USA, 1973. xiii + 251 pp. LCCN JX1974.7 .K24; JZ5665 .K37 1973x. URL http://

catalog.hathitrust.org/api/volumes/oclc/623340.html.

Kuiper:1989:ECc

[KB89] Thomas B. H. (Thomas Bernardus Henricus) Kuiper and Glen David Brin, editors. Extraterrestrial civilization. American Association of Physics Teachers, College Park, MD, USA, 1989. ISBN 0-917853-38-5. 121 pp. LCCN QB54 .E945; QB54 .E97 1989.

Kinoshita:73

[KC73] T. Kinoshita and P. Cvitanovic. Feynman-Dyson rules in parametric space. Report CLNS-209, Cornell Lab. for Nuclear Studies, January 1973.

Kahneman:2012:TCE

[KD12] Daniel Kahneman and Freeman J. Dyson. A ‘thriving collective enterprise’. New York Review of Books, 59(1):56, January 12, 2012. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2012/jan/12/thriving-collective-enterprise/.

Kevles:1987:PHS

[Kev87] Daniel J. Kevles. The physicists: the history of a scientific community in modern America. Harvard University Press, Cambridge, MA, USA, 1987. ISBN 0-674-66655-0 (paperback). xv + 489 pp. LCCN QC9.U5 K48 1987. URL http://catalog.

hathitrust.org/api/volumes/oclc/15860449.html.

Kidder:1989:AC

[Kid89] Rushworth M. Kidder. An Agenda for the 21st Century. MIT Press, Cambridge, MA, USA, 1989. ISBN 0-262-11128-4. xxii + 216 pp. LCCN CB161 .A35 1987.

Klabucar:1998:CSD

[KK98] Dubravko Klabucar and Dalibor Kekez. η and η′ in a coupled Schwinger–Dyson and Bethe–Salpeter approach. Physical Review D (Particles and Fields), 58(??):096003, September 21, 1998. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.58.096003.

Kekez:2002:CSD

[KK02] Dalibor Kekez and Dubravko Klabucar. η and η′ in a coupled Schwinger–Dyson and Bethe–Salpeter approach. II. The γ^{∗}γ transition form factors. Physical Review D (Particles and Fields), 65 (??):057901, January 28, 2002. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.65.057901.

Kisslinger:2002:LCR

[KL02] L. S. Kisslinger and O. Linsuain. Light-cone representation of the quark Schwinger–Dyson equation. Physical Review C (Nuclear Physics), 66(??):045206, October 29, 2002. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://

link.aps.org/doi/10.1103/PhysRevC.66.045206.

Kleinert:1989:SDE

[Kle89] H. Kleinert. Solution of Dyson equation for elastic membranes. Journal of Mathematical Physics, 30(12):2991–2993, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v30/i12/p2991_s1.

Katzir:2013:TTH

[KLR13] Shaul Katzir, Christoph Lehner, and Jürgen Renn, editors. Traditions and transformations in the history of quantum physics: HQ-3, Third International Conference on the History of Quantum Physics, Berlin, June 28–July 2, 2010, volume 5 of Max Planck research library for the history and development of knowledge. Proceedings. Edition Open Access, Berlin, Germany, 2013. ISBN 3-8442-5134-0. LCCN QC173.98. URL http://www.edition-

open-access.de/proceedings/5/.

Katkov:2022:MMH

[KNGT22] Mikhail Katkov, Michelangelo Naim, Antonios Georgiou, and Misha Tsodyks. Mathematical models of human memory. Journal of Mathematical Physics, 63(7):073303, July 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

URL https://pubs.aip.org/aip/jmp/article/63/7/073303/
2843660/Mathematical-models-of-human-memory. Special col-

lection in honor of Freeman Dyson. See erratum [KTNG23].

Kondo:1990:OEO

[Kon90] Yuji Kondo, editor. Observatories in Earth orbit and beyond: proceedings of the 123rd Colloquium of the International Astronomical Union, held in Greenbelt, Maryland, April 24–27, 1990, volume 166 of Astrophysics and space science library. Kluwer Academic Publishers, Norwell, MA, USA, and Dordrecht, The Netherlands, 1990. ISBN 0-7923-1133-7 (hardcover). LCCN QB500.267 .I56 1990.

Kondo:1997:RCC

[Kon97] Kei-Ichi Kondo. Running coupling constant of a gauge theory in the framework of the Schwinger–Dyson equation: Infrared behavior of three-dimensional quantum electrodynamics. Physical Review D (Particles and Fields), 55(??):7826–??, June 15, 1997. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.55.7826.

Kornfeld:1994:STR

[Kor94] R. Kornfeld. Science in trouble + response to Freeman Dyson article. American Scholar, 63(1):157, ???? 1994. ISSN 0003-0937 (print), 2162-2892 (electronic).

Kizilersu:2009:BFF

[KP09] Ayse Kizilersü and Michael R. Pennington. Building the full fermion-photon vertex of QED by imposing multiplicative renormalizability of the Schwinger–Dyson equations for the fermion and photon propagators. Physical Review D (Particles and Fields), 79(??):125020, June 22, 2009. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
79.125020.

Korpa:1990:TSD

[KPS90] C. L. Korpa, Jitendra C. Parikh, and Philip J. Siemens. Truncation of Schwinger–Dyson equations and the 1/N expansion in the O(N) model. Physical Review D (Particles and Fields), 41 (??):1276–??, February 15, 1990. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.41.1276.

Krauss:2001:AOB

[Kra01] Lawrence Maxwell Krauss. Atom: an odyssey from the Big Bang to life on earth — and beyond. Little, Brown, London, UK, 2001. ISBN 0-316-64877-9. 305 pp. LCCN ????

Krauss:2011:QMR

[Kra11] Lawrence Maxwell Krauss. Quantum man: Richard Feynman’s life in science. Great discoveries. W. W. Norton & Co., New York, NY, USA, 2011. ISBN 0-393-06471-9. xvii + 350 pp. LCCN QC16.F49 K73 2011. URL http://

www.scientificcomputing.com/news-DS-Illuminating-the-
Life-and-Legacy-of-Richard-Feynman-032211.aspx.

Krishnaswami:2008:SDO

[Kri08] Govind S. Krishnaswami. Schwinger–Dyson operators as invariant vector fields on a matrix model analog of the group of loops. Journal of Mathematical Physics, 49(6):062303, June 2008. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic), 1527-2427. URL http://jmp.aip.org/resource/
1/jmapaq/v49/i6/p062303_s1.

Krivoruchenko:2010:ROC

[Kri10] M. I. Krivoruchenko. Remarks on the origin of Castillejo– Dalitz–Dyson poles. Physical Review C (Nuclear Physics), 82 (??):018201, July 8, 2010. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.org/

doi/10.1103/PhysRevC.82.018201.

Kizilersu:2013:SCU

[KSW13] Ayse Kizilersü, Tom Sizer, and Anthony G. Williams. Stronglycoupled unquenched QED_{4} propagators using Schwinger–Dyson equations. Physical Review D (Particles and Fields), 88(??):

045008, August 8, 2013. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.88.045008.

Katkov:2023:EMM

[KTNG23] Mikhail Katkov, Misha Tsodyks, Michelangelo Naim, and Antonis Georgiou. Erratum: “Mathematical models of human memory” [J. Math. Phys. 63, 073303 (2022)]. Journal of Mathematical Physics, 64(2):029901, February 2023. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

URL https://pubs.aip.org/aip/jmp/article/64/2/029901/
2874149/Erratum-Mathematical-models-of-human-memory-J-

Math. See [KNGT22].

Kim:1985:IFP

[KV85] G.-K. Kim and C. M. Vincent. Inadequacy of four-particle tests of Dyson boson mapping. Physical Review C (Nuclear Physics), 32(??):1776–??, November 1, 1985. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.

aps.org/doi/10.1103/PhysRevC.32.1776.

Lacroix-A-Chez-Toine:2022:SRP

[LACTFF22] Bertrand Lacroix-A-Chez-Toine, Yan V. Fyodorov, and Sirio Belga Fedeli. Superposition of random plane waves in high spatial dimensions: Random matrix approach to landscape complexity. Journal of Mathematical Physics, 63(9):093301, September 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic). URL https://pubs.aip.org/aip/jmp/article/63/
9/093301/2844320/Superposition-of-random-plane-waves-

in-high. Special collection in honor of Freeman Dyson.

Lamb:2006:WSE

[Lam06] Gregory M. Lamb. Where science and ethics meet: Physicist Freeman Dyson explores the ability of science to help us make sense of the world. The Christian Science Monitor, ??(??):??, December 19, 2006. ISSN 0882-7729 (print), 1540-4617 (electronic). URL

http://www.csmonitor.com/2006/1219/p14s01-bogn.html.

Largeault:1989:NCB

[Lar89] J. Largeault. Note critique: Les dérangeurs d’univers par F. Dyson. (French) [Book review: The disturbers of the Universe by F. Dyson]. Revue de métaphysique et de morale, 94(3):421, July/September 1989. CODEN ???? ISSN 0035-1571 (print),

2102-5177 (electronic). URL http://www.jstor.org/stable/
40903066.

Liou:1972:NRS

[LCW^{+}72] H. I. Liou, H. S. Camarda, S. Wynchank, M. Slagowitz, G. Hacken, F. Rahn, and J. Rainwater. Neutron-resonance spectroscopy. VIII. The separated isotopes of erbium: Evidence for Dyson’s theory concerning level spacings. Physical Review C (Nuclear Physics), 5(??):974–??, March 1, 1972. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://

link.aps.org/doi/10.1103/PhysRevC.5.974.

Loly:1966:RAD

[LD66] P. D. Loly and S. Doniach. Removal of an apparent discrepancy between calculations of Dyson and of Oguchi for the Heisenberg ferromagnet. Physical Review, 144(??):319–??, April 8, 1966. CO-DEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic).

URL http://link.aps.org/doi/10.1103/PhysRev.144.319.

Lenard:1968:SMI

[LD68] A. Lenard and Freeman J. Dyson. Stability of matter. II. Journal of Mathematical Physics, 9(5):698–711, May 1968. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v9/i5/
p698_s1;   http://link.aip.org/link/?JMP/9/698/1.

Lipkin:1990:LES

[LD90] Harry J. Lipkin and Freeman J. Dyson. Letters to the Editor: Sakharov and Aruri: An unapt comparison. Physics Today, 43(5): 119–120, May 1990. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
43/119/1.

Lytton:2008:LEB

[LD08] Bernard Lytton and Freeman J. Dyson. Letter to the Editor: von Braun’s bargain. New York Review of Books, 55(2): 55, February 14, 2008. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2008/feb/14/von-brauns-bargain/. Comments on [Dys08b].

Leonard:1977:BET

[LDB77] H. Jeffrey Leonard, J. Clarence Davies, and Gordon Binder, editors. Business and environment: toward common ground. The Conservation Foundation, Washington, DC, USA, 1977. ISBN 0-89164-046-0. xi + 434 pp. LCCN HD60 B979e.

Llanes-Estrada:2003:QSD

[LEdAB03] F. J. Llanes-Estrada and P. de A. Bicudo. Quark Schwinger– Dyson evaluation of the l_{1}, l_{2} coefficients in the chiral Lagrangian. Physical Review D (Particles and Fields), 68(??):094014, November 19, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.68.094014.

Lee:2007:BPC

[Lee07] Sabine Lee. The Bethe–Peierls correspondence. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2007. ISBN 981-277-135-2. xi + 506 pp. LCCN QC16.B46 A4 2007eb. URL http:

//site.ebrary.com/lib/yale/Doc?id=10255467;   http://
www.loc.gov/catdir/toc/fy0805/2008274437.html;   http:
//www.worldscientific.com/worldscibooks/10.1142/6595.

Lerche:1971:KDT

[Ler71] Ian Lerche. Kinematic dynamo theory: The Dyson equation and the large-scale field: the Bethe–Salpeter equation and the fluctuation intensity. Journal of Mathematical Physics, 12(8):1538– 1547, August 1971. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v12/i8/p1538_s1.

Lewis:1999:BRB

[Lew99] Arthur O. Lewis. Book review: Beyond Progress: An Interpretive Odyssey to the Future by Hugh DeSantis. Imagined Worlds by Freeman Dyson. The World in 2020: Power, Culture and Prosperity: A Vision of the Future by Hamish McRae. Utopian Studies, 10(1):189–196, ???? 1999. ISSN 1045-991X (print), 2154-9648

(electronic). URL http://www.jstor.org/stable/20718034.

Lewin:2022:CRG

[Lew22] Mathieu Lewin. Coulomb and Riesz gases: The known and the unknown. Journal of Mathematical Physics, 63(6):061101, June 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658

(electronic). URL https://pubs.aip.org/aip/jmp/article/
63/6/061101/2845978/Coulomb-and-Riesz-gases-The-known-

and-the-unknown. Special collection in honor of Freeman Dyson.

Li:2013:CSS

[LFJ^{+}13] Jian-Feng Li, Hong-Tao Feng, Yu Jiang, Wei-Min Sun, and Hong-Shi Zong. Calculation of the staggered spin correlation in the framework of the Dyson–Schwinger approach. Physical Review D (Particles and Fields), 87(??):116008, June 20, 2013. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.87.116008.

Luecker:2010:VBQ

[LFW10] Jan Luecker, Christian S. Fischer, and Richard Williams. Volume behavior of quark condensate, pion mass, and decay constant from Dyson–Schwinger equations. Physical Review D (Particles and Fields), 81(??):094005, May 5, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
81.094005.

Lieb:2004:SAF

[Lie04] E. Lieb. A short appreciation of Freeman Dyson’s long career in mathematical physics. Communications in Mathematical Physics, 252(1–3):3–5, ???? 2004. CODEN CMPHAY. ISSN 0010-3616 (print), 1432-0916 (electronic).

Livesay:1954:TFJ

[Liv54] George R. Livesay. On a theorem of F. J. Dyson. Annals of Mathematics, 59(2):227–229, March 1954. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://

www.jstor.org/stable/1969689.

Li:2009:URM

[LLW^{+}09] D. Li, X. Li, F. Wang, H. Huang, X. Li, and L. C. Kwek. Uncertainty relation of mixed states by means of Wigner–Yanase–Dyson information. Physical Review A (Atomic, Molecular, and Optical Physics), 79(??):052106, May 6, 2009. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL

http://link.aps.org/doi/10.1103/PhysRevA.79.052106.

Lieb:1966:MPO

[LM66] Elliott H. Lieb and Daniel C. Mattis, editors. Mathematical Physics in One Dimension: exactly soluble models of interacting particles. Academic Press, New York, USA, 1966. xiv + 565 pp. LCCN QC174.5 .L47.

Leung:1996:SDE

[LNA96] C. N. Leung, Y. J. Ng, and A. W. Ackley. Schwinger–Dyson equation approach to chiral symmetry breaking in an external magnetic field. Physical Review D (Particles and Fields), 54(??): 4181–??, September 15, 1996. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.54.4181.

Lieb:2022:EFI

[LS22] Elliott H. Lieb and Siddhartha Sahi. On the extension of the FKG inequality to n functions. Journal of Mathematical Physics, 63(4):043301, April 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Lieb:1976:SMP

[LSW76] Elliott H. Lieb, Barry Simon, and A. S. Wightman, editors. Studies in mathematical physics: essays in honor of Valentine Bargmann. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1976. xii + 460 pp. LCCN ????

Lieb:1991:SMA

[LT91] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-53039-8 (Berlin), 0-387-53039-8 (New York). viii + 565 pp. LCCN QC173.4.T48 L54 1991. With a preface by F. Dyson.

Lieb:1997:SMA

[LT97] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997. ISBN 3-540-61565-2 (Berlin: hardcover). xi + 675 pp. LCCN QC173.4.T48 L54 1997. With a preface by F. Dyson.

Lieb:2001:SMA

[LT01] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2001. ISBN 3-540-42083-5. xiii + 812 pp. LCCN QC173.4.T48 L54 2001. With a preface by F. Dyson.

Lieb:2005:SMA

[LT05] Elliott H. Lieb and Walter E. Thirring, editors. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2005. ISBN 3-540-22212-X. xv + 932 pp. LCCN QC173.4.T48 L54 2005. URL http://www.loc.

gov/catdir/enhancements/fy0663/2004108033-d.html. With

a preface by F. Dyson.

Lannutti:1978:CTT

[LW78] J. E. Lannutti and P. K. Williams, editors. Current trends in the theory of fields: (Tallahassee 1978): a symposium in honor of P. A. M. Dirac, volume 48 of AIP conference proceedings. American Institute of Physics, Woodbury, NY, USA, 1978. ISBN 0-88318-147-9. LCCN QC793.3.F5 C87.

Liu:1995:CNY

[LY95] C. S. (Chao Shiuan) Liu and Shing-Tung Yau, editors. Chen Ning Yang: a great physicist of the Twentieth Century. International Press, Cambridge, MA, 1995. ISBN 1-57146-001-2. LCCN QC16.Y364 A3 1995.

Maas:2006:SST

[Maa06] Axel Maas. Solving a set of truncated Dyson–Schwinger equations with a globally converging method. Computer Physics Communications, 175(3):167–179, August 1, 2006. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0010465506001111.

Mahlburg:2005:PCA

[Mah05] Karl Mahlburg. Partition congruences and the Andrews–Garvan– Dyson crank. Proceedings of the National Academy of Sciences of the United States of America, 102(43):15373–15376, October 25, 2005. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://www.jstor.org/stable/4143442.

Mahoney:2011:CMP

[Mah11] Michael S. (Michael Sean) Mahoney. Computing and mathematics at Princeton in the 1950s. In Mahoney and Haigh [MH11], page ?? ISBN 0-674-05568-3. LCCN QA76.17 .M34 2011. Edited and with an introduction by Thomas Haigh. See the Mahoney obituary [?] for comments on the author and this book.

Manoukian:1976:EGI

[Man76a] Edward B. Manoukian. Erratum: Generalization and improvement of the Dyson–Salam renormalization scheme and equivalence with other schemes. Physical Review D (Particles and Fields), 14(??):2202–??, October 15, 1976. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.

14.2202. See [Man76b].

Manoukian:1976:GID

[Man76b] Edward B. Manoukian. Generalization and improvement of the Dyson–Salam renormalization scheme and equivalence with other schemes. Physical Review D (Particles and Fields), 14 (??):966–??, August 15, 1976. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.14.966. See er-

ratum [Man76a].

Manoukian:1977:CGI

[Man77] Edward B. Manoukian. Convergence of the generalized and improved Dyson–Salam renormalization scheme. Physical Review D (Particles and Fields), 15(??):535–??, January 15, 1977. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/

PhysRevD.15.535. See erratum [Man82].

Manoukian:1978:HEB

[Man78] Edward B. Manoukian. High-energy behavior of renormalized Feynman amplitudes. Journal of Mathematical Physics, 19(5): 917–923, May 1978. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v19/i5/p917_s1.

Manoukian:1982:ECG

[Man82] E. B. Manoukian. Erratum: Convergence of the generalized and improved Dyson–Salam renormalization scheme. Physical Review D (Particles and Fields), 25(??):1157–??, February 15, 1982. CO-DEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic),

1538-4500 (CD-ROM). URL http://link.aps.org/doi/10.
1103/PhysRevD.25.1157.2. See [Man77].

Manin:2007:MMS

[Man07] Yuri Ivanovich Manin. Mathematics as metaphor: selected essays of Yuri I. Manin. American Mathematical Society, Providence, RI, USA, 2007. ISBN 0-8218-4331-1. xvii + 232 pp. LCCN QA7 .M279 2007. With foreword by Freeman J. Dyson.

Marshall:1956:INM

[Mar56] Andrew W. Marshall. An introductory note [on Monte Carlo method]. In Meyer [Mey56], pages 14–?? LCCN QA273 U577.

Marshalek:1988:FHA

[Mar88] E. R. Marshalek. Finite Hermitian alternatives to the Dyson Hamiltonian. Physical Review C (Nuclear Physics), 38(??):2961– ??, December 1, 1988. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.org/

doi/10.1103/PhysRevC.38.2961.

Maddox:1960:AB

[MASD60] John Maddox, Poul Anderson, Eugene A. Sloane, and Freeman J. Dyson. Artificial biosphere. Science, 132(3421):250–253, July 22, 1960. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.jstor.org/stable/1705166.

Matthews:1949:ADM

[Mat49a] P. T. Matthews. The application of Dyson’s methods to meson interactions. Physical Review, 76(??):684–??, September 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/

PhysRev.76.684.2. See erratum [Mat49b].

Matthews:1949:EAD

[Mat49b] P. T. Matthews. Erratum: The application of Dyson’s methods to meson interactions. Physical Review, 76(??):1419–??, November 1, 1949. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.76.1419.2. See [Mat49a].

Mavrodes:1975:LRC

[Mav75] George I. Mavrodes. Letters: Risk and consent. Bulletin of the Atomic Scientists, 31(8):3–4, October 1975. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys75b].

Mayer:2007:EKS

[May07] Elizabeth Lloyd Mayer. Extraordinary Knowing: Science, Skepticism, and the Inexplicable Powers of the Human Mind. Bantam Books, New York, NY, USA, 2007. ISBN 0-553-80335-2 (hardcover). xiii + 302 pp. LCCN BF315 .M36 2007. URL http:/

/www.loc.gov/catdir/enhancements/fy0738/2006025661-b.
html;   http://www.loc.gov/catdir/enhancements/fy0738/
2006025661-d.html;   http://www.loc.gov/catdir/enhancements/
fy0738/2006025661-s.html;   http://www.loc.gov/catdir/
toc/ecip0619/2006025661.html. Forewords by Freeman Dyson

and Carol Gilligan.

Marshak:1966:PMP

[MB66] R. E. Marshak and J. Warren Blaker, editors. Perspectives in modern physics. Essays in honor of Hans A. Bethe on the occasion of his 60th birthday, July, 1966. Interscience Publishers, New York, NY, USA, 1966. xii + 673 pp. LCCN QC774.B4 M3.

Mccutchen:1994:DFS

[Mcc94] C. W. Mccutchen. Freeman Dyson science in trouble (American Scholar, the 1994 winter issue). American Scholar, 63(2):320, ???? 1994. ISSN 0003-0937 (print), 2162-2892 (electronic).

McLeod:2013:CDD

[McL13] R. J. McLeod. Castillejo, Dalitz, and Dyson ambiguity in field theoretic models of scattering. Physical Review D (Particles and Fields), 87(??):125016, June 10, 2013. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
87.125016.

McMahan:1965:RBM

[McM65] Richard H. McMahan, Jr. Rationales for ballistic missile defense policy. Bulletin of the Atomic Scientists, 21(3):37–40, March 1965. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys64d, Dys65b].

McPhee:1974:CBE

[McP74] John A. McPhee. The curve of binding energy. Farrar, Strauss, and Giroux, New York, NY, USA, 1974. ISBN 0-374-13373-5. 232 pp. LCCN UF767 .M215 1974. URL http://

www.loc.gov/catdir/bios/hol059/74001226.html;   http:
//www.loc.gov/catdir/description/hol054/74001226.html.

McVoy:1957:LPB

[MD57] Kirk W. McVoy and Freeman J. Dyson. Longitudinal polarization of Bremsstrahlung and pair production at relativistic energies. Physical Review, 106(6):1360–1361, June 15, 1957. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.106.1360.

Mehta:1963:STE

[MD63] Madan Lal Mehta and Freeman J. Dyson. Statistical theory of the energy levels of complex systems. V. Journal of Mathematical Physics, 4(5):713–719, May 1963. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v4/i5/
p713_s1;   http://link.aip.org/link/?JMP/4/713/1.

MacLane:1995:MTR

[MD95a] Saunders Mac Lane and Freeman Dyson. A matter of temperament. New York Review of Books, 42(15):56, October 5, 1995. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/1995/oct/05/
a-matter-of-temperament/.

MacLane:1995:LEM

[MD95b] Saunders Mac Lane and Freeman J. Dyson. Letter to the Editor: ‘a matter of temperament’. New York Review of Books, 42(15):??, October 5, 1995. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/1995/oct/05/a-matter-of-temperament/. In re-

sponse to [Dys95i].

May:2008:LEH

[MD08] Robert M. May and Freeman J. Dyson. Letter to the Editor: How long will they stay? New York Review of Books, 55(15):50, October 9, 2008. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2008/oct/09/how-long-will-they-stay/. In re-

sponse to [Dys08d].

Mader:2011:HDM

[MEBK11] V. Mader, G. Eichmann, M. Blank, and A. Krassnigg. Hadronic decays of mesons and baryons in the Dyson–Schwinger approach. Physical Review D (Particles and Fields), 84(??):034012, August 8, 2011. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.84.034012.

Metz:1973:PTP

[Met73] William D. Metz. Physics at a turning point? — Interview with Freeman Dyson. Science, 179(4078):1114, March 16, 1973. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203

(electronic). URL http://www.sciencemag.org/content/179/
4078/1114.full.pdf.

Meurice:2001:SNC

[Meu01] Y. Meurice. Small numerators canceling small denominators: Is Dyson’s hierarchical model solvable? Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 63(??):055101, April 9, 2001. CODEN PLEEE8. ISSN 1539-3755 (print), 1550-2376 (electronic). URL http://link.aps.

org/doi/10.1103/PhysRevE.63.055101.

Meurice:2009:DIN

[Meu09] Y. Meurice. Dyson instability for 2D nonlinear O(N) sigma models. Physical Review D (Particles and Fields), 80(??):054020, September 21, 2009. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://

link.aps.org/doi/10.1103/PhysRevD.80.054020.

Meyer:1956:SMC

[Mey56] Herbert A. Meyer, editor. Symposium on Monte Carlo Methods: held at the University of Florida, March 16 and 17, 1954. Wiley, New York, NY, USA, 1956. LCCN QA273 U577.

Mahoney:2011:HC

[MH11] Michael S. (Michael Sean) Mahoney and Thomas Haigh, editors. Histories of computing. Harvard University Press, Cambridge, MA, USA, 2011. ISBN 0-674-05568-3. 250 pp. LCCN QA76.17 .M34 2011. Edited and with an introduction by Thomas Haigh. See the Mahoney obituary [?] for comments on the author and this book.

Morse:1972:NPM

[MHFW72] Philip M. (Philip McCord) Morse, Bernard T. Held, Herman Feshbach, and Richard Wilson, editors. Nuclear, particle and many body physics. Academic Press, New York, USA, 1972.

ISBN 0-12-508201-0 (vol. 1), 0-323-14303-2 (e-book). xxi + 693 pp. LCCN QC173 .N8837 1972. URL http://lccn.loc.

gov/79183529;   http://www.sciencedirect.com/science/

book/9780125082013. Two volume tribute to Amos de-Shalit (1926–1969).

Mitchell:1980:CSI

[Mit80] Janet A. Mitchell, editor. A community of scholars: the Institute for Advanced Study, faculty and members 1930–1980. The Institute for Advanced Study, Princeton, NJ, USA, 1980. xxiii + 549 pp. LCCN LD2535.I31 I67 1980. With a foreword by Harry Woolf. Biographies and bibliographies of The Institute for Advanced Study faculty and members, in celebration of the 50th anniversary of the Institute.

Munczek:1990:SDE

[MM90] Herman J. Munczek and Douglas W. McKay. Schwinger–Dyson equation in QCD: Comparison of some approximations. Physical Review D (Particles and Fields), 42(??):3548–??, November 15, 1990. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.42.3548.

McKay:1997:SQP

[MM97] Douglas W. McKay and Herman J. Munczek. Study of quark propagator solutions to the Dyson–Schwinger equation in a confining model. Physical Review D (Particles and Fields), 55(??): 2455–??, February 15, 1997. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.55.2455.

Magarshak:1991:DTS

[MMJ91] Y. Magarshak, J. Malinsky, and A. D. Joran. Diagram techniques for solving Schwinger–Dyson equations: Electron transfer pathways in biological molecules. Journal of Chemical Physics, 95(1):418–432, 1991. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://link.aip.org/link/?JCP/
95/418/1.

Marhauser:2007:CSL

[MNBW07] Florian Marhauser, Dominik Nickel, Michael Buballa, and Jochen Wambach. Color-spin locking in a self-consistent Dyson– Schwinger approach. Physical Review D (Particles and Fields), 75(??):054022, March 16, 2007. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.75.054022.

Miao:2011:CDO

[MND11] Y. R. Miao, C. G. Ning, and J. K. Deng. Calculation of Dyson orbitals using a symmetry-adapted-cluster configuration-interaction method for electron momentum spectroscopy: N_{2} and H_{2}O. Physical Review A (Atomic, Molecular, and Optical Physics), 83(??): 062706, June 15, 2011. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.

aps.org/doi/10.1103/PhysRevA.83.062706.

Mitsumori:1997:QCN

[MNK^{+}97] Tomohiro Mitsumori, Nobuo Noda, Hiroaki Kouno, Akira Hasegawa, and Masahiro Nakano. Quark condensate in nuclear matter based on nuclear Schwinger–Dyson formalism. Physical Review C (Nuclear Physics), 55(??):1577–??, March 1, 1997. CO-DEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

URL http://link.aps.org/doi/10.1103/PhysRevC.55.1577.

Miao:2011:DOE

[MNLD11] Y. R. Miao, C. G. Ning, K. Liu, and J. K. Deng. Dyson orbitals of N_{2}O: Electron momentum spectroscopy and symmetry adapted cluster-configuration interaction calculations. Journal of Chemical Physics, 134(20):204304, 2011. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://

link.aip.org/link/?JCP/134/204304/1.

Medikeri:1994:OPS

[MNM94] Milan N. Medikeri, Jayraman Nair, and Manoj K. Mishra. On the orbital picture of shape resonances using Feynman–Dyson amplitudes from different decouplings of the dilated electron propagator. Journal of Chemical Physics, 100(3):2044–2051, 1994. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://link.aip.org/link/?JCP/100/2044/1.

Madhavan:2013:PS

[MOG^{+}13] Guruprasad Madhavan, Barbara Oakley, David Green, David Koon, and Penny Low, editors. Practicing Sustainability. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2013. ISBN 1-4614-4348-2, 1-4614-4349-0 (ebook). xxxvi + 244 pp. LCCN HC79.E5 P6765 2013. URL http:/

/www.loc.gov/catdir/enhancements/fy1316/2012947954-b.
html;   http://www.loc.gov/catdir/enhancements/fy1316/
2012947954-d.html;   http://www.loc.gov/catdir/enhancements/

fy1316/2012947954-t.html. Foreword by Michael Spence. Editorials by Klaus Schwab, Robert Rubin, and George Whitesides. Afterword by M. S. Swaminathan.

Monastyrskii:1987:RTP

[Mon87] Mikhail Ilyich Monastyrskiĭ, editor. Riemann, topology, and physics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1987. ISBN 0-8176-3262-X. xiii + 158 pp. LCCN QA29.R425 M6613 1987. US$34.95. With a foreword by Freeman J. Dyson. Translated from Russian by James King and Victoria King.

Montesano:1994:FDR

[Mon94] Craig Montesano. Freeman Dyson receives 12th Wright Prize. The Scientist (Philadelphia, PA), 8(8):23, April 1994. ISSN 0890-3670 (print), 1945-5127 (electronic). URL http://www.

the-scientist.com/?articles.view/articleNo/27992/.

Monastyrsky:1999:RTP

[Mon99] Michael Monastyrsky. Riemann, topology, and physics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, second edition, 1999. ISBN 0-8176-3789-3, 3-7643-3789-3. xiii + 215 pp. LCCN QA29.R425 M6613 1999.

URL http://link.springer.com/book/10.1007/978-0-8176-

4779-7/page/1. Foreword by Freeman J. Dyson. Translated by Roger Cooke, James King, Victoria King.

Morrison:1958:GRA

[Mor58] Philip Morrison. On gamma-ray astronomy. Il Nuovo Cimento (10), 7(6):858–865, March 16, 1958. CODEN NUCIAD. ISSN 0029-6341 (print), 1827-6121 (electronic). URL http://link.

springer.com/article/10.1007/BF02745590.

Moravcsik:1962:LEF

[Mor62] Michael J. Moravcsik. Letter to the Editor: Further thoughts on shelters. Bulletin of the Atomic Scientists, 18(7), September 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys62j].

Meurice:1995:ECS

[MOR95] Y. Meurice, G. Ordaz, and V. G. J. Rodgers. Evidence for complex subleading exponents from the high-temperature expansion of Dyson’s hierarchical Ising model. Physical Review Letters, 75 (??):4555–??, December 18, 1995. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL http://

link.aps.org/doi/10.1103/PhysRevLett.75.4555.

Moser:1950:APSb

[MS50] Leo Moser and Robert Steinberg. Advanced problems and solutions: 4317. American Mathematical Monthly, 57(5):345, May 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wil48].

Munczek:1995:DCS

[Mun95] H. J. Munczek. Dynamical chiral symmetry breaking, Goldstone’s theorem, and the consistency of the Schwinger–Dyson and Bethe– Salpeter equations. Physical Review D (Particles and Fields), 52 (??):4736–??, October 15, 1995. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.52.4736.

Nakano:1987:DRC

[Nak87] K. Nakano. Dispersion relations and Castillejo–Dalitz–Dyson pole in the πNP wave scattering amplitude. Physical Review C (Nuclear Physics), 36(??):714–??, August 1, 1987. CODEN PRV-CAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http:

//link.aps.org/doi/10.1103/PhysRevC.36.714.

Narcowich:1986:DLE

[Nar86] Francis J. Narcowich. A Dyson-like expansion for solutions to the quantum Liouville equation. Journal of Mathematical Physics, 27 (10):2502–2510, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://

jmp.aip.org/resource/1/jmapaq/v27/i10/p2502_s1.

Navarro:1999:BRF

[Nav99] Jesús Navarro. Book review: Futuro e ideología: Mundos del futuro, by Freeman J. Dyson. Revista de libros, 67(31–32):37, July/

August 1999. URL http://www.jstor.org/stable/30229020.

Nickel:2008:NCF

[NAW08] D. Nickel, R. Alkofer, and J. Wambach. Neutrality of the colorflavor-locked phase in a Dyson–Schwinger approach. Physical Review D (Particles and Fields), 77(??):114010, June 10, 2008. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.77.114010.

Nicmorus:2010:DOE

[NEA10] D. Nicmorus, G. Eichmann, and R. Alkofer. Delta and omega electromagnetic form factors in a Dyson–Schwinger/ Bethe–Salpeter approach. Physical Review D (Particles and Fields), 82(??):114017, December 20, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
82.114017.

Needell:1983:FYS

[Nee83] Allan A. Needell, editor. The First 25 years in space: a symposium. Smithsonian Institution Press, Washington, DC, USA, 1983. ISBN 0-87474-668-X. LCCN TL787 .F5 1983.

Neuenschwander:2016:DPD

[Neu16] Dwight E. Neuenschwander, editor. Dear Professor Dyson: twenty years of correspondence between Freeman Dyson and undergraduate students on science, technology, society and life. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2016. ISBN 981-4675-84-9 (hardcover), 981-4675-85-7 (paperback), 981-4675-86-5 (e-book), 981-4759-09-0 (e-book). xiv + 420 pp. LCCN QC16.D95 A4 2016.

Ng:1996:JSP

[Ng96] Yee Jack Ng, editor. Julian Schwinger: the physicist, the teacher, and the man. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1996. ISBN 981-02-2531-8 (hardcover), 981-02-2532-6 (paperback), 981-283-044-8 (e-book). xiv + 195 pp. LCCN QC16.S29 J85 1996.

URL http://site.ebrary.com/id/10691945;   http://www.
myilibrary.com?id=494808.

Nakano:1991:RMB

[NH91] Masahiro Nakano and Akira Hasegawa. Relativistic many-body theory of finite nuclei and the Schwinger–Dyson formalism. Physical Review C (Nuclear Physics), 43(??):618–??, February 1, 1991. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-

4497. URL http://link.aps.org/doi/10.1103/PhysRevC.43.
618.

Nakano:1994:NSDa

[NHKK94] Masahiro Nakano, Akira Hasegawa, Hiroaki Kouno, and Kazuharu Koide. Nuclear Schwinger–Dyson formalism applied to finite baryon density. I. Formulation. Physical Review C (Nuclear Physics), 49(??):3061–??, June 1, 1994. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://

link.aps.org/doi/10.1103/PhysRevC.49.3061.

Nakano:1994:NSDb

[NMM^{+}94] Masahiro Nakano, Tomohiro Mitsumori, Masaru Muraki, Kazuharu Koide, Hiroaki Kouno, and Akira Hasegawa. Nuclear Schwinger–Dyson formalism applied to finite baryon density. II. Numerical calculations and medium effects of meson selfenergies. Physical Review C (Nuclear Physics), 49(??):3076–??, June 1, 1994. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-

490X, 1538-4497. URL http://link.aps.org/doi/10.1103/
PhysRevC.49.3076.

Nakano:1986:LEP

[NN86] K. Nakano and Y. Nogami. Low equation, pion-nucleon scattering, and Castillejo–Dalitz–Dyson pole. Physical Review C (Nuclear Physics), 33(??):2087–??, June 1, 1986. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://

link.aps.org/doi/10.1103/PhysRevC.33.2087.

Naito:1999:ASD

[NO99] K. Naito and M. Oka. Approximation of the Schwinger–Dyson and the Bethe–Salpeter equations and chiral symmetry of QCD. Physical Review C (Nuclear Physics), 59(??):542–??, January 1, 1999. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-

490X, 1538-4497. URL http://link.aps.org/doi/10.1103/
PhysRevC.59.542.

Nocera:2008:TBE

[Noc08] Joe Nocera. Talking business: At Exxon’s can’t-miss meeting. The New York Times Magazine, ??(??):??, May 31, 2008.

ISSN 0362-1308. URL http://www.nytimes.com/2008/05/31/
business/31nocera.html.

Nam:2022:GSE

[NRT22] Phan Thành Nam, Julien Ricaud, and Arnaud Triay. Ground state energy of the low density Bose gas with three-body interactions. Journal of Mathematical Physics, 63(7):071903, July 2022. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic). URL https://pubs.aip.org/aip/jmp/article/63/7/
071903/2843539/Ground-state-energy-of-the-low-density-

Bose-gas. Special collection in honor of Freeman Dyson.

Narayan:1993:DBM

[NS93] Onuttom Narayan and B. Sriram Shastry. Dyson’s Brownian motion and universal dynamics of quantum systems. Physical Review Letters, 71(??):2106–??, September 27, 1993. CODEN PRLTAO. ISSN 0031-9007 (print), 1079-7114 (electronic), 1092-0145. URL

http://link.aps.org/doi/10.1103/PhysRevLett.71.2106.

Nordhaus:2008:QGW

[NZSD08] William D. Nordhaus, Dimitri Zenghelis, Leigh Sullivan, and Freeman J. Dyson. ‘The question of global warming’: An exchange. New York Review of Books, 55(14):92–96, September 25, 2008. ISSN 0028-7504 (print), 1944-7744 (electronic). URL

http://www.nybooks.com/articles/archives/2008/sep/25/
the-question-of-global-warming-an-exchange/. In response

to [Dys08d].

Odifreddi:1996:KAA

[Odi96] Piergiorgio Odifreddi, editor. Kreiseliana: about and around Georg Kreisel. A. K. Peters, Ltd., Wellesley, MA, USA, 1996. ISBN 1-56881-061-X. xiii + 495 pp. LCCN QA29.K73 K74 1996.

Oana:2007:DOI

[OK07] C. Melania Oana and Anna I. Krylov. Dyson orbitals for ionization from the ground and electronically excited states within equation-of-motion coupled-cluster formalism: Theory, implementation, and examples. Journal of Chemical Physics, 127(23): 234106, 2007. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-

7690 (electronic). URL http://link.aip.org/link/?JCP/127/
234106/1.

Oana:2009:CSP

[OK09] C. Melania Oana and Anna I. Krylov. Cross sections and photoelectron angular distributions in photodetachment from negative ions using equation-of-motion coupled-cluster Dyson orbitals.

Journal of Chemical Physics, 131(12):124114, 2009. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL

http://link.aip.org/link/?JCP/131/124114/1.

Okopinska:1991:SSD

[Oko91] Anna Okopińska. Solving Schwinger–Dyson equations by truncation in zero-dimensional scalar quantum field theory. Physical Review D (Particles and Fields), 43(??):3561–??, May 15, 1991. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.43.3561.

Oltmans:1973:IFD

[Olt73a] W. Oltmans. Interview with Freeman Dyson. In On Growth: The Crisis of Exploding Population and Resource Depletion [Olt73b], page ?? ISBN ???? Japanese translation, Japan Publications, Inc. Tokyo 1973. Spanish translation, Fondo de Cultura Economica, Mexico City, 1974.

Oltmans:1973:GCE

[Olt73b] Willem L. Oltmans. On Growth: The Crisis of Exploding Population and Resource Depletion. Bruna Publishers, Utrecht, The Netherlands, 1973. ISBN ???? ???? pp.

Oltmans:1974:IFD

[Olt74a] Willem L. Oltmans. Interview with Freeman Dyson. In On growth [Olt74b], chapter 60, pages 405–413. ISBN 0-399-11233-2.

LCCN HC59 .O563. URL http://catalog.hathitrust.org/
api/volumes/oclc/979019.html.

Oltmans:1974:G

[Olt74b] Willem L. Oltmans, editor. On growth. Capricorn Books, New York, 1974. ISBN 0-399-11233-2. xii + 493 pp. LCCN HC59

.O563. URL http://catalog.hathitrust.org/api/volumes/
oclc/979019.html.

Osada:2023:EUD

[OO23] Hirofumi Osada and Shota Osada. Ergodicity of unlabeled dynamics of Dyson’s model in infinite dimensions. Journal of Mathematical Physics, 64(3):043505, March 2023. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic). URL https://pubs.aip.org/aip/jmp/article/64/
4/043505/2877773/Ergodicity-of-unlabeled-dynamics-of-

Dyson-s-model?searchresult=1. Special collection in honor of

Freeman Dyson.

Oppenheimer:1989:AVE

[Opp89] J. Robert Oppenheimer. Atom and void: essays on science and community. Princeton science library. Princeton University Press, Princeton, NJ, USA, 1989. ISBN 0-691-08547-1, 0-691-02434-0 (paperback). xi + 155 pp. LCCN Q175

.O65 1989. URL http://www.loc.gov/catdir/description/

prin031/89010413.html. Preface by Freeman J. Dyson.

Ortiz:2004:BOD

[Ort04] J. V. Ortiz. Brueckner orbitals, Dyson orbitals, and correlation potentials. International Journal of Quantum Chemistry, 100(6): 1131–1135, December 20, 2004. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Overbye:2008:JWP

[Ove08] Dennis Overbye. John A. Wheeler, physicist who coined the term ‘black hole,’ is dead at 96. New York Times, ??(??):B7, April 14, 2008. CODEN NYTIAO. ISSN 0362-4331 (print),

1542-667X, 1553-8095. URL https://search.proquest.com/
hnpnewyorktimes/docview/897768871/.

Park:1987:AAP

[Par87] P. Park. Algebraic analysis of physical and spurious states in Dyson boson mapping. Physical Review C (Nuclear Physics), 35 (??):807–??, February 1, 1987. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.35.807. See comment [HGE88].

Pastur:2022:EDL

[Pas22] L. Pastur. Eigenvalue distribution of large random matrices arising in deep neural networks: Orthogonal case. Journal of Mathematical Physics, 63(6):063505, June 2022. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

URL https://pubs.aip.org/aip/jmp/article/63/6/063505/
2846100/Eigenvalue-distribution-of-large-random-matrices.

Special collection in honor of Freeman Dyson.

Patton:2018:M

[Pat18] Lydia Patton. In memoriam. HOPOS: Journal of the International Society for the History of Philosophy of Science, 8(1):??, Spring 2018. CODEN ???? ISSN 2152-5188 (print), 2156-6240 (electronic).

Pais:2006:JRO

[PC06] Abraham Pais and Robert P. Crease. J. Robert Oppenheimer: a life. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 2006. ISBN 0-19-516673-6. xxii + 353 + 16 pp. LCCN QC16.O62 P35 2006. URL http://www.loc.gov/

catdir/enhancements/fy0636/2005002173-d.html;   http:
//www.loc.gov/catdir/enhancements/fy0723/2005002173-b.
html;   http://www.loc.gov/catdir/toc/ecip056/2005002173.
html.

Pratt:2008:LEB

[PD08] D. A. Pratt and Freeman J. Dyson. Letter to the Editor: The brief life of a molecule. New York Review of Books, 55(12): 53, July 17, 2008. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2008/jul/17/the-brief-life-of-a-molecule/.

Press:2012:IPD

[PD12] William H. Press and Freeman J. Dyson. Iterated Prisoner’s Dilemma contains strategies that dominate any evolutionary opponent. Proceedings of the National Academy of Sciences of the United States of America, 109(26):10409–10413, June 2012. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490

(electronic). URL http://labs.adsabs.harvard.edu/ui/abs/
2012PNAS..10910409P.

Patterson:2013:BVR

[PD13] James T. Patterson and Freeman Dyson. Birmingham and voting rights. New York Review of Books, 60(19):??, December 5, 2013. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
2013/dec/05/birmingham-and-voting-rights/. See [Dys13b].

Postol:1980:RU

[PDMS80] Theodore A. Postol, Freeman J. Dyson, E. Paul McClain, and William R. Stratton. Radioactivity in the Urals. Science, 208 (4445):652, 654–655, May 16, 1980. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http://www.jstor.

org/stable/1683824;   http://www.sciencemag.org/content/
208/4445/652.2.full.pdf.

Peierls:1980:NFD

[Pei80] Rudolf Peierls. Nuclear family: Disturbing the Universe by Freeman Dyson Harper and Row, 283 pp, £6.95, November 1979, ISBN 0-06-011108-9 [from Winchester to Orion]. London Review of Books, 2(12):9–10, June 19, 1980. ISSN

0260-9592. URL https://www.lrb.co.uk/v02/n12/rudolf-
peierls/nuclear-family.

Peierls:1985:BPR

[Pei85] Sir Rudolf E. (Rudolf Ernst) Peierls. Bird of Passage: Recollections of a Physicist. Princeton legacy library. Princeton University Press, Princeton, NJ, USA, 1985. ISBN 0-691-08390-8, 0-691-02416-2 (paperback), 0-691-60220-4, 1-4008-5461-X (e-book). xii + 350 + 12 pp. LCCN QC16.P375 A32

1985. URL http://site.ebrary.com/id/10897402;   https:
//www.jstor.org/stable/j.ctt7ztn4b.

Perrin:1968:UDR

[Per68] Robert Perrin. Use of the Dyson representation in currentalgebra calculations. Physical Review, 170(??):1365–??, June 25, 1968. CODEN PHRVAO. ISSN 0031-899X (print), 1536-

6065 (electronic). URL http://link.aps.org/doi/10.1103/
PhysRev.170.1365.

Pierre:1984:BRB

[Pie84] Andrew J. Pierre. Book review: Weapons and Hope by Freeman Dyson. Foreign Affairs, 62(5):1248, Summer 1984. CODEN FR-

NAA3. ISSN 0015-7120. URL http://www.jstor.org/stable/
20042015.

Pitici:2011:BWM

[Pit11] Mircea Pitici, editor. The best writing on mathematics 2010. Princeton University Press, Princeton, NJ, USA, 2011. ISBN 0-691-14841-4. xxviii + 440 pp. LCCN ???? Foreword by William P. Thurston.

Pitici:2012:BWM

[Pit12] Mircea Pitici, editor. The best writing on mathematics 2011. Princeton University Press, Princeton, NJ, USA, 2012. ISBN 0-691-15315-9 (paperback). xxx + 383 pp. LCCN QA8.6.

Phua:2014:PCH

[PKpCH14] K. K. Phua, Leong Chuan Kwek, Ngee pong Chang, and Chan A. H., editors. Proceedings of the conference in honour of the 90th birthday of Freeman Dyson: Nanyang Technological University, Singapore, 26–29 August 2013. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2014. ISBN 981-4590-10-X (hardcover), 981-4590-70-3 (paperback), 981-4590-12-6 (e-book). LCCN ????

URL http://www.worldscientific.com/doi/abs/10.1142/
9789814590112_0001;   http://www.worldscientific.com/
worldscibooks/10.1142/9098.

Peterson:1983:TLH

[PM^{+}83] Grethe B. Peterson, Sterling M. McCurrin, et al., editors. The Tanner Lectures on human values, volume 4. University of Utah Press, Salt Lake City, UT, USA, 1983. ISBN 0-585-19772-5 (ebook), 0-87480-216-4. 254 pp. LCCN BD232 .T24 1983eb.

Pombo:2009:NCD

[Pom09] Claudia Pombo. A new comment on Dyson’s exposition of Feynman’s proof of Maxwell equations. AIP Conference Proceedings, 1101(1):363–367, 2009. CODEN APCPCS. ISSN 0094-243X (print), 1551-7616 (electronic), 1935-0465. URL http://link.

aip.org/link/?APC/1101/363/1.

Pool:1972:MAS

[Poo72] James C. T. Pool, editor. Mathematical aspects of statistical mechanics, volume 5 of SIAM-AMS proceedings. American Mathematical Society, Providence, RI, USA, 1972. ISBN 0-8218-1324-2. LCCN QC174.7 .S92 1971.

Pak:2009:WLD

[PR09] M. Pak and H. Reinhardt. Wilson loop from a Dyson equation. Physical Review D (Particles and Fields), 80(??):125022, December 21, 2009. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.80.125022.

Praveczki:1964:DUF

[Pra64] E. Praveczki. Definition and use of ferromagnon annihilation operators in the Bloch–Dyson theory of ferromagnons. Physica Status Solidi. B, Basic Research, 5(3):481–494, 1964. CODEN PSS-BBD. ISSN 0370-1972 (print), 1521-3951 (electronic).

Pino:2004:LTD

[PS04] Ramiro Pino and Gustavo E. Scuseria. Laplace-transformed diagonal Dyson correction to quasiparticle energies in periodic systems. Journal of Chemical Physics, 121(6):2553–2557, 2004. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://link.aip.org/link/?JCP/121/2553/1.

Pei:2007:CDW

[PSP^{+}07] Minxin Pei, Qiao Shi, Henry Paulson, Bobby Ray Inman, Anthony Lake, Benazir Bhutto, Richard Holbrooke, Dennis Ross, Brent Scowcroft, David M. Walker, Leonard Dicaprio, and Freeman Dyson. China: From Democracy Wall to the shopping mall and back — Fall 2007. New Perspectives Quarterly: NPQ, 25(1): 86–90, Fall 2007. ISSN 0893-7850 (print), 1540-5842 (electronic).

Pupin:1960:II

[Pup60] Michael Pupin. From immigrant to inventor, volume SL26 of The Scribner library. Scribner, New York, NY, USA, 1960. 396 pp. LCCN T40.P8 A3 1960. With a foreword by Freeman J. Dyson.

Peirs:2002:SCS

[PVW02] K. Peirs, D. Van Neck, and M. Waroquier. Self-consistent solution of Dyson’s equation up to second order for open-shell atomic systems. Journal of Chemical Physics, 117(9):4095–4105, 2002. CO-DEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic).

URL http://link.aip.org/link/?JCP/117/4095/1.

Peirs:2003:SCS

[PVW03] K. Peirs, D. Van Neck, and M. Waroquier. Self-consistent solution of Dyson’s equation up to second order for closed- and open-shell atomic systems. International Journal of Quantum Chemistry, 91(2):113–118, 2003. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Rabi:1978:CFA

[Rab78] I. I. (Isidor Isaac) Rabi, editor. Celebration of the fiftieth anniversary of the Pupin Laboratories, Columbia University in the City of New York. Columbia University, New York, NY, USA, 1978. LCCN ????

Radford:2020:FDO

[Rad20] Tim Radford. Freeman Dyson obituary: Brilliant theoretical physicist and mathematician whose far-fetched ideas for the future verged on the bizarre. The Guardian, ??(??):??, March 1, 2020. ISSN 0261-3077 (print), 1756-3224 (elec-

tronic). URL https://www.theguardian.com/science/2020/
mar/01/freeman-dyson-obituary.

Radozycki:1995:FNS

[RBB95] Tomasz Radozycki and Iwo Bialynicki-Birula. Finite nonperturbative solutions of Dyson–Schwinger equations in QED in the infrared domain. Physical Review D (Particles and Fields), 52 (??):2439–??, August 15, 1995. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.52.2439.

Rees:1999:OCC

[RD99] Martin J. Rees and Freeman J. Dyson. Our concepts of the cosmos: progress, prospects and mysteries. In Bearn [Bea99], page ?? ISBN 0-87169-234-1 (hardcover). ISSN 0065-9738. LCCN Q11.P612 vol. 234.

Rembiesa:1990:GIB

[Rem90] Peter Rembiesa. Gauge-independent bifurcation to the chiralsymmetry-breaking solution of the Dyson–Schwinger equation in continuum QED. Physical Review D (Particles and Fields), 41 (??):2009–??, March 15, 1990. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.41.2009.

Revkin:2009:SIT

[Rev09] Andrew C. Revkin. Some inconvenient thinkers. The New York Times Magazine, ??(??):??, March 25, 2009. ISSN 0362-

1308. URL http://dotearth.blogs.nytimes.com/2009/03/
25/some-inconvenient-thinkers/.

Rabinowitch:1969:MM

[RL69] E. Rabinowitch and R. S. Lewis, editors. Man on the Moon. Basic Books, New York, NY, USA, 1969. ???? pp.

Rindorf:2006:COW

[RM06] Lars Rindorf and Niels Asger Mortensen. Calculation of opticalwaveguide grating characteristics using Green’s functions and Dyson’s equation. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 74(??):036616, September 18, 2006. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.74.036616.

Rabson:2004:CPW

[RNM04] D. A. Rabson, B. N. Narozhny, and A. J. Millis. Crossover from Poisson to Wigner–Dyson level statistics in spin chains with integrability breaking. Physical Review B: Condensed Matter and Materials Physics, 69(??):054403, February 9, 2004. CODEN

PRBMDO. ISSN 1098-0121. URL http://link.aps.org/doi/
10.1103/PhysRevB.69.054403.

Rockmore:1971:UCD

[Roc71] Ronald Rockmore. Unitarization, Castillejo–Dalitz–Dyson poles, and scalar companions. Physical Review D (Particles and Fields), 4(??):900–??, August 1, 1971. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.4.900.

Rogers:1997:EBH

[Rog97] Pattiann Rogers. Eating bread and honey. Milkweed Editions, Minneapolis, MN, 1997. ISBN 1-57131-406-7. 95 pp. LCCN PS3568.O454 E38 1997.

Romm:1989:BRB

[Rom89] Joseph J. Romm. Book review: Infinite in All Directions, by Freeman Dyson. Bulletin of the Atomic Scientists, 45(8):35–36, October 1989. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic).

Rosa:1962:LEB

[Ros62] Nicholas Rosa. Letter to the Editor: Buying time. Bulletin of the Atomic Scientists, 18(9):32, November 1962. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys62j].

Rosenblatt:1963:PST

[Ros63] Murray Rosenblatt, editor. Proceedings of the Symposium on Time Series Analysis, held at Brown University, June 11–14, 1962. Wiley, New York, NY, USA, 1963. LCCN QA280.

Rodriguez-Quintero:2011:BRS

[RQ11] J. Rodríguez-Quintero. Brief remarks on the similarities of the infrared solutions for the ghost propagator Dyson–Schwinger equation in Landau and Coulomb gauges. Physical Review D (Particles and Fields), 83(??):097501, May 5, 2011. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
83.097501.

Rosenbluth:1973:NFA

[RRD^{+}73] M. N. Rosenbluth, M. Ruderman, F. Dyson, J. N. Bahcall, J. Shaham, and J. Ostriker. Nuclear fusion in accreting neutron stars. Astrophysical Journal, 184(??):907, ???? 1973. CODEN ASJOAB. ISSN 0004-637X (print), 1538-4357 (electronic). URL

http://adsabs.harvard.edu/abs/1973ApJ...184..907R.

Recht:1948:APS

[RRS48] Leon Recht, Martin Rosenbaum, and E. P. Starke. Advanced problems and solutions: 4247. American Mathematical Monthly, 55(9):588–592, November 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Thé47].

Ring:1977:MCD

[RS77] P. Ring and P. Schuck. Mode coupling and Dyson’s finite boson expansion in nuclei. Physical Review C (Nuclear Physics), 16 (??):801–??, August 1, 1977. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.16.801.

Rastelli:1982:DMT

[RT82] E. Rastelli and A. Tassi. Dyson Maleev transformation in ferromagnets with single ion planar anisotropy. Journal of Applied Physics, 53(3):1867–1869, 1982. CODEN JAPIAU. ISSN 0021-8979 (print), 1089-7550 (electronic), 1520-8850. URL http://

link.aip.org/link/?JAP/53/1867/1.

Ruse:2012:PSU

[Rus12] Michael Ruse. Philosophy’s strain of unevolved thinking. The Chronicle of Higher Education, ??(??):??, November 26, 2012. ISSN 0009-5982 (print), 1931-1362 (electronic). URL http://

chronicle.com/article/Philosophys-Strain-of/135872/.

Rapport:1964:P

[RW64] Samuel Rapport and Helen Wright, editors. Physics. New York University Press, New York City, NY, USA, 1964. xiii + 333 pp.

Rudnick:2021:RPR

[RW21] Zeév Rudnick and Igor Wigman. The Robin problem on rectangles. Journal of Mathematical Physics, 62(11):113503, November 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Richter:1978:OTF

[RWP78] Wolfgang Richter, R. N. (Ray N.) Wilson, and F. (Franco) Pacini, editors. Optical telescopes of the future: ESO conference, Geneva, 12–15 December 1977: proceedings. Southern European Observatory, Geneva, Switzerland, 1978. LCCN QB88 .O685

1977. URL http://catalog.hathitrust.org/api/volumes/
oclc/4374966.html.

Ruskai:2023:LAR

[RY23] Mary Beth Ruskai and Jon Yard. Local additivity revisited. Journal of Mathematical Physics, 64(3):032201, March 2023. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic).

URL https://pubs.aip.org/aip/jmp/article/64/3/032201/
2881610/Local-additivity-revisited. Special collection in

honor of Freeman Dyson.

Sauli:2003:SRB

[SA03] Vladimír Sauli and J. Adam, Jr. Study of relativistic bound states for scalar theories in the Bethe–Salpeter and Dyson–Schwinger formalism. Physical Review D (Particles and Fields), 67(??): 085007, April 21, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.67.085007.

Sagan:1973:CEI

[Sag73] Carl Sagan, editor. Communication with Extraterrestrial Intelligence (CETI). MIT Press, Cambridge, MA, USA, 1973. ISBN 0-262-19106-7. LCCN QB54 .S68 1971a.

Sagan:2000:CSC

[Sag00] Carl Sagan, editor. Carl Sagan’s cosmic connection: an extraterrestrial perspective. Cambridge University Press, Cambridge, UK, 2000. ISBN 0-521-78303-8 (hardback). xxxi + 302 pp. LCCN QB54 .S24 2000. URL http://www.loc.

gov/catdir/description/cam021/00020378.html;   http:/

/www.loc.gov/catdir/enhancements/fy0731/00020378-b.html;
 http://www.loc.gov/catdir/samples/cam031/00020378.html;
 http://www.loc.gov/catdir/toc/cam021/00020378.html.

Sandham:1950:APSc

[San50] H. F. Sandham. Advanced problems and solutions: 4318. American Mathematical Monthly, 57(5):345–346, May 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wil48].

Sankaran:1994:PGN

[San94] Neeraja Sankaran. Physicist and geneticist are named winners of Enrico Fermi Award — Freeman Dyson and Liane B. Russell. The Scientist (Philadelphia, PA), 8(21):4–5, October 1994. ISSN 0890-3670 (print), 1945-5127 (electronic). URL http://

www.the-scientist.com/?articles.view/articleNo/28066/.

Sarkar:2005:TTG

[Sar05] Anoop Sarkar. The tragic tale of a genius by Freeman Dyson. Web glob article, July 20, 2005. URL http://www.cs.sfu.ca/~

anoop/weblog/archives/000158.html. Comment on [Dys05g].

Saxon:1994:BRR

[Sax94] David S. Saxon. Book review: A reshaping in physics: QED and the Men Who Made It, Dyson, Feynman, Schwinger, and Tomonaga. Science, 266(5192):1888–1890, December 16, 1994. CODEN SCIEAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL

http://www.jstor.org/stable/2885047.

Solovej:2021:SCH

[SBZ21] Jan Philip Solovej, Rafael Benguria, and Martin Zirnbauer. Special collection in honor of Freeman Dyson. American Institute of Physics Web site., 2021. URL https:/

/pubs.aip.org/jmp/collection/1381/Special-collection-

in-honor-of-Freeman-Dyson. Special collection in honor of Freeman Dyson.

Schwinger:1958:SPQ

[Sch58] Julian Schwinger, editor. Selected Papers on Quantum Electrodynamics. Dover books on engineering and engineering physics. Dover Publications, Inc., New York, NY, USA, 1958. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35.

Schweber:1994:QMW

[Sch94a] S. S. (Silvan S.) Schweber. QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga. Princeton series in physics. Princeton University Press, Princeton, NJ, USA, 1994. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). xxviii + 732 pp. LCCN QC680 .S34 1994. US$72.50. URL http://www.loc.

gov/catdir/description/prin021/93033550.html;   http:
//www.loc.gov/catdir/toc/prin031/93033550.html.

Schweber:1994:PTS

[Sch94b] Sylvan S. Schweber. A postscript: Tomonaga, Schwinger, Feynman, and Dyson. In QED and the men who made it: Dyson, Feynman, Schwinger, and Tomonaga [Sch94a], pages 572–575. ISBN 0-691-03685-3, 0-691-03327-7 (paperback). LCCN QC680 .S34 1994. US$72.50. URL http://www.loc.

gov/catdir/description/prin021/93033550.html;   http:
//www.loc.gov/catdir/toc/prin031/93033550.html.

Schwinger:2003:SPQ

[Sch03] Julian Schwinger. Selected Papers on Quantum Electrodynamics. Dover Publications, Inc., New York, NY, USA, 2003. ISBN 0-486-60444-6. xvii + 424 pp. LCCN QC680 .S35. Reprint of [Sch58] with ISBN.

Schewe:2013:MGP

[Sch13] Phillip F. Schewe. Maverick genius: the pioneering odyssey of Freeman Dyson. Thomas Dunn Books, St. Martin’s Press, New York, NY, USA, 2013. ISBN 0-312-64235-0 (hardcover), 1-250-02101-4 (e-book). x + 339 pp. LCCN QC16.D95 S34 2013.

Smith:1998:LEW

[SD98] Huston Smith and Freeman J. Dyson. Letter to the Editor: Words and things. New York Review of Books, 45(12): 57, July 16, 1998. ISSN 0028-7504 (print), 1944-7744 (elec-

tronic). URL http://www.nybooks.com/articles/archives/
1998/jul/16/words-and-things/. In response to [Dys98b].

Schwartz:2006:LER

[SD06] Stephen P. Schwartz and Freeman J. Dyson. Letter to the Editor: Respect for our enemies. New York Review of Books, 53(17):63, November 2, 2006. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/

archives/2006/nov/02/respect-for-our-enemies/. In re-

sponse to [Dys06d].

Strawson:2015:WYG

[SD15] Galen Strawson and Freeman Dyson. When you got to go. New York Review of Books, 62(10):??, June 4, 2015. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2015/jun/04/when-you-got-go/. Com-

ments on [Dys15a].

Susskind:2006:LES

[SDGK06] Leonard Susskind, Freeman Dyson, David Gross, and Walter Kohn. Letter to the Editor: Scientists speak out about Guantanamo. New York Times, ??(??):C13, April 30, 2006. CO-DEN NYTIAO. ISSN 0362-4331 (print), 1542-667X, 1553-

8095. URL http://search.proquest.com/hnpnewyorktimes/
docview/93226596/.

Stuckey:2001:LES

[SDL^{+}01] W. Mark Stuckey, Freeman Dyson, Arthur S. Lodge, Matthew Housley, and Mark Friesel. Letters to the Editor: Science, religion, Templeton Prize. Physics Today, 54(8): 74, August 2001. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://labs.adsabs.

harvard.edu/ui/abs/2001PhT....54h..74D;   http://www.
physicstoday.org/resource/1/phtoad/v54/i8/p72_s2.

Sheikh:1986:ATS

[SG86] J. A. Sheikh and Y. K. Gambhir. Analysis of the truncation schemes for the physical boson states with Dyson’s description. Physical Review C (Nuclear Physics), 34(??):2344–??, December 1, 1986. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-

490X, 1538-4497. URL http://link.aps.org/doi/10.1103/
PhysRevC.34.2344.

Sheikh:1993:NND

[SG93] J. A. Sheikh and M. Gupta. Nonunitary nature of the Dyson boson mapping revisited. Physical Review C (Nuclear Physics), 47(??):2998–??, June 1, 1993. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.47.2998.

Snyman:2006:SCL

[SG06] I. Snyman and H. B. Geyer. Strong-coupling limit of the Richardson Hamiltonian analyzed using Dyson mapping. Physical Review B: Condensed Matter and Materials Physics, 73(??):144516, April 26, 2006. CODEN PRBMDO. ISSN 1098-0121. URL http://

link.aps.org/doi/10.1103/PhysRevB.73.144516.

Soffel:1995:DER

[SGT^{+}95] M. Soffel, B. Geyer, H. Teichler, W. Eberhardt, F. Krull, H. Jungblut, H. Kolanoski, H. Fritzsch, and G. Münster. d’Inverno: Einführung in die Relativitätstheorie/Wong: Mathematische Physik/Sutton: Electronic Structure of Materials/ Mönch: Semiconductor Surfaces and Interfaces/Meadows and Hancock-Beaulieu: Front Page Physics/Markvart: Solar Electricity/Zell: Simulation Neuronaler Netze/Enz u. Meÿenn: W. Pauli: Writings on Physics and Philosophy/Mehra: The Beat of a different Drum The Life and Science of Richard Feynman/Miller: Early Quantum Electrodynamics: A Sourcebook/ Schweber: QED and the Men who made it: Dyson, Feynman, Schwinger and Tomonaga. Physikalische Blätter, 51(5):428–434, May 1995. CODEN PHBLAG. ISSN 0031-9279 (print), 1521-3722

(electronic). URL http://onlinelibrary.wiley.com/doi/10.
1002/phbl.19950510519/abstract.

Shapiro:1987:BRB

[Sha87] Robert Shapiro. Book review: Origins of Life, by Freeman Dyson. Quarterly Review of Biology, 62(3):305, September 1987. CODEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic). URL

http://www.jstor.org/stable/2828989.

Shapiro:2001:BRB

[Sha01] Robert Shapiro. Book review: Origins of Life by Freeman Dyson. Origins of Life on the Earth and in the Cosmos by Geoffrey Zubay. Quarterly Review of Biology, 76(1):62, March 2001. CO-DEN QRBIAK. ISSN 0033-5770 (print), 1539-7718 (electronic).

URL http://www.jstor.org/stable/2664141.

Stoller:1991:TSR

[SHD91] S. D. Stoller, W. Happer, and F. J. Dyson. Transverse spin relaxation in inhomogeneous magnetic fields. Physical Review A (Atomic, Molecular, and Optical Physics), 44(11):7459–7477, December 1, 1991. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.aps.org/

doi/10.1103/PhysRevA.44.7459.

Sherman:1982:OWT

[She82] Michael Sherman. Oppenheimer: What a trouble-maker!. book reviews: Reminiscences of Los Alamos, 1943–1945 by Lawrence Badash, Joseph O. Hirschfelder, and Herbert P. Broida. Disturbing the Universe by Freeman Dyson. J. Robert Oppenheimer, Shatterer of Worlds by Peter Goodchild. Robert Oppenheimer, Letters and Recollections by Alice Kimball Smith and Charles Weiner. The Public Historian, 4(4):97–117, Autumn 1982. ISSN 0272-3433 (print), 1533-8576 (electronic). URL http://www.

jstor.org/stable/3377050.

Sheikh:1988:FCB

[She88] J. A. Sheikh. Fermion calculations in the boson space using the Dyson boson mapping. Physical Review C (Nuclear Physics), 37 (??):1295–??, March 1, 1988. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.37.1295.

Sheikh:1989:IBF

[She89] J. A. Sheikh. Intrinsic basis function in the Dyson boson mapping. Physical Review C (Nuclear Physics), 39(??):1641–??, April 1, 1989. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-

490X, 1538-4497. URL http://link.aps.org/doi/10.1103/
PhysRevC.39.1641.

Simon:2022:SGL

[Sim22] Barry Simon. The strong Gauss–Lucas theorem and analyticity of correlation functions via the Lee–Yang theorem. Journal of Mathematical Physics, 63(3):033302, March 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Simons:1993:EPC

[SLA93] B. D. Simons, P. A. Lee, and B. L. Altshuler. Exact pair correlation of the one-dimensional quantum gas with 1/r^{2} repulsion derived from the sympletic Dyson ensemble. Physical Review B: Condensed Matter and Materials Physics, 48(??):11450, October 15, 1993. CODEN PRBMDO. ISSN 1098-0121. URL http://

link.aps.org/doi/10.1103/PhysRevB.48.11450.

Stapledon:2004:SM

[SM04] Olaf Stapledon and Patrick A. McCarthy, editors. Star maker. The Wesleyan early classics of science fiction series. Wesleyan University Press, Middletown, CT, USA, 2004. ISBN 0-8195-6692-6 (hardcover), 0-8195-6693-4 (paperback). xxxiii + 314 pp. LCCN

PR6037.T18 S7 2004. URL http://www.loc.gov/catdir/toc/

fy045/2004299753.html. Original edition 1937. New edition introduced and Edited by Patrick A. McCarthy. Foreword by Freeman Dyson.

Smith:2006:IFD

[Smi06] David Smith. Interview: Freeman Dyson spills the beans. Physics Education, 41(2):180–182, March 2006. CODEN PHEDA7. ISSN 0031-9120 (print), 1361-6552 (electronic). URL http://stacks.

iop.org/0031-9120/41/i=2/a=M02.

Solomon:2013:FCF

[Sol13] Lawrence Solomon. Fighting climate ‘fluff’. National Post, ??

(??):??, ???? 2013. URL http://www.nationalpost.com/news/
story.html?id=985641c9-8594-43c2-802d-947d65555e8e.

Singh:2010:TFA

[SOM10] Raman K. Singh, J. V. Ortiz, and Manoj K. Mishra. Tautomeric forms of adenine: Vertical ionization energies and Dyson orbitals. International Journal of Quantum Chemistry, 110(10): 1901–1915, August 15, 2010. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Spears:1991:BRC

[Spe91] Monroe K. Spears. Book review: Cosmology and the common reader: What Do You Care What Other People Think? by Richard P. Feynman. Infinite in All Directions by Freeman J. Dyson. A Brief History of Time from the Big Bang to Black Holes by Stephen W. Hawking. The Big Bang by Joseph Silk. The Anthropic Cosmological Principle by John D. Barrow and Frank J. Tipler. The Sewanee Review, 99(1):113–121, Winter 1991. ISSN 0037-3052 (print), 1934-421X (electronic). URL http://www.

jstor.org/stable/27546317.

Spohn:2022:HEA

[Spo22] Herbert Spohn. Hydrodynamic equations for the Ablowitz–Ladik discretization of the nonlinear Schrödinger equation. Journal of Mathematical Physics, 63(3):033305, March 2022. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Spanner:2012:DNX

[SPZ^{+}12] Michael Spanner, Serguei Patchkovskii, Congyi Zhou, Spiridoula Matsika, Marija Kotur, and Thomas C. Weinacht. Dyson norms in XUV and strong-field ionization of polyatomics: Cytosine and uracil. Physical Review A (Atomic, Molecular, and Optical Physics), 86(??):053406, November 5, 2012. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL

http://link.aps.org/doi/10.1103/PhysRevA.86.053406.

Szczepaniak:2011:SDE

[SR11] Adam P. Szczepaniak and Hugo Reinhardt. Schwinger–Dyson equations and disorder. Physical Review D (Particles and Fields), 84(??):056011, September 16, 2011. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-

ROM). URL http://link.aps.org/doi/10.1103/PhysRevD.
84.056011.

Stampfer:2005:PSA

[SRBW05] Christoph Stampfer, Stefan Rotter, Joachim Burgdörfer, and Ludger Wirtz. Pseudopath semiclassical approximation to transport through open quantum billiards: Dyson equation for diffractive scattering. Physical Review E (Statistical physics, plasmas, fluids, and related interdisciplinary topics), 72(??):036223, September 30, 2005. CODEN PLEEE8. ISSN 1539-3755 (print),

1550-2376 (electronic). URL http://link.aps.org/doi/10.
1103/PhysRevE.72.036223.

Schroer:1967:RAJ

[SS67] B. Schroer and P. Stichel. Remarks on the application of the Jost–Lehmann–Dyson representation to equal-time commutators. Physical Review, 162(??):1394–??, October 25, 1967. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.162.1394.

Surjan:1998:DCO

[SS98] Péter R. Surján and Ágnes Szabados. Dyson-corrected orbital energies for the perturbative treatment of electron correlation. International Journal of Quantum Chemistry, 69(6):713–719, ????

1998. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-
bin/abstract?ID=74988;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=74988&PLACEBO=IE.pdf.

Schafer:1999:DEA

[SS99] Steffen Schäfer and Peter Schuck. Dyson equation approach to many-body Green’s functions and self-consistent RPA: Application to the Hubbard model. Physical Review B: Condensed Matter and Materials Physics, 59(??):1712–??, January 15, 1999. CO-DEN PRBMDO. ISSN 1098-0121. URL http://link.aps.org/

doi/10.1103/PhysRevB.59.1712.

Sheth:1967:ECD

[ST67] Bidhanchandra I. Sheth and Arnold Tubis. Effect of a Castillejo– Dalitz–Dyson pole in the πN I = 1/2, J = (1/2)^{+} amplitude. Physical Review, 154(??):1322–??, February 25, 1967. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.154.1322.

Saikia:2011:FFB

[ST11] D. J. Saikia and Virginia Trimble, editors. Fluid flows to black holes: a tribute to S. Chandrasekhar on his birth centenary. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2011. ISBN 981-4374-76-8 (hardcover), 981-4374-77-6 (e-book). LCCN QB461 .F58 2011. URL http://ebooks.

worldscinet.com/ISBN/9789814374774/9789814374774.html.

Sasagawa:2012:EIP

[ST12] S. Sasagawa and H. Tanaka. Effect of an imaginary part of the Schwinger–Dyson equation at finite temperature and density. Physical Review C (Nuclear Physics), 85(??):045201, April 4, 2012. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-

490X, 1538-4497. URL http://link.aps.org/doi/10.1103/
PhysRevC.85.045201.

Shen:2020:VMP

[ST20] Allan Shen and Edward Tian. Venerated mathematical physicist Freeman J. Dyson dies at 96. The Daily Princetonian, ??(??):??,

March 5, 2020. URL https://www.dailyprincetonian.com/
article/2020/03/venerated-mathematical-physicist-and-
technology-visionary-freeman-j-dyson-dies-at-96.

Staal:1998:STR

[Sta98] F. Staal. Science & theology (response to Freeman Dyson’s review of books by Feynman and Polkinghorne). New York Review of Books, 45(18):76–77, November 19, 1998. ISSN 0028-7504 (print), 1944-7744 (electronic).

Staley:2003:BRI

[Sta03] Richard Staley. Book reviews: Interdisciplinary atomism? Exploring Twentieth-Century culture through Einstein Marcia Bartusiak, Einstein’s Unfinished Symphony: Listening to the Sounds of Space–Time. Washington, DC: Joseph Henry Press, 2000. Pp. xii + 249. ISBN 0-309-06987-4. £17.95 (hardback). Alice Calaprice (ed.), The Expanded Quotable Einstein. With a foreword by Freeman Dyson. Princeton and Oxford: Princeton University Press, 2000. Pp. xliii+407. ISBN 0-691-07021-0. £11.95, $18.95 (cloth). Klaus Hentschel (ed.), Physics and National Socialism: An Anthology of Primary Sources. Ann M. Hentschel, Editorial Assistant and Translator. Erwin Hiebert and Hans Wussing (eds.), Science Networks: Historical Studies, 18. Basel, Boston and Berlin: Birkhäuser, 1996. Pp. ci+406+civ. ISBN 3-7643-5312-0. DM 178.00, SFR 148.00, EUR98.00 (cloth). Gerald Holton, Einstein, History, and Other Passions: The Rebellion Against Science at the End of the Twentieth Century. Cambridge and London: Harvard University Press, 2000. Pp. xii+240. ISBN 0-674-00433-7. £12.50 (paperback). Don Howard and John Stachel (eds.), Einstein: The Formative Years, 1879–1909. Einstein Studies, 8. Boston, Basel and Berlin: Birkhäuser, 2000. Pp. xi+258. ISBN 0-8176-4030-4. DM 128.00, EUR83.00, SFR 124.00. Arthur I. Miller, Einstein, Picasso: Space, Time and the Beauty that Causes Havoc. New York: Basic Books, 2001. Pp. x+357. ISBN 0-465-01859-9. $30.00, CAN$44.95 (cloth). British Journal for the History of Science, 36(2):221–230, June 2003. CODEN BJHSAT. ISSN 0007-0874 (print), 1474-001X (electronic). URL

http://www.jstor.org/stable/4028234.

Stewart:1992:BRB

[Ste92] Albert B. Stewart. Book review: From Eros to Gaia by Freeman Dyson. The Antioch Review, 50(4):774–775, Autumn 1992. ISSN

0003-5769. URL http://www.jstor.org/stable/4612639.

Steinberg:2004:TDF

[Ste04] Daniel H. Steinberg. Two degrees of freedom. Web document,

July 30, 2004. URL http://www.onlamp.com/pub/a/onlamp/

2004/07/30/dysons.html. Description of talk by George Dyson and Freeman Dyson.

Stone:2020:OFD

[Sto20] Andrea Stone. Obituary: Freeman Dyson, legendary theoretical physicist, dies at 96. National Geographic, ??(??):??, February 28,

2020. URL https://www.nationalgeographic.com/science/
2020/02/freeman-dyson-legendary-theoretical-physicist-
dies-at-96/.

Strick:2000:BRC

[Str00] James Strick. Book review: The Cambrian explosion (of books on the origin of life) Life on Other Worlds: The 20th Century Extraterrestrial Life Debate by Steven J. Dick. Origins of Life by Freeman Dyson. The Emergence of Life on Earth: A Historical and Scientific Overview by Iris Fry. Biogenesis: Theories of Life’s Origin by Noam Lahav. The Origins of Life: From the Birth of Life to the Origin of Language by John Maynard Smith and Eörs Szathmáry. Journal of the History of Biology, 33(2):371–384,

Autumn 2000. URL http://www.jstor.org/stable/4331589.

Steiner:2008:SNL

[SV^{+}08] George Steiner, Emilio Rui Vilar, et al., editors. Is science nearing its limits? Carcanet Press Ltd., Manchester. UK, 2008. ISBN 1-

84777-007-X. LCCN ???? URL http://www.carcanet.co.uk/
cgi-bin/indexer?product=9781847770073.

Storchi:2009:IUD

[SVT09] Loriano Storchi, Giuseppe Vitillaro, and Francesco Tarantelli. Implementation and use of a direct, partially integral-driven non-Dyson propagator method for molecular ionization. Journal of Computational Chemistry, 30(5):818–825, April 15, 2009. CO-DEN JCCHDD. ISSN 0192-8651 (print), 1096-987X (electronic).

Sagan:1966:IDD

[SW66] Carl Sagan and Russell G. Walker. The infrared detectability of Dyson civilizations. Astrophysical Journal, 144(??):1216– 1218, ???? 1966. CODEN ASJOAB. ISSN 0004-637X (print),

1538-4357 (electronic). URL http://adsabs.harvard.edu/abs/
1966ApJ...144.1216S. See [Dys60k].

Salam:1972:AQT

[SW72] Abdus Salam and Eugene Paul Wigner, editors. Aspects of quantum theory. Cambridge University Press, Cambridge, UK, 1972. ISBN 0-521-08600-0. xvi + 268 pp.

LCCN QC174.1 .A85 1972. URL http://hooke.lib.cam.ac.
uk/cgi-bin/bib_seek.cgi?cat=ul&bib=1733506;   http://
www.loc.gov/catdir/enhancements/fy1001/72075298-d.html;
 http://www.loc.gov/catdir/enhancements/fy1001/72075298-
t.html.

Sweet:1988:ACA

[Swe88] William Sweet. AIP Corporate Associates meet at IBM; Dyson and Riordan are honored. Physics Today, 41(12):91–94, 1988. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (elec-

tronic). URL http://link.aip.org/link/?PTO/41/91/1.

Saha:2005:CBC

[SWFB05] Saumitra Saha, Feng Wang, Chantal T. Falzon, and Michael J. Brunger. Coexistence of 1,3-butadiene conformers in ionization energies and Dyson orbitals. Journal of Chemical Physics, 123 (12):124315, 2005. CODEN JCPSA6. ISSN 0021-9606 (print),

1089-7690 (electronic). URL http://link.aip.org/link/?JCP/
123/124315/1.

Xue:1987:PAS

[sXcHmW87] She sheng Xue, Ting chang Hsien, and Chi min Wu. Plaquette average and Schwinger–Dyson equation of a Wilson loop in SU(2) lattice gauge theory. Physical Review D (Particles and Fields), 36(??):3203–??, November 15, 1987. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM).

URL http://link.aps.org/doi/10.1103/PhysRevD.36.3203.

Sills:2006:DDC

[SZ06] Andrew V. Sills and Doron Zeilberger. Disturbing the Dyson conjecture (in a GOOD way). Experimental Mathematics, 15(2): 187–191, ???? 2006. CODEN ???? ISSN 1058-6458 (print), 1944-

950X (electronic). URL http://projecteuclid.org/euclid.
em/1175789739.

Subag:2021:CCS

[SZ21] Eliran Subag and Ofer Zeitouni. Concentration of the complexity of spherical pure p-spin models at arbitrary energies. Journal of Mathematical Physics, 62(12):123301, December 2021. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. Special collection in honor of Freeman Dyson.

Takada:1986:HTD

[Tak86] K. Takada. Hermitian treatment of Dyson boson theory. Physical Review C (Nuclear Physics), 34(??):750–??, August 1, 1986. CO-DEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

URL http://link.aps.org/doi/10.1103/PhysRevC.34.750.

Takada:1988:HTD

[Tak88] K. Takada. Hermitian treatment of Dyson boson theory. Physical Review C (Nuclear Physics), 38(??):2450–??, November 1, 1988. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-

4497. URL http://link.aps.org/doi/10.1103/PhysRevC.38.
2450.

Takeuchi:1989:ANS

[Tak89] Tatsu Takeuchi. Analytical and numerical study of the Schwinger–Dyson equation with four-fermion coupling. Physical Review D (Particles and Fields), 40(??):2697–??, October 15, 1989. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.40.2697.

Tamura:1983:RDB

[Tam83] Taro Tamura. Reformulation of Dyson’s boson expansion theory. Physical Review C (Nuclear Physics), 28(??):2480–??, December 1, 1983. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-

490X, 1538-4497. URL http://link.aps.org/doi/10.1103/
PhysRevC.28.2480.

Taylor:1969:LED

[Tay69] William Palmer Taylor. Letter to the Editor: Dyson views challenged. Bulletin of the Atomic Scientists, 25(9):46, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys69a].

Thebault:1949:APSr

[TB49] Victor Thébault and Robert Bouvaist. Advanced problems and solutions: Solutions: 4248. American Mathematical Monthly, 56 (10):696–697, December 1949. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTD^{+}47, Thé47].

Tarantelli:1992:CIF

[TC92] A. Tarantelli and L. S. Cederbaum. Configuration-interaction formulation of the Dyson equation. Physical Review A (Atomic, Molecular, and Optical Physics), 45(??):2790–??, March 1, 1992. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-

4446, 1538-4519. URL http://link.aps.org/doi/10.1103/
PhysRevA.45.2790.

Troutman:2005:LEN

[TD05] Bill Troutman and Freeman J. Dyson. Letter to the Editor: Norbert Wiener at MIT. New York Review of Books, 52 (20):93, December 15, 2005. ISSN 0028-7504 (print), 1944-

7744 (electronic). URL http://www.nybooks.com/articles/
archives/2005/dec/15/norbert-wiener-at-mit/. In response

to [Dys05g].

Topper:2014:EB

[TD14] David Topper and Freeman Dyson. Einstein’s ‘blunder’. New York Review of Books, 61(8):??, May 8, 2014. ISSN 0028-7504 (print), 1944-7744 (electronic). URL http://www.nybooks.com/

articles/archives/2014/may/08/einsteins-blunder/. See

[Dys14a].

Teller:1948:CPC

[Tel48] Edward Teller. On the change of physical constants. Physical Review, 73(7):801–802, April 1, 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://prola.

aps.org/abstract/PR/v73/i7/p801_1.

Tenney:1969:LEW

[Ten69] Fred H. Tenney. Letter to the Editor: What does this mean? Bulletin of the Atomic Scientists, 25(9):45–46, November 1969. CO-DEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys69a].

Thebault:1947:APSg

[Thé47] Victor Thébault. Advanced problems and solutions: Problems for solution: 4244. American Mathematical Monthly, 54(4):232– 233, April 1947. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BTD^{+}47, BG49, GB48, RRS48, TL48, TB49].

Thebault:1950:APSj

[Thé50] Victor Thébault. Advanced problems and solutions: Problems for solution: 4386. American Mathematical Monthly, 57(3):188–189, March 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Bat51, DTS51, ES51, TK51b, UL51].

Thirring:1991:SMA

[Thi91] Walter E. Thirring, editor. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 3-540-53039-8 (Berlin), 0-387-53039-8 (New York). viii + 565 pp. LCCN QC173.4.T48 L54 1991. With a preface by F. Dyson.

Thirring:2001:SMA

[Thi01] Walter E. Thirring, editor. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., third edition, 2001. ISBN 3-540-42083-5. xiii + 812 pp. LCCN QC173.4.T48 L54 2001. With a preface by F. Dyson.

Thirring:2005:SMA

[Thi05] Walter E. Thirring, editor. The stability of matter: from atoms to stars: selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., fourth edition, 2005. ISBN 3-540-22212-X. xv + 932 pp. LCCN QC173.4.T48

L54 2005. URL http://www.loc.gov/catdir/enhancements/
fy0663/2004108033-d.html.

Tierney:2009:FDG

[Tie09a] John Tierney. Freeman Dyson’s 4th-grade math puzzle. The New York Times Magazine, ??(??):??, April 6, 2009. ISSN 0362-

1308. URL http://tierneylab.blogs.nytimes.com/2009/04/
06/freeman-dysons-4th-grade-math-puzzle/.

Tierney:2009:PAG

[Tie09b] John Tierney. Puzzle answers from a 4th-grader and Freeman Dyson. The New York Times Magazine, ??(??):??, April 10, 2009.

ISSN 0362-1308. URL http://tierneylab.blogs.nytimes.
com/2009/04/10/puzzle-answers-from-freeman-dyson-and-
a-fourth-grader/.

Tierney:2009:TFD

[Tie09c] John Tierney. Tragedy is not Freeman Dyson’s business. The New York Times Magazine, ??(??):??, March 30, 2009. ISSN 0362-

1308. URL http://tierneylab.blogs.nytimes.com/2009/03/
30/tragedy-is-not-freeman-dysons-business/.

Tippett:2010:EGC

[Tip10] Krista Tippett. Einstein’s God: conversations about science and the human spirit. Penguin Books, Harmondsworth, UK, 2010. ISBN 0-14-311677-0 (paperback), 1-4487-2130-X, 1-101-19541-X (e-book), 1-101-19583-5 (e-book). 286 pp. LCCN BL240.3 .T57 2010.

Thebault:1951:APSa

[TK51a] Victor Thébault and L. M. Kelly. Advanced problems and solutions: 4316. American Mathematical Monthly, 58(1):42–44, January 1951. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wil48].

Thebault:1951:APSi

[TK51b] Victor Thébault and L. M. Kelly. Advanced problems and solutions: 4386. American Mathematical Monthly, 58(9):637–638, November 1951. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UTB^{+}50].

Tahir-Kheli:1962:UGF

[TKH62] R. A. Tahir-Kheli and D. Ter Haar. Use of Green functions in the theory of ferromagnetism. II. Dyson spin waves. Physical Review, 127(??):95–??, July 1, 1962. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL http://link.

aps.org/doi/10.1103/PhysRev.127.95.

Thebault:1948:APSn

[TL48] Victor Thébault and Roger Lessard. Advanced problems and solutions: 4244. American Mathematical Monthly, 55(8):511–512, October 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Thé47].

Li:1984:APP

[tL84] Ching teh Li. Approximate projection of physical states in the Dyson boson description of nuclear collective motion. Physical Review C (Nuclear Physics), 29(??):2309–??, June 1, 1984. CO-DEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497.

URL http://link.aps.org/doi/10.1103/PhysRevC.29.2309.

Thirring:1997:SMA

[TL97] W. Thirring and Elliott H. Lieb, editors. The stability of matter: From atoms to stars: Selecta of Elliott H. Lieb. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1997. ISBN 3-540-61565-2. 675 pp. LCCN ????

URL http://www.gbv.de/dms/goettingen/215927141.pdf.

Tong:1996:ACBa

[TMD96a] W. Tong, R. C. Morrison, and O. W. Day. Analysis of chemical bonding in C_{2} using Dyson orbitals. International Journal of Quantum Chemistry, 60(7):411–??, 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-461X (electronic).

Tong:1996:ACBc

[TMD96b] W. Tong, R. C. Morrison, and O. W. Day, Jr. Analysis of chemical bonding in C_{2} using Dyson orbitals. International Journal of Quantum Chemistry. Quantum Chemistry Symposium, 30(??): 411–??, ???? 1996. CODEN IJQSDI. ISSN 0161-3642.

Tong:1996:ACBb

[TMD96c] Wei Tong, Robert C. Morrison, and Orville W. Day Jr. Analysis of chemical bonding in C_{2} using Dyson orbitals. International Journal of Quantum Chemistry, 60(7):1623–1631, ???? 1996. CODEN IJQCB2. ISSN 0020-7608 (print), 1097-

461X (electronic). URL http://www3.interscience.wiley.
com/cgi-bin/abstract?ID=60723.

Takagaki:2001:PWD

[TP01] Y. Takagaki and K. H. Ploog. Poisson and Wigner–Dyson distributions of conductance fluctuations in quantum cavities. Physical Review B: Condensed Matter and Materials Physics, 64(??): 245336, December 10, 2001. CODEN PRBMDO. ISSN 1098-

0121. URL http://link.aps.org/doi/10.1103/PhysRevB.64.
245336.

Treiman:1994:BRB

[Tre94] Sam Treiman. Book review: QED and the Men Who Made It: Dyson, Feynman, Schwinger and Tomonaga, by Silvan S. Schweber. Physics Today, 47(12):59–60, 1994. CODEN PHTOAD.

ISSN 0031-9228 (print), 1945-0699 (electronic). URL http://

link.aip.org/link/?PTO/47/59/1.

Ungar:1951:APS

[UL51] P. Ungar and Roger Lessard. Advanced problems and solutions: Solutions: 4385. American Mathematical Monthly, 58(8):573– 575, October 1951. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [UTB^{+}50].

Ullmann-Margalit:1986:PSI

[UM86] Edna Ullmann-Margalit, editor. The Prism of Science: The Israel Colloquium: Studies in History, Philosophy, and Sociology of Science. Volume 2, volume 95(2) of Boston Studies in the Philosophy of Science. D. Reidel, Dordrecht, The Netherlands; Boston, MA, USA; Lancaster, UK; Tokyo, Japan, 1986. ISBN 90-277-2160-2, 90-277-2161-0 (paperback), 94-009-4566-3 (e-book). ISSN 0068-0346. ix + 250 pp. LCCN Q174 .B67 vol. 95

Q175. URL http://www.springerlink.com/content/978-94-
009-4566-1;   https://link.springer.com/book/10.1007/
978-94-009-4566-1.

Ulam:1986:SCP

[URR86] Stanisław M. Ulam, Mark C. Reynolds, and Gian-Carlo Rota. Science, computers, and people: from the tree of mathematics. Birkhäuser Boston Inc., Cambridge, MA, USA, 1986. ISBN 0-8176-3276-X. xxii + 264 pp. LCCN QA7 .U431 1986; QA7 .U43 1986. From the tree of mathematics, With a preface by Martin Gardner, With an introduction by Françoise Ulam.

Ungar:1950:APS

[UTB^{+}50] P. Ungar, Victor Thebault, Paul Bateman, Paul Erdos, and F. J. Dyson. Advanced Problems and Solutions: Problems For Solution: 4385-4389. American Mathematical Monthly, 57(3):188–189, 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

vonHippel:1984:BRF

[vH84] Frank von Hippel. Book review: Freeman Dyson, Weapons and Hope. Physics Today, 37(11):105–106, November 1984. CODEN PHTOAD. ISSN 0031-9228 (print), 1945-0699 (electronic). URL

http://link.aip.org/link/?PTO/37/105/1.

Vincent:1989:SFT

[VHK89] C. M. Vincent, L. J. Henry, and G.-K. Kim. Six-fermion tests of Dyson boson mappings. Physical Review C (Nuclear Physics), 39(??):1583–??, April 1, 1989. CODEN PRVCAN. ISSN 0556-2813 (print), 1089-490X, 1538-4497. URL http://link.aps.

org/doi/10.1103/PhysRevC.39.1583.

Villars:1995:BRB

[Vil95] Felix M. H. Villars. Book review: QED and the Men Who Made It: Dyson, Feynman, Schwinger and Tomonaga, by Silvan S. Schweber. American Journal of Physics, 63(4):383–384, 1995. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (elec-

tronic). URL http://link.aip.org/link/?AJP/63/383/1.

vonNeumann:1998:CGF

[vN98] John von Neumann. Continuous geometry. Princeton Landmarks in Mathematics. Princeton University Press, Princeton, NJ, USA, 1998. ISBN 0-691-05893-8. xi + 299 pp. LCCN QA611 .V6 1998. US$24.95, UK£18.95. With a foreword by Israel Halperin. New edition.

Vogel:1999:PSE

[Vog99] Gretchen Vogel. Planetary systems: Expanding the habitable zone. Science, 286(5437):70–71, October 1, 1999. CODEN SCI-EAS. ISSN 0036-8075 (print), 1095-9203 (electronic). URL http:

//www.sciencemag.org/content/286/5437/70.full. From the

article: “At a meeting this summer [1999], theoretical physicist Freeman Dyson of the Institute for Advanced Study in Princeton, New Jersey, offered to bet $100 that the first extraterrestrial life would be found on an asteroid or even in a cloud of space dust, rather than on a planet. Smaller objects account for much of the solar system’s mass, and so have ‘simply so much more real estate’ for life to colonize, both at and below the surface.”.

VanNeck:2001:SCS

[VPW01] D. Van Neck, K. Peirs, and M. Waroquier. Self-consistent solution of Dyson’s equation up to second order for atomic systems. Journal of Chemical Physics, 115(1):15–25, 2001. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://

link.aip.org/link/?JCP/115/15/1.

Vernes:2008:TRD

[VWS08] A. Vernes, P. Weinberger, and L. Szunyogh. Time-resolved Dyson equations in the context of time-dependent density-functional theory: Extension to solid systems. Physical Review B: Condensed Matter and Materials Physics, 78(??):155129, October 29, 2008. CODEN PRBMDO. ISSN 1098-0121. URL http://link.

aps.org/doi/10.1103/PhysRevB.78.155129.

Wadia:1981:DSE

[Wad81] Spenta R. Wadia. Dyson–Schwinger equations approach to the large-N limit: Model systems and string representation of Yang– Mills theory. Physical Review D (Particles and Fields), 24(??): 970–??, August 15, 1981. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.24.970.

Walhout:1999:SRH

[Wal99] T. S. Walhout. Similarity renormalization, Hamiltonian flow equations, and Dyson’s intermediate representation. Physical Review D (Particles and Fields), 59(??):065009, February 12, 1999. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.aps.org/

doi/10.1103/PhysRevD.59.065009.

Wataghin:1948:FCE

[Wat48] Gleb Wataghin. On the formation of chemical elements inside the stars. Physical Review, 73(1):79, January 1948. CODEN PHRVAO. ISSN 0031-899X (print), 1536-6065 (electronic). URL

http://link.aps.org/doi/10.1103/PhysRev.73.79. See re-

marks in [Dys93a] about the relation of this work to [?], and the subsequent incorrect neglect of Wataghin’s work. See also related papers [?, ?, DW48].

Wayne:1995:BRB

[Way95] Andrew Wayne. Book review: QED and the Men Who Made It: Dyson, Feynman, Schwinger, and Tomonaga by Silvan S. Schweber. British Journal for the Philosophy of Science, 46(4):624–627, December 1995. CODEN BJPIA5. ISSN 0007-0882 (print), 1464-

3537 (electronic). URL http://www.jstor.org/stable/687905.

Wassermann:2015:MMB

[WD15] Henry Wassermann and Freeman Dyson. Murders & massacres from both sides. New York Review of Books, 62(11):??, June 25, 2015. ISSN 0028-7504 (print), 1944-7744 (electronic). Comments on [Dys15a].

Witten:1989:LEA

[WDW89] Edward Witten, Freeman J. Dyson, and Richard Wilson. Letters to the Editor: Appeals for Tayseer Aruri. Physics Today, 42(5): 15, 118, May 1989. CODEN PHTOAD. ISSN 0031-9228 (print),

1945-0699 (electronic). URL http://link.aip.org/link/?PTO/
42/15/1.

Wilansky:1950:APSa

[WG50] Albert Wilansky and S. H. Gould. Advanced problems and solutions: Solutions: 4315. American Mathematical Monthly, 57(4): 272, April 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Wil48].

Wheeler:1991:AAP

[Whe91] Gerald F. Wheeler. American Association of Physics Teachers 1991 Oersted Medalist: Freeman J. Dyson. American Journal of Physics, 59(6):490, 1991. CODEN AJPIAS. ISSN 0002-9505 (print), 1943-2909 (electronic). URL http://link.aip.

org/link/?AJP/59/490/1.

White:1980:CVP

[Whi80] Guy K. (Guy Kendall) White, editor. Changing views of the physical world, 1954–1979: 2nd Silver Jubilee Symposium: Australian Academy of science: (proceedings: Canberra, 27–28 March 1979), volume 2 of Silver jubilee symposium. Australian Academy of Science, Canberra, ACT, Australia, 1980. ISBN 0-85847-058-6 (v. 1), 0-85847-060-8 (v. 2), 0-85847-064-0 (v. 3). x + 141 pp. LCCN QC6.9 .C48 1980.

Wightman:1990:BRB

[Wig90] Arthur S. Wightman. Book review: Riemann, Topology, and Physics, by Michael Monastyrsky, Freeman J. Dyson, James King, Victoria King, R. O. Wells, Jr. SIAM Review, 32(1):192, March 1990. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (elec-

tronic). URL http://www.jstor.org/stable/2030418.

Wilansky:1948:APS

[Wil48] Albert Wilansky. Advanced problems and solutions: Problems for solution: 4315. American Mathematical Monthly, 55(9):586, November 1948. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [EW50, MS50, San50, TK51a, WG50].

Wilansky:1950:APSb

[Wil50] Albert Wilansky. Advanced problems and solutions: Problems for solution: 4397. American Mathematical Monthly, 57(5):342–343, May 1950. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [BS51, DC52, FZ52, Gal52, WL51].

Wilkinson:1958:DCN

[Wil58] D. H. Wilkinson. Do the ‘constants of nature’ change with time? Philosophical Magazine, 3(30):582–585, 1958. CODEN PHMAA4.

ISSN 0031-8086. URL http://www.tandfonline.com/doi/abs/
10.1080/14786435808565799.

Wilson:1962:PCD

[Wil62] Kenneth G. Wilson. Proof of a conjecture by Dyson. Journal of Mathematical Physics, 3(5):1040–1043, May 1962. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v3/i5/
p1040_s1.

White:1974:CDD

[WK74] A. R. White and Kang Kyungsik. Castillejo–Dalitz–Dyson zeros in the Pomeranchukon scattering amplitude. Physical Review D (Particles and Fields), 10(??):983–??, August 1, 1974. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.10.983.

Willow:2013:SES

[WKH13] Soohaeng Yoo Willow, Kwang S. Kim, and So Hirata. Stochastic evaluation of second-order Dyson self-energies. Journal of Chemical Physics, 138(16):164111, 2013. CODEN JCPSA6. ISSN 0021-9606 (print), 1089-7690 (electronic). URL http://link.aip.

org/link/?JCP/138/164111/1.

Wilansky:1951:APSc

[WL51] Albert Wilansky and G. Lumer. Advanced problems and solutions: Solutions: 4397. American Mathematical Monthly, 58(10): 706–708, December 1951. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). See also [Gal50].

Woolf:1980:SSP

[Woo80] Harry Woolf, editor. Some strangeness in the proportion: a centennial symposium to celebrate the achievements of Albert Einstein. Addison-Wesley, Reading, MA, USA, 1980. ISBN 0-201-09924-1. LCCN QC16.E5 S63.

Woodruff:1992:DEA

[Woo92] S. L. Woodruff. Dyson equation analysis of inertial-range turbulence. Physics of Fluids A: Fluid Dynamics (1989–1993), 4(5): 1077–1079, 1992. CODEN PFADEB. ISSN 0899-8213 (print),

2163-5013 (electronic). URL http://link.aip.org/link/?PFA/
4/1077/1.

Watson:2007:PDS

[WR07] P. Watson and H. Reinhardt. Propagator Dyson–Schwinger equations of Coulomb gauge Yang–Mills theory within the first order formalism. Physical Review D (Particles and Fields), 75 (??):045021, February 26, 2007. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.75.045021.

Watson:2010:CGG

[WR10] P. Watson and H. Reinhardt. Coulomb gauge ghost Dyson– Schwinger equation. Physical Review D (Particles and Fields), 82 (??):125010, December 8, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.82.125010.

Wuthrich:2011:GFD

[Wüt11] Adrian Wüthrich. The Genesis of Feynman Diagrams, volume 26 of Archimedes: New Studies in the History and Philosophy of Science and Technology. Springer Netherlands, Dordrecht, The Netherlands, 2011. ISBN 90-481-9227-7 (hardcover), 90-481-9228-5 (e-book). ISSN 1385-0180 (print), 2215-0064 (electronic). xvii + 208 pp. LCCN QC794.6.F4 W88 2010.

Wuthrich:2013:FSD

[Wüt13] Adrian Wüthrich. Feynman’s struggle and Dyson’s surprise: the development and early application of a new means of representation. In Katzir et al. [KLR13], pages 271–289. ISBN 3-8442-5134-

0. LCCN QC173.98. URL http://www.edition-open-access.
de/proceedings/5/.

Winkler:1981:SDE

[WYL81] Peter Winkler, Robert Yaris, and Ronald Lovett. Solution of Dyson’s equation employing Siegert boundary conditions. Physical Review A (Atomic, Molecular, and Optical Physics), 23(??): 1787–??, April 1, 1981. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-4446, 1538-4519. URL http://link.

aps.org/doi/10.1103/PhysRevA.23.1787.

Xue:1988:STS

[Xue88] She-Sheng Xue. String tension from the Schwinger–Dyson equation for a Wilson loop. Physical Review D (Particles and Fields), 37(??):493–??, January 15, 1988. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL

http://link.aps.org/doi/10.1103/PhysRevD.37.493.

Yang:1954:TBU

[Yan54] Chung-Tao Yang. On theorems of Borsuk–Ulam, Kakutani– Yamabe–Yujobo and Dyson, I. Annals of Mathematics, 60(2): 262–282, September 1954. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://www.jstor.org/

stable/1969632.

Yang:1955:TBU

[Yan55] Chung-Tao Yang. On theorems of Borsuk–Ulam, Kakutani– Yamabe–Yujobo and Dyson, II. Annals of Mathematics, 62(2): 271–283, September 1955. CODEN ANMAAH. ISSN 0003-486X (print), 1939-8980 (electronic). URL http://www.jstor.org/

stable/1969681.

Yanagi:2010:URG

[Yan10] Kenjiro Yanagi. Uncertainty relation on generalized Wigner– Yanase–Dyson skew information. Linear Algebra and its Applications, 433(8–10):1524–1532, December 15, 2010. CODEN LAA-PAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Yang:2013:SPI

[Yan13] Chen Ning Yang, editor. Selected papers II, with commentaries. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 2013. ISBN 981-4449-00-8 (hardcover), 981-4449-01-6 (paperback), 981-4449-02-4 (e-book). x +

34 pp. LCCN QC21.3. URL http://www.worldscientific.
com/worldscibooks/10.1142/8640.

Ye:1992:DEF

[Ye92] Zhen Ye. Dyson equation at finite temperature for the vibrational mode of an admolecule on a crystal surface. Physical Review B: Condensed Matter and Materials Physics, 46(??):2628–??, July 15, 1992. CODEN PRBMDO. ISSN 1098-0121. URL http://

link.aps.org/doi/10.1103/PhysRevB.46.2628.

Yeung:1980:FDS

[Yeu80] Wai Bong Yeung. Is the Feynman–Dyson series adequate for the asymptotic expansion of the Green’s function of the quantum mechanical anharmonic oscillator? Journal of Mathematical Physics, 21(3):477–485, March 1980. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v21/i3/p477_s1.

Yamazaki:2007:DOM

[YHKO07] Masakazu Yamazaki, Takuya Horio, Naoki Kishimoto, and Koichi Ohno. Determination of outer molecular orbitals by collisional ionization experiments and comparison with Hartree–Fock, Kohn–Sham, and Dyson orbitals. Physical Review A (Atomic, Molecular, and Optical Physics), 75(??):032721, March 28, 2007. CODEN PLRAAN. ISSN 1050-2947 (print), 1094-1622, 1538-

4446, 1538-4519. URL http://link.aps.org/doi/10.1103/
PhysRevA.75.032721.

York:1987:MWT

[Yor87] Herbert F. (Herbert Frank) York. Making weapons, talking peace: a physicist’s odyssey from Hiroshima to Geneva. Alfred P. Sloan Foundation series. Basic Books, New York, NY, USA, 1987. ISBN 0-465-04338-0. xiv + 359 + 8 pp. LCCN JX1974.7 .Y575 1987. US$22.95.

Yoshimura:1982:USD

[Yos82] Tetz Yoshimura. Unrenormalized Schwinger–Dyson equations and dynamical mass generation. Journal of Mathematical Physics, 23(5):822–829, May 1982. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:

//jmp.aip.org/resource/1/jmapaq/v23/i5/p822_s1.

Young:1969:LED

[You69] Elizabeth Young. Letter to the Editor: Dyson may be right. Bulletin of the Atomic Scientists, 25(9):46, November 1969. CODEN BASIAP. ISSN 0096-3402 (print), 1938-3282 (electronic). See [Dys69a].

Wu:1998:JTP

[yWHH98] Ta you Wu, J. P. (Jong-Ping) Hsu, and Leonardo Hsu, editors. JingShin Theoretical Physics Symposium in Honor of Professor Ta-You Wu. World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1998. ISBN 981-02-3372-8. LCCN QC20 .J554 1997.

Yuan:2010:DSE

[YxQCxL10] Wei Yuan, Si xue Qin, Huan Chen, and Yu xin Liu. Dyson– Schwinger equations with a parametrized metric. Physical Review D (Particles and Fields), 81(??):114022, June 14, 2010. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-

4500 (CD-ROM). URL http://link.aps.org/doi/10.1103/
PhysRevD.81.114022.

Zenghilis:2008:PDC

[ZDSN08] Dimitri Zenghilis, Freeman J. Dyson, Leigh Sullivan, and William D. Nordhaus. El problema del calentamiento global. (Spanish) [The problem of global warming]. Estudios Públicos, 112(??):83–118, Spring 2008. ISSN 0716-1115 (print), 0718-3089

(electronic). URL http://www.cepchile.cl/1_4298/doc/el_
problema_del_calentamiento_global.html;   http://www.
cepchile.cl/2_4298/doc/the_question_of_global_warming.
html;   http://www.cepchile.cl/dms/lang_2/cat_928_inicio.

html. Spanish translation of debate about global warming [Dys08d].

Zeilberger:1987:PCM

[Zei87] Doron Zeilberger. A proof of the G_{2} case of Macdonald’s root system-Dyson conjecture. SIAM Journal on Mathematical Analysis, 18(3):880–883, May 1987. CODEN SJMAAH. ISSN 0036-1410 (print), 1095-7154 (electronic).

Zuckerman:1995:EWT

[ZH95] Ben Zuckerman and Michael H. Hart, editors. Extraterrestrials — where are they? Cambridge University Press, Cambridge, UK, 1995. ISBN 0-521-44335-0 (hardback), 0-521-44803-4 (paperback). LCCN QB54 .E95 1995. URL http://www.loc.

gov/catdir/description/cam026/94043739.html;   http:/
/www.loc.gov/catdir/toc/cam021/94043739.html.

Zhou:1997:ESL

[Zho97] Bang-Rong Zhou. Exact solutions of linearized Schwinger–Dyson equation of fermion self-energy. Journal of Mathematical Physics, 38(2):809–820, February 1997. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Zwanziger:2003:TIS

[Zwa03] Daniel Zwanziger. Time-independent stochastic quantization, Dyson–Schwinger equations, and infrared critical exponents in QCD. Physical Review D (Particles and Fields), 67(??):105001, May 2, 2003. CODEN PRVDAQ. ISSN 0556-2821 (print), 1089-4918 (electronic), 1538-4500 (CD-ROM). URL http://link.

aps.org/doi/10.1103/PhysRevD.67.105001.